### v2

# Learning Safe Strategies for Value Maximizing Buyers in Uniform Price Auctions

Abstract
We study the bidding problem in repeated uniform price multi-unit auctions from the perspective of a single value-maximizing buyer who aims to maximize their cumulative value over $T$ rounds while adhering to return-on-investment (RoI) constraints in each round. Buyers adopt $m$-uniform bidding format, where they submit $m$ bid-quantity pairs $(b_{i},q_{i})$ to demand $q_{i}$ units at bid $b_{i}$. We introduce safe bidding strategies as those that satisfy RoI constraints in every auction, regardless of competing bids. We show that these strategies depend only on the bidder’s valuation curve, and the bidder can focus on a finite subset of this class without loss of generality. While the number of strategies in this subset is exponential in $m$, we develop a polynomial-time algorithm to learn the optimal safe strategy that achieves sublinear regret in the online setting, where regret is measured against a clairvoyant benchmark that knows the competing bids a priori and selects a fixed hindsight optimal safe strategy. We then evaluate the performance of safe strategies against a clairvoyant that selects the optimal strategy from a richer class of strategies in the online setting. In this scenario, we compute the richness ratio, $\alpha\in(0,1]$ for the class of strategies chosen by the clairvoyant and show that our algorithm, designed to learn safe strategies, achieves $\alpha$-approximate sublinear regret against these stronger benchmarks. Experiments on semi-synthetic data from real-world auctions show that safe strategies substantially outperform the derived theoretical bounds, making them quite appealing in practice.
1.  Introduction
In a uniform price multi-unit auction, the auctioneer sells $K$ identical units of a single good to buyers who may demand multiple units, with the per-unit price set at the $K^{th}$ highest bid. These auctions are widely used to allocate scarce resources in critical markets such as emissions permits, energy markets, and Treasury auctions. To succeed in these markets, bidders must develop effective bidding strategies that balance long-term value maximization with financial risk management, all while operating under limited information, uncertainty in competing bids and strategic (or adversarial) behavior of other bidders, among other challenges.
In this paper, we model the bidders as a value-maximizing agent with per-round return-on-investment (RoI) constraints. This model reflects real-world decision-making in industries where managing financial risks is as critical as maximizing value. Unlike traditional mechanism design, which assumes bidders with quasilinear utility (i.e., utility decreases linearly with payments), many practical settings involve bidders who optimize total value while adhering to financial constraints such as RoI or budgets. These scenarios frequently arise in industries where agents or algorithms bid on behalf of clients, optimizing for high-level objectives within strict constraints—a context akin to the principal-agent framework (Fadaei and Bichler, 2016; Aggarwal et al., 2024). Examples include autobidders in online advertising, where advertisers seek to maximize clicks while keeping the average cost per click below a threshold (Lucier et al., 2024; Balseiro et al., 2021a; Deng et al., 2023b), and consultants bidding for small firms in emission permit auctions (EEX, 2024).
Building on this, we study the bidding problem in repeated uniform price auctions over $T$ rounds, from the perspective of a single bidder who seeks to maximize cumulative value while adhering to per-round RoI constraints—that is, ensuring that the value obtained in each round is at least a fixed multiple of the corresponding payment (see Eq. 2).
To address practical bidding interfaces, we assume bidders adopt $m$-uniform bidding format, a generalization of the uniform bidding format (De Keijzer et al., 2013; Birmpas et al., 2019) for some $m\in\mathbb{N}$. In a $m$-uniform strategy $\mathbf{b}:=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$, bidders bid $b_{1}$ for the first $q_{1}$ units, $b_{2}$ for the next $q_{2}$ units and so on (see Definition 1).
1.1.  Our Contributions
Safe Bidding Strategies (Section 3). To ensure that the bidder satisfies the RoI constraint without knowing the competing bids, we introduce the concept of safe bidding strategies. These strategies guarantee that the RoI constraint is met regardless of how the other participants bid. We then identify the class of $m$-uniform safe undominated bidding strategies, denoted by $\mathscr{S}_{m}^{\star}$ (Theorem 3.2), and demonstrate that the strategies within this class depend solely on the bidder’s valuation vector $[v_{1},\dots,v_{K}]$ and exhibit a “nested” structure (illustrated in Fig. 1).
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$v_{9}$$v_{10}$${w_{3}}$${w_{7}}$${w_{9}}$
Figure 1: Nested structure of the bidding strategies in $\mathscr{S}_{3}^{\star}$. Consider the strategy $\mathbf{b}=\langle(w_{3},3),(w_{7},4),(w_{9},2)\rangle\in\mathscr{S}_{3}^{\star}$, where we note that $Q_{1}=3$, $Q_{2}=3+4=7$ and $Q_{3}=3+4+2=9$. The $j^{th}$ highest bid (i.e., $b_{j}$) is the average of the first $Q_{j}$ entries of the valuation vector, i.e., $b_{j}=w_{Q_{j}}$, where $w_{j}=\frac{1}{j}\sum_{\ell\leq j}v_{\ell}$.
Learning Safe Bidding Strategies (Section 4). Designing an algorithm to learn the optimal $m$-uniform safe bidding strategy poses significant challenges due to the exponential size of the decision space in $m$. To tackle this, we first consider an offline setting where the competing bids are known in advance. In this setting, we construct a polynomial-sized directed acyclic graph (DAG) with carefully assigned edge weights and show that determining the maximum weight path in the DAG is equivalent to computing the optimal $m$-uniform safe bidding strategy (Theorem 4.2).
Building on this, we develop an online learning algorithm when the competing bids are unknown. [^1] For a detailed discussion on leveraging offline algorithms for no-regret learning, we refer readers to Roughgarden and Wang (2019); Niazadeh et al. (2022); Brânzei et al. (2023). Our algorithm runs in polynomial time and achieves a regret of $\widetilde{O}(M\sqrt{mT})$ in a full-information setting and $\widetilde{O}(M^{3/2}m^{3/2}\sqrt{T})$ regret in a bandit setting, where the clairvoyant selects the fixed hindsight optimal safe strategy (Theorem 4.3). Additionally, we establish a regret lower bound of $O(M\sqrt{T})$ (Theorem 4.4).
The problem of learning to bid in multi-unit uniform price and pay-as-bid auctions has been explored recently by Brânzei et al. (2023); Galgana and Golrezaei (2024); Potfer et al. (2024), assuming the bidders are quasilinear utility maximizers. Our work differs from them in two key ways: (a) we consider value-maximizing buyers with RoI constraints, a fundamentally different behavioral model, and (b) from a technical perspective, unlike prior works that require bid spaces to be discretized, our approach does not, due to the structure of safe strategies. Consequently, the time complexity becomes independent of $T$ in both online and offline settings  [^2] Assuming the discretization level in the offline setting is same as that in the online setting. In the online setting, time complexity refers to per-round running time., and under bandit feedback, the dependence on $T$ in the regret bound improves from $T^{2/3}$ to $\sqrt{T}$ implying that it is easier to learn safe strategies for value maximizers compared to bidding strategies for quasilinear utility maximizers.
Richer Classes of Strategies for the Clairvoyant (Section 5). A key novel aspect of our work is to evaluate the performance of safe bidding strategies with at most $m$ bid-quantity pairs and the online learning algorithm from Section 4 against a clairvoyant that selects the optimal strategy from a richer class of strategies. In this case, the bidder (learner) achieves $\alpha$-approximate sublinear regret (same as Theorem 4.3) for a parameter $\alpha\in(0,1]$—defined as the richness ratio—thereby demonstrating the robustness of the safe bidding strategies against a broad range of benchmarks (Corollary 5.2). One of our main contributions is to compute the richness ratio, $\alpha$, for various classes of bidding strategies. Specifically, when the clairvoyant selects the optimal bidding from the class of
(a) strategies that are RoI-feasible, not necessarily safe and have at most $m$ bid-quantity pairs, we get $\alpha=\frac{1}{2}$, notably independent of $m$ (Theorem 5.3).
(b) safe strategies with at most $m^{\prime}$ bid-quantity pairs, where $m^{\prime}\geq m$, we obtain $\alpha=\frac{m}{m^{\prime}}$ (Theorem 5.4).
(c) strategies that are RoI-feasible, not necessarily safe and have at most $m^{\prime}$ bid-quantity pairs, where $m^{\prime}\geq m$, we get that $\alpha=\frac{m}{2m^{\prime}}$ (Theorem 5.5).
Computing $\alpha$ involves two key challenges: (a) deriving an upper bound on the ratio between the value obtained by the clairvoyant’s optimal strategy and that of the optimal safe strategy in the worst case (see Definition 4 for details), and (b) proving that this upper bound is tight. The key ideas for deriving the upper bound are outlined in Section 5. The construction of problem instances, each tailored to a specific strategy class, is non-trivial due to their exponential size in $m$ and the need for a careful choice of competing bids and the valuation vector, as detailed in Sections D.1 to D.4. Simulations using semi-synthetic data show that the empirical richness ratios in practical scenarios are significantly better than the theoretical bounds, making safe strategies highly appealing in practice (Section 5.3).
1.2.  Related Work
Value Maximizers and RoI Constraints. The concept of agents as value maximizers within financial constraints is a well-established notion in microeconomic theory (Mas-Colell et al., 1995, pp. 50). In mechanism design literature, one of the earliest explorations of value-maximizing agents was conducted by Wilkens et al. (2016). Their work primarily delved into the single-parameter setting, characterizing truthful auctions for value maximizers. Similarly, Fadaei and Bichler (2016) and Lu et al. (2023) studied truthful (approximate) revenue-maximizing mechanisms in combinatorial markets tailored for such agents.
In recent years, there has been growing interest in RoI-constrained value maximizers, particularly in the context of autobidding and online advertising. One of the earliest works in this area is Golrezaei et al. (2021), which studied auction design for RoI-constrained buyers and validated the presence of such soft financial constraints using data from online ad auctions. Broadly, the literature in this space can be divided into two categories: (i) works that design optimal auctions under RoI constraints (Balseiro et al., 2021a, b, 2022; Deng et al., 2021, 2023b), and (ii) works that characterize optimal bidding strategies and/or develop learning algorithms in repeated auction settings (Aggarwal et al., 2019; Deng et al., 2023a; Golrezaei et al., 2023; Castiglioni et al., 2024; Aggarwal et al., 2025; Lucier et al., 2024). Our work primarily aligns with the latter. We study value-maximizing buyers in uniform price auctions under RoI constraints and demonstrate that these buyers can employ safe bidding strategies that are efficiently learnable and robust against a variety of strong benchmarks.
Multi-unit Auctions. In this work, we focus on a subset of combinatorial auctions termed as multi-unit auctions in which multiple identical goods are sold to a group of buyers. [^3] For a comprehensive survey on combinatorial auctions, we refer the readers to several excellent works by De Vries and Vohra (2003); Blumrosen and Nisan (2007); Palacios-Huerta et al. (2022). These auctions find widespread application in various practical scenarios, including Treasury auctions (Hortaçsu and McAdams, 2010), procurement auctions (Cramton and Ausubel, 2006), electricity markets (Tierney et al., 2008), and emissions permit auctions (Schmalensee and Stavins, 2017). While several works have focused on studying the equlibria properties in these auctions (Ausubel et al., 2014; Markakis and Telelis, 2015; De Keijzer et al., 2013), computing equilibrium strategies is usually intractable in general multi-unit auctions due to multi-dimensional valuations. As a result, a growing body of work has emphasized on designing optimal bidding strategies in a prior-free setting, i.e., without any assumptions of the competing bidders’ bids or valuations (Galgana and Golrezaei, 2024; Potfer et al., 2024; Brânzei et al., 2023). Our work contributes to this literature by learning optimal strategies for value-maximizing buyers in repeated uniform price auctions.
2.  Model
Preliminaries. There are $n$ buyers (bidders) indexed by $i\in[n]$, and $K$ identical units of a single good. Each bidder $i$ has a fixed, private valuation curve, denoted by $\mathbf{v}_{i}\in\mathbb{R}^{K}_{+}$ that has diminishing marginal returns property, i.e., $v_{i,1}\geq v_{i,2}\geq\dots\geq v_{i,K}$ which is standard in literature (Brânzei et al., 2023; Goldner et al., 2020). As we study optimal bidding strategies from the perspective of a single bidder, we drop the index $i$ when the context is clear. The maximum total demand for bidder $i$, denoted by $M\in[K]$, is defined as $\min\{j\in[K-1]:v_{j+1}=0\}$. If such an index does not exist, we set $M$ as $K$. Hence, without loss of generality, we assume $\mathbf{v}\in\mathbb{R}_{+}^{M}$. For each $\mathbf{v}=[v_{1},\dots,v_{M}]$, we define the average cumulative valuation vector as $\mathbf{w}=[w_{1},\dots,w_{M}]$,
$\text{where}\quad w_{j}$ $=\frac{1}{j}\sum_{\ell\leq j}v_{\ell},\forall j\in[M]\,.$ (1)
As $v_{1}\geq\dots\geq v_{M}$, we also have $w_{1}\geq\dots\geq w_{M}$.
2.1.  Auction Format and Bidders’ Behavior
Allocation and Payment Rule. In a uniform price auction, each bidder $i$ submits a sorted bid vector $\mathbf{b}$ using $m$-uniform bidding language (see Definition 1). The auctioneer collects the bids (entries of the bid vector) from all the bidders, sorts them in non-increasing order, and allocates units to the bidders with the top $K$ bids (also termed as ‘winning’ bids). That is, if bidder $i$ has $j$ bids in the top $K$ positions, they are allocated $j$ units. For ease of exposition, we assume there are no ties (or ties are always broken in the favor of the bidder in consideration). See Appendix F for our discussion on handling ties. We assume the auction follows the last-accepted-bid payment rule (bidders pay the $K^{th}$ highest bid per unit) which is widely used for uniform price auctions in practice (Regulations, 2019; Garbade and Ingber, 2005).
Let $\bm{\beta}_{-}$ denote the bids submitted by all bidders except bidder $i$ and $\bm{\beta}_{-}^{-(j)}$ be the $j^{th}$ smallest among the top $K$ competing bids (i.e., bids from all bidders except bidder $i$). Suppose bidder $i$ submits $\mathbf{b}$, such that the bid profile is $\bm{\beta}:=(\mathbf{b};\bm{\beta}_{-})$. Let $x(\bm{\beta})$ and $p(\bm{\beta})$ denote the number of units allocated to bidder $i$ and the clearing price (i.e., the per-unit price which is the $K^{th}$ highest submitted bid), respectively. The total value obtained by the bidder is $V(\bm{\beta})=\sum_{j\leq x(\bm{\beta})}v_{j},$ while the total payment made is $P(\bm{\beta})=p(\bm{\beta})\cdot x(\bm{\beta}).$
Bidding Language. Multi-unit auctions allocate a large number of identical units, requiring efficient ways for the bidders to express preferences. A common approach is standard bidding, where bidders submit a vector of bids, one for each unit (Brânzei et al., 2023; Galgana and Golrezaei, 2024; Babaioff et al., 2023; Birmpas et al., 2019; Potfer et al., 2024). Although expressive, this becomes computationally impractical when the number of units, $K$, is large, as in EU ETS emission permit auctions and Treasury auctions. To address this, we consider a bidding language called $m$-uniform bidding, for any $m\in\mathbb{N}$ [^4] This format generalizes the uniform bidding format (De Keijzer et al., 2013; Birmpas et al., 2019) and aligns with practical languages like those in product-mix auctions (Klemperer, 2009) and piecewise-linear bidding (Nisan, 2015). In $m$-uniform bidding, bidders submit $m$ bid-quantity pairs $(b_{i},q_{i})$, where $b_{i}$ is the bid value per unit and $q_{i}$ is the quantity demanded:
Definition 1 ($m$-Uniform Bidding).
For a fixed $m\in\mathbb{N}$, a $m$-uniform bidding strategy is characterized by $m$ bid-quantity pairs, denoted as $\mathbf{b}:=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$, where $b_{1}>b_{2}>\dots>b_{m}>0$ and $q_{j}>0$, $j\in[m]$. This $m$-uniform bid can be equivalently expressed as a vector (similar to the standard bidding format) in which the first $q_{1}$ bids are $b_{1}$, followed by $q_{2}$ bids of $b_{2}$, and so on.
We define $\mathbf{b}[1:\ell]=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell})\rangle$, for all $\ell<m$, to represent the first $\ell$ bid-quantity pairs within a $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$. We further define $Q_{j}=\sum_{\ell=1}^{j}q_{\ell}$ for all $j\in[m]$ as the total quantity demanded in the first $j$ bid-quantity pairs, with $Q_{0}=0$, where we assume, without loss of generality that $Q_{m}\leq M$ [^5] Suppose the bidder bids for, and wins more than $M$ units. There is no additional value being allocated over $M$ units, but the total payment increases (assuming the clearing price is positive), potentially violating the RoI constraint. If $m=M$, the bidding format is equivalent to standard bidding but in practice, bidders often submit only a few bid-quantity pairs. For instance, in the 2023 EU ETS auctions, bidders submitted $\sim 4.35$ bid-quantity pairs per auction on average (EEX, 2023).
Bidders’ Behavior. The bidders maximize their total value obtained while adhering to a constraint that ensures the total value obtained in an auction is at least a constant multiple of the payment in that auction. This can be equivalently expressed as a return-on-investment (RoI) constraint:
$V(\mathbf{b};\bm{\beta}_{-})\geq(1+\gamma)P(\mathbf{b};\bm{\beta}_{-})\,.$ (2)
Here, $\gamma$ is defined as the target RoI which is private and fixed. Without loss of generality, we assume $\gamma=0$ (or equivalently the valuation curve $\mathbf{v}$ is scaled by $\frac{1}{1+\gamma}$) for the rest of this work. For $\gamma=0$, the RoI constraint implies the value obtained in an auction is at least the payment. We illustrate how the auction operates in Example 1.
Example 1.
Consider an auction with $n=2$ bidders, and $K=5$ identical units. The valuations are: $\mathbf{v}_{1}=[6,4,3,1,1]$ and $\mathbf{v}_{2}=[5,3,1,1,0]$. Both the bidders are value maximizing buyers with target RoI, $\gamma_{1}=\gamma_{2}=0$. Suppose $m=2$ and the bids submitted by the bidders are $\mathbf{b}_{1}=\langle(5,2),(3,3)\rangle$ and $\mathbf{b}_{2}=\langle(4,2),(2,2)\rangle$. The bid profile: $\bm{\beta}=[5,5,4,4,3,3,3,2,2]$ and top $K=5$ winning bids are $[\underline{5},\underline{5},4,4,\underline{3}]$. Bidder 1 is allocated $3$ units as they have $3$ bids (underlined) in the winning bids, and bidder 2 gets the remaining $2$ units. The clearing price $p(\bm{\beta})=3$, $V_{1}(\bm{\beta})=6+4+3=13,V_{2}(\bm{\beta})=5+3=8,P_{1}(\bm{\beta})=3\cdot 3=9$, and $P_{2}(\bm{\beta})=2\cdot 3=6$. The RoI constraint is satisfied for both the bidders as $13>9$ and $8>6$.
2.2.  Learning to Bid in Repeated Settings
In practice, most multi-unit auctions, such as emission permit auctions and Treasury auctions, are conducted in a repeated setting. Formally, the auction described in the previous section takes place sequentially over $T$ rounds indexed by $t\in[T]$. We assume that the valuation vector $\mathbf{v}$ is fixed over the $T$ rounds which is a standard assumption in the literature (Brânzei et al., 2023; Galgana and Golrezaei, 2024; Potfer et al., 2024). Similarly, the target RoI is also assumed to be fixed over the $T$ rounds.
We now extend the notations from the previous section to the repeated setting. Formally, in this setting, $\bm{\beta}_{-}^{t}$ denotes the bids submitted by all bidders except bidder $i$ in round $t$ and $\bm{\beta}_{-,t}^{-(j)}$ is the $j^{th}$ smallest among the top $K$ competing bids in round $t$. In round $t$, if bidder $i$ submits a bid $\mathbf{b}^{t}$, the bid profile is $\bm{\beta}^{t}:=(\mathbf{b}^{t};\bm{\beta}_{-}^{t})$. Let $x(\bm{\beta}^{t})$ and $p(\bm{\beta}^{t})$ denote the number of units allocated to bidder $i$ and the clearing price, respectively, in round $t$. The total value obtained by the bidder is $V(\bm{\beta}^{t})=\sum_{j\leq x(\bm{\beta}^{t})}v_{j},$ while the total payment made is $P(\bm{\beta}^{t})=p(\bm{\beta}^{t})\cdot x(\bm{\beta}^{t}).$
RoI Constraints in Repeated Setting. In the repeated setting, we require the bidders to satisfy the RoI constraint described in Eq. 2 in every round. A bidding strategy $\mathbf{b}$ is called feasible for a sequence of competing bids $[\bm{\beta}_{-}^{t}]_{t\in[T]}$, if the RoI constraint is satisfied for every round $t\in[T]$.
Remark 2.1 (RoI Constraints).
Our notion of RoI constraints in the repeated setting aligns with the definitions in Wilkens et al. (2016, 2017); Lv et al. (2023). Similar constraints are considered by Lucier et al. (2024) and Gaitonde et al. (2023), termed as marginal RoI (or value) and maximum bid constraints (up to scaling factors), respectively. Unlike the aggregate constraints over $T$ rounds typically assumed in online ad auction literature (Deng et al., 2021, 2024; Feng et al., 2023; Deng et al., 2023a), we enforce RoI constraints for each auction individually. This distinction reflects the fundamental differences between the two settings: ad auctions often occur simultaneously and frequently, with values in the order of cents, whereas Treasury and emission permit auctions occur sequentially over longer horizons, with units valued in millions. Hence, bidders in such auctions prioritize profitability in each auction rather than waiting for an indefinite period, leading to per-round RoI constraints. [^6] Although EU ETS emission permit auctions are scheduled to occur regularly, regulations stipulate that an auction may be canceled if the bidders’ demand falls short of the supply of permits or if the clearing price of the auction does not meet the reserve price (Regulations, 2019). Hence, the bidders are more likely to ensure RoI feasibility in each round.
2.3.  Objective and Performance Metric
We consider an online setting where a bidder seeks to maximize cumulative value while satisfying RoI constraints and managing uncertainty about competing bids. In this setting, bidders privately submit their bids, and the auctioneer allocates items and sets prices based on these bids. In round $t$, the learning algorithm maps the feedback information set from the first $t-1$ rounds, denoted by $\mathcal{I}^{t-1}$, to a bid $\mathbf{b}^{t}$, where this mapping may be deterministic or random. Under full information feedback, the information set $\mathcal{I}^{t-1}=(\bm{\beta}^{1},\bm{\beta}^{2},\ldots,\bm{\beta}^{t-1}),$ which includes all bids from previous rounds. In the bandit setting, the information set $\mathcal{I}^{t-1}=(V(\bm{\beta}^{1}),P(\bm{\beta}^{1}),\ldots,V(\bm{\beta}^{t-1}),P(\bm{\beta}^{t-1}))$ includes only the value and price outcomes of the previous rounds.
We design learning algorithms to minimize regret, the difference in the value obtained by a fixed hindsight-optimal strategy chosen by a clairvoyant with a priori knowledge of competing bids and that by the learner over time. Formally,
$\mathsf{REG}=\max_{\mathbf{b}\in\mathscr{B}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})],$ (3)
where the expectation is with respect to any randomness in the learning algorithm. Here, $\mathscr{B}$ is the class of bidding strategies (formally characterized in Section 3). We require the bidding strategies in $\mathscr{B}$ to be RoI feasible for the sequence of competing bids $[\bm{\beta}_{-}^{t}]_{t\in[T]}$. In Section 4, we consider the case when the both the clairvoyant and the learner choose strategies from the same class and aim to obtain sublinear regret. Later, we consider more challenging settings where the clairvoyant can choose the optimal strategy from much richer classes of strategies compared to the learner. In these cases, we obtain sublinear $\alpha$-approximate regret, where $\alpha\in(0,1]$ is the richness ratio (see details in Section 5).
3.  Safe Bidding Strategies
Recall that strategies chosen by the learner must be causal (mapping the history to a bidding strategy) while remaining RoI feasible for all rounds $t\in[T]$, even under adversarially generated competing bids. This creates an obvious challenge: a bidding strategy that satisfies RoI feasibility in previous rounds may become infeasible even under the slightest change in the competing bids e.g., if $\mathbf{v}=[0.9,0.5,0.1]$, the bidding strategy $(0.6,3)$ is RoI feasible for competing bids $\bm{\beta}_{-}^{1}=[0.61,0.59,0.59]$ but not for $\bm{\beta}_{-}^{2}=[0.59,0.59,0.59]$. To address this, we focus on safe bidding strategies that inherently guarantee RoI feasibility, regardless of the adversarial behavior of competing bids:
Definition 2 (Safe Strategies).
A $m$-uniform bidding strategy, $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$, is called a safe strategy if it is feasible irrespective of the competing bids, i.e.,
$V(\mathbf{b};\bm{\beta}_{-})\geq P({\mathbf{b}};\bm{\beta}_{-}),\quad\forall\bm{\beta}_{-}\,.$
The class of all $m$-uniform safe bidding strategies is denoted as $\mathscr{S}_{m}$. The union of classes of safe bidding strategies with at most $m$ bid-quantity pairs is denoted as $\mathscr{U}_{m}=\bigcup_{k\in[m]}\mathscr{S}_{k}$.
3.1.  Characterizing Safe Bidding Strategies
We begin by defining underbidding (overbidding) under the $m$-uniform bidding format in the given context. Recall that in a single-item auction, if the value of the item is $v$, then a bid $b$ is an underbid (overbid) if $b<v~{}(b>v)$.
Definition 3 (Underbid and Overbid).
A $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an
(a) underbidding strategy if $b_{j}\leq w_{Q_{j}},\forall j\in[m]$ and $\exists\ell\in[m]$ such that $b_{\ell}<w_{Q_{\ell}}$.
(b) overbidding strategy if $\exists\ell\in[m]$ such that $b_{\ell}>w_{Q_{\ell}}$.
Figure 2: The solid line represents the average cumulative valuation curve, $\mathbf{w}$, and the dotted line represents the valuation curve, $\mathbf{v}$. The figure in the left (resp. right) illustrates underbidding (resp. overbidding) for a $2$-uniform bidding strategy. Note that the notions of underbidding and overbidding in Definition 3 are defined with respect to $\mathbf{w}$ and not $\mathbf{v}$ [^8] In Definition 3 (and throughout the paper), we assume $\gamma=0$ without loss of generality for ease of exposition. Footnote 8 illustrates how different values of $\gamma$ impacts the results. Specifically, for any $\gamma\geq 0$, underbidding and overbidding are defined with respect to the scaled average cumulative valuation curve, $\frac{\mathbf{w}}{1+\gamma}$.Here, the plots of $\mathbf{v}$ and $\mathbf{w}$ are shown to be linear for illustrative purposes only.
In other words, if $\gamma=0$, $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an underbidding strategy if, for all $j\in[m]$, the bid $b_{j}$ is at most the average of the first $Q_{j}$ entries of the valuation vector, denoted as $w_{Q_{j}}$, and there exists $\ell\in[m]$ where the inequality is strict. Recall that $Q_{j}=\sum_{\ell\leq j}q_{\ell}$ denotes the maximum number of demanded units in the first $j$ bid-quantity pairs. Similarly, $\mathbf{b}$ is an overbidding strategy if there exists some $\ell\in[m]$ such that $b_{\ell}$ is strictly greater than the average of the first $Q_{\ell}$ entries of the valuation vector. Having defined the notions of overbidding and underbidding, we characterize the class of safe bidding strategies, $\mathscr{S}_{m}$:
Theorem 3.1.
For any $m\in\mathbb{N}$, no overbidding is allowed in $\mathscr{S}_{m}$. So, the collection of all $m$-uniform safe strategies is
$\mathscr{S}_{m}=\Big{\{}\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}\leq w_{Q_{\ell}},\forall\ell\in[m]\Big{\}}\,,$
where $\bf w$ is defined in Eq. (1).
The proofs of this section are presented in Appendix A. Since there are infinitely many safe strategies in $\mathscr{S}_{m}$, we focus on the subset of safe undominated strategies, $\mathscr{S}_{m}^{\star}$, obtained by removing very weakly dominated strategies. A safe strategy $\mathbf{b}$ is said to be very weakly dominated if there exists another safe strategy $\mathbf{b}^{\prime}$ such that, for any competing bid $\bm{\beta}_{-}$, $V(\mathbf{b},\bm{\beta}_{-})\leq V(\mathbf{b}^{\prime},\bm{\beta}_{-})$ (Shoham and Leyton-Brown, 2008, pp. 79). Keeping these strategies does not improve the bidder’s performance, but removing them yields a finite strategy class (as shown below), which significantly simplifies the design of online learning algorithms. We now present the main result of this section:
Theorem 3.2.
The class of $m$-uniform safe undominated bidding strategies, $\mathscr{S}_{m}^{\star}$, is given as follows:
$\mathscr{S}_{m}^{\star}=\Big{\{}\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}},\ell\in[m]\Big{\}}\,.$
The union of classes of safe undominated strategies with at most $m$ bid-quantity pairs is denoted by $\mathscr{U}_{m}^{\star}=\bigcup_{k\in[m]}\mathscr{S}_{k}^{\star}$.
Observe that $\mathscr{S}_{m}^{\star}$ is a finite class of bidding strategies that depend only the valuation curve and is independent of the competing bids. The strategies in $\mathscr{S}_{m}^{\star}$ exhibit a “nested” structure, referred to as nested $m$-uniform bidding strategies hereafter, where the $j^{th}$ highest bid ($b_{j}$) is the average of the first $Q_{j}=\sum_{\ell=1}^{j}q_{\ell}$ entries of the valuation curve (illustrated in Fig. 1). Importantly, fixing $Q_{j}$’s uniquely determines the bidding strategy, as $b_{j}=w_{Q_{j}}$ for strategies in $\mathscr{S}_{m}^{\star}$ which will be crucial in learning the optimal safe bidding strategy, as discussed in the following section.
4.  Learning Safe Bidding Strategies
Here, the clairvoyant and the learner choose strategies from the class of safe bidding strategies with at most $m$ bid-quantity pairs, $\mathscr{U}_{m}^{\star}$. Thus, the regret defined in Eq. 3 is
$\mathsf{REG}=\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\,.$
To design learning algorithms, we first consider the offline setting, where we aim to solve the following problem given the bid history, $[\bm{\beta}_{-}^{t}]_{t\in[T]}$.
$\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (Offline)
Since $\mathscr{U}_{m}^{\star}$ contains $O(M^{m})$ strategies, evaluating each strategy individually is intractable. To overcome this, we reduce the offline problem to finding the maximum weight path in a carefully constructed edge-weighted directed acyclic graph (DAG), which allows for the computation of the optimal offline bidding strategy in poly($m,M$) time.
4.1.  Offline Problem
Fix a bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, and consider (Offline). The following lemma shows that the objective function can be decomposed across the bid-quantity pairs. Formally,
Lemma 4.1.
Problem (Offline) can be formulated as
$\max_{\ell\in[m]}\max_{Q_{1},\dots,Q_{\ell}}\sum_{j=1}^{\ell}\sum_{t=1}^{T}\sum_{k=Q_{j-1}+1}^{Q_{j}}v_{k}\cdot\mathbb{I}{\left[w_{Q_{j}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$
The proofs of this section are presented in Appendix B. Building on the decomposition in Lemma 4.1, we now construct an edge-weighted DAG $\mathcal{G}(V,E)$ as proposed earlier.
Vertices. The DAG has a ‘layered’ structure with a source ($s$), destination ($d$) and $m$ intermediate layers each with $M$ nodes, denoted by $(\ell,j)$ where $\ell\in[m],j\in[M]$ as shown in Fig. 3. For convenience, we set $s=(0,0)$.
Edges and Edge Weights. A directed edge exists from vertex $x$ to $y$ under the following conditions:
(i) $x=(\ell-1,j)$ and $y=(\ell,j^{\prime})$, $\forall\ell\in[m]$ and $j<j^{\prime}$. This edge connects consecutive layers in the graph, and its weight is given by:
$\mathsf{w}(e)=\sum_{t=1}^{T}\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$ (4)
(ii) $x=(\ell,j)$ and $y=d$, $\forall\ell\in[m],j\in[M]$. This edge connects the current layer to the destination node $d$, and its weight is $\mathsf{w}(e)=0$.
Figure 3: In this DAG, $M=3$, $m=2$. The red path refers to $\mathbf{b}=\langle(w_{1},1),(w_{3},2)\rangle\in\mathscr{U}_{2}^{\star}.$ The values in red are the corresponding $Q_{j}$’s. Similarly, the blue path refers to $\mathbf{b}=(w_{3},3)\in\mathscr{U}_{2}^{\star}$.
Theorem 4.2.
  • There exists a bijective mapping between the $s$-$d$ paths in $\mathcal{G}(V,E)$ and bidding strategies in $\mathscr{U}_{m}^{\star}$, i.e., the path $s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for $k\in[m]$ refers to the strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
    $b_{\ell}=w_{z_{\ell}}\quad\text{and}\quad q_{\ell}=z_{\ell}-z_{\ell-1},\forall\ell\in[k]\,.$
    where $w_{z_{\ell}}$ is defined as per Eq. 1, $\forall\ell\in[k]$. Conversely, the strategy $\mathbf{b}=\langle(w_{Q_{1}},q_{1}),\dots,(w_{Q_{k}},q_{k})\rangle$ maps to the path $s\to(1,Q_{1})\to\dots\to(k,Q_{k})\to d$.
  • The value obtained by a bidding strategy is the weight of the corresponding $s$-$d$ path. Thus, the Problem (Offline) is equivalent to finding the maximum weight $s$-$d$ path in $\mathcal{G}(V,E)$ which can be computed in $\text{poly}(m,M)$ time.
4.2.  Online Setting
In the online setting, in each round $t\in[T]$, the learner submits a strategy $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}$ and receives feedback from the auction. We consider two feedback models: (a) full information setting and (b) bandit setting, with the latter discussed in Section B.3.1. Leveraging the DAG formulation, we design an online learning algorithm such that $\mathsf{REG}=o(T)$. Without loss of generaity, assume $v_{j}\in[0,1],\forall j\in[M]$.
In the full information setting, after receiving feedback at the end of round $t$, a DAG $\mathcal{G}^{t}(V,E)$ is constructed as described earlier. The edge weights are set as follows:
(i) If $x=(\ell-1,j)$ and $y=(\ell,j^{\prime})$ with $\ell\in[m]$ and $j<j^{\prime}$,
$\mathsf{w}^{t}(e)=\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$ (5)
(ii) If $x=(\ell,j)$ and $y=d$, then $\mathsf{w}^{t}(e)=0,\forall\ell\in[m],j\in[M]$.
The online algorithm in Algorithm 1 consists of three main steps: (1) $\mathsf{UPDATE}$, (2) $\mathsf{SAMPLE}$, and (3) $\mathsf{MAP}$. In the first two steps, the algorithm maintains and updates the probabilities over the edges in the constructed DAG and selects a $s$-$d$ path in the DAG by sampling from these probabilities. These steps implement exponential weight updates using the weight-pushing method proposed by Takimoto and Warmuth (2003). Naively implementing exponential weight updates by treating each $s$-$d$ path (safe strategy) as an expert is intractable as there are $O(M^{m})$ such experts. However, following the decomposition in Lemma 4.1, Algorithm 1 only needs to maintain weights for individual edges and then sample a path based on these weights, avoiding the need to track weights for each path separately. The $\mathsf{MAP}$ step is the key component of the algorithm that maps the sampled path to a safe strategy (by Theorem 4.2).
Algorithm 1 Compute Bidding Strategies (Full Information)
0:  valuation curve $\mathbf{v}$, time horizon $T$, learning rate $\eta>0$.
1:  for $t=1,2,\dots,T$ do
2:     Construct $\mathcal{G}^{t}(V,E)$ similar to $\mathcal{G}(V,E)$ without weights.
3:     $\mathsf{UPDATE}:$ Obtain edge probabilities $\varphi^{t}(\cdot)$ by Algorithm 2. $\mathsf{SAMPLE}$: Define initial node $u=s$ and path $\mathfrak{p}=s$.
4:     while $u\neq d$ do
5:        Sample $v$ with probability $\varphi^{t}(u\to v)$.
6:        Append $v$ to the path $\mathfrak{p}$; set $u\leftarrow v$.
7:     end while
8:     $\mathsf{MAP}$: If $\mathfrak{p}=s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for some $k\in[m]$, submit $\mathbf{b}^{t}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
$b_{\ell}=w_{z_{\ell}}\quad\text{and}\quad q_{\ell}=z_{\ell}-z_{\ell-1},~{}\forall\ell\in[k]\,.$
where $w_{z_{\ell}}$ is defined as per Eq. 1 for all $\ell\in[k]$.
9:     The bidder observes $\bm{\beta}_{-}^{t}$ and sets $\mathsf{w}^{t}(e)$ as per Eq. 5.
10:  end for
Theorem 4.3.
Algorithm 1 runs in $\text{poly}(m,M)$ time per round and achieves $\mathsf{REG}\leq O(M\sqrt{mT\log M})$ under full information feedback, and $\mathsf{REG}\leq O(M\sqrt{m^{3}MT\log M})$ in the bandit setting.
Algorithm 2 Update Edge Probabilities
0:  For $t\geq 1$, $\mathcal{G}^{t}(V,E)$. For $t\geq 2$, additionally require $\eta$, $\Gamma^{t-1}(v),\varphi^{t-1}(e),$ and $\mathsf{w}^{t-1}(e)$, $\forall e\in E$ and $\forall v\in V$.
1:  If $t=1$, – Initialize the edge probability of each edge $e=x\to y$ as $\varphi^{1}(e)=\nicefrac{{1}}{{|\mathcal{N}_{x}|}}$ where $\mathcal{N}_{x}$ is the set of out-neighbors of $x$.
2:  If $t\geq 2$: – Set $\Gamma^{t-1}(d)=1$ and recursively compute in bottom-to-top fashion for every node $u$ in $\mathcal{G}^{t}(V,E)$:
$\Gamma^{t-1}(u)=\sum_{v:u\to v=e\ni E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))$
– For edge $e=u\to v$ in $\mathcal{G}^{t}(V,E)$, update edge probability: $\varphi^{t}(e)=\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)}.$
We prove Theorem 4.3 by showing that Algorithm 1 is an efficient (polynomial time) and equivalent implementation of the exponential weight updates algorithm. We complement Theorem 4.3 by an establishing matching lower bound.
Theorem 4.4.
Suppose $M\geq 2$ and $m=1$. Then, there exist competing bids, $[\bm{\beta}_{-}^{t}]_{t\in[T]}$, such that, under any learning algorithm, the expected regret $\mathbb{E}[\mathsf{REG}]=\Omega(M\sqrt{T})$ in both full information and bandit settings.
5.  Competing Against Richer Classes of Bidding Strategies
In previous sections, we characterized the class of safe bidding strategies and proposed a learning algorithm that obtain sublinear regret, where the regret is computed against a clairvoyant that also selects the optimal safe bidding strategy. In this section, we consider the cases where the clairvoyant can choose the optimal strategy from richer bidding classes, which we will describe shortly. We show that under Algorithm 1, the class of safe bidding strategies achieves sublinear (approximate) regret when competing against these stronger benchmarks. The performance metric we consider in this section is
$\alpha\text{-}\mathsf{REG}_{\mathscr{B}_{c}}=\alpha\cdot\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\,.$
Here, $\mathscr{B}_{c}$ is the class of bidding strategies from which the clairvoyant (which serves as our benchmark) chooses the optimal bidding strategy whereas the bidder (learner) submits bids, $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}$, in each round unless stated otherwise.
Remark 5.1 (Richness Ratio $\alpha$).
In prior works on approximate regret, $\alpha$ typically measures the hardness of the offline problem, such as $\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (see Streeter and Golovin (2008); Niazadeh et al. (2022)). In contrast, our work reinterprets $\alpha$ as richness ratio, capturing the relative richness of the clairvoyant’s strategy class compared to the learner’s. Specifically, it quantifies how closely the value achieved by the optimal strategy in $\mathscr{U}_{m}^{\star}$ approximates that by the optimal strategy in $\mathscr{B}_{c}$. If $\mathscr{B}_{c}=\mathscr{U}_{m}^{\star}$, then $\alpha=1$ (see Section 4). By considering a richer class of strategies for the clairvoyant, we aim to quantify the robustness and performance of safe strategies against stronger benchmarks.
Our main contribution in this section is to compute the richness ratio $\alpha$ for different classes of bidding strategies. To this end, for any $\mathscr{B}_{c}$, define:
$\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}$ $:=\max_{\mathbf{v},\mathcal{H}_{-}}\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v}),\text{where}$
$\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ $=\frac{\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}^{\prime}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime};\bm{\beta}_{-}^{t})}$ (6)
for $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$. In words, $\Lambda_{\mathscr{B}_{c}}$ is the maximum ratio of the value obtained by the optimal bidding strategy in $\mathscr{B}_{c}$ and that by the optimal strategy in $\mathscr{U}_{m}^{\star}$ over all valuation curves and bid histories in an offline setting. We drop the argument $\mathbf{v}$ from $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ when the context is clear.
Definition 4 (Richness Ratio).
Suppose the clairvoyant chooses the optimal strategy from $\mathscr{B}_{c}$ (which may depend on the bid history, $\mathcal{H}_{-}$) such that $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}\leq\lambda$, and the bound is tight, i.e., there exist a bid history $\mathcal{H}_{-}$, a valuation curve $\mathbf{v}$, and $\overline{\delta}>0$ such that $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\lambda-\delta$ for any $\delta\in(0,\overline{\delta}]$. Then, the richness ratio of $\mathscr{B}_{c}$ is $\alpha=\nicefrac{{1}}{{\lambda}}$.
The key idea is that if, in the worst-case scenario for the offline problem, the optimal safe strategy achieves no more than a $\frac{1}{\lambda}$-fraction of the value obtained by the optimal bidding strategy in $\mathscr{B}_{c}$, then the learner can achieve at most a $\frac{1}{\lambda}$-fraction of that value in the online setting as well. In the subsequent sections, we consider different classes of bidding strategies for the clairvoyant and compute its richness ratio $\alpha$. An immediate corollary of Definition 4 is:
Corollary 5.2.
Suppose the clairvoyant chooses the optimal strategy from the class $\mathscr{B}_{c}$ such that $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}\leq\lambda$ and this bound is tight. Then, Algorithm 1 obtains $\frac{1}{\lambda}$-$\mathsf{REG}_{\mathscr{B}_{c}}\leq O(M\sqrt{mT\log M})$ under full information feedback and $\frac{1}{\lambda}$-$\mathsf{REG}_{\mathscr{B}_{c}}\leq O(M\sqrt{m^{3}MT\log M})$ under bandit feedback.
Proof.
As $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}\leq\lambda$,
$\frac{1}{\lambda}\text{-}\mathsf{REG}_{\mathscr{B}_{c}}$ $=\frac{1}{\lambda}\cdot\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]$
$\leq\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]=\mathsf{REG}\,.\vspace{-0.5cm}$
By Theorem 4.3, we get the stated regret upper bound. ∎
5.1.  Richness Ratio of $m$-uniform Non-Safe Strategies
Here, we consider the clairvoyant class $\mathscr{B}_{c}$ to be the class of bidding strategies containing at most $m$ bid-quantity pairs, which are feasible only for the given bid history $\mathcal{H}_{-}$, rather than for every possible sequence of competing bids as in the case of safe strategies. The following theorem characterizes the richness ratio of this class, denoted as $\mathscr{F}^{\mathcal{H}_{-}}_{m}$:
Theorem 5.3.
For any $m\in\mathbb{N}$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}\leq 2$. Furthermore, there exist a bid history and valuation curve such that for any $\delta\in(0,\frac{1}{2}]$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq 2-\delta$. Thus, the richness ratio of the class $\mathscr{F}^{\mathcal{H}_{-}}_{m}$ is $\alpha=1/2$.
Theorem 5.3 implies that restricting the learner to safe bidding strategies has a bounded cost and does not lead to an arbitrary loss in the obtained value as the upper bound is independent of $m$. Building on this, Corollary 5.2 shows that Algorithm 1 results in $\frac{1}{2}$-approximate sublinear regret in the online setting when the safe strategies compete against $\mathscr{F}^{\mathcal{H}_{-}}_{m}$. Additionally, the factor-of-two loss represents a worst-case scenario, occurring in a highly non-trivial setting (see Section D.2). In practice, we expect these strategies to perform near-optimally, a claim further supported by the numerical simulations presented in Section 5.3.
Proof Sketch of Theorem 5.3 (Upper Bound). We now present the central ideas to prove the upper bound, which are also used to show the subsequent theorems, albeit with some changes specific to each case. Fixing a bid history, $\mathcal{H}_{-}$, we consider the expression in Section 5. Maximizing over all bid histories and valuation curves, we get the desired result.
Firstly, we bridge the we bridge the gap between the bidding strategies of the clairvoyant and the learner by constructing a restricted class of safe strategies, $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})\subseteq\mathscr{U}_{m}^{\star}$. This restricted class is exponentially smaller in size and is derived based on the optimal strategy selected by the clairvoyant (see exact characterization in Eq. 17 for Theorem 5.3 and Algorithm 3 for Theorem 5.4). We demonstrate that considering this restricted class of safe strategies is sufficient to derive tight upper bounds.
Second, as a crucial step, we show that for any $m$-uniform safe strategy $\mathbf{b}=\langle(w_{Q_{1}},q_{1}),\dots,(w_{Q_{m}},q_{m})\rangle$,
$V(\mathbf{b};\bm{\beta}_{-})=\max_{\ell\in[m]}V((w_{Q_{\ell}},Q_{\ell});\bm{\beta}_{-})$
for all competing bids $\bm{\beta}_{-}$, where, $Q_{\ell}=\sum_{j=1}^{\ell}q_{j}$. Thus, the value obtained by any $m$-uniform safe strategy can be determined by knowing the value obtained by the strategies in $\mathscr{U}_{1}^{\star}$, allowing us to focus solely on 1-uniform strategies.
As the third key idea, we partition the $T$ rounds into sets $T_{j}$, based on the least winning bid of the optimal bidding strategy chosen by the clairvoyant (see details in Eq. 15) and derive lower bounds on the value obtained by 1-uniform strategies in $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$ in terms of the value obtained by optimal strategy chosen by the clairvoyant in each partition $T_{j}$. With these three key ideas, we bound the ratio in Section 5. The detailed proof is presented in Section C.1.
Proof Sketch of Theorem 5.3 (Tight Lower Bound). For any $\delta\in(0,1]$, we construct a tight lower-bound instance for any $m\in\mathbb{N}$ by considering an offline problem where $M=T=N^{2m}$ and $K=M+1$, with $N=O(1/\delta)$. We set $\mathbf{v}=[1,v,\dots,v]\in\mathbb{R}^{M}$, where $v=1-O(\delta)$. The sequence of $T$ auctions is divided into $2m$ partitions of exponentially varying sizes, each with an identical competing bid profile. In each bid profile, there are two types of bids: ‘big’ and ‘small’. All ‘big’ bids are set to $C\gg 1$, while the ‘small’ bids remain identical within a partition. The partition sizes, the number of ‘small’ competing bids, and the exact values of these bids are carefully chosen to ensure that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-};\mathbf{v})\geq 2-\delta$ (see Sections D.1 to D.2).
5.2.  Richness Ratio of $m^{\prime}$-uniform Strategies
We first consider the case when the clairvoyant selects the optimal strategy from the class of safe bidding strategies with at most $m^{\prime}$ bid-quantity pairs where $m^{\prime}\geq m$.
Theorem 5.4.
For any $m,m^{\prime}\in\mathbb{N}$ such that $m^{\prime}\geq m$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}\leq\frac{m^{\prime}}{m}$. Additionally, there exist $\mathcal{H}_{-}$ and $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{m^{\prime}}{m}-\delta$. Thus, the richness ratio of the class $\mathscr{U}_{m^{\prime}}^{\star}$ is $\alpha=m/m^{\prime}$.
Suppose the clairvoyant can now select the optimal strategy from $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$, the class of bidding strategies with at most $m^{\prime}$ bid-quantity pairs ($m^{\prime}\geq m$) that are feasible for the bid history $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, but not necessarily safe. Then,
Theorem 5.5.
For any $m,m^{\prime}\in\mathbb{N}$ such that $m^{\prime}\geq m$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}\leq\frac{2m^{\prime}}{m}$. Moreover, there exist $\mathcal{H}_{-}$ and $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{2m^{\prime}}{m}-\delta$. Thus, the richness ratio of the class $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$ is $\alpha=m/2m^{\prime}$.
By Corollary 5.2, for $m^{\prime}\geq m$, Algorithm 1 obtains $\frac{m}{m^{\prime}}$-approximate (resp. $\frac{m}{2m^{\prime}}$-approximate) sublinear regret in the online setting when the clairvoyant draws the optimal strategy from the class of safe (resp. feasible but not necessarily safe) strategies with at most $m^{\prime}$ bid-quantity pairs.
The richness ratio in Theorem 5.4 and Theorem 5.5 can theoretically become arbitrarily small as $m^{\prime}$ grows, for a fixed $m$. However, the bidder has the flexibility to choose $m$; increasing $m$, though, incurs higher space and time complexity for Algorithm 1. Moreover, for any given pair $(m,m^{\prime})$, the bound is tight under a highly non-trivial setting (see details in Sections D.3 to D.4). In practice, as shown in Section 5.3, the observed richness ratios are substantially better than the theoretical bounds, suggesting that the worst-case scenarios are unlikely to arise in real-world applications.
5.3.  Estimating The Richness Ratio: A Case Study
In this section, we estimate the richness ratio (equivalently $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}$) for the discussed classes of bidding strategies by experiments using EU ETS emission permit auction data from 2022 and 2023. Detailed methods are provided in Appendix E, and the results are shown in Fig. 4.
Figure 4: The left (resp. right) figure refers to (the upper bound on) $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ (resp. $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-})$) as a function of $m$.
Estimating $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$. The left plot of Fig. 4 shows that (the upper bound for) $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ is significantly better than the theoretical bound of 2. The decaying trend of the plot with increasing $m$ can be attributed to the fact that we are computing an upper bound for the value achieved by the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m}$, which is independent of $m$, rather than the exact value (see Section E.2). For $m\geq 4$, even the upper bound of $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})\sim 1.05$ indicating that the safe bidding strategies are near-optimal.
Estimating $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-})$. From the right plot of Fig. 4, we observe that the empirical values of $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-})$ is significantly better than the worst-case bound $m$ (Theorem 5.4). In fact, the gain obtained by increasing the number of bid-quantity pairs plateaus after $m=4$ and even for $m=10$, the ratio of the value obtained by the optimal bidding strategy with at most 10 bid-quantity pairs to that by optimal 1-uniform bidding strategy is $\sim 1.25$.
Both plots show that the richness ratio is significantly better in practice than the theoretical bounds, suggesting that safe strategies with small $m$ are near-optimal in practice.
6.  Conclusion and Open Problems
We studied bidding in repeated uniform price auctions for a value maximizing buyer with per-round RoI constraints and characterized safe bidding strategies that can be efficiently learnt and are robust against various strong benchmarks. This study suggests several interesting future research directions: (1) analyzing bidding strategies in discriminatory-price (“pay-as-bid”) auctions; (2) modeling bidders with time-varying valuations and cumulative RoI constraints; (3) incorporating both RoI and budget constraints for a more realistic model of bidder behavior; and (4) extending the analysis to the more complex setting of combinatorial auctions with non-identical items.
Appendix A Omitted Proofs from Section 3
A.1.  Proof of Theorem 3.1
Before proving Theorem 3.1, we present the following lemma about the bidder’s per unit payments, which we will also use to prove several other results.
Lemma A.1 (Per-unit Payments).
Suppose the bidder bids $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$. Recall that $Q_{j}=\sum_{\ell=1}^{j}q_{\ell},\forall j\in[m]$. Recall that, $\bm{\beta}_{-,t}^{-(j)}$ is the $j^{th}$ smallest winning bid in the absence of bids from bidder $i$ for round $t$. If $j=0$, $\bm{\beta}_{-,t}^{-(j)}=0$ and $j>K$, $\bm{\beta}_{-,t}^{-(j)}=\infty$. Then, the per-unit payments by the bidder in round $t$ is
$p(\bm{\beta}^{t})=\begin{cases}0,&~{}\text{if }x(\bm{\beta}^{t})=0\\ b_{\ell},&~{}\text{if }Q_{\ell-1}<x(\bm{\beta}^{t})<Q_{\ell}\\ \min(b_{\ell},\bm{\beta}_{-,t}^{-(Q_{\ell}+1)}),&~{}\text{if }x(\bm{\beta}^{t})=Q_{\ell}\end{cases}\,.$ (7)
Define the following class of no-overbidding (NOB) strategies with $m$ bid-quantity pairs:
$\mathscr{S}_{m}^{\mathbf{NOB}}:=\Big{\{}\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}\leq w_{Q_{\ell}},\forall\ell\in[m]\Big{\}}\,.$
We first show that $\mathscr{S}_{m}\subseteq\mathscr{S}_{m}^{\mathbf{NOB}}$, and then complete the proof by showing $\mathscr{S}_{m}^{\mathbf{NOB}}\subseteq\mathscr{S}_{m}$.
Proof of $\mathscr{S}_{m}\subseteq\mathscr{S}_{m}^{\mathbf{NOB}}$.
Observation 1.
Overbidding is not a safe bidding strategy. To see this, let $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell}),\dots,(b_{m},q_{m})\rangle$ be an overbid such that $b_{\ell}>w_{Q_{\ell}}$. Now, consider an auction in which the competing bids are:
$\bm{\beta}_{-,}^{-(j)}=\begin{cases}\epsilon,&~{}\text{if }1\leq j\leq Q_{\ell}\\ 2b_{1},&~{}\text{if }Q_{\ell}<j\leq K\end{cases}\,,$
where $\epsilon<\frac{b_{m}}{2}$. Submitting $\mathbf{b}$ gets allocated $Q_{\ell}$ units and from Lemma A.1, we conclude that the clearing price of auction is $b_{\ell}$. As $b_{\ell}>w_{Q_{\ell}}$ by assumption, the RoI constraint is violated.
As overbidding is not a safe strategy, every safe bidding strategy is a NOB strategy, i.e., $\mathscr{S}_{m}\subseteq\mathscr{S}_{m}^{\mathbf{NOB}}$.
Proof of $\mathscr{S}_{m}^{\mathbf{NOB}}\subseteq\mathscr{S}_{m}$. We now prove that the converse is also true, i.e., every NOB strategy is also a safe strategy. To show this, fix any competing bid, $\bm{\beta}_{-}$, and consider any $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle\in\mathscr{S}_{m}^{\mathbf{NOB}}$. Suppose bidding $\mathbf{b}$ wins $x(\bm{\beta})$ units. So, from Lemma A.1, if $x(\bm{\beta})=0$, trivially, $P(\bm{\beta})=0=V(\bm{\beta})$. If $Q_{\ell-1}<x(\bm{\beta})\leq Q_{\ell}$, for some $\ell\in[m]$,
$P(\bm{\beta})$ $=x(\bm{\beta})\cdot p(\bm{\beta})\leq x(\bm{\beta})\cdot b_{\ell}\leq x(\bm{\beta})\cdot w_{Q_{\ell}}\leq x(\bm{\beta})\cdot w_{x(\bm{\beta})}=V(\bm{\beta})\,.$
The first inequality holds true because bidders’ per-unit payment is at most their least winning bid (individual rationality of the auction format), the second is true by definition, and the third is true because the $w_{j}$ is non-decreasing in $j$ and $x(\bm{\beta})\leq Q_{\ell}$. As the choice of competing bids and bidding strategy $\mathbf{b}$ was arbitrary, we conclude that every strategy in $\mathscr{S}_{m}^{\mathbf{NOB}}$ is safe, i.e., $\mathscr{S}_{m}^{\mathbf{NOB}}\subseteq\mathscr{S}_{m}$, which completes the proof.
A.1.1.  Proof of Lemma A.1
Consider the following three cases:
(1) If $x(\bm{\beta}^{t})=0$, trivially, $p(\bm{\beta}^{t})=0$.
(2) Let $Q_{\ell-1}<x(\bm{\beta}^{t})<Q_{\ell}$. Let $b$ be the last accepted bid after including $\mathbf{b}$, i.e., the smallest bid in $\bm{\beta}^{t}=(\mathbf{b},\bm{\beta}_{-}^{t})$. Then $b_{\ell}\stackrel{{\scriptstyle(a)}}{{\geq}}b\stackrel{{\scriptstyle(b)}}{{\geq}}b_{\ell}$.
  1. I.
    $(a)$ holds true because the bidder is allocated at least one unit for bid $b_{\ell}$ and
  2. II.
    $(b)$ is correct because they do not acquire at least one unit for bid $b_{\ell}$. Hence, $p(\bm{\beta}^{t})=b_{\ell}$.
(3) Suppose $x(\bm{\beta}^{t})=Q_{\ell}$. If $b_{\ell}>\bm{\beta}_{-,t}^{-(Q_{\ell}+1)}$, then $p(\bm{\beta}^{t})=\bm{\beta}_{-,t}^{-(Q_{\ell}+1)}$. However, if $b_{\ell}\leq\bm{\beta}_{-,t}^{-(Q_{\ell}+1)}$, $p(\bm{\beta}^{t})=b_{\ell}$. So, $p(\bm{\beta}^{t})=\min(b_{\ell},\bm{\beta}_{-,t}^{-(Q_{\ell}+1)})$.
A.2.  Proof of Theorem 3.2
We show that for any $m\in\mathbb{N}$, most safe bidding strategies are very weakly dominated for which they can removed from $\mathscr{S}_{m}$, resulting in $\mathscr{S}_{m}^{\star}$. We begin by establishing a general result regarding the monotonocity of feasible bid vectors (not necessarily $m$-uniform strategies) for value maximizing bidders. As the result holds for any bid vector, it is also true for $m$-uniform bidding strategies.
Lemma A.2 (Monotonocity of feasible bids).
Consider two sorted bid vectors: $\mathbf{b}=[b_{1},b_{2},\dots,b_{k}]$ and $\mathbf{b}^{\prime}=[b_{1}^{\prime},b_{2}^{\prime},\dots,b_{k}^{\prime}]$ such that $b_{j}\geq b_{j}^{\prime},\forall j\in[k]$. Suppose $\mathbf{b}$ is RoI feasible for some competing bid, $\bm{\beta}_{-}$. Then, $\mathbf{b}^{\prime}$ is also feasible for $\bm{\beta}_{-}$ and $V(\mathbf{b};\bm{\beta}_{-})\geq V(\mathbf{b}^{\prime};\bm{\beta}_{-})$.
Suppose $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an underbidding strategy per Definition 3. Consider $\mathbf{b}^{\prime}=\langle(b_{1}^{\prime},q_{1}^{\prime}),\dots,(b_{m}^{\prime},q_{m}^{\prime})\rangle$ such that $q_{j}^{\prime}=q_{j}$ and $b_{j}^{\prime}=w_{Q_{j}},\forall j\in[m]$. By Theorem 3.1, we establish that $\mathbf{b},\mathbf{b}^{\prime}\in\mathscr{S}_{m}$. Invoking Lemma A.2 gives us that the underbidding strategy is very weakly dominated, and hence can be removed from $\mathscr{S}_{m}$ to obtain $\mathscr{S}_{m}^{\star}$.
A.2.1.  Proof of Lemma A.2
Let $\bm{\beta}=(\mathbf{b},\bm{\beta}_{-})$ and $\bm{\beta}^{\prime}=(\mathbf{b}^{\prime},\bm{\beta}_{-})$. We first prove that $\mathbf{b}^{\prime}$ is also feasible for $\bm{\beta}_{-}$. Contrary to our claim, suppose $\mathbf{b}^{\prime}$ is infeasible, i.e., the value obtained by $\mathbf{b}^{\prime}$ when the competing bids are $\bm{\beta}_{-}$ is strictly less than the payment. Suppose $\mathbf{b}^{\prime}$ is allocated $r^{\prime}$ units with clearing price $p(\bm{\beta}^{\prime})$ such that the RoI constraint is violated:
$p(\bm{\beta}^{\prime})>w_{r^{\prime}}.$ (8)
Suppose $\mathbf{b}$ is allocated $r$ units when the competing bids are $\bm{\beta}_{-}$. By definition of allocation and payment rule, the units allocated and the clearing price in an auction are weakly increasing in bids, so $r^{\prime}\leq r$ and $p(\bm{\beta}^{\prime})\leq p(\bm{\beta})$. As $\mathbf{b}$ is feasible,
$p(\bm{\beta})\leq w_{r}.$ (9)
Combining Equations (8) and (9), we have
$p(\bm{\beta})\leq w_{r}\stackrel{{\scriptstyle(a)}}{{\leq}}w_{r^{\prime}}<p(\bm{\beta}^{\prime})\implies p(\bm{\beta})<p(\bm{\beta}^{\prime}),$
which is a contradiction. Here $(a)$ is true as $w_{j}$ is non-increasing in $j$ and $r^{\prime}\leq r$. So, $\mathbf{b}^{\prime}$ is feasible.
By definition of the allocation rule, the value obtained in an auction is weakly increasing in the bids submitted. As $\mathbf{b}$ and $\mathbf{b}^{\prime}$ are both feasible, $V(\mathbf{b};\bm{\beta}_{-})\geq V(\mathbf{b}^{\prime};\bm{\beta}_{-})$.
Appendix B Omitted Proofs From Section 4
B.1.  Proof of Lemma 4.1
By Theorem 3.2, a strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell})\rangle\in\mathscr{U}_{m}^{\star}$ for some $\ell\in[m]$ can be uniquely identified by $\{Q_{1},\dots,Q_{\ell}\}$ where $Q_{j}=\sum_{k\leq j}q_{k}$ because $b_{j}=w_{Q_{j}}$ for all $j\in[\ell]$. Let the bid history $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$. Then, for any round $t$,
$V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{j=1}^{\ell}\sum_{k=Q_{j-1}+1}^{Q_{j}}v_{k}\cdot\mathbb{I}{\left[w_{Q_{j}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}$
Hence, the problem (Offline) can be equivalently expressed as
$\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})=\max_{\ell\in[m]}\max_{Q_{1},\dots,Q_{\ell}}\sum_{j=1}^{\ell}\sum_{t=1}^{T}\sum_{k=Q_{j-1}+1}^{Q_{j}}v_{k}\cdot\mathbb{I}{\left[w_{Q_{j}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$
B.2.  Proof of Theorem 4.2
$\mathcal{G}(V,E)$ is a DAG. For convenience, let $d=(m+1,0)$. In the constructed graph, any node $(\ell,j)$ where $\ell\in[m]$ has edges edges either to the next layer, i.e. nodes of the form $(\ell+1,j^{\prime})$ where $j^{\prime}>j$ or to the destination node $(m+1,0)$. Hence, the directed graph has a topological sorting of the nodes implying that $\mathcal{G}(V,E)$ is a DAG.
Bijection between $s$-$d$ paths and strategies in $\mathscr{U}_{m}^{\star}$. Consider a path [^9] Here, we assume the path has $m+1$ edges without loss of generality. The same argument follows if $\mathfrak{p}=s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ with $k+1$ edges for some $k\in[m]$ is considered instead.
$\mathfrak{p}=s\to(1,z_{1})\to\dots\to(m,z_{m})\to d\,.$
By construction of $\mathcal{G}(V,E)$, edges $e=s\to(1,z_{1})$ and $e=(m,z_{m})\to d$ always exist. For any $\ell\in[m-1]$, the edge $(\ell,z_{\ell})\to(\ell+1,z_{\ell+1})$ exists if $z_{\ell}<z_{\ell+1}$. With this path $\mathfrak{p}$, we associate the strategy
$\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle\quad\text{where}\quad b_{\ell}=w_{z_{\ell}}\quad\text{and}\quad q_{\ell}=z_{\ell}-z_{\ell-1},$ $\forall\ell\in[m]\,.$
where $z_{0}=0$. By definition, $Q_{j}=\sum_{\ell=1}^{j}q_{\ell}=z_{j}$ and $b_{j}=w_{z_{j}}=w_{Q_{j}},\forall j\in[m]$. Hence, $\mathbf{b}\in\mathscr{U}_{m}^{\star}$.
Conversely, consider any safe strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ with $b_{j}=w_{Q_{j}}$. With this strategy, we associate the $s$-$d$ path $\mathfrak{p}^{\prime}=s\to(1,Q_{1})\to\dots\to(m,Q_{m})\to d$ where $Q_{j}=\sum_{\ell=1}^{j}q_{\ell},\forall j\in[m]$. We claim that $\mathfrak{p}^{\prime}$ is a valid path, i.e., all the edges exist, because by definition $e=s\to(1,Q_{1})$ and $e=(m,Q_{m})\to d$ always exist. Furthermore, for any $\ell\in[m-1]$, the edge $(\ell,Q_{l})\to(\ell+1,Q_{\ell+1})$ also exists as $Q_{\ell+1}-Q_{\ell}=q_{\ell+1}>0$.
Weight of $s$-$d$ paths. By assumption, $s=(0,0)$ and $z_{0}=0$. The weight of $\mathfrak{p}=s\to(1,z_{1})\to\dots\to(m,z_{m})\to d$ is:
$\mathsf{w}(\mathfrak{p})$ $=\sum_{e\in\mathfrak{p}}\mathsf{w}(e)=\sum_{\ell=1}^{m}\mathsf{w}((\ell-1,z_{\ell-1})\to(\ell,z_{\ell}))$
=(4)${{=}}\sum_{\ell=1}^{m}\sum_{t=1}^{T}\sum_{k=z_{\ell-1}+1}^{z_{\ell}}v_{k}\cdot\mathbb{I}{\left[w_{z_{\ell}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}$
$=\sum_{t=1}^{T}\sum_{\ell=1}^{m}\sum_{k=z_{\ell-1}+1}^{z_{\ell}}v_{k}\cdot\mathbb{I}{\left[w_{z_{\ell}}\geq\bm{\beta}_{-,t}^{-(k)}\right]},$
which is the value obtained by the safe strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle\in\mathscr{U}_{m}^{\star}$ where $b_{\ell}=w_{z_{\ell}},\forall\ell\in[m]$.
Computing maximum weight path. As $\mathcal{G}(V,E)$ is a DAG, the edge weights can be negated and the maximum (resp. minimum) weight path problem in the original (resp. ‘negated’) DAG can be solved in space and time complexity of $O(|V|+|E|)=O(mM^{2})$ which is polynomial in the parameters of the problem.
B.3.  Proof of Theorem 4.3
We begin by observing that there is a bijective mapping between $s$-$d$ paths in the DAG $\mathcal{G}^{t}(V,E)$ and bidding strategies $\mathbf{b}\in\mathscr{U}_{m}^{\star}$ (see proof of Theorem 4.2). The core idea of the proof is to show that Algorithm 1 is an equivalent and efficient implementation of the Hedge algorithm (Freund and Schapire, 1997) where every $s$-$d$ path is treated as an individual expert. To this end, we use an approach proposed in Takimoto and Warmuth (2003) and later used by Brânzei et al. (2023); Potfer et al. (2024); Galgana and Golrezaei (2024). For the sake of completeness, we provide the details.
In Algorithm 1, $\varphi^{t}(u\to\cdot)$ denotes the probability distribution over the outgoing neighbors of node $u$. By the recursive sampling of nodes (Line 4-7 of Algorithm 1), we get that the probability of selecting $s$-$d$ path $\mathfrak{p}$ is
$\mathbb{P}^{t}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{t}(e),$ (10)
and for any edge $e=u\to v$ in $\mathcal{G}^{t}(V,E)$, the edge probabilities, $\varphi^{t}(e)$, are updated as
$\varphi^{t}(e)=\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)},$ (11)
where $\Gamma^{t-1}(d)=1$ and $\Gamma^{t-1}(\cdot)$ is computed recursively in bottom-to-top fashion for every node $u$ in $\mathcal{G}^{t}(V,E)$ as follows:
$\Gamma^{t-1}(u)=\sum_{v:u\to v\in E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(u\to v)\cdot\exp(\eta\mathsf{w}^{t-1}(u\to v))\,.$ (12)
Now, consider a naïve implementation of the Hedge algorithm with learning rate $\eta$ in which each $s$-$d$ path is treated as an individual expert. For $t=1$, $\mathbb{P}_{\textsc{hedge}}^{1}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{1}(e)$. For $t\geq 2$, the probability of selecting path $\mathfrak{p}$ in round $t$ is
$\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})=\frac{\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})\exp(\eta\sum_{e\in\mathfrak{p}}\mathsf{w}^{t-1}(e))}{\sum_{\mathfrak{p}^{\prime}}\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p}^{\prime})\exp(\eta\sum_{e\in\mathfrak{p}^{\prime}}\mathsf{w}^{t-1}(e))}\,.$
Now, we show that for any path $\mathfrak{p}$ and $t\in[T]$, $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$. To this end, we first present the following result:
Claim B.1.
For any node $u$ in the graph, let $\mathscr{P}(u)$ be the set of paths starting at $u$ and terminating at $d$. Then,
$\Gamma^{t-1}(u)=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))$
Proof.
We prove the result by induction in a bottom-to-top order. For the base case, $u=d$, $\Gamma^{t-1}(d)=1$. Suppose the result is true for all the nodes in layer $\ell+1$ for some $0\leq\ell\leq m$. By the recursion in Eq. 12, for any node $u$ in layer $\ell$,
$\Gamma^{t-1}(u)$ $=\sum_{v:u\to v\in E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(u\to v)\cdot\exp(\eta\mathsf{w}^{t-1}(u\to v))$
$=\sum_{v:u\to v\in E}\Big{(}\sum_{\mathfrak{p}\in\mathscr{P}(v)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))\Big{)}\varphi^{t-1}(u\to v)\cdot\exp(\eta\mathsf{w}^{t-1}(u\to v))$
$=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e)),$
where the first equality follows from induction hypothesis. ∎
We will show by induction on $t\in[T]$ that for any path $\mathfrak{p}$, $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$. For $t=1$ and any path $\mathfrak{p}$, $\mathbb{P}_{\textsc{hedge}}^{1}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{1}(e)=\mathbb{P}^{1}(\mathfrak{p})$. Here, the second equality holds by Eq. 10. Suppose the result holds for round $t-1$, i.e., $\mathbb{P}^{t-1}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)$. For round $t$,
[…]
where the last equality follows by telescoping product and invoking the induction hypothesis. By definition, $\Gamma^{t-1}(d)=1$. By B.1,
$\Gamma^{t-1}(s)=\sum_{\mathfrak{p}^{\prime}}\prod_{e\in\mathfrak{p}^{\prime}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))=\sum_{\mathfrak{p}^{\prime}}\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p}^{\prime})\cdot\exp\Big{(}\eta\sum_{e\in\mathfrak{p}^{\prime}}\mathsf{w}^{t-1}(e)\Big{)}$
Substituting the values gives $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$ implying Algorithm 1 is a correct implementation of the Hedge algorithm.
B.3.1.  Bandit Feedback Model
In the bandit setting, the bidder learns about their own allocation only for which $\mathsf{w}^{t}(e)$ can not be exactly computed for all the edges. Hence, we use an unbiased estimator, $\widehat{\mathsf{w}}^{t}(e)$, of $\mathsf{w}^{t}(e)$ as follows:
$\widehat{\mathsf{w}}^{t}(e)$ $=\overline{\mathsf{w}}(e)-\frac{\overline{\mathsf{w}}(e)-\mathsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}\right]}\quad\text{where}\quad p^{t}(e)=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\,.$
Here, $p^{t}(e)$ is the probability of selecting edge $e$ in round $t$, $\mathbb{P}^{t}$ is the distribution over all $s$-$d$ paths in $\mathcal{G}^{t}(V,E)$ and $\overline{\mathsf{w}}(e)$ is an edge-dependent upper bound of the edge weight $\mathsf{w}^{t}(e)$. Formally,
$\overline{\mathsf{w}}(e)=\begin{cases}j^{\prime}-j,&\text{if }e=(\ell,j)\to(\ell+1,j^{\prime})\\ 0,&\text{if }e=(\ell,j)\to d\,.\end{cases}$
The algorithm followed by the bidder in the bandit setting is the same as Algorithm 1, except in Line 9, the bidder sets the edge weights as $\widehat{\mathsf{w}}^{t}(e)$ instead of $\mathsf{w}^{t}(e)$. Clearly, $\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)]=\mathsf{w}^{t}(e)$. The unbiased estimator $\widehat{\mathsf{w}}^{t}(e)$ is different from the natural importance-weighted estimator $\frac{\mathsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}\right]}$ for technical reasons (similar to Lattimore and Szepesvári (2020, Equation 11.6)) which we will clarify shortly.
B.3.2.  Regret Upper Bound
Full Information Setting. Having shown that Algorithm 1 emulates the Hedge algorithm, we now recall the its regret bound from Brânzei et al. (2023, Corollary 1) which states that for learning rate $\eta$:
$\mathsf{REG}=-\frac{1}{\eta}\max_{\mathfrak{p}}\log(\mathbb{P}^{1}(\mathfrak{p}))+\frac{\eta TM^{2}}{8}\,.$
Observe that $\mathbb{P}^{1}(\mathfrak{p})\geq M^{-m}$, which implies,
$\mathsf{REG}\leq\frac{m\log M}{\eta}+\frac{\eta TM^{2}}{8}\,.$
Setting $\eta=\frac{1}{M}\sqrt{\frac{8m\log M}{T}}$, we get $\mathsf{REG}\leq O(M\sqrt{mT\log M})$.
Bandit Setting. Using the standard analysis for EXP3 algorithm (see Lattimore and Szepesvári (2020, Chapter 11)), we get
$\mathsf{REG}=-\frac{1}{\eta}\max_{\mathfrak{p}}\log(\mathbb{P}^{1}(\mathfrak{p}))+\sum_{t=1}^{T}\mathbb{E}\left[\frac{1}{\eta}\log\left(\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})e^{\eta\widehat{\mathsf{w}}^{t}(\mathfrak{p})}\right)-\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\widehat{\mathsf{w}}^{t}(\mathfrak{p})\right]$
where $\widehat{\mathsf{w}}^{t}(\mathfrak{p})=\sum_{e\in\mathfrak{p}}\widehat{\mathsf{w}}^{t}(e)$. Note that $\widehat{\mathsf{w}}^{t}(\mathfrak{p})=\sum_{e\in\mathfrak{p}}\widehat{\mathsf{w}}^{t}(e)\leq\sum_{e\in\mathfrak{p}}\overline{\mathsf{w}}(e)\leq M$ and $\mathbb{P}^{1}(\mathfrak{p})\geq M^{-m}$. For $\eta\leq\frac{1}{M}$,
$\mathsf{REG}$ $\leq\frac{m\log M}{\eta}+\eta\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\mathbb{E}[\widehat{\mathsf{w}}^{t}(\mathfrak{p})^{2}]$
$\leq\frac{m\log M}{\eta}+\eta m\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]$
where in the first inequality, we used $e^{x}\leq 1+x+x^{2}$ for $x\leq 1$ and $\log(1+x)\leq x$ for $x\geq 0$ and the second inequality follows from Cauchy-Schwarz. Observe that
$\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]=\sum_{e}\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})=\sum_{e}p^{t}(e)\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}],$
where the last equality holds as $p^{t}(e)=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})$.
$\sum_{e}p^{t}(e)\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]$ $=\sum_{e}p^{t}(e)\Big{[}\overline{\mathsf{w}}(e)^{2}(1-p^{t}(e))+\Big{(}\overline{\mathsf{w}}(e)-\frac{\overline{\mathsf{w}}(e)-\mathsf{w}^{t}(e)}{p^{t}(e)}\Big{)}^{2}p^{t}(e)\Big{]}$
$=\sum_{e}\mathsf{w}^{t}(e)^{2}p^{t}(e)+(\overline{\mathsf{w}}(e)-\mathsf{w}^{t}(e))^{2}(1-p^{t}(e))\leq\sum_{e}\overline{\mathsf{w}}(e)^{2}\,.$
We now aim to upper bound $\sum_{e}\overline{\mathsf{w}}(e)^{2}$.
$\sum_{e}\overline{\mathsf{w}}(e)^{2}=\sum_{\ell=0}^{m-1}\sum_{j<j^{\prime}\leq M}\overline{\mathsf{w}}((\ell,j)\to(\ell+1,j^{\prime}))^{2}=m\sum_{j<j^{\prime}\leq M}(j-j^{\prime})^{2}\leq 4m\sum_{j=1}^{M}j^{2}\leq 4mM^{3}\,.$
Hence,
$\mathsf{REG}$ $\leq\frac{m\log M}{\eta}+\eta m\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]$
$\lesssim\frac{m\log M}{\eta}+\eta Tm^{2}M^{3}$
Setting $\eta=\frac{1}{M}\sqrt{\frac{\log M}{mMT}}$, we get $\mathsf{REG}\leq O(M\sqrt{m^{3}MT\log M})\,.$  [^10] Recall that the analysis assumes that $\eta\leq\nicefrac{{1}}{{M}}$. As $m\geq 1,T\geq 1$ and $\log M\leq M$, the chosen $\eta=\frac{1}{M}\sqrt{\frac{\log M}{mMT}}\leq\frac{1}{M}.$, [^11] The upper bound on $\mathsf{REG}$ has a linear dependence on $M$ which is unavoidable as the reward in each round is in $[0,M]$. If normalized regret (Zhang et al., 2024) is considered instead, i.e., the rewards are scaled such that in each round it lies in $[0,1]$, the dependence on $M$ is logarithmic (or sublinear in the bandit setting).
B.3.3.  Time Complexity
In the full information setting, the running time bottleneck is Eq. 12. Thus, Algorithm 1 runs in $|E|=O(mM^{2})$ time per round. The bottleneck of running Algorithm 1 under the bandit feedback is the efficient computation of the marginal distribution, $p^{t}(e)$. Note that $p^{t}(e)$ has a sum-product structure as follows:
$p^{t}(e)=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\prod_{e^{\prime}\in\mathfrak{p}}\varphi^{t}(e^{\prime})\,.$
Hence, by the message passing algorithm from Wainwright and Jordan (2008, Section 2.5.1), we can compute the marginal probabilities (and hence run Algorithm 1) in $O(M|E|)=O(mM^{3})$ time per round.
B.4.  Proof of Theorem 4.4
For $m=1$, let $K=M$ be an even integer. Let $\mathbf{v}=[1,\dots,1,v,\dots,v]$. Here, the first $\frac{M}{2}$ entries are 1 and the remaining $\frac{M}{2}$ entries are $v$. Let $v=1-\delta$. Define $\delta^{\prime}=\frac{\delta}{2M}$. Consider two scenarios:
Scenario 1. In this scenario, for every $t\in[T]$, the competing bids $\bm{\beta}_{-}^{t}$ are:
$\bm{\beta}_{-}^{t}=\begin{cases}[1-\delta^{\prime},\dots,1-\delta^{\prime}],&\text{ w.p. }\frac{1}{2}+\delta,\\ [\frac{1+v}{2}-\delta^{\prime},\dots,\frac{1+v}{2}-\delta^{\prime}],&\text{ w.p. }\frac{1}{2}-\delta,\\ \end{cases}$
Scenario 2. In this scenario, for every $t\in[T]$, the competing bids $\bm{\beta}_{-}^{t}$ are:
$\bm{\beta}_{-}^{t}=\begin{cases}[1-\delta^{\prime},\dots,1-\delta^{\prime}],&\text{ w.p. }\frac{1}{2}-\delta,\\ [\frac{1+v}{2}-\delta^{\prime},\dots,\frac{1+v}{2}-\delta^{\prime}],&\text{ w.p. }\frac{1}{2}+\delta,\\ \end{cases}$
for some $\delta\in(0,1/4)$ that would be determined shortly. Assume the randomness used in different rounds are independent.
Let $P$ and $Q$ be the distribution of $[\bm{\beta}_{-}^{t}]_{t\in[T]}$ for scenario 1 and 2 respectively. Then, for $\delta\in(0,\frac{1}{4})$,
$\textsc{KL}(P||Q)=T\cdot\textsc{KL}(\textsc{Bern}(0.5+\delta)||\textsc{Bern}(0.5-\delta))=2T\delta\log\Big{(}\frac{1+2\delta}{1-2\delta}\Big{)}\leq\frac{8T\delta^{2}}{1-2\delta}\leq 16T\delta^{2}$
where the first inequality follows from $\log(\frac{1+x}{1-x})\leq\frac{2x}{1-x}$. By Tsybakov (2009, Lemma 2.6),
$1-\textsc{TV}(P,Q)\geq\frac{1}{2}\exp{(-\textsc{KL}(P||Q))}\geq\frac{1}{2}\exp{\left(-16T\delta^{2}\right)}\,.$
Consider the class of 1-uniform safe strategies. We first show that we need to consider only two strategies out of the $M$ possible strategies. Recall that any strategy in this class is of the form $(w_{q},q)$.
Case 1. $1\leq q\leq\frac{M}{2}$: Here, $w_{q}=1,\forall q$. So, $(1,\frac{M}{2})$ weakly dominates all the strategies of the form $(1,q)$ for $1\leq q\leq\frac{M}{2}$.
Case 2. $\frac{M}{2}+1\leq q\leq M$: In this interval, the highest bid value is for $q=\frac{M}{2}+1$ due to diminishing marginal returns property. Note that,
$w_{\frac{M}{2}+1}=\frac{\frac{M}{2}+1-\delta}{\frac{M}{2}+1}=1-\frac{\delta}{\frac{M}{2}+1}\leq 1-\frac{\delta}{M}<1-\delta^{\prime}\,.$
Hence, no strategy of form $(w_{q},q)$ for $q\geq\frac{M}{2}+1$ is allocated any unit in when the competing bids are $[1-\delta^{\prime},\dots,1-\delta^{\prime}]$. The smallest bid value for $\frac{M}{2}+1\leq q\leq M$ is for $q=M$. Note that, $w_{M}=\frac{1+v}{2}>\frac{1+v}{2}-\delta^{\prime}$. Hence, $(w_{M},M)=(\frac{1+v}{2},M)=(1-\frac{\delta}{2},M)$ is allocated all the units when the competing bids are $[\frac{1+v}{2}-\delta^{\prime},\dots,\frac{1+v}{2}-\delta^{\prime}]$ and, by definition, weakly dominates all the strategies of the form $(w_{q},q)$ for $q\geq\frac{M}{2}+1$. Hence, we have two undominated (for the constructed competing bids) strategies: $\mathbf{b}_{1}=(1,\frac{M}{2})$ and $\mathbf{b}_{2}=(1-\frac{\delta}{2},M)$.
Now, we compute the expected value obtained by $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ for scenarios $P$ and $Q$.
$\mathbb{E}_{P}[V(\mathbf{b}_{1};\bm{\beta}_{-}^{t})]$ $=\mathbb{E}_{Q}[V(\mathbf{b}_{1};\bm{\beta}_{-}^{t})]=\frac{M}{2}$
$\mathbb{E}_{P}[V(\mathbf{b}_{2};\bm{\beta}_{-}^{t})]$ $=M\left(\frac{1}{2}-\delta\right)\left(\frac{1+v}{2}\right)=\frac{M}{2}\left(\frac{1}{2}-\delta\right)(2-\delta)<\frac{M}{2}$
$\mathbb{E}_{Q}[V(\mathbf{b}_{2};\bm{\beta}_{-}^{t})]$ $=M\left(\frac{1}{2}+\delta\right)\left(\frac{1+v}{2}\right)=\frac{M}{2}\left(\frac{1}{2}+\delta\right)(2-\delta)>\frac{M}{2}$
So, $\mathbf{b}_{1}$ (resp. $\mathbf{b}_{2}$) is optimal for scenario $P$ (resp. scenario $Q$). Now, consider the distribution $\frac{P+Q}{2}$, i.e., $\textsc{Bern}(0.5)$.
$\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}_{1};\bm{\beta}_{-}^{t})]=\frac{M}{2}~{}~{}\text{ and }~{}~{}$ $\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}_{2};\bm{\beta}_{-}^{t})]=\frac{M}{2}\left(\frac{1+v}{2}\right)\leq\frac{M}{2}$
$\implies\max_{\mathbf{b}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b};\bm{\beta}_{-}^{t})]$ $\leq\frac{M}{2}$ (13)
Hence, for any $\mathbf{b}\in\mathscr{U}_{1}^{\star}$, and any round $t\in[T]$,
$\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{P}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})-V(\mathbf{b};\bm{\beta}_{-}^{t})]+\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{Q}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})-V(\mathbf{b};\bm{\beta}_{-}^{t})]$
$\geq\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{P}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]+\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{Q}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]-2\max_{b^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]$
$=\mathbb{E}_{P}[V(\mathbf{b}_{1};\bm{\beta}_{-}^{t})]+\mathbb{E}_{Q}[V(\mathbf{b}_{2};\bm{\beta}_{-}^{t})]-2\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]$
(13)${{\geq}}\frac{M}{2}+\frac{M}{2}\left(\frac{1}{2}+\delta\right)(2-\delta)-2\cdot\frac{M}{2}=\frac{M}{2}\left(\frac{3\delta}{2}-\delta^{2}\right)\geq\frac{5M\delta}{8},$
where the last inequality follows as $\delta\in(0,\frac{1}{4})$. Hence, any strategy $\mathbf{b}\in\mathscr{U}_{1}^{\star}$ incurs a total regret of $\frac{5MT\delta}{16}$ either under $P$ or under $Q$. By two-point method from Tsybakov (2009, Theorem 2.2),
$\mathbb{E}_{\frac{P+Q}{2}}[\mathsf{REG}]\geq\frac{5MT\delta}{16}\cdot(1-\text{TV}(P,Q))\geq\frac{5MT\delta}{32}\exp{\left(-16T\delta^{2}\right)}$
Setting $\delta=\frac{1}{4\sqrt{2T}}$, we get $\mathbb{E}_{\frac{P+Q}{2}}[\mathsf{REG}]\geq\frac{M\sqrt{T}}{36\sqrt{e}}$. Hence, $\mathbb{E}_{\frac{P+Q}{2}}[\mathsf{REG}]=\Omega(M\sqrt{T})$.
Appendix C Omitted Proofs from Section 5
C.1.  Proof of Theorem 5.3
To prove that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}\leq 2$ for any $m\in\mathbb{N}$, recall from the proof sketch that we first defined the following metric for any bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})=\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$ (14)
Let $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{m}^{*},q_{m}^{*})\rangle$ where $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right):=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ [^12] We assumed $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{m}^{*},q_{m}^{*})\rangle$ without loss of generality. The proof also follows if we considered $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{k}^{*},q_{k}^{*})\rangle$ for some $k\in[m]$ instead.Define $Q_{\ell}^{*}=\sum_{j\leq\ell}q_{j}^{*}$. Suppose $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated $r_{t}^{*}$ units in round $t$. For any $j\in[m]$, let $T_{j}$ be the set of rounds in which the least winning bid is $b_{j}^{*}$, i.e.,
$T_{j}=\big{\{}t\in[T]:Q_{j-1}^{*}<r_{t}^{*}\leq Q_{j}^{*}\big{\}}\,.$ (15)
For any $j\in[m]$, partition $T_{j}$ into $T_{j,0}$ and $T_{j,1}$ such that $T_{j,0}$ is the set of rounds where the bidder gets strictly less than $Q_{j}^{*}$ units and $T_{j,1}$ is the set of rounds when they get exactly $Q_{j}^{*}$ units:
$T_{j,0}=\Big{\{}t\in[T_{j}]:r_{t}^{*}<Q_{j}^{*}\Big{\}},\qquad T_{j,1}=\Big{\{}t\in[T_{j}]:r_{t}^{*}=Q_{j}^{*}\Big{\}}\,.$ (16)
For any $j\in[m]$, let
$\widehat{Q}_{j}=\begin{cases}\max\{r_{t}^{*}:t\in T_{j,0}\},~{}&\text{ if }T_{j,0}\neq\emptyset\\ Q_{j}^{*}~{}&\text{ if }T_{j,0}=\emptyset\,.\end{cases}$
In other words, $\widehat{Q}_{j}$ is the $2^{nd}$ highest number of units allocated to $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ over the rounds in $T_{j}$ (or the highest if $T_{j,0}=\emptyset$ or $T_{j,1}=\emptyset$).
Step 1. Constructing a Restricted Class of Safe Bidding Strategies. Here, as a crucial part of the proof, we construct a restricted class of safe bidding strategies, denoted by $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, where $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})\subset\mathscr{U}_{m}^{\star}$. This construction serves two purposes. First, it reduces the search space for the optimal safe bidding strategy. Second, and more importantly, it enables us to establish a connection between the optimal safe bidding strategy and $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$.
In defining the restricted class, we use the quantities, $\{\widehat{Q}_{j},Q_{j}^{*}\}_{j\in[m]}$ as follows:
$\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})=\Big{\{}\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}},~{}~{}Q_{\ell}\in\{\widehat{Q}_{\ell},Q_{\ell}^{*}\},\quad\forall\ell\in[m]\Big{\}}\,.$ (17)
Recall that any strategy in $\mathscr{U}_{m}^{\star}$ with $m$ bid-quantity pairs takes the form of $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}},\forall\ell\in[m]$. The strategies in $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$ also have the same structure but as an important difference, for any $\ell\in[m]$, we enforce $Q_{\ell}\in\{\widehat{Q}_{\ell},Q_{\ell}^{*}\}$. Observe that for any $\ell\in[m-1]$,
$Q_{\ell}\leq Q_{\ell}^{*}<\widehat{Q}_{\ell+1}\leq Q_{\ell+1},$
where the first and third inequalities follow directly from the definition of $Q_{\ell}$ and the second one follows from the definition of $\widehat{Q}_{\ell+1}$. So, $Q_{\ell}$’s are distinct and ordered. Further observe that the number of bidding strategies in $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$ is $O(2^{m})$; significantly smaller than the number of strategies in $\mathscr{U}_{m}^{\star}$, which is $O(M^{m})$.
With the definition of the restricted class of safe bidding strategies, we have
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})=\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\leq\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$ (18)
Step 2. Value Decomposition for $m$-uniform Strategies. Let $V_{j,k}$ be the time-average value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in the set of rounds in $T_{j,k}$ (as defined in Eq. 16). Formally, $\forall j\in[m],k\in\{0,1\}$,
$V_{j,k}=\dfrac{1}{|T_{j,k}|}\sum_{t\in T_{j,k}}V(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right),\bm{\beta}_{-}^{t})\,.$
Define ${W}_{k}=\sum_{j=1}^{k}v_{j}$. Note that $V_{j,1}={W}_{Q_{j}^{*}}$ because for any $t\in T_{j,1}$, we have $r_{t}^{*}=Q_{j}^{*}$. For any $\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, where $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}}$, and $Q_{\ell}\in\{Q_{\ell}^{*},\widehat{Q}_{\ell}\}$ for any $\ell\in[m]$, define $N^{*},\widehat{N}\subseteq[m]$ as follows:
$N^{*}=\{j:Q_{j}=Q_{j}^{*},j\in[m]\}\quad\text{ and }\quad\widehat{N}=\{j:Q_{j}=\widehat{Q}_{j},j\in[m]\}\,.$
We now present a crucial result that expresses the value obtained by a $m$-uniform bidding strategy as a function of the value obtained by $m$ 1-uniform strategies.
Lemma C.1.
Let $\mathbf{b}=\langle(w_{Q_{1}},q_{1}),\dots,(w_{Q_{m}},q_{m})\rangle$ be a $m$-uniform safe bidding strategy for some $m\in\mathbb{N}$. Then, for any competing bids $\bm{\beta}_{-}$,
$V(\mathbf{b};\bm{\beta}_{-})=\max_{\ell\in[m]}V((w_{Q_{\ell}},Q_{\ell});\bm{\beta}_{-}),$
where we recall that $Q_{\ell}=\sum_{j=1}^{\ell}q_{j},\forall\ell\in[m]$.
A key consequence of Lemma C.1 is that knowing the value obtained by the strategies in $\mathscr{U}_{1}^{\star}$ is sufficient to compute the value obtained any $m$-uniform safe bidding strategy. By Lemma C.1, for any round $t\in T_{j}$, and $\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, we get
$V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\max_{\ell\in[m]}V((w_{Q_{\ell}},Q_{\ell});\bm{\beta}_{-}^{t})\geq V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\,,$ (19)
$\implies\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $\geq\sum_{j\in N^{*}}\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})+\sum_{j\in\widehat{N}}\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\,.$ (20)
Step 3. Allocation Lower Bounds for 1-uniform Strategies. For any $j\in[m]$, we now invoke the following lemma to establish lower bound on $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})$
Lemma C.2.
Let $\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, where $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{j}=w_{Q_{j}}$, and $Q_{j}\in\{Q_{j}^{*},\widehat{Q}_{j}\},\forall j\in[m]$. Then, for any $j\in[m]$,
  • (a)
    if $Q_{j}=\widehat{Q}_{j}$ (i.e., $j\in\widehat{N}$), we have
    $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\geq V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\,.$ (21)
  • (b)
    If $Q_{j}=Q_{j}^{*}$ (i.e., $j\in N^{*}$), we have
    $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\geq{W}_{Q_{j}^{*}}|T_{j,1}|\,.$ (22)
Note that the right hand side of Eq. 20 depends only on the choice of the partitions $N^{*}$ and $\widehat{N}$. Substituting the lower bound from Lemma C.2 in Eq. 20, we establish that,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ (18)${{\leq}}\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}$
$\leq\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{(N^{*},\widehat{N})}\Big{\{}\sum_{j\in N^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}\Big{\}}}\,.$ (23)
Let $(N_{0}^{*},\widehat{N}_{0})=\mathop{\mathrm{argmax}}_{(N^{*},\widehat{N})}\Big{\{}\sum_{j\in N^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}\Big{\}}$. Then,
$\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{t=1}^{T}\sum_{\ell=0}^{1}V_{j,\ell}|T_{j,\ell}|$
$=\sum_{j\in N_{0}^{*}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}\,.$ (24)
Consider a partition of $[m]$ that differs from the maximizing partition $(N_{0}^{*},\widehat{N}_{0})$ by exactly one element, i.e., for any $a\in\widehat{N}_{0}$, consider the following partition: $(N_{0}^{*}\cup\{a\},\widehat{N}_{0}\setminus\{a\})$. By definition,
$\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$ $\geq\sum_{j\in N_{0}^{*}\cup\{a\}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}\setminus\{a\}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$
$\implies V_{a,0}|T_{a,0}|+{W}_{\widehat{Q}_{a}}|T_{a,1}|$ $\geq{W}_{Q_{a}^{*}}|T_{a,1}|\,.$ (25)
Now, for any $b\in N^{*}_{0}$, consider the following partition: $(N_{0}^{*}\setminus\{b\},\widehat{N}_{0}\cup\{b\})$. By definition,
$\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$ $\geq\sum_{j\in N_{0}^{*}\setminus\{b\}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}\cup\{b\}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$
$\implies{W}_{Q_{b}^{*}}|T_{b,1}|-{W}_{\widehat{Q}_{b}}|T_{b,1}|$ $\geq V_{b,0}|T_{b,0}|\,.$ (26)
Plugging in the values,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ (C.1)${{\leq}}\frac{\sum_{j\in N_{0}^{*}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
(25)${{\leq}}\frac{\sum_{j\in N_{0}^{*}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}2V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
(26)${{\leq}}\frac{\sum_{j\in N_{0}^{*}}\Big{(}2{W}_{Q_{j}^{*}}|T_{j,1}|-{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}2V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
$=\frac{2\Big{\{}\sum_{j\in N_{0}^{*}}{W}_{Q_{j}^{*}}|T_{j,1}|+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}\Big{\}}-\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
$=2-\theta_{\mathcal{H}_{-}}\,,$
where
$0<\theta_{\mathcal{H}_{-}}$ $=\frac{\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
(25)${{\leq}}\frac{\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}}$
$=\frac{\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j=1}^{m}{W}_{Q_{j}^{*}}|T_{j,1}|}\stackrel{{\scriptstyle(a)}}{{\leq}}\max_{j\in[m]}\frac{{W}_{\widehat{Q}_{j}}}{{W}_{Q_{j}^{*}}}\leq 1\,,$
and $(a)$ follows from 1. Finally,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}=\max_{\mathcal{H}_{-}}\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})\leq 2-\min_{\mathcal{H}_{-}}\theta_{\mathcal{H}_{-}}=:2-\theta,\quad\text{where}\quad\theta=\min_{\mathcal{H}_{-}}\theta_{\mathcal{H}_{-}}\in(0,1]\,.$
Fact 1 (Williamson and Shmoys (2011, pp. 25)).
For positive numbers $a_{1},\dots,a_{k}$ and $b_{1},\dots,b_{k}$,
$\frac{\sum_{j=1}^{k}a_{j}}{\sum_{j=1}^{k}b_{j}}\leq\max_{j\in[m]}\frac{a_{j}}{b_{j}}\,.$
C.2.  Proof of Lemma C.1
We state and prove a stronger version of the result in Lemma C.1. Formally,
Lemma C.3.
For any $m\in\mathbb{N}$ and competing bid $\bm{\beta}_{-}$, let $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ be a feasible (not necessarily safe) $m$-uniform strategy for $\bm{\beta}_{-}$. Then,
$V(\mathbf{b};\bm{\beta}_{-})=\max_{\ell\in[m]}V((b_{\ell},Q_{\ell});\bm{\beta}_{-}),$
where we recall that $Q_{\ell}=\sum_{j=1}^{\ell}q_{j},\forall\ell\in[m]$.
Lemma C.3 states a similar result as Lemma C.1, except a key difference that the bidding strategies are not necessarily safe.
C.2.1.  Proof of Lemma C.3
We prove the lemma via induction on $m$. The base case is $m=1$ for which the result is trivially true. Now assume that the result holds for any $m$-uniform bidding strategy for any competing bid $\bm{\beta}_{-}$. We now show that the result holds for any $(m+1)$-uniform bidding strategy which is feasible for $\bm{\beta}_{-}$.
Consider any $m+1$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m}),(b_{m+1},q_{m+1})\rangle$ feasible for $\bm{\beta}_{-}$. Then,
Claim C.4.
The bid-quantity pair $(b_{m+1},Q_{m+1})$ is feasible for $\bm{\beta}_{-}$ and $V(\mathbf{b};\bm{\beta}_{-})\geq V((b_{m+1},Q_{m+1});\bm{\beta}_{-})$.
Proof.
By assumption, $\mathbf{b}$ is feasible, the total demand of $\mathbf{b}$ and $(b_{m+1},Q_{m+1})$ are equal and $b_{(j)}\geq b_{m+1},\forall j\in[Q_{m+1}]$, where $b_{(j)}$ denotes the bid value in $j$th position in the sorted bid vector. So, by Lemma A.2, $(b_{m+1},Q_{m+1})$ is feasible and $V(\mathbf{b};\bm{\beta}_{-})\geq V((b_{m+1},Q_{m+1}),\bm{\beta}_{-})$. ∎
Suppose that by bidding $\mathbf{b}$, the bidder is allocated $r$ units. There are two cases: (a) $r\leq Q_{m}$ and (b) $r>Q_{m}$.
Case I. $r\leq Q_{m}$. In this case, we have
$V(\mathbf{b};\bm{\beta}_{-})=V(\mathbf{b}[1:m];\bm{\beta}_{-})\,.$ (27)
Hence, by C.4 and Eq. 27,
$V(\mathbf{b};\bm{\beta}_{-})$ $=\max\Big{\{}V(\mathbf{b}[1:m];\bm{\beta}_{-}),V((b_{m+1},Q_{m+1});\bm{\beta}_{-})\Big{\}}\,.$ (28)
Case II. $r>Q_{m}$. As $r>Q_{m}$, $b_{m+1}$ is the least winning bid which implies $b_{m+1}\geq\bm{\beta}_{-}^{-(r)}$, where we recall that $\bm{\beta}_{-}^{-(r)}$ is the $r^{th}$ smallest competing bid in $\bm{\beta}_{-}$. So, $(b_{m+1},Q_{m+1})$ is allocated at least $r$ units which implies $V((b_{m+1},Q_{m+1}),\bm{\beta}_{-})\geq V(\mathbf{b};\bm{\beta}_{-})$. By C.4, we also have $V(\mathbf{b};\bm{\beta}_{-})\geq V((b_{m+1},Q_{m+1}),\bm{\beta}_{-})$. Hence,
$V(\mathbf{b};\bm{\beta}_{-})=V((b_{m+1},Q_{m+1});\bm{\beta}_{-})\,.$ (29)
As $r>Q_{m}$, $(b_{m+1},Q_{m+1})$ is allocated at least $Q_{m}+1$ units, whereas $\mathbf{b}[1:m]$ has demand for (hence, can be allocated) at most $Q_{m}$ units. So,
$V(\mathbf{b};\bm{\beta}_{-})$ $\geq V(\mathbf{b}[1:m];\bm{\beta}_{-})$
$\implies V(\mathbf{b};\bm{\beta}_{-})$ $=\max\Big{\{}V(\mathbf{b}[1:m];\bm{\beta}_{-}),V((b_{m+1},Q_{m+1});\bm{\beta}_{-})\Big{\}}\,.$ (30)
For both Case I and Case II, we get the same result (cf. (28) and (30)). Hence,
$V(\mathbf{b};\bm{\beta}_{-})$ $=\max\Big{\{}V(\mathbf{b}[1:m];\bm{\beta}_{-}),V((b_{m+1},Q_{m+1});\bm{\beta}_{-})\Big{\}}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\max\Big{\{}\max_{\ell\in[m]}V((b_{\ell},Q_{\ell});\bm{\beta}_{-}),V((b_{m+1},Q_{m+1});\bm{\beta}_{-})\Big{\}}$
$=\max_{\ell\in[m+1]}V((b_{\ell},Q_{\ell});\bm{\beta}_{-})\,.$
Here, $(a)$ holds as $\mathbf{b}[1:m]$ is feasible for $\bm{\beta}_{-}$ allowing us to apply the induction hypothesis for $m$.
C.3.  Proof of Lemma C.2
To prove this result, we use the following key lemma that measures the value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in terms of 1-uniform safe bidding strategies.
Lemma C.5.
Let $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{m}^{*},q_{m}^{*})\rangle$. Suppose $Q_{j}^{*}=\sum_{\ell\leq j}q_{\ell}^{*},\forall j\in[m]$ and $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated $r_{t}^{*}$ units in any round $t\in[T]$. Then,
1. For any $t\in[T]$ and $q\leq r_{t}^{*}$, $(w_{q},q)$ gets exactly $q$ units.
2. For any $j\in[m]$, let $T_{j}\subseteq[T]$ be the rounds in which $b_{j}^{*}$ is the least winning bid, i.e.,
$T_{j}=\big{\{}t\in[T]:Q_{j-1}^{*}<r_{t}^{*}\leq Q_{j}^{*}\big{\}}\,.$
Suppose that $\exists t\in T_{j}$ such that $r_{t}^{*}<Q_{j}^{*}$. Then in any round $t^{\prime}\in T_{j}$ in which $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated at most $r_{t}^{*}$ units, i.e., $r_{t^{\prime}}^{*}\leq r_{t}^{*}$, $(w_{r_{t}^{*}},r_{t}^{*})$ is allocated at least $r_{t^{\prime}}^{*}$ units.
With this lemma, we are ready to prove Lemma C.2.
Case I: $Q_{j}=\widehat{Q}_{j}$. By assumption, $Q_{j}=\widehat{Q}_{j}\leq Q_{j}^{*}$. So, for any $t\in T_{j,1}$, invoking Lemma C.5 (1) with $q=\widehat{Q}_{j}$, we conclude that $(w_{\widehat{Q}_{j}},\widehat{Q}_{j})$ is allocated exactly $\widehat{Q}_{j}$ units. Summing over all rounds in $T_{j,1}$,
$\sum_{t\in T_{j,1}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})={W}_{\widehat{Q}_{j}}|T_{j,1}|\,.$ (31)
Now, suppose $T_{j,0}\neq\emptyset$ and $r_{t}^{*}=\widehat{Q}_{j}$ for some $t\in T_{j,0}$. By definition, $\widehat{Q}_{j}<Q_{j}^{*}$ and $\widehat{Q}_{j}\geq r_{s}^{*}$ for any $s\in T_{j,0}$. So, for any $s\in T_{j,0}$, invoking Lemma C.5 (2) with $r_{t}^{*}=\widehat{Q}_{j}$, we conclude that $(w_{\widehat{Q}_{j}},\widehat{Q}_{j})$ is allocated at least $r_{s}^{*}$ units. Hence, summing over all rounds, $(w_{\widehat{Q}_{j}},\widehat{Q}_{j})$ gets at least the value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ over the rounds in $T_{j,0}$. So,
$\sum_{t\in T_{j,0}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})$ $\geq\sum_{t\in T_{j,0}}V\Big{(}\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right),\bm{\beta}_{-}^{t}\Big{)}=V_{j,0}|T_{j,0}|\,.$ (32)
If $T_{j,0}=\emptyset$, the lower bound is trivially true. Combining Eq. 31 and (32), for $Q_{j}=\widehat{Q}_{j}$,
$\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\geq V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\,.$ (33)
Case II: $Q_{j}=Q_{j}^{*}$. So, for any $t\in T_{j,1}$, using Lemma C.5 (1) with $q=Q_{j}^{*}$, we get that $(w_{Q_{j}^{*}},Q_{j}^{*})$ is allocated exactly $Q_{j}^{*}$ units, which is the same as the allocation for $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$. Summing over all rounds in $T_{j,1}$,
$\sum_{t\in T_{j,1}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})={W}_{Q_{j}^{*}}|T_{j,1}|\,.$ (34)
For the rounds in $T_{j,0}$, trivially,
$\sum_{t\in T_{j,0}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})$ $\geq 0\,.$ (35)
Combining Eq. 34 and (35), for $Q_{j}=Q_{j}^{*}$,
$\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\geq{W}_{Q_{j}^{*}}|T_{j,1}|\,.$ (36)
C.3.1.  Proof of Lemma C.5
(1) First we show that for any $t\in[T]$ and $q\leq r_{t}^{*}$, the 1-uniform bid $(w_{q},q)$ is allocated exactly $q$ units.
As $(w_{q},q)\in\mathscr{U}_{1}^{\star}$, it is a safe strategy. By assumption, $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated $r_{t}^{*}$ units in round $t$. Let $\bm{\beta}^{t}=(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right),\bm{\beta}_{-}^{t})$. Recall that, $\bm{\beta}_{-,t}^{-(j)}$ is the $j^{th}$ smallest winning bid in the absence of bids from bidder $i$ for round $t$. If $r_{t}^{*}=0$, the result is vacuously true. Suppose $r_{t}^{*}>0$, then
$\bm{\beta}_{-,t}^{-(q)}\leq\bm{\beta}_{-,t}^{-(r_{t}^{*})}\leq p(\bm{\beta}^{t})\leq w_{r_{t}^{*}}\leq w_{q}\,,$
where the first inequality holds by definition of $\bm{\beta}_{-,t}^{-(j)}$ and our assumption that $q\leq r_{t}^{*}$. For the second inequality, suppose $Q_{\ell-1}^{*}<r_{t}^{*}\leq Q_{\ell}^{*}$ for some $\ell\in[m]$ which by definition implies that $b_{\ell}^{*}$ is the least winning bid. By Lemma A.1,
  1. 1.
    If $p(\bm{\beta}^{t})=b^{*}_{\ell}$, we have $p(\bm{\beta}^{t})\geq\bm{\beta}_{-,t}^{-(r_{t}^{*})}$ as $b_{\ell}^{*}$ is the least winning bid.
  2. 2.
    If $p(\bm{\beta}^{t})=\bm{\beta}_{-,t}^{-(Q^{*}_{\ell}+1)}$, we have $r_{t}^{*}=Q_{\ell}^{*}$ and by definition of $\bm{\beta}_{-,t}^{-(j)}$, $p(\bm{\beta}^{t})=\bm{\beta}_{-,t}^{-(Q^{*}_{\ell}+1)}\geq\bm{\beta}_{-,t}^{-(r_{t}^{*})}$.
The third inequality is true as RoI constraint is satisfied by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ for round $t$, and the fourth is true as $w_{j}$ is a non-decreasing function of $j$ and $q\leq r_{t}^{*}$. From the first and last expressions, $\bm{\beta}_{-,t}^{-(q)}\leq w_{q}$ which implies that $(w_{q},q)$ is allocated at least $q$ units in round $t$. Moreover, $(w_{q},q)$ can be allocated at most $q$ units. Hence, the 1-uniform bid $(w_{q},q)$ is allocated exactly $q$ units.
(2) For any $j\in[m]$, let $T_{j}\subseteq[T]$, defined in Eq. 15, be the rounds in which $b_{j}^{*}$ is the least winning bid. Suppose that $\exists t\in T_{j}$ such that $r_{t}^{*}<Q_{j}^{*}$. We show that in any round $t^{\prime}\in T_{j}$ in which $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated at most $r_{t}^{*}$ units, i.e., $r_{t^{\prime}}^{*}\leq r_{t}^{*}$, the 1-uniform strategy $(w_{r_{t}^{*}},r_{t}^{*})$ is allocated at least $r_{t^{\prime}}^{*}$ units.
Observe that $(w_{r_{t}^{*}},r_{t}^{*})\in\mathscr{U}_{1}^{\star}$, so it is a safe strategy. If $r_{t}^{*}=0$, the result is trivially true. Hence, suppose $r_{t}^{*}>0$ and consider the set of rounds in $T_{j}$ for any $j\in[m]$. As $r_{t}^{*}<Q_{j}^{*}$, by Lemma A.1, $p(\bm{\beta}^{t})=b_{j}^{*}$. So, for any $t^{\prime}\in T_{j}$ such that $r_{t^{\prime}}^{*}\leq r_{t}^{*}$,
$\bm{\beta}_{-,t^{\prime}}^{-(r_{t^{\prime}}^{*})}\leq b_{j}^{*}\leq w_{r_{t}^{*}}\,.$
Here, the first inequality holds as $b_{j}^{*}$ is the least winning bid for round $t^{\prime}\in T_{j}$ and the second one holds as the RoI constraint is true for $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ for round $t$. Hence, $(w_{r_{t}^{*}},r_{t}^{*})$ is allocated at least $r_{t^{\prime}}^{*}$ units in round $t^{\prime}$.
C.4.  Proof of Theorem 5.4
We state and prove a stronger result which recovers the upper bound in Theorem 5.4 as a corollary and is also crucial to prove the upper bound in Theorem 5.5. Let $\mathscr{F}^{\mathcal{H}_{-}}_{m}$ be the class of bidding strategies with at most $m$ bid-quantity pairs that are feasible for the bid history $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$. Then,
Theorem C.6.
For any $m,m^{\prime}\in\mathbb{N}$ such that $m^{\prime}\geq m$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}\leq\frac{m^{\prime}}{m}$.
Proof.
For $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, we define
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-})=\frac{\max_{\mathbf{b}^{\prime}\in\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$
Then, maximizing over all bid histories, we get the desired result.
Let $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$. If $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is a $k$-uniform bidding strategy where $k\leq m\leq m^{\prime}$, by definition, $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)\in\mathscr{F}^{\mathcal{H}_{-}}_{m}$ implying $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-})=1\leq\frac{m^{\prime}}{m}$.
Without loss of generality, let the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$ be $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b^{*}_{1},q^{*}_{1}),\dots,(b^{*}_{m^{\prime}},q^{*}_{m^{\prime}})\rangle$ and $Q_{\ell}^{*}=\sum_{j\leq\ell}q_{j}^{*}$ for all $\ell\in[m^{\prime}]$ [^13] A similar analysis also follows if $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is a $k$-uniform strategy where $m\leq k\leq m^{\prime}$. Let $r_{t}^{*}$ be the number of units allocated to $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in round $t$. Recall that $T_{j}$ is the set of rounds when $b_{j}^{*}$ is the least winning bid when $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is submitted (see Eq. 15). So,
$\sum_{t=1}^{T}V(\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})=\sum_{j=1}^{m^{\prime}}\sum_{t\in T_{j}}V(\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})\,.$
Observe that, for any $t\in T_{j}$,
$V(\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})=V(\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)[1:j];\bm{\beta}_{-}^{t})=\max_{\ell\in[j]}V((b_{\ell}^{*},Q_{\ell}^{*});\bm{\beta}_{-}^{t})\,,$
where recall that $\mathbf{b}[1:j]$ represents the first $j$ bid-quantity pairs and the last equality holds due to Lemma C.3. For any $t\in T_{j}$, $r_{t}^{*}>Q_{j-1}^{*}$. So, $\forall\ell<j$, $V(\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})>V((b_{\ell}^{*},Q_{\ell}^{*});\bm{\beta}_{-}^{t})$ as the demand of the strategy $(b_{\ell}^{*},Q_{\ell}^{*})$ is strictly less than the number of units obtained by $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in that round. Hence, for any $t\in T_{j}$,
$V(\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})$ $=V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})\implies\sum_{t=1}^{T}V(\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})=\sum_{j=1}^{m^{\prime}}\sum_{t\in T_{j}}V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})\,.$ (37)
Let $\mathscr{G}^{\mathcal{H}_{-}}_{m}\subseteq\mathscr{F}^{\mathcal{H}_{-}}_{m}$ be the class of feasible strategies for $\mathcal{H}_{-}$ that contain exactly $m$ bid-quantity pairs. Then,
[…]
Constructing the class of bidding strategies, $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$. We now construct a class of $m$-uniform bidding strategies, $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$ using $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ as a ‘parent’ bidding strategy, in Algorithm 3.
Algorithm 3 Constructing $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$
0:  $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b^{*}_{1},q^{*}_{1}),\dots,(b^{*}_{m^{\prime}},q^{*}_{m^{\prime}})\rangle$. The collection, $\mathscr{C}$, of all subsets $\mathcal{S}\subseteq[m^{\prime}]$ such that $|S|=m$. Initialize $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}=\emptyset$.
1:  for $\mathcal{S}\in\mathscr{C}$ do
2:     Suppose $\mathcal{S}=\{k_{1},\dots,k_{m}\}$ such that ${k_{1}}<{k_{2}}<\dots<{k_{m}}$.
3:     Construct a $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ where
$b_{j}=b_{k_{j}}^{*}\quad\text{and}\quad q_{j}=Q_{k_{j}}^{*}-Q_{k_{j-1}}^{*},~{}$ $\forall j\in[m]\,.$ (39)
4:     $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\leftarrow\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\cup\{\mathbf{b}\}$
5:  end forReturn: The class of bidding strategies, $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$.
For example: suppose $m^{\prime}=4$ and $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(w_{3},3),(w_{7},4),(w_{8},1),(w_{10},2)\rangle$. Let $m=2$ and $\mathcal{S}=\{2,4\}$. Then, $\mathbf{b}=\langle(w_{7},7),(w_{10},3)\rangle$.
Lemma C.7.
Let $b_{(j)}^{*}$ be the $j^{th}$ entry when $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is expressed as a bid vector. Similarly, for any $\mathbf{b}\in\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$, let $b_{(j)}$ be the $j^{th}$ entry when $\mathbf{b}$ is expressed as a bid vector. Then, for any $j\in[Q_{k_{m}}]$, $b_{(j)}\leq b_{(j)}^{*}$. Thus, by Lemma A.2, $\mathbf{b}$ is also feasible for $\mathcal{H}_{-}$ and as $\mathbf{b}$ was chosen arbitrarily, we conclude that $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\subseteq\mathscr{G}^{\mathcal{H}_{-}}_{m}$.
Proof.
Let $\mathcal{S}$ contain $k_{1}<\dots<{k_{m}}$ and $\mathbf{b}$ be constructed per Algorithm 3. Then, for any $j\in[m]$ and any entry $\ell\in(Q_{k_{j-1}}^{*},Q_{k_{j}}^{*}]$,
$b_{(\ell)}^{*}$ $\geq b_{(Q_{k_{j}}^{*})}^{*}=b_{k_{j}}^{*}=b_{(\ell)},$
where the first inequality follows because entries in bid vector are in non increasing order, the first equality follows as $b_{(Q_{r})}^{*}=b_{r}^{*}$ for all $r\in[m^{\prime}]$ and second equality follows from Eq. 39. As $j$ and $\ell$ were picked arbitrarily, we get that for all $j\in[Q_{k_{m}}]$, $b_{(j)}\leq b_{(j)}^{*}$. ∎
By Lemma C.7,
[…]
Consider any $\mathbf{b}\in\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$ obtained from the set $\mathcal{S}=\{k_{1},\dots,k_{m}\}\subseteq[m^{\prime}]$. As $\mathbf{b}$ is feasible for $\mathcal{H}_{-}$, invoking Lemma C.3 for any round $t\in[T]$, we get:
$V(\mathbf{b};\bm{\beta}_{-}^{t})=\max_{\ell\in[m]}V((b_{k_{\ell}}^{*},Q_{k_{\ell}}^{*});\bm{\beta}_{-}^{t})=\max_{\ell\in\mathcal{S}}V((b_{\ell}^{*},Q_{\ell}^{*});\bm{\beta}_{-}^{t}),$ (41)
because for any $\ell\in[m]$, $b_{\ell}=b_{k_{\ell}}^{*}$ and $Q_{\ell}=\sum_{j=1}^{j}q_{j}=\sum_{j=1}^{j}Q_{k_{j}}^{*}-Q_{k_{j-1}}^{*}=Q_{k_{\ell}}^{*}\,.$
Recall that $T_{j}$s, the set of rounds in which $b_{j}^{*}$ is the least winning bid, form a partition of $[T]$. So,
$\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{t=1}^{m^{\prime}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})=\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})+\sum_{j\in[m^{\prime}]\setminus\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})\geq\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (42)
For any $j\in\mathcal{S}$, consider any round $t\in T_{j}$:
$V(\mathbf{b};\bm{\beta}_{-}^{t})$ =(41)${{=}}\max_{\ell\in\mathcal{S}}V((b_{\ell}^{*},Q_{\ell}^{*});\bm{\beta}_{-}^{t})\geq V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})\implies\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})\geq\sum_{t\in T_{j}}V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})$ ) (43)
which implies,
[…]
Note that a bidding strategy $\mathbf{b}\in\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$ is uniquely determined by the set $\mathcal{S}\subseteq[m^{\prime}]$ that generates $\mathbf{b}$. Hence,
$\max_{\mathbf{b}\in\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\max_{\mathcal{S}\subseteq[m^{\prime}]:|\mathcal{S}|=m}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$
(44)${{\geq}}\max_{\mathcal{S}\subseteq[m^{\prime}]:|\mathcal{S}|=m}\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})$ )
$\geq\frac{m}{m^{\prime}}\sum_{j=1}^{m^{\prime}}\sum_{t\in T_{j}}V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})\,.$ (45)
Hence,
[…]
Maximizing over all bid histories, we get that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}\leq\frac{m^{\prime}}{m}$.
Now, we prove that the result also holds for safe bidding strategies. Formally, we show that for $m,m^{\prime}\in\mathbb{N}$ and $m^{\prime}\geq m$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}\leq\frac{m^{\prime}}{m}$. To this end, we consider the quantity $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ analogous to $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-})$, upper bound it and maximize over all bid histories to get the desired result.
For any bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, let $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right):=\mathop{\mathrm{argmax}}_{\mathbf{b}^{\prime}\in\mathscr{U}_{m^{\prime}}^{\star}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime};\bm{\beta}_{-}^{t})$. Without loss of generality, let $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)=\langle(w_{Q_{1}^{*}},q_{1}^{*}),\dots,(w_{Q_{m^{\prime}}^{*}},q_{m^{\prime}}^{*})\rangle$. The proof to bound $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ is similar to that of Theorem C.6 till Eq. 38, where we have
$\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})\leq\frac{\sum_{j=1}^{m^{\prime}}\sum_{t\in T_{j}}V((w_{Q_{j}}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{S}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$
Recall that $\mathscr{S}_{m}$ is the class of $m$-uniform safe bidding strategies. Suppose $\mathscr{S}_{m}(\mathcal{H}_{-})$ be the class that is constructed analogous to the class $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$ obtained in Algorithm 3. We now show that $\mathscr{S}_{m}(\mathcal{H}_{-})\subseteq\mathscr{S}_{m}$.
Lemma C.8.
For any subset $\mathcal{S}\subseteq[m^{\prime}]$ such that $|S|=m$, suppose $\mathbf{b}$ is the bidding strategy obtained from $\mathcal{S}$. Then, $\mathbf{b}$ is a safe bidding strategy. As $\mathcal{S}$ was chosen arbitrarily, $\mathscr{S}_{m}(\mathcal{H}_{-})\subseteq\mathscr{S}_{m}$.
Proof.
Let $\mathcal{S}$ contain $k_{1}<\dots<{k_{m}}$ and $\mathbf{b}$ be constructed per Algorithm 3. Then, for any $j\in[m]$,
$b_{j}=b_{k_{j}}^{*}=w_{Q_{k_{j}}^{*}},\quad\text{and}\quad Q_{j}=Q_{k_{j}}^{*}\,.$
Thus, $\mathbf{b}$ is a safe bidding strategy. ∎
The rest of the proof follows similar to that of Theorem C.6 which gives the desired result.
C.5.  Proof of Theorem 5.5
We need to establish upper bounds on $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}$. From Theorem C.6 and Theorem 5.3, we have:
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}$ $=\frac{\max_{\mathbf{b}^{\prime}\in\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}$
$=\frac{\max_{\mathbf{b}^{\prime}\in\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}^{\prime\prime}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime\prime};\bm{\beta}_{-}^{t})}\cdot\frac{\max_{\mathbf{b}^{\prime\prime}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime\prime};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}$
$=\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}\cdot\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}\leq\frac{2m^{\prime}}{m}\,.$
Appendix D Tight Lower Bounds for Results in Section 5
D.1.  Tight Lower Bound for Theorem 5.3 (For $m=1$)
In this section, we construct a bid history, $\mathcal{H}_{-}$ and valuation curve $\mathbf{v}$ (equivalently the $\mathscr{U}_{m}^{\star}$) for which the upper bound on $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}$, presented in Theorem 5.3, is tight. Recall that for any choice of $\mathcal{H}_{-}$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})\leq 2-\theta_{\mathcal{H}_{-}}$, where $\theta_{\mathcal{H}_{-}}\leq\max_{j\in[m]}({W}_{\widehat{Q}j}/{W}_{Q_{j}^{*}})$. To minimize $\theta_{\mathcal{H}_{-}}$, we choose a valuation curve that is very weakly decreasing. We then set the competing bids such that $\widehat{Q}_{j}\ll Q_{j}^{*}$ for all $j\in[m]$. Finally, we determine the values of $|T_{j,0}|$ and $|T_{j,1}|$ for which the upper bound is tight. See the definition of these quantities in Section C.1, where the first part of the theorem is proven.
We present the case when $m=1$ below. The case for $m\geq 2$ is deferred to Section D.2. Although the main idea for both the cases are the same, for $m\geq 2$, the construction is more involved, and hence presented separately. Formally, for any $\delta\in(0,1/2]$, we design a bid history, $\mathcal{H}_{-}$, and valuation vector, $\mathbf{v}$ for which $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{1},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq 2-\delta$.
Let $M=2\left\lceil\frac{1}{\delta}\right\rceil$. Suppose $\mathbf{v}=[1,v,\cdots,v]\in\mathbb{R}^{M}$, target RoI $\gamma=0$ and $v=1-4\epsilon$ where $\epsilon=\frac{\delta-1/M}{4(1-1/M)}<\frac{\delta}{4}$. Observe that $\epsilon\in(0,1/8)$ as $\delta\leq 1/2$. Let $T=M$ and $K=M+1$. The bid history is defined as:
$\bm{\beta}_{-,t}^{-(j)}=\begin{cases}1-\epsilon,~{}&\text{ if $t\leq M-1$ and $j=1$}\\ C,~{}&\text{ if $t\leq M-1$ and $2\leq j\leq K$}\\ \epsilon,~{}&\text{ if $t=M$ and $1\leq j\leq K$}\\ \end{cases}\,,$
where $C\gg w_{1}$. The bid history is presented in Table 1.
Table 1: Bid history, $\mathcal{H}_{-}$, that achieves the $2$-approximation asymptotically for $m=1$. Here $C\gg w_{1}$ and $\epsilon>0$ is a small real number.
$t=1$ $t=2$cdots $t=M-1$ $t=M$
$C$ $C$cdots $C$ $\epsilon$
$\vdots$ $\vdots$ $\cdots$ $\vdots$ $\vdots$
$C$ $C$cdots $C$vdots
$1-\epsilon$ $1-\epsilon$cdots $1-\epsilon$ $\epsilon$
Notice that the constructed $\mathcal{H}_{-}$ contains $M-1$ rounds with high competition (where the bidder can acquire at most 1 unit), while there is one round with minimal competition, allowing the bidder to obtain any desired number of units. The 1-uniform safe bidding strategies of the form $(w_{q},q)$ do not perform well universally on all rounds because with increase in $q$, despite the increasing demand (i.e., $q$), the bid value $w_{q}$ decreases, thereby reducing the likelihood of acquiring any units.
Notations. For a given bid history $\mathcal{H}_{-}$, recall that for any $m\in\mathbb{N}$, $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$. We define $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)=\sum_{t=1}^{T}V(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})$, the value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ over the bid history $\mathcal{H}_{-}$. Similarly, $\mathbf{b}_{m}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right):=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ and $V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)=\sum_{t=1}^{T}V(\mathbf{b}_{m}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})$.
Computing the optimal strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{1}$. We observe that no bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{1}$ can obtain more than 1 unit in the first $M-1$ rounds and $M$ units in the final round.
We claim that $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=(1,M)$. Observe that $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated 1 unit in each of the first $M-1$ rounds and $M$ units in the final round. Hence, it is allocated the maximum number of units possible. To verify that $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is feasible for the bid history, notice that for $\forall t\in[M-1]$, $p(\bm{\beta}^{t})=1\leq w_{1}=1$. For round $t=M$, as $K>M$,
$p(\bm{\beta}^{t})=\epsilon\leq 1-4\epsilon<w_{M}=\mu\cdot 1+(1-\mu)\cdot(1-4\epsilon),\quad\text{where}\quad\mu=\frac{1}{M},$
implying that RoI constraint is satisfied by $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in round $t=T$. So, $V^{\mathsf{OPT}}_{1}\left(\mathcal{H}_{-}\right)=M+v(M-1)$.
Computing the optimal strategy in $\mathscr{U}_{1}^{\star}$. Here, we argue that the optimal strategy in $\mathscr{U}_{1}^{\star}$ is $(w_{1},1)=(1,1)$. Observe that the strategy $(1,1)$ is allocated 1 unit in each round. Hence, it obtains a total value of $M$.
As $\bm{\beta}_{-,t}^{-(1)}=1-\epsilon>1-2\epsilon=w_{2}$, for $t=1,2,\dots,M-1$, bidding $\left(w_{q},q\right)$ for $q\geq 2$ does not get any value in the first $M-1$ rounds. Bidding $\left(w_{q},q\right)$ gets exactly $q$ units in round $M$ as $w_{q}>\epsilon$ for any $q\geq 2$. So, for $2\leq q\leq M$, the total value obtained by bidding $(w_{q},q)$ is $1+(q-1)v<q$. Hence, $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)=M$, which is the value obtained by the safe strategy $(w_{1},1)=(1,1)$. So,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{1};\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})=\frac{V^{\mathsf{OPT}}_{1}\left(\mathcal{H}_{-}\right)}{V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}=\frac{M+v(M-1)}{M}=\frac{2M-1-4\epsilon(M-1)}{M}=2-\delta\,.$
D.2.  Tight Lower Bound for Theorem 5.3 (For $m\geq 2$)
In this section, we design a bid history $\mathcal{H}_{-}$ and valuation vector, $\mathbf{v}$ such that for any $\delta\in(0,1/2]$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq 2-\delta$. By definition, for any $\mathcal{H}_{-}$ and $m\geq 2$,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})=\frac{V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)}{V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)}\geq\frac{V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)}{mV^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}\,,$ (46)
where the second inequality follows because $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}\leq m$ (by Theorem 5.4). So, instead of computing $V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)$ directly which requires solving a DP, we obtain $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$ and show that the bound is tight.
D.2.1.  Construction of $\mathcal{H}_{-}$.
We first decide all the parameters.
  • Fix $m\geq 2$ and any integer $N\geq 2\left\lceil\frac{1}{\delta}\right\rceil$.
  • Let $M=N^{2m-1}$. Consider $T=N^{2m-1}$ rounds and $K=N^{2m-1}+1$ units in each auction.
  • Let $\epsilon^{\prime}=\frac{m\delta/(2m-1)-1/N}{2(1-1/N)}<\frac{\delta}{2}\leq\frac{1}{4}$. Set $\epsilon$ such that $\epsilon^{\prime}=\epsilon N^{2m-1}(N^{2m-1}+1)$.
Consider a valuation vector $\mathbf{v}=[1,v,\cdots,v]$ such that $v=1-2\epsilon^{\prime}$, and target RoI $\gamma=0$. Partition the $N^{2m-1}$ rounds into $2m$ partitions such that the first partition has $1$ round and the $j^{th}$ partition has $N^{j-1}-N^{j-2}$ rounds for $2\leq j\leq 2m$. Each partition has identical competing bid profile submitted by other bidders. In particular,
  1. 1.
    The first partition (containing one round) has all the bids submitted by others as $w_{N^{2m-1}+1}+\epsilon$.
  2. 2.
    If $j>1$ and $j$ is odd, for the $j^{th}$ partition (of size $N^{j-1}-N^{j-2}$), the smallest $N^{2m-j}+1$ winning competing bids are $w_{N^{2m-j}+1}+\epsilon$ and the remaining bids are $C\gg w_{1}$.
  3. 3.
    If $j>1$ and $j$ is even, for the $j^{th}$ partition (of size $N^{j-1}-N^{j-2}$), the smallest $N^{2m-j}$ winning competing bids are $w_{N^{2m-j}+1}+\epsilon$ and the remaining bids are $C\gg w_{1}$.
We present an example for such a bid history in Table 2.
Table 2: Example of bid history, $\mathcal{H}_{-}$, achieving 2-approximation asymptotically for $m=2$. Each round in the same partition has identical competing bid profile. Total number of units in each auction is $K=N^{3}+1$.
Partition 1 Partition 2 Partition 3 Partition 4
$t=1$ $t\in[2,N]$ $t\in[N+1,N^{2}]$ $t\in[N^{2}+1,N^{3}]$
00 bids are $C$ $N^{3}-N^{2}+1$ bids are $C$ $N^{3}-N$ bids are $C$ $N^{3}$ bids are $C$
$N^{3}+1$ bids are $w_{N^{3}+1}+\epsilon$ $N^{2}$ bids are $w_{N^{2}+1}+\epsilon$ $N+1$ bids are $w_{N+1}+\epsilon$ 1 bid is $w_{2}+\epsilon$
D.2.2.  Computing $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$.
We make the following claim about the optimal $m$-uniform bidding strategy for the constructed $\mathcal{H}_{-}$.
Lemma D.1.
For the aforementioned $\mathcal{H}_{-}$, $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ where
$(b_{j},q_{j})=\begin{cases}\left(1,N\right),&~{}\text{ if }j=1\\ \left(w_{N^{2j-2}},N^{2j-1}-N^{2j-3}\right),&~{}\text{ if }2\leq j\leq m\,.\end{cases}$ (47)
Furthermore,
$V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)=N^{2m-1}+(2m-1)(N^{2m-1}-N^{2m-2})v\,.$
Proof.
We begin by a crucial observation that the bid history does not allow obtaining more than $N^{2m-j}$ units in the $j^{th}$ partition while satisfying the RoI constraint, irrespective of the number of bids submitted by the bidder. To verify this, note that, the maximum number of units that can be allocated to any bidding strategy in the $j^{th}$ partition is either $N^{2m-j}$ or $N^{2m-j}+1$ (depending on if $j$ is even or odd). Suppose contrary to our claim, the bidder is allocated $N^{2m-j}+1$ units in the some round $t$ in the $j^{th}$ partition by bidding some $\mathbf{b}$. Let $\bm{\beta}^{t}=(\mathbf{b},\bm{\beta}_{-}^{t})$ be the complete bid profile. So, $p(\bm{\beta}^{t})\geq w_{N^{2m-j}+1}+\epsilon$ but $x(\bm{\beta}^{t})=N^{2m-j}+1$ indicating that the RoI constraint is violated, which verifies our claim.
So, the total number of units, $N_{\text{total}}$, that can be obtained by the bidder over all the rounds is:
$N_{\text{total}}\leq N^{2m-1}+\sum_{j=2}^{2m}N^{2m-j}(N^{j-1}-N^{j-2})=2mN^{2m-1}-(2m-1)N^{2m-2}=:N_{\max}.$
Now, we compute the the number of units obtained by bidding $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ and show that it is allocated $N_{\max}$ units for the constructed bid history, demonstrating that it is the optimal bidding strategy.
Consider any auction in the $j^{th}$ partition. The lowest winning bid in the bid profile is $w_{N^{2m-j}+1}+\epsilon$. Note that the unique bid values (ignoring the quantity for the sake of brevity) in $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ are $\mathbf{b}=\left\{1,w_{N^{2}},\dots,w_{N^{2m-2}}\right\}$. We claim that the winning bid values of $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in the $j^{th}$ partition are $\widehat{\mathbf{b}}=\left\{1,w_{N^{2}},\dots,w_{N^{2m+2\left\lfloor-j/2\right\rfloor}}\right\}$. This is true because the least bid value in $\widehat{\mathbf{b}}$ is greater than $w_{N^{m-j}+1}+\epsilon$, i.e.,
$w_{N^{2m+2\left\lfloor-j/2\right\rfloor}}-(w_{N^{2m-j}+1}+\epsilon)$ $\geq w_{N^{2m-j}}-(w_{N^{2m-j}+1}+\epsilon)$
$=\frac{1-v}{N^{2m-j}(N^{2m-j}+1)}-\epsilon=\frac{2\epsilon N^{2m-1}(N^{2m-1}+1)}{N^{2m-j}(N^{2m-j}+1)}-\epsilon\geq\epsilon>0\,.$
Let $N_{j}$ denote the number of units allocated to $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in each auction in the $j^{th}$ partition. There are two cases:
  1. (a)
    for $j$ odd, recall that for the $j^{th}$ partition (of size $N^{j-1}-N^{j-2}$), the smallest $N^{2m-j}+1$ winning competing bids are $w_{N^{2m-j}+1}+\epsilon$ and the remaining bids are $C\gg w_{1}$. Then,
    $N_{j}=N+\sum_{\ell=2}^{m-\frac{j-1}{2}}(N^{2\ell-1}-N^{2\ell-3})=N^{2m-j}\,.$
  2. (b)
    For $j$ even, recall that for the $j^{th}$ partition (of size $N^{j-1}-N^{j-2}$), the smallest $N^{2m-j}$ winning competing bids are $w_{N^{2m-j}+1}+\epsilon$ and the remaining bids are $C\gg w_{1}$.
    $N_{j}=\min\left\{N^{2m-j},N+\sum_{\ell=2}^{m+1-\frac{j}{2}}(N^{2\ell-1}-N^{2\ell-3})\right\}=\min\{N^{2m-j},N^{2m-j+1}\}=N^{2m-j}\,.$
Here, the minimum is taken over two quantities where the first quantity is the number of competing bids less than $C$ in any round $t$ in the $j^{th}$ partition and the second quantity represents the total demand of the winning bids in $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ for that round. So, the total number of units obtained across all rounds is
$N^{2m-1}+\sum_{j=2}^{2m}N^{2m-j}(N^{j-1}-N^{j-2})=2mN^{2m-1}-(2m-1)N^{2m-2}.$
As this is the maximum number of units that can be obtained by the bidder, $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is optimal. The total value obtained by bidding $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is
$V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ $=1+(N^{2m-1}-1)v+\sum_{j=2}^{2m}(N^{j-1}-N^{j-2})(1+(N^{2m-j}-1)v)$
$=N^{2m-1}+(2m-1)(N^{2m-1}-N^{2m-2})v\,.$
D.2.3.  Computing $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$
Recall that we compute $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$ and invoke the bounds on $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}$, instead of directly evaluating $V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)$.
Lemma D.2.
For the aforementioned $\mathcal{H}_{-}$, $\mathbf{b}_{1}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)=(1,1)$ and $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)=N^{2m-1}$.
Proof.
The basic idea is to enumerate the total units (value) that can be obtained by bidding $(w_{q},q)$ for $q\in[N^{2m-1}]$ and then finding the maximum of those values. As $q$ can be exponential in $m$, we exploit the structure of the bid history to compute the objective in an efficient manner.
Suppose $q=1$. The maximum number of units (value) that can be obtained by bidding $(1,1)$ is trivially $N^{2m-1}$, So, $(1,1)$ obtains a total value $N^{2m-1}$.
Suppose $q\geq 2$. Furthermore, assume $N^{2m-j}<q\leq N^{2m-j+1}$, for some $2\leq j\leq 2m$. Consider any bid of the form $(w_{q},q)$. Bidding $(w_{q},q)$ does not obtain any units in the partitions indexed by $j,j+1,\dots,2m$ as $w_{q}$ is strictly less than the least winning competing bids in those partitions, i.e., $w_{q}\leq w_{N^{2m-j}+1}<w_{N^{2m-\ell}+1}+\epsilon$, for any $\ell\in\{j,j+1,\dots,2m\}$. So, if $N^{2m-j}<q\leq N^{2m-j+1}$, $(w_{q},q)$ gets no units in
$\sum_{\ell=j}^{2m}N^{\ell-1}-N^{\ell-2}=N^{2m-1}-N^{j-2}\text{ auctions.}$.
In the remaining $N^{j-2}$ auctions it can win at most $q$ units. So, the maximum value obtained by $(w_{q},q)$ for any $q\geq 2$ is
$N^{j-2}(1+(q-1)v)$ $\leq N^{j-2}(1+(N^{2m-j+1}-1)v)=N^{j-2}+(N^{2m-1}-N^{j-2})v<N^{2m-1},$
where the last inequality holds as $v<1$. Hence, $\mathbf{b}_{1}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)=(1,1)$ and $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)=N^{2m-1}$. ∎
[…]
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)}{mV^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}$ $=\frac{N^{2m-1}+(2m-1)(N^{2m-1}-N^{2m-2})v}{mN^{2m-1}}$
$=\frac{2mN^{2m-1}-(2m-1)N^{2m-2}-2\epsilon^{\prime}(2m-1)(N^{2m-1}-N^{2m-2})}{mN^{2m-1}}$
$=2-\frac{2m-1}{m}\Big{(}\frac{1}{N}+2\epsilon^{\prime}\Big{(}1-\frac{1}{N}\Big{)}\Big{)}$
$=2-\delta\,,$
where the last step follows by substituting the value of $\epsilon^{\prime}$ and thus concludes the proof.
D.3.  Tight Lower Bound for Theorem 5.4
In this section, we construct a bid history $\mathcal{H}_{-}$ and valuation vector, $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{m^{\prime}}{m}-\delta$, i.e., the upper bound is tight for $m^{\prime}\geq m+1$ (If $m^{\prime}=m$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}=1$ implying the bound is tight).
Recall that Theorem 5.4 states that $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}\leq m$ for any $m\geq 1$. So,
$\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})=\frac{V^{\mathsf{SAFE}}_{m^{\prime}}\left(\mathcal{H}_{-}\right)}{V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)}\geq\frac{V^{\mathsf{SAFE}}_{m^{\prime}}\left(\mathcal{H}_{-}\right)}{mV^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}=\frac{1}{m}\cdot\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})\,.$ (48)
We now establish a lower bound on $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$.
D.3.1.  Construction of $\mathcal{H}_{-}$.
We first decide all the parameters.
  • Fix $m^{\prime}\geq 2$ and any integer $N\geq\left\lceil\frac{m^{\prime}}{\delta}\right\rceil$.
  • Let $M=N^{m^{\prime}-1}$. Consider $T=N^{m^{\prime}-1}$ rounds and $K=N^{m^{\prime}-1}$ units in each auction.
  • Let $\epsilon^{\prime}=\frac{m\delta/(m^{\prime}-1)-1/N}{2(1-1/N)}<\frac{m\delta}{2(m^{\prime}-1)}\leq\frac{1}{4}$. Set $\epsilon$ such that $\epsilon^{\prime}=\epsilon N^{m^{\prime}-1}(N^{m^{\prime}-1}+1)$.
Let the valuation vector be $\mathbf{v}=[1,v,\cdots,v]$ such that $v=1-2\epsilon^{\prime}$, and target RoI, $\gamma=0$. Partition the $N^{m^{\prime}-1}$ rounds into $m^{\prime}$ partitions such that the first partition has $1$ round and the $j^{th}$ partition has $N^{j-1}-N^{j-2}$ rounds for $2\leq j\leq m^{\prime}$. Each partition has identical competing bid profile submitted by the other bidders. In particular,
  1. 1.
    The first partition (containing one round) has all the competing bids to be $w_{N^{m^{\prime}-1}+1}+\epsilon$.
  2. 2.
    For $2\leq j\leq m^{\prime}$, the $j^{th}$ partition (of size $N^{j-1}-N^{j-2}$); the smallest $N^{m^{\prime}-j}+1$ competing winning bids are $w_{N^{m^{\prime}-j}+1}+\epsilon$ and the remaining bids are $C\gg w_{1}$.
We present an example for such a bid history in Table 3.
Table 3: Example of bid history, $\mathcal{H}_{-}$, achieving 4-approximation asymptotically for $m^{\prime}=4$ (and $m=1$). Each round in the same partition has identical competing bid profile. Total number of units in each auction is $K=N^{3}$.
Partition 1 Partition 2 Partition 3 Partition 4
$t=1$ $t\in[2,N]$ $t\in[N+1,N^{2}]$ $t\in[N^{2}+1,N^{3}]$
00 bids are $C$ $N^{3}-N^{2}-1$ bids are $C$ $N^{3}-N-1$ bids are $C$ $N^{3}-2$ bids are $C$
$N^{3}$ bids are $w_{N^{3}+1}+\epsilon$ $N^{2}+1$ bids are $w_{N^{2}+1}+\epsilon$ $N+1$ bids are $w_{N+1}+\epsilon$ 2 bids are $w_{2}+\epsilon$
D.3.2.  Computing $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
We make the following claim about the optimal $m^{\prime}$-uniform safe bidding strategy for $\mathcal{H}_{-}$.
Lemma D.3.
For the aforementioned $\mathcal{H}_{-}$, $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)=\langle(b_{1},q_{1}),\dots,(b_{m}^{\prime},q_{m}^{\prime})\rangle$ where
$(b_{j},q_{j})=\begin{cases}\left(1,1\right),&~{}\text{if }j=1\\ \left(w_{N^{j-1}},N^{j-1}-N^{j-2}\right),&~{}\text{if }2\leq j\leq m^{\prime}\end{cases}\,.$ (49)
Furthermore,
$V^{\mathsf{SAFE}}_{m^{\prime}}\left(\mathcal{H}_{-}\right)=N^{m^{\prime}-1}+(m^{\prime}-1)(N^{m^{\prime}-1}-N^{m^{\prime}-2})v\,.$
Proof.
We begin by a crucial observation that the bid history does not allow obtaining more than $N^{m^{\prime}-j}$ units in the $j^{th}$ partition, while satisfying the RoI constraint, irrespective of the number of bids submitted by the bidder. To verify this, suppose contrary to our claim, that the bidder is allocated $N^{m^{\prime}-j}+1$ units in the some round $t$ in the $j^{th}$ partition by bidding some $\mathbf{b}$. Let $\bm{\beta}^{t}=(\mathbf{b},\bm{\beta}_{-}^{t})$ be the complete bid profile. Hence, $p(\bm{\beta}^{t})\geq w_{N^{m^{\prime}-j}+1}+\epsilon$ but $x(\bm{\beta}^{t})=N^{m^{\prime}-j}+1$ indicating that the RoI constraint is violated which verifies our claim.
So, the total number of units, $N_{\text{total}}$, that can be obtained by the bidder across all the rounds is:
$N_{\text{total}}\leq N^{m^{\prime}-1}+\sum_{j=2}^{m^{\prime}}N^{m^{\prime}-j}(N^{j-1}-N^{j-2})=m^{\prime}N^{m^{\prime}-1}-(m^{\prime}-1)N^{m^{\prime}-2}:=N_{\max}.$
Now, we compute the number of units obtained by bidding $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$ and show that it is allocated $N_{\max}$ units across all the auctions, implying that it is the optimal bidding strategy.
To show $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$ is optimal, consider any auction in the $j^{th}$ partition. The lowest winning competing bid is $w_{N^{m^{\prime}-j}+1}+\epsilon$. Note that the unique bid values (ignoring the quantity for the sake of brevity) in $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$, provided in Eq. 49, are $\mathbf{b}=\left\{1,w_{N},\dots,w_{N^{m^{\prime}-1}}\right\}$. We claim that the winning bid values of $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$ in the $j^{th}$ partition are $\widetilde{\mathbf{b}}=\left\{1,w_{N},\dots,w_{N^{m^{\prime}-j}}\right\}$. Again recall that for $2\leq j\leq m^{\prime}$, for the $j^{th}$ partition (of size $N^{j-1}-N^{j-2}$), the smallest $N^{m^{\prime}-j}+1$ competing winning bids are $w_{N^{m^{\prime}-j}+1}+\epsilon$ and the remaining bids are $C\gg w_{1}$. And here, the least bid value in $\widetilde{\mathbf{b}}$ is greater than $w_{N^{m^{\prime}-j}+1}+\epsilon$, i.e.,
$w_{N^{m^{\prime}-j}}-(w_{N^{m^{\prime}-j}+1}+\epsilon)=\frac{1-v}{N^{m^{\prime}-j}(N^{m^{\prime}-j}+1)}-\epsilon=\frac{2\epsilon N^{m^{\prime}-1}(N^{m^{\prime}-1}+1)}{N^{m^{\prime}-j}(N^{m^{\prime}-j}+1)}-\epsilon\geq\epsilon>0\,.$
Moreover, observe that the bidder is allocated the maximum number of units demanded for each of the bid value in $\widetilde{\mathbf{b}}$. So, the number of units in each auction in the $j^{th}$ partition by bidding $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$ is
$N_{j}=1+\sum_{\ell=2}^{m^{\prime}-j+1}(N^{\ell-1}-N^{\ell-2})=N^{m^{\prime}-j}\,.$
Hence, the total number of units obtained across all rounds is
$N^{m^{\prime}-1}+\sum_{j=2}^{m^{\prime}}N^{m^{\prime}-j}(N^{j-1}-N^{j-2})=m^{\prime}N^{m^{\prime}-1}-(m^{\prime}-1)N^{m^{\prime}-2}=N_{\max}.$
As this is the maximum number of units that can be obtained by the bidder, $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$ is optimal. The total value obtained by bidding $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$ is given by
$V^{\mathsf{SAFE}}_{m^{\prime}}\left(\mathcal{H}_{-}\right)$ $=1+(N^{m^{\prime}-1}-1)v+\sum_{j=2}^{m^{\prime}}(N^{j-1}-N^{j-2})(1+(N^{m^{\prime}-j}-1)v)$
$=N^{m^{\prime}-1}+(m^{\prime}-1)(N^{m^{\prime}-1}-N^{m^{\prime}-2})v\,.$
D.3.3.  Computing $\mathbf{b}_{1}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
In this section, we compute the optimal 1-uniform safe bidding strategy for the bid history, $\mathcal{H}_{-}$.
Lemma D.4.
For the constructed $\mathcal{H}_{-}$, $\mathbf{b}_{1}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)=\left(1,1\right)$ and $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)=N^{m-1}$.
Proof.
The proof is similar to that of Lemma D.2 in the sense that it involves enumerating the total units obtaining by submitting $(w_{q},q)$ for all $q\in[N^{m-1}]$ in an efficient manner by leveraging the structure in the bid history. ∎
Substituting values from Lemma D.3, Lemma D.4 and Eq. 48
$\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{1}{m}\cdot\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ $=\frac{V^{\mathsf{SAFE}}_{m^{\prime}}\left(\mathcal{H}_{-}\right)}{mV^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}$
$=\frac{N^{m^{\prime}-1}+(m^{\prime}-1)(N^{m^{\prime}-1}-N^{m^{\prime}-2})v}{mN^{m^{\prime}-1}}$
$=\frac{m^{\prime}N^{m^{\prime}-1}-(m^{\prime}-1)N^{m^{\prime}-2}-2\epsilon^{\prime}(m^{\prime}-1)(N^{m^{\prime}-1}-N^{m^{\prime}-2})}{mN^{m^{\prime}-1}}$
$=\frac{m^{\prime}}{m}-\frac{(m^{\prime}-1)}{m}\Big{(}\frac{1}{N}+2\epsilon^{\prime}\Big{(}1-\frac{1}{N}\Big{)}\Big{)}$
$=\frac{m^{\prime}}{m}-\delta\,,$
where the last step follows by substituting the value of $\epsilon^{\prime}$ and thus concludes the proof.
Corollary D.5.
As $\mathscr{U}_{m}^{\star}\subseteq\mathscr{F}^{\mathcal{H}_{-}}_{m}$ for all $m\in\mathbb{N}$, the upper bound in Theorem C.6 is also tight, i.e., for any $m^{\prime}\geq m$ and $\delta\in(0,\frac{1}{2}]$, there exists a bid history $\mathcal{H}_{-}$ and valuation curve $\mathbf{v}$ such that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{m^{\prime}}{m}-\delta$.
D.4.  Tight Lower Bound for Theorem 5.5
In this section, we present a bid history $\mathcal{H}_{-}$ and valuation vector, $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{2m^{\prime}}{m}-\delta$, i.e., the upper bound is tight for $m^{\prime}\geq m$.
Recall that Theorem 5.4 states that $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}\leq m$ for any $m\geq 1$. So,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})=\frac{V^{\mathsf{OPT}}_{m^{\prime}}\left(\mathcal{H}_{-}\right)}{V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)}\geq\frac{V^{\mathsf{OPT}}_{m^{\prime}}\left(\mathcal{H}_{-}\right)}{mV^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}=\frac{1}{m}\cdot\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})\,.$ (50)
Recall that for showing tightness of the bound for Theorem 5.3 for any $m\geq 2$, we computed a lower bound on $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ (cf. Eq. 46). Thus, we consider a valuation curve and a bid history that has a structure identical to the one presented in Section D.2 but with the following modified parameters,
  • Fix $m^{\prime}\geq 1$ and any integer $N\geq 2\left\lceil\frac{2m^{\prime}}{\delta}\right\rceil$.
  • Let $M=N^{2m^{\prime}-1}$. Consider $T=N^{2m^{\prime}-1}$ rounds and $K=N^{2m^{\prime}-1}+1$ units in each auction.
  • Let $\epsilon^{\prime}=\frac{m\delta/(2m^{\prime}-1)-1/N}{2(1-1/N)}<\frac{m\delta}{2(2m^{\prime}-1)}\leq\frac{1}{4}$. Set $\epsilon$ such that $\epsilon^{\prime}=\epsilon N^{2m^{\prime}-1}(N^{2m^{\prime}-1}+1)$.
[…]
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{1}{m}\cdot\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ $=\frac{N^{2m^{\prime}-1}+(2m^{\prime}-1)(N^{2m^{\prime}-1}-N^{2m^{\prime}-2})v}{mN^{2m^{\prime}-1}}$
$=\frac{2m^{\prime}N^{2m^{\prime}-1}-(2m^{\prime}-1)N^{2m^{\prime}-2}-2\epsilon^{\prime}(2m^{\prime}-1)(N^{2m^{\prime}-1}-N^{2m^{\prime}-2})}{mN^{2m^{\prime}-1}}$
$=\frac{2m^{\prime}}{m}-\frac{(2m^{\prime}-1)}{m}\Big{(}\frac{1}{N}+2\epsilon^{\prime}\Big{(}1-\frac{1}{N}\Big{)}\Big{)}$
$=\frac{2m^{\prime}}{m}-\delta\,,$
where the last step follows by substituting the value of $\epsilon^{\prime}$ and thus concludes the proof.
Appendix E Simulation Details
In Section 5.3, we conducted numerical experiments to estimate the richness ratio (or equivalently $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}$) in practice for the classes of bidding strategies discussed earlier. We now discuss a few details regarding the same.
Dataset. We estimate the richness ratio for the EU ETS emission permit auction data for 2022 and 2023. However, only aggregate statistics of the submitted bids is publicly available for privacy reasons. Hence, we synthesize individual level bid data from these available statistics. The exact procedure to reconstruct the bid data is presented in Section E.1. We sample the values from the $\text{Unif}[0,1]$ distribution. In each simulation, we sample $T\sim\text{Unif}[100,300]$ auctions and let $M\sim\text{Unif}[10,80]$. We vary $m=1$ to $m=10$ and average over 100 simulations to obtain plots in Fig. 4. As computing $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ (the value obtained by the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m}$) can be non-trivial, we obtain a uniform upper bound for $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ that is independent of $m$ (see details in Section E.2).
E.1.  Reconstructing Individual Bid Data
We obtained the publicly available auction data for $T_{\max}=443$ EU ETS emission permit auctions held in 2022 and 2023 (EEX, 2023). For each auction indexed by $t\in[T_{\max}]$, we have the following relevant information: the minimum bid ($b_{\min}^{t}$), the maximum bid ($b_{\max}^{t}$), the average of the bids ($b_{\text{avg}}^{t}$), the median of the bids ($b_{\text{med}}^{t}$), and the number of bid-quantity pairs submitted ($n_{\text{sub}}^{t}$). We normalized the bids to be in $[0,1]$. For all rounds $t$, $b_{\text{avg}}^{t}\approx b_{\text{med}}^{t}$ (linear regression yields coefficient 1.01 and intercept $-0.008$).
Upon further investigation, we observed that, except a few, a significant number of auctions had either $b_{\min}^{t}\approx b_{\text{avg}}^{t}\ll b_{\max}^{t}$ (Type I) or $b_{\min}^{t}\ll b_{\text{avg}}^{t}\approx b_{\max}^{t}$ (Type II). As $b_{\text{avg}}^{t}\approx b_{\text{med}}^{t},\forall t$, we deduce that for Type I, most of the bids are concentrated in the interval $[b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\min}^{t}]$ whereas for Type II, most of the bids are in the interval $[2b_{\text{avg}}^{t}-b_{\max}^{t},b_{\max}^{t}]$. We posit that for Type I (resp. Type II) auctions, $f\in(0,1)$ fraction of the bids ($n_{\text{sub}}^{t}$) are in $[b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\min}^{t}]$ (resp. $[2b_{\text{avg}}^{t}-b_{\max}^{t},b_{\max}^{t}]$) and the $1-f$ fraction of bids are in $[2b_{\text{avg}}^{t}-b_{\min}^{t},b_{\max}^{t}]$ (resp. $[b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\max}^{t}]$). If for Type I (resp. Type II) auctions, $2b_{\text{avg}}^{t}>b_{\min}^{t}+b_{\max}^{t}$ (resp. $2b_{\text{avg}}^{t}<b_{\min}^{t}+b_{\max}^{t}$), we assume that all the bids are uniformly present in the interval $[b_{\min}^{t},b_{\max}^{t}]$. With these assumptions, we generate individual bid data for each auction by sampling uniformly from these intervals.
After generating the individual bid data, we compute the metrics for the reconstructed bids (say $\widehat{b}_{\text{avg}}^{t}$) for each auction and reject those with a relative error of at least $\delta$ (tolerance). For our simulations, we set $\delta=0.05$. We perform a grid search for $f$ to maximize the number of auctions where the metrics of the reconstructed data are within $\delta$ relative error of the actual metrics, and obtain that $f=0.97$. Following this pre-processing, we have reconstructed individual bid data for $T=341$ auctions. The bids are normalized to be in $[0,1]$.
E.2.  An Uniform Upper Bound for $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right).$
For any bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, suppose $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated $r_{t}$ units in any round $t$. Then, by Lemma C.5 (1), we know that $(w_{r_{t}},r_{t})$ also obtains $r_{t}$ units in round $t$. So,
$V(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})$ $=V((w_{r_{t}},r_{t});\bm{\beta}_{-}^{t})\leq\max_{\mathbf{b}\in\mathscr{U}_{1}^{\star}}V(\mathbf{b};\bm{\beta}_{-}^{t})$
$\implies V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ $=\sum_{t=1}^{T}V(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})\leq\sum_{t=1}^{T}\max_{\mathbf{b}\in\mathscr{U}_{1}^{\star}}V(\mathbf{b};\bm{\beta}_{-}^{t})\,.$
Appendix F Resolving Ties
We first emphasize that the effect of resolution of ties on the objective function under the value maximization behavioral model is fundamentally different from the quasilinear utility maximization model. To illustrate this, consider a single-item second price auction with two bidders each valuing the item equally at $v$. Assume that both the bidders bid $v$ and there exists a definitive tie breaking rule (either randomized or deterministic). If the bidders are considered quasilinear utility maximizers, the objective function value (value obtained minus the payments) for both the bidders is 0. However, under the value maximization model (assume $\gamma=0$ for both the bidders), the objective function value for the winning bidder is $v$ and 0 for the losing bidder. [^14] Although, the total (liquid) welfare is $v$ in both the cases. Hence, we need to carefully analyze the key results of our work in the event of ties.
In the context of multi-unit auctions, we can classify ties into two informal types: (a) ‘good ties’—ties occurring at any bid other than the last accepted bid (LAB) and (b) ‘bad ties’, which are the ties occurring at the LAB. All the results in this work are unaffected in case only ‘good ties’ occur. Thus, we focus on ‘bad ties’ for the rest of this section. For this discussion, we consider a public, deterministic tie breaking rule under which ties are always broken in favor of lower indexed bidder. In the presence of ties, the undominated class of safe bidding strategies is still the class of nested strategies (Theorem 3.2). The tie breaking rule is incorporated into the decomposition in Lemma 4.1 and computing the edge weights in the DAG in the offline and online settings in Section 4, ensuring that the maximum weight path in the DAG gives the optimal offline solution and Algorithm 1 achieves sublinear regret in the online setting.

### v4

@@
-Learning Safe Strategies for Value Maximizing Buyers in Uniform Price Auctions
+Negin Golrezaei
 Abstract
-We study the bidding problem in repeated uniform price multi-unit auctions from the perspective of a single value-maximizing buyer who aims to maximize their cumulative value over $T$ rounds while adhering to return-on-investment (RoI) constraints in each round.
-Buyers adopt $m$-uniform bidding format, where they submit $m$ bid-quantity pairs $(b_{i},q_{i})$ to demand $q_{i}$ units at bid $b_{i}$.
-We introduce safe bidding strategies as those that satisfy RoI constraints in every auction, regardless of competing bids.
-We show that these strategies depend only on the bidder’s valuation curve, and the bidder can focus on a finite subset of this class without loss of generality.
-While the number of strategies in this subset is exponential in $m$, we develop a polynomial-time algorithm to learn the optimal safe strategy that achieves sublinear regret in the online setting, where regret is measured against a clairvoyant benchmark that knows the competing bids a priori and selects a fixed hindsight optimal safe strategy.
-We then evaluate the performance of safe strategies against a clairvoyant that selects the optimal strategy from a richer class of strategies in the online setting.
-In this scenario, we compute the richness ratio, $\alpha\in(0,1]$ for the class of strategies chosen by the clairvoyant and show that our algorithm, designed to learn safe strategies, achieves $\alpha$-approximate sublinear regret against these stronger benchmarks.
-Experiments on semi-synthetic data from real-world auctions show that safe strategies substantially outperform the derived theoretical bounds, making them quite appealing in practice.
+We study the bidding problem in repeated uniform price multi-unit auctions, a format prevalent in emission permit auctions, Treasury auctions and energy markets, from the perspective of a value-maximizing buyer.
+The buyer aims to maximize their cumulative value over $T$ rounds while adhering to per-round return-on-investment (RoI) constraints in a strategic (or adversarial) environment.
+The buyer uses $m$-uniform bidding format, submitting $m$ bid-quantity pairs $(b_{i},q_{i})$ to demand $q_{i}$ units at bid $b_{i}$, with $m\ll M$ in practice, where, $M$ is the maximum demand of the buyer.
+We introduce the notion of safe bidding strategies as those that satisfy the RoI constraint irrespective of competing bids.
+Despite the stringent requirement, we show that these strategies satisfy a mild no-overbidding condition, depend only on the bidder’s valuation curve, and the bidder can focus on a finite subset without loss of generality.
+While the number of strategies in this subset is $O(M^{m})$, leveraging a key insight on the decomposition of a strategy’s obtained value across its constituent bid-quantity pairs, we develop a polynomial-time algorithm to learn the optimal safe strategy that achieves a regret of $\widetilde{O}(M\sqrt{mT})$ and $\widetilde{O}(M^{2}m^{3/2}\sqrt{T})$ in the full-information and bandit settings respectively, where regret is measured against a clairvoyant that selects the fixed hindsight optimal safe strategy.
+We complement the upper bounds by establishing a $O(M\sqrt{T})$ regret lower bound.
+We then assess the robustness of safe strategies against the hindsight-optimal strategy from a richer class of strategies.
+We define the richness ratio $\alpha\in(0,1]$ as the minimum ratio of the value of the optimal safe strategy to that of the optimal strategy from richer class and construct hard instances showing the tightness of $\alpha$.
+Our algorithm achieves $\alpha$-approximate sublinear regret against these stronger benchmarks.
+Simulations on semi-synthetic real-world auction data demonstrate that empirical richness ratios are significantly better than the tight theoretical bounds.
+Finally, we show that (variants of) safe strategies and the proposed learning algorithm are well-suited to cater to more nuanced behavioral models of the buyer and the competitors.
+The simplicity, robustness, and adaptability of safe bidding strategies make them highly promising for practitioners.
 1. Introduction
 In a uniform price multi-unit auction, the auctioneer sells $K$ identical units of a single good to buyers who may demand multiple units, with the per-unit price set at the $K^{th}$ highest bid.
-These auctions are widely used to allocate scarce resources in critical markets such as emissions permits, energy markets, and Treasury auctions.
+These auctions are widely used to allocate scarce resources in critical avenues such as emissions permit auctions, wholesale electricity markets, and Treasury auctions.
 To succeed in these markets, bidders must develop effective bidding strategies that balance long-term value maximization with financial risk management, all while operating under limited information, uncertainty in competing bids and strategic (or adversarial) behavior of other bidders, among other challenges.
 In this paper, we model the bidders as a value-maximizing agent with per-round return-on-investment (RoI) constraints.
-This model reflects real-world decision-making in industries where managing financial risks is as critical as maximizing value.
-Unlike traditional mechanism design, which assumes bidders with quasilinear utility (i.e., utility decreases linearly with payments), many practical settings involve bidders who optimize total value while adhering to financial constraints such as RoI or budgets.
-These scenarios frequently arise in industries where agents or algorithms bid on behalf of clients, optimizing for high-level objectives within strict constraints—a context akin to the principal-agent framework (Fadaei and Bichler, 2016; Aggarwal et al., 2024).
-Examples include autobidders in online advertising, where advertisers seek to maximize clicks while keeping the average cost per click below a threshold (Lucier et al., 2024; Balseiro et al., 2021a; Deng et al., 2023b), and consultants bidding for small firms in emission permit auctions (EEX, 2024).
+This model reflects real-world decision-making in industries where managing financial risks is as important as maximizing value.
+Unlike traditional auction theory, which assumes bidders with quasilinear utility (i.e., utility decreases linearly with payments), many practical settings involve bidders who optimize total value while adhering to financial constraints such as RoI or budgets.
+These scenarios frequently arise in industries where agents or algorithms bid on behalf of clients, optimizing for high-level objectives within strict constraints–a context akin to the principal-agent framework (Fadaei and Bichler, 2016; Aggarwal et al., 2024).
+For example, in online advertising markets, performance marketing agencies often manage autobidding systems for clients who seek to maximize conversions or clicks while ensuring the average cost-per-click or return-on-ad-spend remains within specified limits (Lucier et al., 2024; Balseiro et al., 2021a; Deng et al., 2023b).
+Similarly, in the EU Emissions Trading System (EU ETS), many compliant companies participate in emission permits auctions indirectly through third-party firms such as MK Brokers JSC [^2] https://mkb.bg/en/services/sustainability-products/., Act Group [^3] https://www.actgroup.com/products-solutions/compliance/emissions-compliance-products., and AFS Group [^4] https://www.afsgroup.nl/exchange-traded-energy-instruments..
+These intermediaries bid on behalf of clients often under contractual constraints on allowable expenditures or minimum cost-effectiveness of permit acquisition (Ellerman et al., 2010).
 Building on this, we study the bidding problem in repeated uniform price auctions over $T$ rounds, from the perspective of a single bidder who seeks to maximize cumulative value while adhering to per-round RoI constraints—that is, ensuring that the value obtained in each round is at least a fixed multiple of the corresponding payment (see Eq.
@@
 Safe Bidding Strategies (Section 3).
-To ensure that the bidder satisfies the RoI constraint without knowing the competing bids, we introduce the concept of safe bidding strategies.
+To ensure that the bidder satisfies the RoI constraint without knowing the (potentially adversarial) competing bids in the online setting, we introduce the concept of safe bidding strategies.
 These strategies guarantee that the RoI constraint is met regardless of how the other participants bid.
-We then identify the class of $m$-uniform safe undominated bidding strategies, denoted by $\mathscr{S}_{m}^{\star}$ (Theorem 3.2), and demonstrate that the strategies within this class depend solely on the bidder’s valuation vector $[v_{1},\dots,v_{K}]$ and exhibit a “nested” structure (illustrated in Fig.
-1).
+We show that these strategies follow a mild no-overbidding condition (see Definition 3) and the bidder can focus on a finite subset of this class without loss of generality.
+For any $m\in\mathbb{N}$, we characterize this finite subset as the class of $m$-uniform safe undominated bidding strategies, denoted by $\mathscr{S}_{m}^{\star}$ (Theorem 3.2), and demonstrate that the strategies within this class depend solely on the bidder’s valuation vector $v_{1},\dots,v_{K}]$ and exhibit a “nested” structure as illustrated in Fig.
+1.
 $v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$$v_{8}$$v_{9}$$v_{10}$$w_{3}$$w_{7}$$w_{9}$}
@@
 Consider the strategy $\mathbf{b}=\langle(w_{3},3),(w_{7},4),(w_{9},2)\rangle\in\mathscr{S}_{3}^{\star}$, where we note that $Q_{1}=3$, $Q_{2}=3+4=7$ and $Q_{3}=3+4+2=9$.
-The $j^{th}$ highest bid (i.e., $b_{j}$) is the average of the first $Q_{j}$ entries of the valuation vector, i.e., $b_{j}=w_{Q_{j}}$, where $w_{j}=\frac{1}{j}\sum_{\ell\leq j}v_{\ell}$.
+The $j^{th}$ highest bid (i.e., $b_{j}$) is the average of the first $Q_{j}=\sum_{\ell\leq j}q_{\ell}$ entries of the valuation vector, i.e., $b_{j}=w_{Q_{j}}$, where $w_{j}=\frac{1}{j}\sum_{\ell\leq j}v_{\ell}$.
 Learning Safe Bidding Strategies (Section 4).
-Designing an algorithm to learn the optimal $m$-uniform safe bidding strategy poses significant challenges due to the exponential size of the decision space in $m$.
-To tackle this, we first consider an offline setting where the competing bids are known in advance.
-In this setting, we construct a polynomial-sized directed acyclic graph (DAG) with carefully assigned edge weights and show that determining the maximum weight path in the DAG is equivalent to computing the optimal $m$-uniform safe bidding strategy (Theorem 4.2).
-Building on this, we develop an online learning algorithm when the competing bids are unknown.
-[^1] For a detailed discussion on leveraging offline algorithms for no-regret learning, we refer readers to Roughgarden and Wang (2019); Niazadeh et al.
-(2022); Brânzei et al.
-(2023).
-Our algorithm runs in polynomial time and achieves a regret of $\widetilde{O}(M\sqrt{mT})$ in a full-information setting and $\widetilde{O}(M^{3/2}m^{3/2}\sqrt{T})$ regret in a bandit setting, where the clairvoyant selects the fixed hindsight optimal safe strategy (Theorem 4.3).
-Additionally, we establish a regret lower bound of $O(M\sqrt{T})$ (Theorem 4.4).
+Designing an algorithm to learn the optimal safe bidding strategy with at most $m$ bid-quantity pairs poses significant challenges as the size of the decision space in $O(M^{m})$, where $M$ is the maximum number of units demanded by the bidder and in practice, $m\ll M$.
+To this end, we first consider an offline setting where the competing bids are known a priori.
+We show that the value obtained by a $m$-uniform safe strategy can be decomposed across its $m$ constituent bid-quantity pairs.
+Leveraging this decomposition, we construct a directed acyclic graph (DAG), of size $\text{poly}(m,M)$, with carefully assigned edge weights and show that determining the maximum weight path in the DAG is equivalent to computing the optimal safe bidding strategy with at most $m$ bid-quantity pairs (Theorem 4.2).
+This reduction allows us to compute the offline optimal safe strategy in $\text{poly}(m,M)$ time, despite the exponential size of the original decision space.
+[…]
 The problem of learning to bid in multi-unit uniform price and pay-as-bid auctions has been explored recently by Brânzei et al.
@@
 Our work differs from them in two key ways: (a) we consider value-maximizing buyers with RoI constraints, a fundamentally different behavioral model, and (b) from a technical perspective, unlike prior works that require bid spaces to be discretized, our approach does not, due to the structure of safe strategies.
-Consequently, the time complexity becomes independent of $T$ in both online and offline settings  [^2] Assuming the discretization level in the offline setting is same as that in the online setting.
+Consequently, the time complexity becomes independent of $T$ in both online and offline settings [^7] Assuming the discretization level in the offline setting is the same as that in the online setting.
 In the online setting, time complexity refers to per-round running time., and under bandit feedback, the dependence on $T$ in the regret bound improves from $T^{2/3}$ to $\sqrt{T}$ implying that it is easier to learn safe strategies for value maximizers compared to bidding strategies for quasilinear utility maximizers.
 Richer Classes of Strategies for the Clairvoyant (Section 5).
-A key novel aspect of our work is to evaluate the performance of safe bidding strategies with at most $m$ bid-quantity pairs and the online learning algorithm from Section 4 against a clairvoyant that selects the optimal strategy from a richer class of strategies.
-In this case, the bidder (learner) achieves $\alpha$-approximate sublinear regret (same as Theorem 4.3) for a parameter $\alpha\in(0,1]$—defined as the richness ratio—thereby demonstrating the robustness of the safe bidding strategies against a broad range of benchmarks (Corollary 5.2).
-One of our main contributions is to compute the richness ratio, $\alpha$, for various classes of bidding strategies.
-Specifically, when the clairvoyant selects the optimal bidding from the class of
-(a) strategies that are RoI-feasible, not necessarily safe and have at most $m$ bid-quantity pairs, we get $\alpha=\frac{1}{2}$, notably independent of $m$ (Theorem 5.3).
-(b) safe strategies with at most $m^{\prime}$ bid-quantity pairs, where $m^{\prime}\geq m$, we obtain $\alpha=\frac{m}{m^{\prime}}$ (Theorem 5.4).
-(c) strategies that are RoI-feasible, not necessarily safe and have at most $m^{\prime}$ bid-quantity pairs, where $m^{\prime}\geq m$, we get that $\alpha=\frac{m}{2m^{\prime}}$ (Theorem 5.5).
-Computing $\alpha$ involves two key challenges: (a) deriving an upper bound on the ratio between the value obtained by the clairvoyant’s optimal strategy and that of the optimal safe strategy in the worst case (see Definition 4 for details), and (b) proving that this upper bound is tight.
-The key ideas for deriving the upper bound are outlined in Section 5.
-The construction of problem instances, each tailored to a specific strategy class, is non-trivial due to their exponential size in $m$ and the need for a careful choice of competing bids and the valuation vector, as detailed in Sections D.1 to D.4.
-Simulations using semi-synthetic data show that the empirical richness ratios in practical scenarios are significantly better than the theoretical bounds, making safe strategies highly appealing in practice (Section 5.3).
+We then evaluate the performance of the bidder’s strategy class, i.e., safe bidding strategies with at most $m$ bid-quantity pairs and the online learning algorithm from Section 4 against a clairvoyant benchmark that selects the optimal strategy from a richer class.
+Here, a richer class refers to any class of strategies that is a superset of the bidder’s strategy class (see examples below).
+Recall that the bidder chooses to follow the $m$-uniform safe strategies to satisfy the RoI constraints with limited knowledge about the competing bids in any auction in the online setting.
+Comparing against such stronger benchmarks demonstrates the robustness of the bidder’s strategy class.
+We quantify this robustness via the richness ratio $\alpha\in(0,1]$ (Definition 4), analogous to an approximation ratio.
+We prove that for any richer class with richness ratio $\alpha$, the bidder achieves $\alpha$-approximate regret (as defined in Eq.
+8) of the same order as in Theorem 4.3 (see Corollary 5.2).
+One of our main contributions is to compute the richness ratio $\alpha$ for various classes of bidding strategies.
+We define richness along two dimensions: (i) the safety of strategies (safe vs.
+only RoI-feasible), and (ii) the number of bid-quantity pairs allowed.
+A strategy is defined as RoI-feasible (or feasible in short) if it satisfies the RoI constraint for a given sequence of competing bids over $T$ rounds, unlike safe strategies that satisfy the constraint for every such sequence.
+We choose these two dimensions of richness as the bidder follows the class of strategies that are (i) safe and (ii) have at most $m$ bid-quantity pairs.
+When the clairvoyant selects the optimal strategy from the class of
+  • (a)
+    strategies that are RoI-feasible and have at most $m$ bid-quantity pairs, we get $\alpha=\frac{1}{2}$ (Theorem 5.3).
+Notably, for this particular choice of richer class of strategies, $\alpha$ is independent of $m$.
+  • (b)
+    safe strategies with at most $m^{\prime}$ bid-quantity pairs, where $m^{\prime}\geq m$, we obtain $\alpha=\frac{m}{m^{\prime}}$ (Theorem 5.4).
+Here, setting $m^{\prime}=m$ refers to the case considered in Section 4 where the bidder and clairvoyant follow the same class of strategies.
+In this case, Theorem 5.4 correctly implies the optimal value of $\alpha=1$.
+  • (c)
+    strategies that are RoI-feasible and have at most $m^{\prime}$ bid-quantity pairs, where $m^{\prime}\geq m$, we get $\alpha=\frac{m}{2m^{\prime}}$ (Theorem 5.5).
+Setting $m^{\prime}=m$ in this case recovers the result of Theorem 5.3.
+Computing the richness ratio involves two different challenges: (a) deriving an upper bound on the ratio between the value obtained by the clairvoyant’s optimal strategy from the richer class and that of the optimal safe strategy in the worst case (see Definition 4 for details), and (b) proving that this upper bound is tight by constructing a hard problem instance for the given richer class of strategies.
+The key ideas for deriving the upper bounds are outlined in Section 5.
+The hard problem instances that show the tightness of the bounds are highly non-trivial where we need to consider an instance of size that is exponential in $m$, with careful choice of competing bids and the valuation vector (see details in Sections D.1 to D.3).
+Extensions (Section 6).
+Finally, we show that safe strategies and the learning algorithm from Section 4 can be adapted to handle more nuanced modeling assumptions.
+First, as competitors may adapt to the bidder’s past actions in repeated auctions, we consider an adaptive adversary model for generating competing bids (as opposed to an oblivious one).
+In this setting, the learning algorithm with appropriate modifications still yields a high-probability regret bound of $\text{poly}(m,M)\cdot\sqrt{T}$ (Section 6.1).
+Second, when the RoI constraint is enforced over a sliding window of $T_{0}\in\mathbb{N}$ rounds instead of each round, we propose a heuristic based on shifted safe strategies for this setting that shows good empirical performance on semi-synthetic real-world auction data (Section 6.2).
+Third, in the case of time-varying valuation vectors, that are either sampled from a distribution or generated adversarially from a finite set $\mathcal{V}$, we treat the valuation vector as contexts and show that a contextual version of our learning algorithms achieves $\text{poly}(m,M,|\mathcal{V}|)\cdot\sqrt{T}$ regret in this setting (Section 6.3).
+These extensions highlight the practical appeal and flexibility of our approach.
+1.2. Managerial Insights
+Our findings offer several key implications for RoI-constrained value maximizing buyers.
+  1. 1.
+    Safe strategies are an intuitive class of bidding strategies due to their mild no-overbidding conditions.
+As the bidder can focus on a finite subset of the class without any loss of generality, designing and implementing the proposed learning algorithm in repeated auction environments is quite efficient from a practical perspective.
+Furthermore, the theoretical regret upper and lower bounds with $\sqrt{T}$ dependence indicate that the proposed algorithm is near-optimal, which is desirable to practitioners.
+  2. 2.
+    We crucially show the robustness of safe strategies and the learning algorithm by proving that they are approximately optimal against stronger benchmarks.
+A key highlight here is that for any given $m$, the cost of following safe strategies compared to the instance dependent RoI-feasible strategies, which are impossible to know a priori, is uniformly bounded.
+So, a clairvoyant with prior knowledge of the adversarially generated competing bids in this setting can not be arbitrarily better than the bidder following safe strategies.
+  3. 3.
+    Simulations on semi-synthetic data from EU Emission Trading System (ETS) auctions show that the empirical richness ratios in real-world scenarios are significantly better than tight theoretical bounds, which reinforces the fact that tight bounds are attained under highly pathological settings and suggests that safe strategies can perform near-optimal in practice, even for small values $m$ (Section 5.3).
+  4. 4.
+    Finally, we show that variants of safe strategies and the proposed learning algorithm can be appropriately deployed in several practical extensions of the primary model considered in this work.
+This aspect is particularly interesting as it allows practitioners to tweak safe strategies and/or the proposed algorithm to design novel algorithms or heuristics to adapt to their objectives and risk appetite that may change over time.
 1.3. Related Work
 Value Maximizers and RoI Constraints.
-The concept of agents as value maximizers within financial constraints is a well-established notion in microeconomic theory (Mas-Colell et al., 1995, pp.
-50).
+The concept of agents as value maximizers within financial constraints is a well-established notion in microeconomic theory (Mas-Colell et al., 1995).
 In mechanism design literature, one of the earliest explorations of value-maximizing agents was conducted by Wilkens et al.
@@
 (2023) studied truthful (approximate) revenue-maximizing mechanisms in combinatorial markets tailored for such agents.
+More recently, Tang et al.
+(2024) investigated auction design in a setting where buyers may be either quasilinear utility maximizers or RoI-constrained value maximizers.
+Xu et al.
+(2024) proposed the allowance utility maximization model, which generalizes both quasilinear and value-maximizing preferences.
 In recent years, there has been growing interest in RoI-constrained value maximizers, particularly in the context of autobidding and online advertising.
 One of the earliest works in this area is Golrezaei et al.
-(2021), which studied auction design for RoI-constrained buyers and validated the presence of such soft financial constraints using data from online ad auctions.
+(2021c), which studied auction design for RoI-constrained buyers and validated the presence of such soft financial constraints using data from online ad auctions.
 Broadly, the literature in this space can be divided into two categories: (i) works that design optimal auctions under RoI constraints (Balseiro et al., 2021a, b, 2022; Deng et al., 2021, 2023b), and (ii) works that characterize optimal bidding strategies and/or develop learning algorithms in repeated auction settings (Aggarwal et al., 2019; Deng et al., 2023a; Golrezaei et al., 2023; Castiglioni et al., 2024; Aggarwal et al., 2025; Lucier et al., 2024).
 Our work primarily aligns with the latter.
-We study value-maximizing buyers in uniform price auctions under RoI constraints and demonstrate that these buyers can employ safe bidding strategies that are efficiently learnable and robust against a variety of strong benchmarks.
+We study value-maximizing buyers in uniform price auctions under RoI constraints and show that they can employ safe strategies that are efficiently learnable and robust against a variety of strong benchmarks.
 Multi-unit Auctions.
 In this work, we focus on a subset of combinatorial auctions termed as multi-unit auctions in which multiple identical goods are sold to a group of buyers.
-[^3] For a comprehensive survey on combinatorial auctions, we refer the readers to several excellent works by De Vries and Vohra (2003); Blumrosen and Nisan (2007); Palacios-Huerta et al.
+[^8] For a comprehensive survey on combinatorial auctions, we refer readers to several excellent works by De Vries and Vohra (2003); Blumrosen and Nisan (2007); Palacios-Huerta et al.
 (2024).
-These auctions find widespread application in various practical scenarios, including Treasury auctions (Hortaçsu and McAdams, 2010), procurement auctions (Cramton and Ausubel, 2006), electricity markets (Tierney et al., 2008), and emissions permit auctions (Schmalensee and Stavins, 2017).
-While several works have focused on studying the equlibria properties in these auctions (Ausubel et al., 2014; Markakis and Telelis, 2015; De Keijzer et al., 2013), computing equilibrium strategies is usually intractable in general multi-unit auctions due to multi-dimensional valuations.
-As a result, a growing body of work has emphasized on designing optimal bidding strategies in a prior-free setting, i.e., without any assumptions of the competing bidders’ bids or valuations (Galgana and Golrezaei, 2024; Potfer et al., 2024; Brânzei et al., 2023).
+These auctions find widespread application in various practical scenarios, including Treasury auctions (Nyborg et al., 2002; Hortaçsu and McAdams, 2010; Hortaçsu and Kastl, 2012), procurement auctions (Cramton and Ausubel, 2006), electricity markets (Tierney et al., 2008; Fabra et al., 2006), and emissions permit auctions (Goulder and Schein, 2013; Schmalensee and Stavins, 2017; Goldner et al., 2020).
+While several works have focused on studying the equilibria properties in these auctions (Ausubel et al., 2014; Engelbrecht-Wiggans and Kahn, 1998; Noussair, 1995; Markakis and Telelis, 2015; De Keijzer et al., 2013), computing equilibrium strategies is usually intractable in general multi-unit auctions due to multi-dimensional valuations (Kasberger and Woodward, 2025).
+As a result, a growing body of work has emphasized on designing optimal bidding strategies in a prior-free setting, i.e., without any assumptions of the competing bidders’ bids or valuations–in a repeated setting (Galgana and Golrezaei, 2024; Potfer et al., 2024; Brânzei et al., 2023) or under a minimax loss framework (Kasberger and Woodward, 2025).
 Our work contributes to this literature by learning optimal strategies for value-maximizing buyers in repeated uniform price auctions.
+Data-Driven Methods in Auctions.
+Traditionally, the problems of bidding in auctions and determining parameters such as reserve prices have been studied in the Bayesian framework (Myerson, 1981; Riley and Samuelson, 1981; Hartline and Roughgarden, 2009; Beyhaghi et al., 2021).
+However, these approaches rely on full knowledge of bidders’ valuation distributions, which is often unavailable or unreliable in practice.
+Moreover, they become computationally intractable in multi-unit auctions, as discussed earlier.
+In practice, historical bidding data is crucial for bidders to calibrate their strategies.
+For instance, Kasberger and Schlag (2024) note that none of the auction consultants surveyed could offer viable solutions without access to sufficient real-world data.
+These challenges, along with the increasing availability of auction data, have led to a growing interest in data-driven approaches.
+Beyond learning optimal bidding strategies in a repeated setting, as discussed in the previous section, this line of work includes data-driven reserve price optimization (Roughgarden and Wang, 2019; Derakhshan et al., 2022, 2021; Golrezaei et al., 2021a; Feng et al., 2021), boost value optimization in second-price auctions (Golrezaei et al., 2021b) and deep learning methods for auction design (Feng et al., 2018; Dütting et al., 2024; Wang et al., 2024, 2025; Liu et al., 2025).
 2. Model
@@
 Hence, without loss of generality, we assume $\mathbf{v}\in\mathbb{R}_{+}^{M}$.
-For each $\mathbf{v}=[v_{1},\dots,v_{M}]$, we define the average cumulative valuation vector as $\mathbf{w}=[w_{1},\dots,w_{M}]$,
-$\text{where}\quad w_{j}$ $=\frac{1}{j}\sum_{\ell\leq j}v_{\ell},\forall j\in[M]\,.$ (1)
+For each $\mathbf{v}=[v_{1},\dots,v_{M}]$, we define the average cumulative valuation vector as $\mathbf{w}=[w_{1},\dots,w_{M}]$, where
+$w_{j}$ $=\frac{1}{j}\sum_{\ell\leq j}v_{\ell},\forall j\in[M]\,.$ (1)
 As $v_{1}\geq\dots\geq v_{M}$, we also have $w_{1}\geq\dots\geq w_{M}$.
@@
 For ease of exposition, we assume there are no ties (or ties are always broken in the favor of the bidder in consideration).
-See Appendix F for our discussion on handling ties.
-We assume the auction follows the last-accepted-bid payment rule (bidders pay the $K^{th}$ highest bid per unit) which is widely used for uniform price auctions in practice (Regulations, 2019; Garbade and Ingber, 2005).
-Let $\bm{\beta}_{-}$ denote the bids submitted by all bidders except bidder $i$ and $\bm{\beta}_{-}^{-(j)}$ be the $j^{th}$ smallest among the top $K$ competing bids (i.e., bids from all bidders except bidder $i$).
+See Appendix G for our discussion on handling ties.
+We assume the auction follows the last-accepted-bid payment rule (bidders pay the $K^{th}$ highest bid per unit), which is widely used for uniform price auctions in practice (Regulations, 2019; Garbade and Ingber, 2005; Burkett and Woodward, 2020).
+Let $\bm{\beta}_{-}$ denote the vector of bids submitted by all bidders except bidder $i$.
+Let $\textsc{top}_{K}(\bm{\beta}_{-})\subseteq\bm{\beta}_{-}$ be the multiset of the top $K$ highest bids among these competing bids.
+Furthermore, we define $\bm{\beta}_{-}^{-(j)}$ to be the $j^{\text{th}}$ smallest element in $\textsc{top}_{K}(\bm{\beta}_{-})$.
 Suppose bidder $i$ submits $\mathbf{b}$, such that the bid profile is $\bm{\beta}:=(\mathbf{b};\bm{\beta}_{-})$.
@@
 Although expressive, this becomes computationally impractical when the number of units, $K$, is large, as in EU ETS emission permit auctions and Treasury auctions.
-To address this, we consider a bidding language called $m$-uniform bidding, for any $m\in\mathbb{N}$ [^4] This format generalizes the uniform bidding format (De Keijzer et al., 2013; Birmpas et al., 2019) and aligns with practical languages like those in product-mix auctions (Klemperer, 2009) and piecewise-linear bidding (Nisan, 2015).
-In $m$-uniform bidding, bidders submit $m$ bid-quantity pairs $(b_{i},q_{i})$, where $b_{i}$ is the bid value per unit and $q_{i}$ is the quantity demanded:
+To address this, we consider a bidding language called $m$-uniform bidding, for any $m\in\mathbb{N}$.
+[^9] This format generalizes the uniform bidding format (De Keijzer et al., 2013; Birmpas et al., 2019) and aligns with practical languages like those in product-mix auctions (Klemperer, 2009) and piecewise-linear bidding (Nisan, 2015).
+In $m$-uniform bidding format, bidders submit $m$ bid-quantity pairs $(b_{i},q_{i})$, where $b_{i}$ is the bid value per unit and $q_{i}$ is the quantity demanded:
 Definition 1 ($m$-Uniform Bidding).
-For a fixed $m\in\mathbb{N}$, a $m$-uniform bidding strategy is characterized by $m$ bid-quantity pairs, denoted as $\mathbf{b}:=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$, where $b_{1}>b_{2}>\dots>b_{m}>0$ and $q_{j}>0$, $j\in[m]$.
-This $m$-uniform bid can be equivalently expressed as a vector (similar to the standard bidding format) in which the first $q_{1}$ bids are $b_{1}$, followed by $q_{2}$ bids of $b_{2}$, and so on.
+For a fixed $m\in\mathbb{N}$, a $m$-uniform bidding strategy is characterized by $m$ bid-quantity pairs, denoted as
+$\mathbf{b}:=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle\,$,
+where $b_{1}>b_{2}>\dots>b_{m}>0$ and $q_{j}>0$, $j\in[m]$.
+This $m$-uniform bidding strategy can be equivalently expressed as a vector (similar to the standard bidding format) in which the first $q_{1}$ bids are $b_{1}$, followed by $q_{2}$ bids of $b_{2}$, and so on.
 We define $\mathbf{b}[1:\ell]=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell})\rangle$, for all $\ell<m$, to represent the first $\ell$ bid-quantity pairs within a $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$.
-We further define $Q_{j}=\sum_{\ell=1}^{j}q_{\ell}$ for all $j\in[m]$ as the total quantity demanded in the first $j$ bid-quantity pairs, with $Q_{0}=0$, where we assume, without loss of generality that $Q_{m}\leq M$ [^5] Suppose the bidder bids for, and wins more than $M$ units.
+We further define
+$Q_{j}=\sum_{\ell=1}^{j}q_{\ell},\qquad\text{ for all $j\in[m]$}$
+as the total quantity demanded in the first $j$ bid-quantity pairs, with $Q_{0}=0$, and we assume, without loss of generality that $Q_{m}\leq M$.
+[^10] Suppose the bidder bids for, and wins more than $M$ units.
 There is no additional value being allocated over $M$ units, but the total payment increases (assuming the clearing price is positive), potentially violating the RoI constraint.
-If $m=M$, the bidding format is equivalent to standard bidding but in practice, bidders often submit only a few bid-quantity pairs.
-For instance, in the 2023 EU ETS auctions, bidders submitted $\sim 4.35$ bid-quantity pairs per auction on average (EEX, 2023).
+If $m=M$, the bidding format is equivalent to standard bidding, but in practice, bidders often submit only a few bid-quantity pairs.
+For instance, in the 2023 EU ETS auctions, bidders submitted $\sim 4.35$ pairs per auction on average (EEX, 2023).
 Bidders’ Behavior.
@@
 Here, $\gamma$ is defined as the target RoI which is private and fixed.
-Without loss of generality, we assume $\gamma=0$ (or equivalently the valuation curve $\mathbf{v}$ is scaled by $\frac{1}{1+\gamma}$) for the rest of this work.
+Without loss of generality, we assume $\gamma=0$ (or equivalently, the valuation curve $\mathbf{v}$ is scaled by $\frac{1}{1+\gamma}$) for the rest of this work.
 For $\gamma=0$, the RoI constraint implies the value obtained in an auction is at least the payment.
-We illustrate how the auction operates in Example 1.
 Example 1.
@@
 Suppose $m=2$ and the bids submitted by the bidders are $\mathbf{b}_{1}=\langle(5,2),(3,3)\rangle$ and $\mathbf{b}_{2}=\langle(4,2),(2,2)\rangle$.
-The bid profile: $\bm{\beta}=[5,5,4,4,3,3,3,2,2]$ and top $K=5$ winning bids are $[\underline{5},\underline{5},4,4,\underline{3}]$.
+The bid profile: $\bm{\beta}=[5,5,4,4,3,3,3,2,2]$ and top $K=5$ winning bids are $\underline{5},\underline{5},4,4,\underline{3}]$.
 Bidder 1 is allocated $3$ units as they have $3$ bids (underlined) in the winning bids, and bidder 2 gets the remaining $2$ units.
@@
 Formally, the auction described in the previous section takes place sequentially over $T$ rounds indexed by $t\in[T]$.
+Throughout this work, we assume that $T$ is known to the bidder.
+[^11] In case $T$ is unknown, the bidder can use the standard doubling trick (Auer et al., 2002; Cesa-Bianchi and Lugosi, 2006) while designing learning algorithms.
 We assume that the valuation vector $\mathbf{v}$ is fixed over the $T$ rounds which is a standard assumption in the literature (Brânzei et al., 2023; Galgana and Golrezaei, 2024; Potfer et al., 2024).
-Similarly, the target RoI is also assumed to be fixed over the $T$ rounds.
+Nonetheless, we discuss the case of time-varying valuation vectors in Section 6.3 and show that the techniques developed for the scenario when the valuation vector is fixed can be leveraged to learn to bid when the valuation vectors change across rounds.
 We now extend the notations from the previous section to the repeated setting.
@@
 Let $x(\bm{\beta}^{t})$ and $p(\bm{\beta}^{t})$ denote the number of units allocated to bidder $i$ and the clearing price, respectively, in round $t$.
-The total value obtained by the bidder is $V(\bm{\beta}^{t})=\sum_{j\leq x(\bm{\beta}^{t})}v_{j},$ while the total payment made is $P(\bm{\beta}^{t})=p(\bm{\beta}^{t})\cdot x(\bm{\beta}^{t}).$
+For the $t^{th}$ auction, the total value obtained by the bidder is $V(\bm{\beta}^{t})=\sum_{j\leq x(\bm{\beta}^{t})}v_{j},$ while the total payment made is $P(\bm{\beta}^{t})=p(\bm{\beta}^{t})\cdot x(\bm{\beta}^{t}).$
+Nature of Competing Bids.
+We model the competing bids in each round as being chosen adversarially to capture realistic scenarios where the bidder has partial or no information about their competitors’ strategy.
+We assume that the competing bids are selected by an oblivious adversary, i.e., the bids in round $t$ are chosen adversarially but independently of the bidder’s strategies in earlier rounds $\tau=1,\dots,t-1$.
+This is equivalent to the case where all competing bids are fixed in advance, prior to the first round, and without any knowledge of the bidder’s realized strategies.
+In our setting, an oblivious adversary models a sufficiently large market where the bids submitted by a single bidder do not materially change the allocation, for which the competing bids could be viewed as independent of the strategies submitted by the bidder.
+In Section 6.1, we extend our analysis for an adaptive adversary, [^12] Also called as non-oblivious adversary (Cesa-Bianchi and Lugosi, 2006) or reactive adversary (Maillard and Munos, 2011).
+where the competing bids in any round may depend on the bidder’s strategies in the previous rounds.
 RoI Constraints in Repeated Setting.
@@
 2 in every round.
-A bidding strategy $\mathbf{b}$ is called feasible for a sequence of competing bids $[\bm{\beta}_{-}^{t}]_{t\in[T]}$, if the RoI constraint is satisfied for every round $t\in[T]$.
+A bidding strategy $\mathbf{b}$ is called feasible for a sequence of competing bids $\bm{\beta}_{-}^{t}]_{t\in[T]}$, if the RoI constraint is satisfied for every round $t\in[T]$.
 Remark 2.1 (RoI Constraints).
@@
 (2023).
-Similar constraints are considered by Lucier et al.
+Similar constraints (up to scaling factors) are considered by Lucier et al.
 (2024) and Gaitonde et al.
-(2023), termed as marginal RoI (or value) and maximum bid constraints (up to scaling factors), respectively.
+(2023), which they term as maximum bid constraints.
+[^13] An earlier version of Lucier et al.
+(2024) termed this as marginal RoI (or value) constraints.
 Unlike the aggregate constraints over $T$ rounds typically assumed in online ad auction literature (Deng et al., 2021, 2024; Feng et al., 2023; Deng et al., 2023a), we enforce RoI constraints for each auction individually.
 This distinction reflects the fundamental differences between the two settings: ad auctions often occur simultaneously and frequently, with values in the order of cents, whereas Treasury and emission permit auctions occur sequentially over longer horizons, with units valued in millions.
+Furthermore, we consider a prior-free setting that allows for strategic (possibly adversarial) behavior by the competitors.
 Hence, bidders in such auctions prioritize profitability in each auction rather than waiting for an indefinite period, leading to per-round RoI constraints.
@@
 Hence, the bidders are more likely to ensure RoI feasibility in each round.
+Nonetheless, in Section 6.2, we present a heuristic for the setting where the RoI constraints are enforced only in aggregate over a sliding window of $T_{0}\geq 1$ rounds.
 2.3. Objective and Performance Metric
@@
 In this setting, bidders privately submit their bids, and the auctioneer allocates items and sets prices based on these bids.
-In round $t$, the learning algorithm maps the feedback information set from the first $t-1$ rounds, denoted by $\mathcal{I}^{t-1}$, to a bid $\mathbf{b}^{t}$, where this mapping may be deterministic or random.
+In round $t$, the learning algorithm maps the feedback information set from the first t1t-1 rounds, denoted by $\mathcal{I}^{t-1}$, to a bid $\mathbf{b}^{t}$, where this mapping may be deterministic or random.
 Under full information feedback, the information set $\mathcal{I}^{t-1}=(\bm{\beta}^{1},\bm{\beta}^{2},\ldots,\bm{\beta}^{t-1}),$ which includes all bids from previous rounds.
-In the bandit setting, the information set $\mathcal{I}^{t-1}=(V(\bm{\beta}^{1}),P(\bm{\beta}^{1}),\ldots,V(\bm{\beta}^{t-1}),P(\bm{\beta}^{t-1}))$ includes only the value and price outcomes of the previous rounds.
+In the bandit setting, the information set $\mathcal{I}^{t-1}=(x(\bm{\beta}^{1}),P(\bm{\beta}^{1}),\ldots,x(\bm{\beta}^{t-1}),P(\bm{\beta}^{t-1}))$ includes only the allocations and prices of the previous rounds.
 We design learning algorithms to minimize regret, the difference in the value obtained by a fixed hindsight-optimal strategy chosen by a clairvoyant with a priori knowledge of competing bids and that by the learner over time.
 Formally,
-$\mathsf{REG}=\max_{\mathbf{b}\in\mathscr{B}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})],$ (3)
+$\textsf{REG}=\max_{\mathbf{b}\in\mathscr{B}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})],$ (3)
 where the expectation is with respect to any randomness in the learning algorithm.
 Here, $\mathscr{B}$ is the class of bidding strategies (formally characterized in Section 3).
-We require the bidding strategies in $\mathscr{B}$ to be RoI feasible for the sequence of competing bids $[\bm{\beta}_{-}^{t}]_{t\in[T]}$.
-In Section 4, we consider the case when the both the clairvoyant and the learner choose strategies from the same class and aim to obtain sublinear regret.
+We require the bidding strategies in $\mathscr{B}$ to be RoI feasible for the sequence of competing bids $\bm{\beta}_{-}^{t}]_{t\in[T]}$.
+In Section 4, we consider the case when both the clairvoyant and the learner choose strategies from the same class and aim to obtain sublinear regret.
 Later, we consider more challenging settings where the clairvoyant can choose the optimal strategy from much richer classes of strategies compared to the learner.
-In these cases, we obtain sublinear $\alpha$-approximate regret, where $\alpha\in(0,1]$ is the richness ratio (see details in Section 5).
+In these cases, we obtain sublinear $\alpha$-approximate regret, where $\alpha\in(0,1]$ is the richness ratio (see Section 5).
 3. Safe Bidding Strategies
-Recall that strategies chosen by the learner must be causal (mapping the history to a bidding strategy) while remaining RoI feasible for all rounds $t\in[T]$, even under adversarially generated competing bids.
+Recall that strategies chosen by the learner must be non-anticipating (mapping the history to a bidding strategy) while remaining RoI feasible for all rounds $t\in[T]$, even under adversarially generated competing bids.
 This creates an obvious challenge: a bidding strategy that satisfies RoI feasibility in previous rounds may become infeasible even under the slightest change in the competing bids e.g., if $\mathbf{v}=[0.9,0.5,0.1]$, the bidding strategy $(0.6,3)$ is RoI feasible for competing bids $\bm{\beta}_{-}^{1}=[0.61,0.59,0.59]$ but not for $\bm{\beta}_{-}^{2}=[0.59,0.59,0.59]$.
@@
 The union of classes of safe bidding strategies with at most $m$ bid-quantity pairs is denoted as $\mathscr{U}_{m}=\bigcup_{k\in[m]}\mathscr{S}_{k}$.
+Recall that we assume $\gamma=0$ throughout; otherwise the constraint in Definition 2 is identical to Eq. 2.
 3.1. Characterizing Safe Bidding Strategies
 We begin by defining underbidding (overbidding) under the $m$-uniform bidding format in the given context.
-Recall that in a single-item auction, if the value of the item is $v$, then a bid $b$ is an underbid (overbid) if $b<v~{}(b>v)$.
 Definition 3 (Underbid and Overbid).
-A $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an
-(a) underbidding strategy if $b_{j}\leq w_{Q_{j}},\forall j\in[m]$ and $\exists\ell\in[m]$ such that $b_{\ell}<w_{Q_{\ell}}$.
-(b) overbidding strategy if $\exists\ell\in[m]$ such that $b_{\ell}>w_{Q_{\ell}}$.
+A $m$-uniform strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an
+(a) underbidding strategy if $b_{j}\leq w_{Q_{j}},\forall j\in[m]$ and $\exists\ell\in[m]$ such that $b_{\ell}<w_{Q_{\ell}}$, and
+(b) overbidding strategy if $\exists\ell\in[m]$ such that $b_{\ell}>w_{Q_{\ell}}$,
+where we recall that $Q_{j}=\sum_{\ell\leq j}q_{\ell},\forall j\in[m]$.
 Figure 2: The solid line represents the average cumulative valuation curve, $\mathbf{w}$, and the dotted line represents the valuation curve, $\mathbf{v}$.
@@
 overbidding) for a $2$-uniform bidding strategy.
-Note that the notions of underbidding and overbidding in Definition 3 are defined with respect to $\mathbf{w}$ and not $\mathbf{v}$ [^8] In Definition 3 (and throughout the paper), we assume $\gamma=0$ without loss of generality for ease of exposition.
-Footnote 8 illustrates how different values of $\gamma$ impacts the results.
-Specifically, for any $\gamma\geq 0$, underbidding and overbidding are defined with respect to the scaled average cumulative valuation curve, $\frac{\mathbf{w}}{1+\gamma}$.Here, the plots of $\mathbf{v}$ and $\mathbf{w}$ are shown to be linear for illustrative purposes only.
-In other words, if $\gamma=0$, $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an underbidding strategy if, for all $j\in[m]$, the bid $b_{j}$ is at most the average of the first $Q_{j}$ entries of the valuation vector, denoted as $w_{Q_{j}}$, and there exists $\ell\in[m]$ where the inequality is strict.
-Recall that $Q_{j}=\sum_{\ell\leq j}q_{\ell}$ denotes the maximum number of demanded units in the first $j$ bid-quantity pairs.
+Note that the notions of underbidding and overbidding in Definition 3 are defined with respect to $\mathbf{w}$ and not $\mathbf{v}$.
+[^16] In Definition 3 (and throughout the paper), we assume $\gamma=0$ without loss of generality for ease of exposition.
+Footnote 16 illustrates how different values of $\gamma$ impacts the results.
+Specifically, for any $\gamma\geq 0$, underbidding and overbidding are defined with respect to the scaled average cumulative valuation curve, $\frac{\mathbf{w}}{1+\gamma}$.
+Here, the plots of $\mathbf{v}$ and $\mathbf{w}$ are shown to be linear for illustrative purposes only.
+In other words, after normalizing $\gamma=0$, $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ is an underbidding strategy if, for all $j\in[m]$, the bid $b_{j}$ is at most the average of the first $Q_{j}$ entries of the valuation vector, denoted as $w_{Q_{j}}$, and there exists $\ell\in[m]$ where the inequality is strict.
 Similarly, $\mathbf{b}$ is an overbidding strategy if there exists some $\ell\in[m]$ such that $b_{\ell}$ is strictly greater than the average of the first $Q_{\ell}$ entries of the valuation vector.
@@
 where $\bf w$ is defined in Eq. (1).
-The proofs of this section are presented in Appendix A.
-Since there are infinitely many safe strategies in $\mathscr{S}_{m}$, we focus on the subset of safe undominated strategies, $\mathscr{S}_{m}^{\star}$, obtained by removing very weakly dominated strategies.
+Since there are infinitely many safe strategies in $\mathscr{S}_{m}$, we focus on the subset of safe undominated strategies, $\mathscr{S}_{m}^{\star}$, obtained by eliminating very weakly dominated strategies.
 A safe strategy $\mathbf{b}$ is said to be very weakly dominated if there exists another safe strategy $\mathbf{b}^{\prime}$ such that, for any competing bid $\bm{\beta}_{-}$, $V(\mathbf{b},\bm{\beta}_{-})\leq V(\mathbf{b}^{\prime},\bm{\beta}_{-})$ (Shoham and Leyton-Brown, 2008, pp.
@@
 Keeping these strategies does not improve the bidder’s performance, but removing them yields a finite strategy class (as shown below), which significantly simplifies the design of online learning algorithms.
-We now present the main result of this section:
+We show that for every underbidding strategy (as per Definition 3) in $\mathscr{S}_{m}$, there exists a non-underbidding strategy in the same class that very weakly dominates it, allowing the former to be safely eliminated.
+Hence,
 Theorem 3.2.
@@
 Observe that $\mathscr{S}_{m}^{\star}$ is a finite class of bidding strategies that depend only the valuation curve and is independent of the competing bids.
-The strategies in $\mathscr{S}_{m}^{\star}$ exhibit a “nested” structure, referred to as nested $m$-uniform bidding strategies hereafter, where the $j^{th}$ highest bid ($b_{j}$) is the average of the first $Q_{j}=\sum_{\ell=1}^{j}q_{\ell}$ entries of the valuation curve (illustrated in Fig.
+The strategies in $\mathscr{S}_{m}^{\star}$ exhibit a “nested” structure, where the $j^{th}$ highest bid ($b_{j}$) is the average of the first $Q_{j}=\sum_{\ell=1}^{j}q_{\ell}$ entries of the valuation curve (illustrated in Fig.
 1).
-Importantly, fixing $Q_{j}$’s uniquely determines the bidding strategy, as $b_{j}=w_{Q_{j}}$ for strategies in $\mathscr{S}_{m}^{\star}$ which will be crucial in learning the optimal safe bidding strategy, as discussed in the following section.
+Importantly, fixing $Q_{j}$’s uniquely determines the bidding strategy, as $b_{j}=w_{Q_{j}}$ for strategies in $\mathscr{S}_{m}^{\star}$, which will be crucial in learning the optimal safe bidding strategy, as discussed in the following section.
 4. Learning Safe Bidding Strategies
-Here, the clairvoyant and the learner choose strategies from the class of safe bidding strategies with at most $m$ bid-quantity pairs, $\mathscr{U}_{m}^{\star}$. Thus, the regret defined in Eq. 3 is
-$\mathsf{REG}=\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\,.$
-To design learning algorithms, we first consider the offline setting, where we aim to solve the following problem given the bid history, $[\bm{\beta}_{-}^{t}]_{t\in[T]}$.
+Here, the clairvoyant and the learner choose strategies from the class of safe bidding strategies with at most $m$ bid-quantity pairs, $\mathscr{U}_{m}^{\star}$ (defined in Theorem 3.2).
+Thus, in this case, the regret defined in Eq.
+3 becomes
+$\textsf{REG}=\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\,.$ (4)
+To design learning algorithms, we first consider the offline setting, where we aim to solve the following problem for a given bid history, $\bm{\beta}_{-}^{t}]_{t\in[T]}$.
 $\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (Offline)
 Since $\mathscr{U}_{m}^{\star}$ contains $O(M^{m})$ strategies, evaluating each strategy individually is intractable.
-To overcome this, we reduce the offline problem to finding the maximum weight path in a carefully constructed edge-weighted directed acyclic graph (DAG), which allows for the computation of the optimal offline bidding strategy in poly($m,M$) time.
-4.1.  Offline Problem
-Fix a bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, and consider (Offline).
+To overcome this, we reduce the offline problem to finding the maximum weight path in an appropriate edge-weighted directed acyclic graph (DAG), which allows for the computation of the optimal offline strategy in poly($m,M$) time.
+4.1. Offline Setting
+Fix a bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, and consider the Problem (Offline).
 The following lemma shows that the objective function can be decomposed across the bid-quantity pairs.
@@
 $\max_{\ell\in[m]}\max_{Q_{1},\dots,Q_{\ell}}\sum_{j=1}^{\ell}\sum_{t=1}^{T}\sum_{k=Q_{j-1}+1}^{Q_{j}}v_{k}\cdot\mathbb{I}{\left[w_{Q_{j}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$
-The proofs of this section are presented in Appendix B. Building on the decomposition in Lemma 4.1, we now construct an edge-weighted DAG $\mathcal{G}(V,E)$ as proposed earlier.
+Proof.
+By Theorem 3.2, a bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell})\rangle\in\mathscr{U}_{m}^{\star}$ for some $\ell\in[m]$ can be uniquely identified by $\{Q_{1},\dots,Q_{\ell}\}$ where $Q_{j}=\sum_{k\leq j}q_{k}$ because $b_{j}=w_{Q_{j}}$ for all $j\in[\ell]$.
+Let the bid history $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$.
+Observe that if the bidder wins $r$ units in any auction, then the $r^{th}$ highest bid value in its submitted strategy must be at least the $r^{th}$ smallest bid in the top $K$ competing bids, i.e., $\bm{\beta}_{-,t}^{-(r)}$.
+Hence, for any round $t$,
+$V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{j=1}^{\ell}\sum_{k=Q_{j-1}+1}^{Q_{j}}v_{k}\cdot\mathbb{I}{\left[w_{Q_{j}}\geq\bm{\beta}_{-,t}^{-(k)}\right]},$
+which implies that the problem (Offline) can be equivalently expressed as
+$\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})=\max_{\ell\in[m]}\max_{Q_{1},\dots,Q_{\ell}}\sum_{j=1}^{\ell}\sum_{t=1}^{T}\sum_{k=Q_{j-1}+1}^{Q_{j}}v_{k}\cdot\mathbb{I}{\left[w_{Q_{j}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$
+Building on the decomposition in Lemma 4.1, we now construct an edge-weighted DAG $\mathcal{G}(V,E)$.
 Vertices.
-The DAG has a ‘layered’ structure with a source ($s$), destination ($d$) and $m$ intermediate layers each with $M$ nodes, denoted by $(\ell,j)$ where $\ell\in[m],j\in[M]$ as shown in Fig.
+The DAG has a ‘layered’ structure with a source node ($s$), destination node ($d$) and $m$ intermediate layers.
+The $\ell^{th}$ intermediate layer contains $M+1$-$\ell$ nodes, denoted by $(\ell,j)$ where $\ell\in[m],j\in\{\ell,\ell+1,\dots,M\}$ as shown in Fig.
 3.
@@
 (i) $x=(\ell-1,j)$ and $y=(\ell,j^{\prime})$, $\forall\ell\in[m]$ and $j<j^{\prime}$. This edge connects consecutive layers in the graph, and its weight is given by:
-$\mathsf{w}(e)=\sum_{t=1}^{T}\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$ (4)
-(ii) $x=(\ell,j)$ and $y=d$, $\forall\ell\in[m],j\in[M]$. This edge connects the current layer to the destination node $d$, and its weight is $\mathsf{w}(e)=0$.
+$\textsf{w}(e)=\sum_{t=1}^{T}\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$ (5)
+(ii) $x=(\ell,j)$ and $y=d$, $\forall\ell\in[m],j\in[M]$. This edge connects the current layer to the destination node $d$, and its weight is $\textsf{w}(e)=0$.
 Figure 3: In this DAG, $M=3$, $m=2$.
@@
 Theorem 4.2.
-  • There exists a bijective mapping between the $s$-$d$ paths in $\mathcal{G}(V,E)$ and bidding strategies in $\mathscr{U}_{m}^{\star}$, i.e., the path $s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for $k\in[m]$ refers to the strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
+There exists a bijective mapping between the $s$-$d$ paths in $\mathcal{G}(V,E)$ and strategies in $\mathscr{U}_{m}^{\star}$, i.e., the path $s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for $k\in[m]$ refers to the strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
 $b_{\ell}=w_{z_{\ell}}\quad\text{and}\quad q_{\ell}=z_{\ell}-z_{\ell-1},\forall\ell\in[k]\,.$
 where $w_{z_{\ell}}$ is defined as per Eq.
-1, $\forall\ell\in[k]$.
+1 and $z_{0}=0$.
 Conversely, the strategy $\mathbf{b}=\langle(w_{Q_{1}},q_{1}),\dots,(w_{Q_{k}},q_{k})\rangle$ maps to the path $s\to(1,Q_{1})\to\dots\to(k,Q_{k})\to d$.
@@
 In the online setting, in each round $t\in[T]$, the learner submits a strategy $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}$ and receives feedback from the auction.
-We consider two feedback models: (a) full information setting and (b) bandit setting, with the latter discussed in Section B.3.1.
-Leveraging the DAG formulation, we design an online learning algorithm such that $\mathsf{REG}=o(T)$.
-Without loss of generaity, assume $v_{j}\in[0,1],\forall j\in[M]$.
-In the full information setting, after receiving feedback at the end of round $t$, a DAG $\mathcal{G}^{t}(V,E)$ is constructed as described earlier. The edge weights are set as follows:
+We consider two feedback models: (a) full information setting and (b) bandit setting.
+Leveraging the DAG formulation, we design a polynomial-time online learning algorithm such that $\textsf{REG}=\widetilde{O}(\text{poly}(m,M)\cdot\sqrt{T})$.
+For the rest of the work, we assume $v_{j}\in[0,1],\forall j\in[M],$ without loss of generality.
+In Theorem 4.2, we established a bijection mapping between ssdd paths in $\mathcal{G}(V,E)$ and safe strategies in $\mathscr{U}_{m}^{\star}$.
+While naïvely applying the Hedge algorithm (Freund and Schapire, 1997) in the full-information setting by treating each path as an expert yields optimal regret bounds, it is computationally intractable due to the exponential number of paths: $O(M^{m})$.
+The situation further worsens in the bandit setting, as implementing the EXP3 algorithm (Auer et al., 2002), treating each path as an expert, also yields unsatisfactory regret bounds (recall that EXP3 obtains $O(\sqrt{NT\log N})$ regret where $N$ is the number of arms and per-round rewards belong to $[0,1]$.
+In this setting, $N=O(M^{m})$.).
+Our proposed algorithm retains the core idea of treating paths as experts but does so efficiently.
+Leveraging the decomposition in Lemma 4.1, it maintains probabilities over edges rather than paths, enabling efficient path sampling as a Markov chain without explicitly maintaining a distribution over all the paths.
+Specifically, in every round $t$, a DAG $\mathcal{G}^{t}(V,E)$, as described earlier, is constructed.
+The proposed algorithm consists of three main steps: (1) UPDATE, (2) SAMPLE, and (3) MAP.
+In the UPDATE step, the bidder uses the feedback from the previous round to update the edge probabilities, $\varphi^{t}(u\to v)$, for every edge $e=u\to v$ in the DAG.
+Using the dynamic programming-based weight-pushing method from Takimoto and Warmuth (2003), this step can be performed in $\text{poly}(m,M)$ time.
+In the SAMPLE step, the algorithm samples an ssdd path in a Markovian manner: starting at the source node $s$, it repeatedly samples the next node $v$ from the set of neighbors of the current node $u$ according to probabilities $\varphi^{t}(u\to v)$, until the destination node $d$ is reached.
+Once a $s$-$d$ path is sampled, in the MAP step, the bidder maps it to a safe strategy $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}$ and submits $\mathbf{b}^{t}$ in round $t$.
+After receiving feedback at the end of the round, in the full information setting, the edge weights are set as follows:
 (i) If $x=(\ell-1,j)$ and $y=(\ell,j^{\prime})$ with $\ell\in[m]$ and $j<j^{\prime}$,
-$\mathsf{w}^{t}(e)=\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$ (5)
-(ii) If $x=(\ell,j)$ and $y=d$, then $\mathsf{w}^{t}(e)=0,\forall\ell\in[m],j\in[M]$.
-The online algorithm in Algorithm 1 consists of three main steps: (1) $\mathsf{UPDATE}$, (2) $\mathsf{SAMPLE}$, and (3) $\mathsf{MAP}$.
-In the first two steps, the algorithm maintains and updates the probabilities over the edges in the constructed DAG and selects a $s$-$d$ path in the DAG by sampling from these probabilities.
-These steps implement exponential weight updates using the weight-pushing method proposed by Takimoto and Warmuth (2003).
-Naively implementing exponential weight updates by treating each $s$-$d$ path (safe strategy) as an expert is intractable as there are $O(M^{m})$ such experts.
-However, following the decomposition in Lemma 4.1, Algorithm 1 only needs to maintain weights for individual edges and then sample a path based on these weights, avoiding the need to track weights for each path separately.
-The $\mathsf{MAP}$ step is the key component of the algorithm that maps the sampled path to a safe strategy (by Theorem 4.2).
-Algorithm 1 Compute Bidding Strategies (Full Information)
-0:  valuation curve $\mathbf{v}$, time horizon $T$, learning rate $\eta>0$.
-1:  for $t=1,2,\dots,T$ do
-2:     Construct $\mathcal{G}^{t}(V,E)$ similar to $\mathcal{G}(V,E)$ without weights.
-3:     $\mathsf{UPDATE}:$ Obtain edge probabilities $\varphi^{t}(\cdot)$ by Algorithm 2. $\mathsf{SAMPLE}$: Define initial node $u=s$ and path $\mathfrak{p}=s$.
-4:     while $u\neq d$ do
-5:        Sample $v$ with probability $\varphi^{t}(u\to v)$.
-6:        Append $v$ to the path $\mathfrak{p}$; set $u\leftarrow v$.
-7:     end while
-8:     $\mathsf{MAP}$: If $\mathfrak{p}=s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for some $k\in[m]$, submit $\mathbf{b}^{t}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
-$b_{\ell}=w_{z_{\ell}}\quad\text{and}\quad q_{\ell}=z_{\ell}-z_{\ell-1},~{}\forall\ell\in[k]\,.$
-where $w_{z_{\ell}}$ is defined as per Eq. 1 for all $\ell\in[k]$.
-9:     The bidder observes $\bm{\beta}_{-}^{t}$ and sets $\mathsf{w}^{t}(e)$ as per Eq. 5.
-10:  end for
+$\textsf{w}^{t}(e)=\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$ (6)
+(ii) If $x=(\ell,j)$ and $y=d$, then $\textsf{w}^{t}(e)=0,\forall\ell\in[m],j\in[M]$.
+The proposed algorithm is formally presented in Algorithm 1. The main result of this section is:
+Algorithm 1 Learning Safe Bidding Strategies (Full Information)
+1:valuation curve $\mathbf{v}$, time horizon $T$, learning rate $\eta\in(0,\frac{1}{M}]$, $\textsf{w}^{0}(e)=0,\text{and}~\varphi^{0}(e)=1,\forall e\in E$.
+2:for $t=1,2,\dots,T$ do
+3:  Construct $\mathcal{G}^{t}(V,E)$ similar to $\mathcal{G}(V,E)$ without weights.
+4:  $\textsf{UPDATE}:$ Set $\Gamma^{t-1}(d)=1$ and recursively compute in bottom-to-top fashion for every node $u$ in $\mathcal{G}^{t}(V,E)$:
+$\Gamma^{t-1}(u)=\sum_{v:u\to v=e\ni E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))\,.$
+5:  For edge $e=u\to v$ in $\mathcal{G}^{t}(V,E)$, update edge probability:
+$\varphi^{t}(e)=\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)}.$
+6:  SAMPLE: Define initial node $u=s$ and path $\mathfrak{p}^{t}=s$.
+7:  while $u\neq d$ do
+8:    Sample $v$ with probability $\varphi^{t}(u\to v)$.
+9:    Append $v$ to the path $\mathfrak{p}^{t}$; set $u\leftarrow v$.
+10:  end while
+11:  MAP: If $\mathfrak{p}^{t}=s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for some $k\in[m]$, submit $\mathbf{b}^{t}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
+$b_{\ell}=w_{z_{\ell}}\qquad\text{and}\qquad q_{\ell}=z_{\ell}-z_{\ell-1},~\forall\ell\in[k]\,.$
+where $w_{z_{\ell}}$ is defined as per Eq. 1 for all $\ell\in[k]$ and $z_{0}=0$.
+12:  The bidder observes $\bm{\beta}_{-}^{t}$ and sets $\textsf{w}^{t}(e)$ as per Eq. 6.
+13:end for
 Theorem 4.3.
-Algorithm 1 runs in $\text{poly}(m,M)$ time per round and achieves $\mathsf{REG}\leq O(M\sqrt{mT\log M})$ under full information feedback, and $\mathsf{REG}\leq O(M\sqrt{m^{3}MT\log M})$ in the bandit setting.
-Algorithm 2 Update Edge Probabilities
-0:  For $t\geq 1$, $\mathcal{G}^{t}(V,E)$. For $t\geq 2$, additionally require $\eta$, $\Gamma^{t-1}(v),\varphi^{t-1}(e),$ and $\mathsf{w}^{t-1}(e)$, $\forall e\in E$ and $\forall v\in V$.
-1:  If $t=1$, – Initialize the edge probability of each edge $e=x\to y$ as $\varphi^{1}(e)=\nicefrac{{1}}{{|\mathcal{N}_{x}|}}$ where $\mathcal{N}_{x}$ is the set of out-neighbors of $x$.
-2:  If $t\geq 2$: – Set $\Gamma^{t-1}(d)=1$ and recursively compute in bottom-to-top fashion for every node $u$ in $\mathcal{G}^{t}(V,E)$:
-$\Gamma^{t-1}(u)=\sum_{v:u\to v=e\ni E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))$
-– For edge $e=u\to v$ in $\mathcal{G}^{t}(V,E)$, update edge probability: $\varphi^{t}(e)=\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)}.$
-We prove Theorem 4.3 by showing that Algorithm 1 is an efficient (polynomial time) and equivalent implementation of the exponential weight updates algorithm.
-We complement Theorem 4.3 by an establishing matching lower bound.
+In the full information setting, Algorithm 1 achieves $\textsf{REG}\leq O(M\sqrt{mT\log M})$ in $\text{poly}(m,M)$ space and time per round.
+In the bandit setting, as the bidder receives only $x(\bm{\beta}^{t})$ and $P(\bm{\beta}^{t})$ as feedback in round $t$, they can not compute Eq.
+6 for all the edges.
+Hence, we use an unbiased estimator, $\widehat{\textsf{w}}^{t}(e)$, of $\textsf{w}^{t}(e)$ as follows:
+$\widehat{\textsf{w}}^{t}(e)$ $=\overline{\textsf{w}}(e)-\frac{\overline{\textsf{w}}(e)-\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}\quad\text{where}\quad p^{t}(e)=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\,.$ (7)
+Here, $p^{t}(e)$ is the probability of selecting edge $e$ in round $t$, $\mathbb{P}^{t}$ is the distribution over all $s$-$d$ paths in $\mathcal{G}^{t}(V,E)$.
+Observe that $\widehat{\textsf{w}}^{t}(e)$ is well defined as $p^{t}(e)>0,\forall e\in E,\forall t\in[T]$.
+Here, $\overline{\textsf{w}}(e)$ is an edge-dependent upper bound of the edge weight $\textsf{w}^{t}(e)$.
+Formally, for $\ell\in[m]$,
+$\overline{\textsf{w}}(e)=\begin{cases}j^{\prime}-j,&\text{if }e=(\ell-1,j)\to(\ell,j^{\prime})\\ 0,&\text{if }e=(\ell,j)\to d\,.\end{cases}$
+The algorithm followed by the bidder in the bandit setting is identical to Algorithm 1, with each instance of $\textsf{w}^{t}(e)$ replaced by $\widehat{\textsf{w}}^{t}(e),\forall t\in[T],\forall e\in E$.
+Furthermore, define $\widehat{\textsf{w}}^{0}(e)=0,\forall e\in E$.
+Clearly, $\mathbb{E}[\widehat{\textsf{w}}^{t}(e)]=\textsf{w}^{t}(e)$.
+The unbiased estimator $\widehat{\textsf{w}}^{t}(e)$ is different from the natural importance-weighted estimator $\frac{\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}$ for technical reasons (similar to Lattimore and Szepesvári (2020, Eq.
+11.6)) which we will clarify in the regret analysis (see details in Section B.3).
+Hence,
+Theorem 4.4.
+In the bandit setting, Algorithm 1 with the aforementioned changes achieves $\textsf{REG}\leq O(M^{2}\sqrt{m^{3}T\log M})$ in $\text{poly}(m,M)$ space and time per round.
+Note that the per-round reward $V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})\in[0,M]$ and number of experts is $N=O(M^{m})$.
+In the bandit setting, although we compute unbiased estimators of the edge weights (equivalently path weights) and perform Hedge-like updates efficiently, the resulting regret bound is tighter than that of a naïve EXP3 implementation.
+This improvement stems from two key advantages: (a) after each round, the bidder can observe the rewards of the constituent edges of the submitted path, rather than just the aggregate reward, [^17] Strictly speaking, this corresponds to the semi-bandit feedback model in online combinatorial optimization (Neu and Bartók, 2013; Jourdan et al., 2021; Brandt et al., 2022), whereas the full bandit model reveals only the total reward.
+However, in our setting, the full bandit model lacks a natural interpretation.
+For simplicity, we refer to this partial feedback setting as the bandit feedback model.
+and (b) the structure of the combinatorial action space reveals partial information about unselected paths as well.
+We complement Theorem 4.3 and Theorem 4.4 by establishing the following lower bound:
 Theorem 4.5.
 Suppose $M\geq 2$ and $m=1$.
-Then, there exist competing bids, $[\bm{\beta}_{-}^{t}]_{t\in[T]}$, such that, under any learning algorithm, the expected regret $\mathbb{E}[\mathsf{REG}]=\Omega(M\sqrt{T})$ in both full information and bandit settings.
+Then, there exist competing bids, $\bm{\beta}_{-}^{t}]_{t\in[T]}$, such that, under any learning algorithm, $\mathbb{E}[\textsf{REG}]=\Omega(M\sqrt{T})$ in the full information setting.
+This implies that the lower bound also holds under bandit feedback.
 5. Competing Against Richer Classes of Bidding Strategies
-In previous sections, we characterized the class of safe bidding strategies and proposed a learning algorithm that obtain sublinear regret, where the regret is computed against a clairvoyant that also selects the optimal safe bidding strategy.
+In previous sections, we characterized the class of safe bidding strategies and proposed a learning algorithm that obtains sublinear regret, where the regret is computed against a clairvoyant that also selects the optimal safe bidding strategy.
 In this section, we consider the cases where the clairvoyant can choose the optimal strategy from richer bidding classes, which we will describe shortly.
-We show that under Algorithm 1, the class of safe bidding strategies achieves sublinear (approximate) regret when competing against these stronger benchmarks.
+We show that the bidder who follows safe bidding strategies and runs Algorithm 1 in the full information setting achieves sublinear (approximate) regret when competing against these stronger benchmarks.
+A similar result also holds in the bandit setting, mutatis mutandis.
 The performance metric we consider in this section is
-$\alpha\text{-}\mathsf{REG}_{\mathscr{B}_{c}}=\alpha\cdot\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\,.$
+$\alpha\text{-}\textsf{REG}_{\mathscr{B}_{c}}=\alpha\cdot\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\,.$ (8)
 Here, $\mathscr{B}_{c}$ is the class of bidding strategies from which the clairvoyant (which serves as our benchmark) chooses the optimal bidding strategy whereas the bidder (learner) submits bids, $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}$, in each round unless stated otherwise.
+Here, $\alpha\in(0,1]$ is defined as the richness ratio.
 Remark 5.1 (Richness Ratio $\alpha$).
-In prior works on approximate regret, $\alpha$ typically measures the hardness of the offline problem, such as $\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (see Streeter and Golovin (2008); Niazadeh et al.
+In prior works on approximate regret, the parameter $\alpha$ typically measures the hardness of the offline problem, such as $\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (see Streeter and Golovin (2008); Niazadeh et al.
 (2022)).
 In contrast, our work reinterprets $\alpha$ as richness ratio, capturing the relative richness of the clairvoyant’s strategy class compared to the learner’s.
-Specifically, it quantifies how closely the value achieved by the optimal strategy in $\mathscr{U}_{m}^{\star}$ approximates that by the optimal strategy in $\mathscr{B}_{c}$.
+Specifically, it quantifies how closely the value achieved by the optimal safe strategy in $\mathscr{U}_{m}^{\star}$ approximates that by the optimal strategy in $\mathscr{B}_{c}$.
 If $\mathscr{B}_{c}=\mathscr{U}_{m}^{\star}$, then $\alpha=1$ (see Section 4).
@@
 for $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$.
-In words, $\Lambda_{\mathscr{B}_{c}}$ is the maximum ratio of the value obtained by the optimal bidding strategy in $\mathscr{B}_{c}$ and that by the optimal strategy in $\mathscr{U}_{m}^{\star}$ over all valuation curves and bid histories in an offline setting.
-We drop the argument $\mathbf{v}$ from $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ when the context is clear.
+In words, $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}$ is the maximum ratio of the value obtained by the optimal bidding strategy in $\mathscr{B}_{c}$ and that by the optimal strategy in $\mathscr{U}_{m}^{\star}$ over all valuation curves and all bid histories of arbitrary length.
 Definition 4 (Richness Ratio).
@@
 Corollary 5.2.
-Suppose the clairvoyant chooses the optimal strategy from the class $\mathscr{B}_{c}$ such that $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}\leq\lambda$ and this bound is tight.
-Then, Algorithm 1 obtains $\frac{1}{\lambda}$-$\mathsf{REG}_{\mathscr{B}_{c}}\leq O(M\sqrt{mT\log M})$ under full information feedback and $\frac{1}{\lambda}$-$\mathsf{REG}_{\mathscr{B}_{c}}\leq O(M\sqrt{m^{3}MT\log M})$ under bandit feedback.
+Suppose the clairvoyant chooses the optimal strategy from the class $\mathscr{B}_{c}$ such that $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}\leq\lambda$ and this bound is tight implying the richness ratio of $\mathscr{B}_{c}$ is $\alpha=\frac{1}{\lambda}$.
+Then, Algorithm 1 obtains $\alpha$-$\textsf{REG}_{\mathscr{B}_{c}}\leq O(M\sqrt{mT\log M})$ in the full information setting.
 Proof.
 As $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}\leq\lambda$,
-$\frac{1}{\lambda}\text{-}\mathsf{REG}_{\mathscr{B}_{c}}$ $=\frac{1}{\lambda}\cdot\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]$
-$\leq\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]=\mathsf{REG}\,.\vspace{-0.5cm}$
-By Theorem 4.3, we get the stated regret upper bound. ∎
-5.1.  Richness Ratio of $m$-uniform Non-Safe Strategies
-Here, we consider the clairvoyant class $\mathscr{B}_{c}$ to be the class of bidding strategies containing at most $m$ bid-quantity pairs, which are feasible only for the given bid history $\mathcal{H}_{-}$, rather than for every possible sequence of competing bids as in the case of safe strategies.
-The following theorem characterizes the richness ratio of this class, denoted as $\mathscr{F}^{\mathcal{H}_{-}}_{m}$:
+$\alpha\text{-}\textsf{REG}_{\mathscr{B}_{c}}$ $=\frac{1}{\lambda}\cdot\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]$
+$\leq\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]=\textsf{REG},$
+where the first equality holds as $\alpha=\frac{1}{\lambda}$. By Theorem 4.3, we get the stated regret upper bound. ∎
+Recall that we defined richness of a class of bidding strategies along two dimensions: (i) the safety of strategies (safe vs.
+only RoI-feasible), and (ii) expressiveness, i.e., the number of bid-quantity pairs allowed.
+In the subsequent sections, we consider classes of strategies that are richer along either one or both of the said dimensions and compute the corresponding richness ratios, $\alpha$.
+5.1. The Class of RoI-feasible Strategies with $\leq m$ Bid-Quantity Pairs
+First, we consider the clairvoyant class $\mathscr{B}_{c}$ to be the class of strategies containing at most $m$ bid-quantity pairs, which are RoI-feasible only for the given bid history $\mathcal{H}_{-}$, rather than for every possible sequence of competing bids as in the case of safe strategies.
+We denote this class by $\mathscr{F}^{\mathcal{H}_{-}}_{m}$.
+Then,
 Theorem 5.3.
@@
 Thus, the richness ratio of the class $\mathscr{F}^{\mathcal{H}_{-}}_{m}$ is $\alpha=1/2$.
+We present the proof of Theorem 5.3 in Section 5.4.
 Theorem 5.3 implies that restricting the learner to safe bidding strategies has a bounded cost and does not lead to an arbitrary loss in the obtained value as the upper bound is independent of $m$.
@@
 In practice, we expect these strategies to perform near-optimally, a claim further supported by the numerical simulations presented in Section 5.3.
-Proof Sketch of Theorem 5.3 (Upper Bound).
-We now present the central ideas to prove the upper bound, which are also used to show the subsequent theorems, albeit with some changes specific to each case.
-Fixing a bid history, $\mathcal{H}_{-}$, we consider the expression in Section 5.
-Maximizing over all bid histories and valuation curves, we get the desired result.
-Firstly, we bridge the we bridge the gap between the bidding strategies of the clairvoyant and the learner by constructing a restricted class of safe strategies, $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})\subseteq\mathscr{U}_{m}^{\star}$.
-This restricted class is exponentially smaller in size and is derived based on the optimal strategy selected by the clairvoyant (see exact characterization in Eq.
-17 for Theorem 5.3 and Algorithm 3 for Theorem 5.4).
-We demonstrate that considering this restricted class of safe strategies is sufficient to derive tight upper bounds.
-Second, as a crucial step, we show that for any $m$-uniform safe strategy $\mathbf{b}=\langle(w_{Q_{1}},q_{1}),\dots,(w_{Q_{m}},q_{m})\rangle$,
-$V(\mathbf{b};\bm{\beta}_{-})=\max_{\ell\in[m]}V((w_{Q_{\ell}},Q_{\ell});\bm{\beta}_{-})$
-for all competing bids $\bm{\beta}_{-}$, where, $Q_{\ell}=\sum_{j=1}^{\ell}q_{j}$.
-Thus, the value obtained by any $m$-uniform safe strategy can be determined by knowing the value obtained by the strategies in $\mathscr{U}_{1}^{\star}$, allowing us to focus solely on 1-uniform strategies.
-As the third key idea, we partition the $T$ rounds into sets $T_{j}$, based on the least winning bid of the optimal bidding strategy chosen by the clairvoyant (see details in Eq.
-15) and derive lower bounds on the value obtained by 1-uniform strategies in $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$ in terms of the value obtained by optimal strategy chosen by the clairvoyant in each partition $T_{j}$.
-With these three key ideas, we bound the ratio in Section 5.
-The detailed proof is presented in Section C.1.
-Proof Sketch of Theorem 5.3 (Tight Lower Bound).
-For any $\delta\in(0,1]$, we construct a tight lower-bound instance for any $m\in\mathbb{N}$ by considering an offline problem where $M=T=N^{2m}$ and $K=M+1$, with $N=O(1/\delta)$.
-We set $\mathbf{v}=[1,v,\dots,v]\in\mathbb{R}^{M}$, where $v=1-O(\delta)$.
-The sequence of $T$ auctions is divided into $2m$ partitions of exponentially varying sizes, each with an identical competing bid profile.
-In each bid profile, there are two types of bids: ‘big’ and ‘small’.
-All ‘big’ bids are set to $C\gg 1$, while the ‘small’ bids remain identical within a partition.
-The partition sizes, the number of ‘small’ competing bids, and the exact values of these bids are carefully chosen to ensure that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-};\mathbf{v})\geq 2-\delta$ (see Sections D.1 to D.2).
-5.2.  Richness Ratio of $m^{\prime}$-uniform Strategies
-We first consider the case when the clairvoyant selects the optimal strategy from the class of safe bidding strategies with at most $m^{\prime}$ bid-quantity pairs where $m^{\prime}\geq m$.
+5.2. Classes of Strategies with $\leq m^{\prime}$ Bid-Quantity Pairs For $m^{\prime}\geq m$
+We now consider the setting where the clairvoyant selects the optimal strategy from the class of safe strategies with at most $m$’ bid-quantity pairs, where $m’\geq m$.
+This strategy class, $\mathscr{U}_{m’}^{\star}$, is richer than the bidder’s class $\mathscr{U}_{m}^{\star}$, although, along the expressiveness dimension compared to the class $\mathscr{F}^{\mathcal{H}_{-}}_{m}$ discussed in the previous section, which is richer along the safety dimension.
+Then,
 Theorem 5.4.
 For any $m,m^{\prime}\in\mathbb{N}$ such that $m^{\prime}\geq m$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}\leq\frac{m^{\prime}}{m}$.
-Additionally, there exist $\mathcal{H}_{-}$ and $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{m^{\prime}}{m}-\delta$.
+Additionally, there exist $\mathcal{H}_{-}$ and $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$,
+$\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{m^{\prime}}{m}-\delta\,.$
 Thus, the richness ratio of the class $\mathscr{U}_{m^{\prime}}^{\star}$ is $\alpha=m/m^{\prime}$.
-Suppose the clairvoyant can now select the optimal strategy from $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$, the class of bidding strategies with at most $m^{\prime}$ bid-quantity pairs ($m^{\prime}\geq m$) that are feasible for the bid history $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, but not necessarily safe.
+Finally, suppose the clairvoyant can now select the optimal strategy from $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$, the class of bidding strategies with at most $m^{\prime}$ bid-quantity pairs ($m^{\prime}\geq m$) that are only RoI-feasible for the bid history $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$.
+This strategy class is richer than the bidder’s strategy class along both dimensions.
 Then,
@@
 For any $m,m^{\prime}\in\mathbb{N}$ such that $m^{\prime}\geq m$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}\leq\frac{2m^{\prime}}{m}$.
-Moreover, there exist $\mathcal{H}_{-}$ and $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{2m^{\prime}}{m}-\delta$.
+Moreover, there exist $\mathcal{H}_{-}$ and $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, we have
+$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{2m^{\prime}}{m}-\delta\,.$
 Thus, the richness ratio of the class $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$ is $\alpha=m/2m^{\prime}$.
 By Corollary 5.2, for $m^{\prime}\geq m$, Algorithm 1 obtains $\frac{m}{m^{\prime}}$-approximate (resp.
-$\frac{m}{2m^{\prime}}$-approximate) sublinear regret in the online setting when the clairvoyant draws the optimal strategy from the class of safe (resp.
-feasible but not necessarily safe) strategies with at most $m^{\prime}$ bid-quantity pairs.
-The richness ratio in Theorem 5.4 and Theorem 5.5 can theoretically become arbitrarily small as $m^{\prime}$ grows, for a fixed $m$.
-However, the bidder has the flexibility to choose $m$; increasing $m$, though, incurs higher space and time complexity for Algorithm 1.
-Moreover, for any given pair $(m,m^{\prime})$, the bound is tight under a highly non-trivial setting (see details in Sections D.3 to D.4).
-In practice, as shown in Section 5.3, the observed richness ratios are substantially better than the theoretical bounds, suggesting that the worst-case scenarios are unlikely to arise in real-world applications.
-5.3.  Estimating The Richness Ratio: A Case Study
-In this section, we estimate the richness ratio (equivalently $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}$) for the discussed classes of bidding strategies by experiments using EU ETS emission permit auction data from 2022 and 2023.
-Detailed methods are provided in Appendix E, and the results are shown in Fig.
-4.
+$\frac{m}{2m^{\prime}}$-approximate) sublinear regret (same as Theorem 4.3) in the online setting when the clairvoyant draws the optimal strategy from the class of safe (resp.
+RoI-feasible) strategies with at most $m^{\prime}$ bid-quantity pairs.
+The richness ratio in Theorem 5.4 and Theorem 5.5 can theoretically become arbitrarily small as $m^{\prime}$ increases, for a fixed $m$.
+However, the bidder retains the flexibility to choose $m$: increasing $m$, though, incurs higher space and time complexity for Algorithm 1.
+Moreover, for any given pair $(m,m^{\prime})$, the bounds are tight for carefully constructed non-trivial hard instances (see details in Sections D.2 to D.3).
+In practice, as shown in Section 5.3, the empirical richness ratios are substantially better than the worst-case bounds, suggesting that such pathological scenarios are unlikely to arise in real-world settings.
+We conclude the section with two key observations.
+First, when $m=m$’, Theorem 5.4 implies $\alpha=1$, which corresponds to the setting where both the bidder and the clairvoyant follow the class of safe strategies with at most $m$ bid-quantity pairs, the case considered in Section 4.
+Similarly, when $m=m$’, Theorem 5.5 yields $\alpha=\frac{1}{2}$, recovering the result of Theorem 5.3.
+Second, recall that we defined the richness of a strategy class along two dimensions: safety and expressiveness.
+The results in this section imply that when a benchmark strategy class is richer along both dimensions (Theorem 5.5), its overall richness ratio is the product of the richness ratios along each individual dimension (Theorem 5.3 and Theorem 5.4).
+5.3. Computing Empirical Richness Ratio: A Case Study
+As discussed in the previous section, constructing worst-case instances for showing tight bounds are significantly intricate that may not reflect real world settings.
+This motivates us to empirically estimate how well do the class of safe strategies approximate the richer strategy classes in practice.
+To this end, with slight abuse of notation, we define the empirical richness ratio as
+$\alpha_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})=\frac{1}{\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})},$
+where $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ is defined in Section 5.
+We assess the tightness of the theoretical bounds in practice by computing $\alpha_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ using the EU ETS emission permit auction data from 2022 and 2023.
+However, only aggregate statistics of the submitted bids is publicly available for privacy reasons.
+Hence, we synthesize individual level bid data from these available statistics.
+The exact procedure to reconstruct the bid data is presented in Section F.1.
 Figure 4: The left (resp.
-right) figure refers to (the upper bound on) $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ (resp.
-$\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-})$) as a function of $m$.
-Estimating $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$.
+right) figure refers to (the lower bound on) $\alpha_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ (resp.
+$\alpha_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$) as a function of $m$.
+The shaded area corresponds to one standard deviation.
+Estimating $\alpha_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$.
+Since computing the exact value attained by the optimal strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m}$ is non-trivial, we instead compute a uniform upper bound that is independent of $m$ (see Section F.2).
 The left plot of Fig.
-4 shows that (the upper bound for) $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ is significantly better than the theoretical bound of 2.
-The decaying trend of the plot with increasing $m$ can be attributed to the fact that we are computing an upper bound for the value achieved by the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m}$, which is independent of $m$, rather than the exact value (see Section E.2).
-For $m\geq 4$, even the upper bound of $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})\sim 1.05$ indicating that the safe bidding strategies are near-optimal.
-Estimating $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-})$.
-From the right plot of Fig.
-4, we observe that the empirical values of $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-})$ is significantly better than the worst-case bound $m$ (Theorem 5.4).
-In fact, the gain obtained by increasing the number of bid-quantity pairs plateaus after $m=4$ and even for $m=10$, the ratio of the value obtained by the optimal bidding strategy with at most 10 bid-quantity pairs to that by optimal 1-uniform bidding strategy is $\sim 1.25$.
-Both plots show that the richness ratio is significantly better in practice than the theoretical bounds, suggesting that safe strategies with small $m$ are near-optimal in practice.
+4 shows that the empirical lower bound of $\alpha_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ is substantially better than the theoretical worst-case bound of $1/2$.
+In particular, for $m\geq 4$, even the lower bound exceeds $0.95$, indicating that safe bidding strategies are nearly optimal in practice.
+Estimating $\alpha_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$.
+As shown in the right plot of Fig.
+4, the empirical values of $\alpha_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ significantly exceed the worst-case theoretical bound of $1/m$ (Theorem 5.4).
+The marginal gain from increasing the number of bid-quantity pairs plateaus beyond $m=4$; even at $m=10$, the ratio of the value achieved by the optimal 1-uniform safe strategy to that of the optimal safe strategy with at most 10 bid-quantity pairs is $\sim 0.8$, in contrast to the worst-case bound of $0.1$.
+Both these results suggest that empirical richness ratios are substantially better than the worst-case guarantees, implying that safe strategies with small $m$ offer near-optimal performance in realistic settings.
+5.4. Proof of Theorem 5.3
+We present the proof of Theorem 5.3 below.
+While Theorem 5.4 follows a similar high-level approach, it requires different techniques, which we detail in Section C.4.
+Finally, Theorem 5.5 follows directly from the proofs of Theorem 5.3 and Theorem 5.4; see Section C.5 for details.
+To prove that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}\leq 2$ for any $m\in\mathbb{N}$, consider the metric presented in Section 5 for any bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$,
+$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})=\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$ (10)
+To prove the upper bound in Theorem 5.3, we show that for any $\mathcal{H}_{-}$ and $\mathbf{v}$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\leq 2$.
+Maximizing over all bid histories and valuation vectors, we get the desired result.
+Let $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{m}^{*},q_{m}^{*})\rangle$ where $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right):=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$.
+[^18] We assumed $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{m}^{*},q_{m}^{*})\rangle$ without loss of generality.
+The proof also follows if we considered $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{k}^{*},q_{k}^{*})\rangle$ for some $k\in[m]$ instead.
+Define $Q_{\ell}^{*}=\sum_{j\leq\ell}q_{j}^{*}$.
+Suppose $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated $r_{t}^{*}$ units in round $t$.
+For any $j\in[m]$, let $T_{j}$ be the set of rounds in which the least winning bid is $b_{j}^{*}$, i.e.,
+$T_{j}=\big\{t\in[T]:Q_{j-1}^{*}<r_{t}^{*}\leq Q_{j}^{*}\big\}\,.$ (11)
+For any $j\in[m]$, partition $T_{j}$ into $T_{j,0}$ and $T_{j,1}$ such that $T_{j,0}$ is the set of rounds where the bidder gets strictly less than $Q_{j}^{*}$ units and $T_{j,1}$ is the set of rounds when they get exactly $Q_{j}^{*}$ units:
+$T_{j,0}=\{t\in T_{j}:r_{t}^{*}<Q_{j}^{*}\},\qquad T_{j,1}=\{t\in T_{j}:r_{t}^{*}=Q_{j}^{*}\}\,.$ (12)
+For any $j\in[m]$, let
+[…]
+In other words, $\widehat{Q}_{j}$ is the $2^{nd}$ highest number of units allocated to $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ over the rounds in $T_{j}$ (or the highest if $T_{j,0}=\emptyset$ or $T_{j,1}=\emptyset$).
+Step 1.
+Constructing a Restricted Class of Safe Bidding Strategies.
+Here, as a crucial part of the proof, we construct a restricted class of safe bidding strategies, denoted by $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, where $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})\subset\mathscr{U}_{m}^{\star}$.
+This construction serves two purposes.
+First, it reduces the search space for the optimal safe bidding strategy.
+Second, and more importantly, it enables us to establish a connection between the optimal safe bidding strategy and $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$.
+In defining the restricted class, we use the quantities, $\{\widehat{Q}_{j},Q_{j}^{*}\}_{j\in[m]}$ as follows:
+$\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})=\Big\{\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}},~~Q_{\ell}\in\{\widehat{Q}_{\ell},Q_{\ell}^{*}\},\quad\forall\ell\in[m]\Big\}\,.$ (13)
+Recall that any strategy in $\mathscr{U}_{m}^{\star}$ with $m$ bid-quantity pairs takes the form of $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}},\forall\ell\in[m]$.
+The strategies in $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$ also have the same structure, but as an important difference, for any $\ell\in[m]$, we enforce $Q_{\ell}\in\{\widehat{Q}_{\ell},Q_{\ell}^{*}\}$.
+Observe that for any $\ell\in[m-1]$,
+$Q_{\ell}\leq Q_{\ell}^{*}<\widehat{Q}_{\ell+1}\leq Q_{\ell+1},$
+where the first and third inequalities follow directly from the definition of $Q_{\ell}$ and the second one follows from the definition of $\widehat{Q}_{\ell+1}$.
+So, $Q_{\ell}$’s are distinct and ordered.
+Further observe that the number of bidding strategies in $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$ is $O(2^{m})$; significantly smaller than the number of strategies in $\mathscr{U}_{m}^{\star}$, which is $O(M^{m})$.
+With the definition of the restricted class of safe bidding strategies, we have
+$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})=\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\leq\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$ (14)
+Step 2.
+Value Decomposition for $m$-uniform Strategies.
+We now present a crucial result that expresses the value obtained by a $m$-uniform strategy as a function of the value obtained by $m$ 1-uniform strategies.
+Lemma 5.6.
+Let $\mathbf{b}=\langle(w_{Q_{1}},q_{1}),\dots,(w_{Q_{m}},q_{m})\rangle$ be a $m$-uniform safe bidding strategy for some $m\in\mathbb{N}$. Then, for any competing bids $\bm{\beta}_{-}$,
+$V(\mathbf{b};\bm{\beta}_{-})=\max_{\ell\in[m]}V((w_{Q_{\ell}},Q_{\ell});\bm{\beta}_{-}),$
+where we recall that $Q_{\ell}=\sum_{j=1}^{\ell}q_{j},\forall\ell\in[m]$.
+A key consequence of Lemma 5.6 is that knowing the value obtained by the strategies in $\mathscr{U}_{1}^{\star}$ is sufficient to compute the value obtained any $m$-uniform safe bidding strategy.
+For any $\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, where $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}}$, and $Q_{\ell}\in\{Q_{\ell}^{*},\widehat{Q}_{\ell}\}$ for any $\ell\in[m]$, partition the set $[m]$ into $\mathcal{S}^{*}$ and $\widehat{\mathcal{S}}$ as follows:
+$\mathcal{S}^{*}=\{j:Q_{j}=Q_{j}^{*},j\in[m]\}\quad\text{ and }\quad\widehat{\mathcal{S}}=\{j:Q_{j}=\widehat{Q}_{j}<Q_{j}^{*},j\in[m]\}\,.$
+Hence, by Lemma 5.6, for any round $t\in T_{j}$, and $\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, we get
+$V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\max_{\ell\in[m]}V((w_{Q_{\ell}},Q_{\ell});\bm{\beta}_{-}^{t})\geq V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\,$, (15)
+$\implies\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $\geq\sum_{j\in\mathcal{S}^{*}}\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})+\sum_{j\in\widehat{\mathcal{S}}}\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\,.$ (16)
+Step 3.
+Allocation Lower Bounds for 1-uniform Strategies.
+Let $V_{j,k}$ be the time-average value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in the set of rounds in $T_{j,k}$ (as defined in Eq.
+12).
+Formally, $\forall j\in[m],k\in\{0,1\}$,
+$V_{j,k}=\dfrac{1}{|T_{j,k}|}\sum_{t\in T_{j,k}}V(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right),\bm{\beta}_{-}^{t})\,.$
+Define ${W}_{k}=\sum_{j=1}^{k}v_{j}$. Note that $V_{j,1}={W}_{Q_{j}^{*}}$ because for any $t\in T_{j,1}$, we have $r_{t}^{*}=Q_{j}^{*}$.
+For any $j\in[m]$, we present the following lemma to establish lower bound on $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})$.
+Lemma 5.7.
+Let $\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, where $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{j}=w_{Q_{j}}$, and $Q_{j}\in\{Q_{j}^{*},\widehat{Q}_{j}\},\forall j\in[m]$.
+Then, for any $j\in[m]$,
+  • (a)
+    if $Q_{j}=\widehat{Q}_{j}<Q_{j}^{*}$ (i.e., $j\in\widehat{\mathcal{S}}$), we have
+    $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\geq V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\,.$ (17)
+  • (b)
+    If $Q_{j}=Q_{j}^{*}$ (i.e., $j\in\mathcal{S}^{*}$), we have
+    $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\geq{W}_{Q_{j}^{*}}|T_{j,1}|\,.$ (18)
+Note that the right hand side of Eq.
+16 depends only on the choice of the partitions $\mathcal{S}^{*}$ and $\widehat{\mathcal{S}}$.
+Substituting the lower bound from Lemma 5.7 in Eq.
+16, we establish that,
+[…]
+$\leq\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{(\mathcal{S}^{*},\widehat{\mathcal{S}})}\Big\{\sum_{j\in\mathcal{S}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)\Big\}}\,.$ (19)
+We complete the proof by invoking the following result:
+Lemma 5.8.
+Let $(\mathcal{S}_{0}^{*},\widehat{\mathcal{S}}_{0})=\mathop{\mathrm{argmax}}_{(\mathcal{S}^{*},\widehat{\mathcal{S}})}\Big\{\sum_{j\in\mathcal{S}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)\Big\}$.
+Then,
+$\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)}\leq 2\,.$
+From Eq. 19 and Lemma 5.8, we get $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\leq 2$ as desired.
+5.4.1. Tight Lower Bound for $m=1$.
+We complement the proof of the upper bound of Theorem 5.3 by constructing a tight lower bound instance for $m=1$.
+The case for general $m\geq 2$ is deferred to Section D.1.
+While the key ideas behind the hard instance construction remain the same for both cases, the $m\geq 2$ case is substantially more intricate and is therefore presented separately.
+Formally, for any $\delta\in(0,\frac{1}{2}]$, we construct a bid history, $\mathcal{H}_{-}$, and valuation vector, $\mathbf{v}$ for which $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{1},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq 2-\delta$.
+Let $M=T=2\left\lceil\frac{1}{\delta}\right\rceil$ and $K=M+1$.
+Suppose $\mathbf{v}=[1,v,\cdots,v]\in\mathbb{R}^{M}$, target RoI $\gamma=0$ and $v=1-4\epsilon$ where $\epsilon=\frac{\delta-1/M}{4(1-1/M)}<\frac{\delta}{4}\leq\frac{1}{8}$.
+The bid history is defined as:
+[…]
+where $C\gg 1$. The bid history is presented in Table 1.
+Table 1: Bid history for tight lower bound for $m=1$. Here, $C\gg 1$ and $\epsilon>0$ is a small real number.
+$t=1$ $t=2$ $\cdots$ t=M1t=M-1 $t=M$
+$C$ $C$ $\cdots$ $C$ $\epsilon$
+$\vdots$ $\vdots$ $\cdots$ $\vdots$ $\vdots$
+$C$ $C$ $\cdots$ $C$ $\vdots$
+$1-\epsilon$ $1-\epsilon$ $\cdots$ $1-\epsilon$ $\epsilon$
+Notice that the constructed $\mathcal{H}_{-}$ contains M1M-1 rounds with high competition (where the bidder can acquire at most 1 unit), while there is one round with minimal competition, allowing the bidder to obtain any desired number of units.
+The 1-uniform safe bidding strategies of the form $(w_{q},q)$ do not perform well universally on all rounds because with increase in $q$, despite the increasing demand (i.e., $q$), the bid value $w_{q}$ decreases, thereby reducing the likelihood of acquiring any units.
+For a given $\mathcal{H}_{-}$, recall that for any $m\in\mathbb{N}$, $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$.
+We define
+$V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)=\sum_{t=1}^{T}V(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t}),$
+the value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ over the bid history $\mathcal{H}_{-}$. Similarly, define
+$\mathbf{b}_{m}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})~\quad\text{and}\quad V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)=\sum_{t=1}^{T}V(\mathbf{b}_{m}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})\,.$
+Computing $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$.
+Observe that no strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{1}$ can obtain more than 1 unit in the first M1M-1 rounds and $M$ units in the final round.
+It is easy to verify that $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=(1,M)$ as it obtains the maximum number of units possible for the constructed $\mathcal{H}_{-}$.
+So, $V^{\mathsf{OPT}}_{1}\left(\mathcal{H}_{-}\right)=M+v(M-1)$.
+Computing $\mathbf{b}_{1}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$.
+We claim that $\mathbf{b}_{1}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)=(w_{1},1)=(1,1)\implies V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)=M$.
+As $\bm{\beta}_{-,t}^{-(1)}=1-\epsilon>1-2\epsilon=w_{2}$ for $t\in[M-1]$, the strategy $\left(w_{q},q\right)$ for $2\leq q\leq M$ does not get any value in the first M1M-1 rounds.
+Bidding $\left(w_{q},q\right)$ gets exactly $q$ units in round $M$ as $w_{q}>\epsilon$ for any $q\geq 2$.
+So, for $2\leq q\leq M$, the total value obtained by bidding $(w_{q},q)$ is 1+(q1)v<q1+(q-1)v<q which proves that $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)=M$.
+So,
+$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{1},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})=\frac{V^{\mathsf{OPT}}_{1}\left(\mathcal{H}_{-}\right)}{V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}=\frac{M+v(M-1)}{M}=\frac{2M-1-4\epsilon(M-1)}{M}=2-\delta\,.$
+6. Extensions
+In this section, we consider several important extensions of the model introduced in Section 2 by relaxing some of its assumptions.
+We first consider the setting where the competing bids are generated by an adaptive adversary instead of an oblivious adversary.
+Next, we present an efficient heuristic for the bidder to follow when RoI constraints are enforced over a sliding window of $T_{0}\in\mathbb{N}$ rounds.
+Finally, we examine the case where the valuation vectors vary across rounds—either stochastically or adversarially.
+In each case, we build upon the safe bidding strategies and learning algorithm introduced in Section 4.
+6.1. Adaptive Adversary
+In many practical marketplaces, particularly in repeated auctions, the actions of a bidder may influence the behavior of competitors over time.
+Competing bidders can observe past auction outcomes, infer bidding patterns, and adapt their strategies accordingly.
+This motivates extending our model to an adaptive adversary setting, where the competing bids in each round may depend on the bidder’s past submitted strategies.
+Formally, we consider the model presented in Section 2 with the crucial change that the competing bids are now generated by an adaptive adversary, i.e., the competing bids in any round $t$ can depend on the strategies submitted by the bidder in previous rounds, $\mathbf{b}^{1},\dots,\mathbf{b}^{t-1}$.
+For an adaptive adversary, we consider the following performance metric:
+$\textsf{REG}_{\text{adap}}=\mathbb{E}\left[\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})\right]\,$, (20)
+where the expectation is with respect to the randomness of the learning algorithm of the bidder.
+[^19] Without loss of generality, we assume that the mapping from past bidding strategies, $\mathbf{b}^{1},\dots,\mathbf{b}^{t-1}$, to current competing bids, $\bm{\beta}_{-}^{t}$, is deterministic (Dani and Hayes, 2006, Remark 3.1).
+Observe that unlike the case for an oblivious adversary (see Eq.
+4), the first term in Eq.
+20 is not fixed in advance for an adaptive adversary as the competing bids $\bm{\beta}_{-}^{t}$ can adaptively change as a function of historical bids.
+In the remaining portion of this section, we propose learning algorithms that yield $\textsf{REG}_{\text{adap}}=o(T)$ both under the full-information and bandit feedback model, when the competing bids are generated by an adaptive adversary.
+All the remaining modeling assumptions from Section 2 continue to hold without any change.
+Full Information Setting.
+In our setting, a learning algorithm is said to be forgetful (Dani and Hayes, 2006), if the probability distribution for choosing a strategy $\mathbf{b}^{t}$ in round $t$ is fully determined by the past competing bids $\bm{\beta}_{-}^{1},\dots,\bm{\beta}_{-}^{t-1}$, and does not explicitly depend on the bidder’s actions $\mathbf{b}^{1},\dots,\mathbf{b}^{t-1}$.
+In particular,
+$\mathbb{P}[\mathbf{b}^{t}=\mathbf{b}|\bm{\beta}_{-}^{1},\dots,\bm{\beta}_{-}^{t-1},\mathbf{b}^{1},\dots,\mathbf{b}^{t-1}]=\mathbb{P}[\mathbf{b}^{t}=\mathbf{b}|\bm{\beta}_{-}^{1},\dots,\bm{\beta}_{-}^{t-1}],\quad\forall\mathbf{b}\in\mathscr{U}_{m}^{\star}\,.$ (21)
+Now, we state the following result that rephrases (Cesa-Bianchi and Lugosi, 2006, Lemma 4.1):
+Lemma 6.1.
+Suppose a learning algorithm satisfies Eq.
+21 and achieves $\textsf{REG}\leq B$ against an oblivious adversary.
+Then, the same algorithm also yields $\textsf{REG}_{\text{adap}}\leq B$ when deployed against an adaptive adversary.
+From Section 4, recall that in the full-information setting, Algorithm 1 is an equivalent implementation of the Hedge algorithm. Hence, for any learning rate $\eta>0$,
+$\mathbb{P}[\mathbf{b}^{t}=\mathbf{b}|\bm{\beta}_{-}^{1},\dots,\bm{\beta}_{-}^{t-1},\mathbf{b}^{1},\dots,\mathbf{b}^{t-1}]=\frac{\exp(\eta\sum_{\tau=1}^{t-1}V(\mathbf{b};\bm{\beta}_{-}^{\tau}))}{\sum_{\mathbf{b}^{\prime}\in\mathscr{U}_{m}^{\star}}\exp(\eta\sum_{\tau=1}^{t-1}V(\mathbf{b}^{\prime};\bm{\beta}_{-}^{\tau}))}=\mathbb{P}[\mathbf{b}^{t}=\mathbf{b}|\bm{\beta}_{-}^{1},\dots,\bm{\beta}_{-}^{t-1}]\,.$
+This implies Algorithm 1 is forgetful and hence yields $\textsf{REG}_{\text{adap}}\leq O(M\sqrt{mT\log M})$ in the full-information setting even when the competing bids are generated by an adaptive adversary.
+Bandit Setting.
+In the bandit feedback model, the forgetfulness property of Algorithm 1 no longer holds, because the unbiased estimator of the value of a strategy $\mathbf{b}$ depends on the bidder’s past actions.
+This dependency necessitates a different approach for controlling regret.
+To bound $\textsf{REG}_{\text{adap}}$, we begin by analyzing the random regret, defined as
+$\widehat{\textsf{REG}}_{\text{adap}}=\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}V(\mathbf{b}^{t};\bm{\beta}_{-}^{t}),$
+and we show that it is bounded with high probability. We then relate this high-probability bound to the expected regret using the following lemma:
+Lemma 6.2.
+For any $\delta\in(0,1]$, if $\widehat{\textsf{REG}}_{\text{adap}}\leq B(\delta)$ with probability at least $1-\delta$, then
+$\textsf{REG}_{\text{adap}}\leq B(\delta)+MT\delta.$
+To address the challenge posed by adaptive adversaries, we develop a new algorithm inspired by the EXP3.P algorithm (Auer et al., 2002), which provides high-probability regret guarantees.
+Our method builds on Algorithm 1 and incorporates ideas from György et al.
+(2007).
+The algorithm maintains a covering path set $\mathcal{C}$, which is a collection of ssdd paths in the DAG $\mathcal{G}(V,E)$ such that every edge $e\in E$ appears in at least one path $\mathfrak{p}\in\mathcal{C}$.
+At each round $t$, the algorithm samples a path using a mixture strategy: with probability $\lambda>0$, it samples uniformly at random from $\mathcal{C}$; with probability $1-\lambda$, it updates edge weights and samples a path as a Markov chain, following the procedure of Algorithm 1.
+The resulting path distribution is thus a convex combination of the uniform distribution over $\mathcal{C}$ and the distribution induced by Hedge-like updates.
+See Algorithm 3 in Section E.3 for implementation details.
+This yields the following regret bound:
+Theorem 6.3.
+Against an adaptive adversary, there exists a learning algorithm that requires $\text{poly}(m,M)$ space, runs in $\text{poly}(m,M)$ time per round and, for any $\delta\in(0,1]$ and sufficiently large $T$, guarantees
+$\widehat{\textsf{REG}}_{\text{adap}}\leq O\left(m^{3/2}M^{2}\sqrt{T\log M}+mM^{2}\sqrt{T\log(M/\delta)}\right)$
+with probability at least $1-\delta$ under bandit feedback. Setting $\delta=1/T$ gives the bound
+$\textsf{REG}_{\text{adap}}\leq O\left(m^{3/2}M^{2}\sqrt{T\log(MT)}\right).$
+6.2. Cumulative RoI Constraints
+In this section, we consider a natural relaxation of the per-round RoI constraints introduced in Section 2.
+Specifically, we assume that bidders aim to satisfy the RoI constraint in aggregate over the entire horizon of $T$ rounds, which is a common assumption in online ad auction literature (Lucier et al., 2024; Feng et al., 2023; Deng et al., 2021, 2023a).
+In this setting, we define regret as
+$\textsf{REG}_{\text{cumul}}=\max_{\mathbf{b}\in\mathscr{B}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})],$ (22)
+where the expectation is with respect to any randomness in the learning algorithm.
+Here, $\mathscr{B}$ is the class of strategies that satisfy cumulative RoI constraints over $T$ rounds for the sequence of competing bids $\bm{\beta}_{-}^{t}]_{t\in[T]}$.
+In this setting, we consider the following assumption on the bidder’s behavior:
+Assumption 6.4.
+For adversarial competing bids, a value maximizing bidder aiming to satisfy cumulative RoI constraints over $T$ rounds does not submit non-safe strategies, i.e., overbidding strategies per Definition 3, in any round until they have accrued a non-zero value from previous auctions.
+Formally, for any round $t\in[T]$, if $\sum_{\tau=1}^{t-1}V(\mathbf{b}^{\tau};\bm{\beta}_{-}^{\tau})=0$, then the bidder submits $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}$.
+6.4 is a natural assumption in adversarial settings.
+We argue that if a bidder does not adhere to 6.4, then there exists a sequence of competing bids under which their cumulative RoI constraint will necessarily be violated.
+To illustrate this, consider a bidder, not having won any units in rounds $1,\dots,t-1$, who overbids in round $t$ and wins at least one unit violating the round-wise RoI constraint.
+[^20] For any overbidding strategy per Definition 3, there exists a competing bid profile for which the RoI constraint in that round is violated.
+See proof of Theorem 3.1 for details.
+Suppose now that in all subsequent rounds $t+1,\dots,T$, the competing bids are fixed at $v_{1}+\epsilon$, irrespective of the bidder’s strategy, the initial RoI violation cannot be made up for in subsequent rounds, and the cumulative RoI constraint is violated.
+This highlights the risk of early overbidding in an adversarial environment and motivates the practicality of 6.4.
+Under 6.4, we have the following strong impossibility result:
+Theorem 6.5.
+There exists a sequence of competing bids for which there exists no constant $c\in(0,1]$ for which the $c$-approximate regret is sublinear in $T$, i.e., $c$-$\textsf{REG}_{\text{cumul}}=\Omega(T)$ for every $c\in(0,1]$.
+Despite this impossibility result, we propose a flexible heuristic based on safe strategies and Algorithm 1 for practitioners seeking to satisfy cumulative RoI constraints in practice.
+To design our heuristic, we adopt a general behavioral model in which bidders aim to satisfy the RoI constraint over a sliding window of $T_{0}\geq 1$ rounds.
+Formally, for any sequence of competing bids $\bm{\beta}_{-}^{t}]_{t\in[T]}$, the bidder seeks to ensure that
+$\sum_{\tau=\max(1,t-T_{0}+1)}^{t}\left(V(\mathbf{b}^{\tau};\bm{\beta}_{-}^{\tau})-P(\mathbf{b}^{\tau};\bm{\beta}_{-}^{\tau})\right)\geq 0,\quad\forall t\in[T]\,.$ (23)
+An immediate consequence of the condition in Eq.
+23 is that it guarantees the RoI constraint is satisfied in aggregate over the interval $[1,t]$ for all $t\in[T]$.
+To see this, let $n=\left\lfloor t/T_{0}\right\rfloor$.
+Then, the interval $[1,t]$ can be partitioned into the following subintervals: $1,t-nT_{0}]$, $t-nT_{0}+1,t-(n-1)T_{0}]$, $\dots$, $t-T_{0}+1,t]$.
+As Eq.
+23 ensures that the constraint holds in each disjoint interval, summing the RoI constraints over these intervals yields the desired result.
+In this setting, the per-round RoI constraint is relaxed and evaluated over sliding windows of length $T_{0}$.
+This allows the bidder some flexibility to temporarily overbid in individual auctions, as long as the cumulative value collected over the preceding $T_{0}-1$ rounds sufficiently exceeds the total payment incurred.
+Intuitively, this creates a temporary “surplus” that can be used to justify more aggressive bidding in the current round.
+To capture this effect, we introduce the notion of $\delta$-shifted $m$-uniform undominated safe strategies for $\delta\geq 0$. Specifically, we define:
+$\mathscr{S}_{m}^{\star}(\delta)=\left\{\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}}+\delta,,\ell\in[m]\right\}.$
+We denote the union of these classes over all $k\in[m]$ as $\mathscr{U}_{m}^{\star}(\delta)=\bigcup_{k\in[m]}\mathscr{S}_{k}^{\star}(\delta)$.
+This class corresponds to the set of safe bidding strategies with at most $m$ bid-quantity pairs under a shifted valuation vector $\mathbf{v}(\delta)=\mathbf{v}+\delta\cdot\bm{1}$, where $\bm{1}$ is the all-ones vector.
+The tunable parameter $\delta$ captures the bidder’s willingness to overbid relative to their valuation, reflecting the additional flexibility afforded by the relaxed RoI constraint.
+The Heuristic.
+When the bidder aims to satisfy the RoI constraints according to Eq.
+23, we propose a heuristic in the online setting that builds on Algorithm 1.
+Specifically, in round $t$, the bidder constructs a DAG, $\mathcal{G}^{t}(V,E)$ similar to $\mathcal{G}(V,E)$, and follows the UPDATE and SAMPLE step from Algorithm 1 to select a path.
+Then, the selected path is mapped to a safe bidding strategy within the class $\mathscr{U}_{m}^{\star}(\delta^{t})$, where
+$\delta^{t}:=\frac{1}{M}\sum_{\tau=\max(1,t-T_{0}+1)}^{t-1}(V(\mathbf{b}^{\tau};\bm{\beta}_{-}^{\tau})-P(\mathbf{b}^{\tau};\bm{\beta}_{-}^{\tau})),\quad\text{and}\quad\delta^{1}=0\,.$ (24)
+Note that $\delta^{t}$ can be computed in both the full-information and bandit settings.
+If the sampled path $\mathfrak{p}^{t}=s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for some $k\in[m]$, the bidder submits $\mathbf{b}^{t}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
+$b_{\ell}=w_{z_{\ell}}+\delta^{t}\quad\text{and}\quad q_{\ell}=z_{\ell}-z_{\ell-1},~\forall\ell\in[k]\,$,
+where $w_{z_{\ell}}$ is defined as per Eq. 1 for all $\ell\in[k]$. Finally, the bidder sets the edge weights in the DAG. In the full-information setting, the edge weights are:
+(i) If $x=(\ell-1,j)$ and $y=(\ell,j^{\prime})$ with $\ell\in[m]$ and $j<j^{\prime}$,
+$\textsf{w}^{t}(e)=\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}+\delta^{t}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$
+(ii) If $x=(\ell,j)$ and $y=d$, then $\textsf{w}^{t}(e)=0,\forall\ell\in[m],j\in[M]$.
+In the bandit setting, $\textsf{w}^{t}(\cdot)$ is replaced by its unbiased estimator, $\widehat{\textsf{w}}^{t}(\cdot)$, as stated in Eq. 7.
+We claim that a bidder following the $\delta^{t}$-shifted safe strategies satisfies the RoI constraints described in Eq.
+23.
+To see this, fix any round $t$.
+The maximum RoI violation in round $t$ under a $\delta^{t}$-shifted safe strategy is at most $M\delta^{t}$.
+Thus, the maximum violation of the RoI constraints over the window of $T_{0}$ rounds in the interval $\max(1,t-T_{0}+1),t]$ is $\sum_{\tau=\max(1,t-T_{0}+1)}^{t-1}\big(P(\mathbf{b}^{\tau};\bm{\beta}_{-}^{\tau})-V(\mathbf{b}^{\tau};\bm{\beta}_{-}^{\tau})\big)+M\delta^{t},$ which equals zero by construction.
+Hence, the RoI constraints are satisfied as desired.
+Experiments.
+We conduct experiments using EU ETS emission permit data to evaluate the proposed heuristic.
+First, we quantify the improvement in the bidder’s cumulative value over $T$ rounds when RoI constraints are enforced over a sliding window of $T_{0}$ rounds, compared to the stricter case of per-round RoI enforcement.
+We measure this via the relative gain as a function of $m$.
+Next, we compute the degree of per-round RoI feasibility by computing the relative feasibility in each round $t\in[T]$, defined as $\frac{V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})}{P(\mathbf{b}^{t};\bm{\beta}_{-}^{t})}-1.$ When per-round RoI constraints are enforced, this quantity is guaranteed to be at least the target RoI $\gamma$ for all tt.
+[^21] In our experiments, we ensure that the smallest competing bid in each auction is positive, so that the relative feasibility is well-defined whenever the bidder wins at least one unit.
+Thus, $P(\mathbf{b}^{t};\bm{\beta}_{-}^{t})=0$ if and only if $V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})=0$.
+In such cases, we define the relative feasibility to be the target RoI $\gamma$.
+In our experiments, we set $T=200$ and $T_{0}\in\{1,8,16,50,\infty\}$.
+These values are chosen to reflect real-world scenarios: approximately 200 EU ETS auctions are conducted per year, and $T_{0}$ values of 8, 16, and 50 correspond to the bidder enforcing RoI constraints roughly biweekly, monthly, and quarterly, respectively.
+In the second experiment, we fix $m=4$, which reflects the average number of bid-quantity pairs submitted by bidders in EU ETS auctions (EEX, 2023).
+Each experiment is repeated over 50 simulations.
+Figure 5: The left figure shows the relative gain obtained as a function of $m$.
+Here, relative gain is the ratio of the value obtained by the bidder over $T$ rounds when RoI constraints are enforced over $T_{0}$ rounds to that when RoI constraints are enforced in each round, i.e., $T_{0}=1$.
+The box plot in the right figure shows the distribution of per-round relative feasibility $\frac{V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})}{P(\mathbf{b}^{t};\bm{\beta}_{-}^{t})}-1.$
+Table 2: Performance metrics of shifted safe strategies as a function of $T_{0}$. Here, $\gamma=0$ and $m=4$. Here, RF stands for relative feasibility.
+$T_{0}=1$ $T_{0}=8$ $T_{0}=16$ $T_{0}=50$ $T_{0}=\infty$
+Fraction of rounds with RoI violations $(\text{RF}\leq\gamma)$ $0.0$ $0.3225$ $0.353$ $0.3815$ $0.4235$
+Median RF $(\frac{V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})}{P(\mathbf{b}^{t};\bm{\beta}_{-}^{t})}-1)$ $0.0786$ $0.0231$ $0.0127$ $0.0042$ $0.0$
+Discussion.
+Our experiments reveal a few interesting trends.
+In Fig.
+5, we observe that while the relative gain increases with $T_{0}$ for a fixed $m$ as expected, for a fixed $T_{0}$, it decreases with $m$.
+The fraction of rounds with RoI violations steadily increases as $T_{0}$ increases (see Table 2).
+Hence, the buyers face an inherent trade- off between the increase in obtained value and per-round RoI feasibility.
+For example, fixing $m=4$, for $T_{0}=\infty$, the shifted safe strategies improve the obtained value by $\sim 21\%$ compared to the safe strategies, but the RoI constraint is violated in $\sim 42\%$ of the rounds.
+Similarly, for $T_{0}=8$, the same metrics are $\sim 10\%$ and $\sim 32\%$ respectively.
+Thus, based on their requirements and risk profile, bidders can employ shifted safe strategies with appropriate values of $T_{0}$.
+The median relative feasibility also steadily decreases with larger $T_{0}$, approaching zero as the strategy becomes more aggressive.
+This indicates that while occasional high-value rounds may drive up cumulative gains, most rounds are closer to the feasibility barrier or even violate it.
+6.3. Time Varying Valuations
+Recall that we had assumed the valuation vector remains fixed over the $T$ rounds.
+Here, we consider the case where the valuation vector in each round is chosen from a finite set $\mathcal{V}$ in each round $t\in[T]$ and revealed to the bidder before they play their bidding strategy.
+As earlier, the competing bids are chosen by an oblivious adversary and the RoI constraints are enforced for each round.
+To satisfy the RoI constraint in round $t$, the bidder submits a safe bidding strategy corresponding to the valuation vector $\mathbf{v}^{t}$, i.e., $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}(\mathbf{v}^{t})$, where
+$\mathscr{U}_{m}^{\star}(\mathbf{v}^{t})=\bigcup_{k\in[m]}\mathscr{S}_{m}^{\star}(\mathbf{v}^{t})\quad\text{and}\quad\mathscr{S}_{m}^{\star}(\mathbf{v}^{t})=\Big\{\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}}^{t},\forall\ell\in[m]\Big\}\,.$
+Here, $\mathbf{v}^{t}=[v^{t}_{1},\dots,v^{t}_{M}]$ and $w_{j}^{t}=\frac{1}{j}\sum_{\ell\leq j}v_{\ell}^{t}$ for all $j\in[M],t\in[T]$.
+For time varying valuations, we frame the learning task as an instance of contextual online learning problem with a finite number of contexts, where the contexts are the valuation vectors.
+These contexts can be either sampled in each round from an (unknown) distribution $F_{\mathbf{v}}$ or they can be chosen adversarially at the beginning of the first round.
+The case of adversarially generated contexts is discussed in Section E.6.
+In both cases, we assume full information feedback for ease of exposition.
+Suppose the contexts are sampled from an (unknown) distribution $F_{\mathbf{v}}$ with a finite support. The regret in this setting is defined as
+$\textsf{REG}_{sto}=\max_{\pi^{*}\in\Pi}\sum_{t=1}^{T}\mathbb{E}_{\mathbf{v}^{t}\sim F_{\mathbf{v}}}[V(\pi^{*}({\mathbf{v}^{t}});\bm{\beta}_{-}^{t})]-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\,$,
+where $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}(\mathbf{v}^{t}),\forall t\in[T]$. Here, $\pi^{*}$ is a stationary policy belonging to the class
+$\Pi=\left\{\pi:\mathcal{V}\to\bigcup_{\mathbf{v}\in\mathcal{V}}\mathscr{U}_{m}^{\star}(\mathbf{v}),~~\text{such that}~~\pi(\mathbf{v})\in\mathscr{U}_{m}^{\star}(\mathbf{v}),\forall\mathbf{v}\in\mathcal{V}\right\}\,.$ (25)
+The expectation in the second term in the regret is with respect to the randomness in the contexts as well as any internal randomness of the learning algorithm.
+Our benchmark in this setting is consistent with existing literature on adversarial rewards with stochastic contexts (Balseiro et al., 2019; Schneider and Zimmert, 2023; Galgana and Golrezaei, 2024).
+Suppose $\mathbb{P}[\mathbf{v}^{t}=\mathbf{v}]=p_{\mathbf{v}},\forall\mathbf{v}\in\mathcal{V},\forall t\in[T]$.
+As $\sum_{t=1}^{T}\mathbb{E}_{\mathbf{v}^{t}\sim F_{\mathbf{v}}}[V(\pi({\mathbf{v}^{t}});\bm{\beta}_{-}^{t})]=\sum_{\mathbf{v}\in\mathcal{V}}p_{\mathbf{v}}\cdot\sum_{t=1}^{T}V(\pi(\mathbf{v});\bm{\beta}_{-}^{t})$, the optimal stationary policy is
+$\pi^{*}(\mathbf{v})=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t}),\quad\forall\mathbf{v}\in\mathcal{V}.$
+In this setting, the bidder follows an algorithm that generalizes Algorithm 1 by maintaining $|\mathcal{V}|$ copies of the DAG, one for each context in $\mathcal{V}$.
+In each round $t\in[T]$, the context $\mathbf{v}^{t}\sim F_{\mathbf{v}}$ is observed, and the edge probabilities for every DAG copy are updated.
+The bidder then samples a bidding strategy $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}(\mathbf{v}^{t})$ and observes the competing bids $\bm{\beta}_{-}^{t}$ ex post.
+Finally, edge weights are assigned for all edges in all DAGs, which are used to compute edge probabilities for the next round (see Algorithm 4 in Section E.5 for details).
+Theorem 6.6.
+When the contexts are stochastic, there exists an algorithm that runs in $\text{poly}(m,M,|\mathcal{V}|)$ space and per-round time and achieves $\textsf{REG}_{sto}\leq O(M\sqrt{mT\log M})$ under full information feedback.
 7. Conclusion and Open Problems
-We studied bidding in repeated uniform price auctions for a value maximizing buyer with per-round RoI constraints and characterized safe bidding strategies that can be efficiently learnt and are robust against various strong benchmarks.
-This study suggests several interesting future research directions: (1) analyzing bidding strategies in discriminatory-price (“pay-as-bid”) auctions; (2) modeling bidders with time-varying valuations and cumulative RoI constraints; (3) incorporating both RoI and budget constraints for a more realistic model of bidder behavior; and (4) extending the analysis to the more complex setting of combinatorial auctions with non-identical items.
+In this work, we studied the bidding problem in repeated uniform-price auctions for a value-maximizing buyer subject to per-round return-on-investment (RoI) constraints.
+To address this, we introduced the notion of safe bidding strategies as those that satisfy RoI constraints regardless of competing bids and characterized the class of $m$-uniform safe strategies.
+These strategies depend solely on the buyer’s valuation curve, and the buyer can focus on a finite subset of this class without loss of generality.
+We proposed a polynomial-time online learning algorithm to efficiently learn optimal strategies within this class.
+We then analyzed the robustness of safe strategies and the proposed algorithm against stronger benchmarks, showing that our approach achieves $\alpha$-approximate sublinear regret, where $\alpha$ is the richness ratio.
+We complemented the theoretical results by simulations on semi-synthetic EU ETS auction data which demonstrate that empirical richness ratios are substantially better than the worst-case bounds.
+We also showed the flexibility and adaptability of safe strategies and our learning algorithm across several model extensions such as adaptive adversaries, sliding-window RoI constraints, and time-varying valuations highlighting their practical relevance.
+From a practitioner’s perspective, safe bidding strategies are intuitive, robust, and adaptable to a wide range of risk preferences.
+The proposed algorithm is simple to implement and backed by strong theoretical guarantees, making it well-suited for real-world deployment.
+This study opens up several exciting avenues for future investigation.
+A natural extension is to characterize safe strategies for value-maximizing buyers in other multi-unit auction formats, such as discriminatory-price (pay-as-bid) auctions.
+From a modeling standpoint, incorporating both RoI and budget constraints could lead to richer and more realistic representations of bidder behavior.
+While our analysis focuses on a single buyer, it would be valuable to explore market dynamics, equilibrium outcomes, and convergence properties when all participants adopt safe strategies—paralleling recent developments under different market conditions and bidders’ behavior (Conitzer et al., 2022; Li and Tang, 2024; Lucier et al., 2024; Fikioris and Tardos, 2023; Paes Leme et al., 2024).
+Finally, extending this framework to combinatorial auctions with heterogeneous items presents a host of new algorithmic and design challenges.
+Acknowledgements
+N.G. and S.S. are partially supported by the MIT Junior Faculty Research Assistance Grant and the Young Investigator Award from the Office of Naval Research (ONR), Award No. N00014-21-1-2776.
 Appendix A Omitted Proofs from Section 3
@@
 Observation 1.
-Overbidding is not a safe bidding strategy.
-To see this, let $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell}),\dots,(b_{m},q_{m})\rangle$ be an overbid such that $b_{\ell}>w_{Q_{\ell}}$.
-Now, consider an auction in which the competing bids are:
+Overbidding is not a safe bidding strategy. To see this, let
+$\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell}),\dots,(b_{m},q_{m})\rangle$
+be an overbid such that $b_{\ell}>w_{Q_{\ell}}$. Now, consider an auction in which the competing bids are:
 $\bm{\beta}_{-,}^{-(j)}=\begin{cases}\epsilon,&~\text{if }1\leq j\leq Q_{\ell}\\ 2b_{1},&~\text{if }Q_{\ell}<j\leq K\end{cases}\,$,
@@
 A.2. Proof of Theorem 3.2
-We show that for any $m\in\mathbb{N}$, most safe bidding strategies are very weakly dominated for which they can removed from $\mathscr{S}_{m}$, resulting in $\mathscr{S}_{m}^{\star}$.
+We show that for any $m\in\mathbb{N}$, most safe bidding strategies are very weakly dominated for which they can eliminated from $\mathscr{S}_{m}$, resulting in $\mathscr{S}_{m}^{\star}$.
 We begin by establishing a general result regarding the monotonocity of feasible bid vectors (not necessarily $m$-uniform strategies) for value maximizing bidders.
@@
 Appendix B Omitted Proofs From Section 4
-B.1.  Proof of Lemma 4.1
-By Theorem 3.2, a strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell})\rangle\in\mathscr{U}_{m}^{\star}$ for some $\ell\in[m]$ can be uniquely identified by $\{Q_{1},\dots,Q_{\ell}\}$ where $Q_{j}=\sum_{k\leq j}q_{k}$ because $b_{j}=w_{Q_{j}}$ for all $j\in[\ell]$.
-Let the bid history $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$.
-Then, for any round $t$,
-$V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{j=1}^{\ell}\sum_{k=Q_{j-1}+1}^{Q_{j}}v_{k}\cdot\mathbb{I}{\left[w_{Q_{j}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}$
-Hence, the problem (Offline) can be equivalently expressed as
-$\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})=\max_{\ell\in[m]}\max_{Q_{1},\dots,Q_{\ell}}\sum_{j=1}^{\ell}\sum_{t=1}^{T}\sum_{k=Q_{j-1}+1}^{Q_{j}}v_{k}\cdot\mathbb{I}{\left[w_{Q_{j}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$
-B.2.  Proof of Theorem 4.2
+B.1. Proof of Theorem 4.2
 $\mathcal{G}(V,E)$ is a DAG.
-For convenience, let $d=(m+1,0)$.
+For convenience, let $d=(m+1,\infty)$.
 In the constructed graph, any node $(\ell,j)$ where $\ell\in[m]$ has edges edges either to the next layer, i.e.
-nodes of the form $(\ell+1,j^{\prime})$ where $j^{\prime}>j$ or to the destination node $(m+1,0)$.
+nodes of the form $(\ell+1,j^{\prime})$ where $j^{\prime}>j$ or to the destination node $(m+1,\infty)$.
 Hence, the directed graph has a topological sorting of the nodes implying that $\mathcal{G}(V,E)$ is a DAG.
@@
 Weight of $s$-$d$ paths. By assumption, $s=(0,0)$ and $z_{0}=0$. The weight of $\mathfrak{p}=s\to(1,z_{1})\to\dots\to(m,z_{m})\to d$ is:
-$\mathsf{w}(\mathfrak{p})$ $=\sum_{e\in\mathfrak{p}}\mathsf{w}(e)=\sum_{\ell=1}^{m}\mathsf{w}((\ell-1,z_{\ell-1})\to(\ell,z_{\ell}))$
-=(4)${{=}}\sum_{\ell=1}^{m}\sum_{t=1}^{T}\sum_{k=z_{\ell-1}+1}^{z_{\ell}}v_{k}\cdot\mathbb{I}{\left[w_{z_{\ell}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}$
+$\textsf{w}(\mathfrak{p})$ $=\sum_{e\in\mathfrak{p}}\textsf{w}(e)=\sum_{\ell=1}^{m}\textsf{w}((\ell-1,z_{\ell-1})\to(\ell,z_{\ell}))$
+[…]
 $=\sum_{t=1}^{T}\sum_{\ell=1}^{m}\sum_{k=z_{\ell-1}+1}^{z_{\ell}}v_{k}\cdot\mathbb{I}{\left[w_{z_{\ell}}\geq\bm{\beta}_{-,t}^{-(k)}\right]},$
@@
 ‘negated’) DAG can be solved in space and time complexity of $O(|V|+|E|)=O(mM^{2})$ which is polynomial in the parameters of the problem.
-B.3.  Proof of Theorem 4.3
+B.2. Proof of Theorem 4.3
 We begin by observing that there is a bijective mapping between $s$-$d$ paths in the DAG $\mathcal{G}^{t}(V,E)$ and bidding strategies $\mathbf{b}\in\mathscr{U}_{m}^{\star}$ (see proof of Theorem 4.2).
-The core idea of the proof is to show that Algorithm 1 is an equivalent and efficient implementation of the Hedge algorithm (Freund and Schapire, 1997) where every $s$-$d$ path is treated as an individual expert.
-To this end, we use an approach proposed in Takimoto and Warmuth (2003) and later used by Brânzei et al.
-(2023); Potfer et al.
-(2024); Galgana and Golrezaei (2024).
-For the sake of completeness, we provide the details.
+The core idea of the proof is to show that Algorithm 1 is an equivalent and efficient implementation of the Hedge algorithm (Freund and Schapire, 1997) where every $s$-$d$ path is treated as an expert.
 In Algorithm 1, $\varphi^{t}(u\to\cdot)$ denotes the probability distribution over the outgoing neighbors of node $u$.
-By the recursive sampling of nodes (Line 4-7 of Algorithm 1), we get that the probability of selecting $s$-$d$ path $\mathfrak{p}$ is
+By the recursive sampling of nodes, we get that the probability of selecting $s$-$d$ path $\mathfrak{p}$ is
 $\mathbb{P}^{t}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{t}(e),$ (29)
 and for any edge $e=u\to v$ in $\mathcal{G}^{t}(V,E)$, the edge probabilities, $\varphi^{t}(e)$, are updated as
-$\varphi^{t}(e)=\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)},$ (11)
-where $\Gamma^{t-1}(d)=1$ and $\Gamma^{t-1}(\cdot)$ is computed recursively in bottom-to-top fashion for every node $u$ in $\mathcal{G}^{t}(V,E)$ as follows:
-$\Gamma^{t-1}(u)=\sum_{v:u\to v\in E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(u\to v)\cdot\exp(\eta\mathsf{w}^{t-1}(u\to v))\,.$ (12)
-Now, consider a naïve implementation of the Hedge algorithm with learning rate $\eta$ in which each $s$-$d$ path is treated as an individual expert.
-For $t=1$, $\mathbb{P}_{\textsc{hedge}}^{1}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{1}(e)$.
+$\varphi^{t}(e)=\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)},$ (30)
+where $\Gamma^{t-1}(d)=1$ and $\Gamma^{t-1}(\cdot)$ is computed recursively in bottom-to-top fashion for every node $u\in V$ as follows:
+$\Gamma^{t-1}(u)=\sum_{v:u\to v\in E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(u\to v)\cdot\exp(\eta\textsf{w}^{t-1}(u\to v))\,.$ (31)
+Now, consider a naïve version of the Hedge algorithm with learning rate $\eta$ in which each $s$-$d$ path is treated as an individual expert.
+For $t=1$, the algorithm samples a path uniformly at random from $\mathscr{P}$, the collection of all $s$-$d$ paths in the DAG.
 For $t\geq 2$, the probability of selecting path $\mathfrak{p}$ in round $t$ is
-$\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})=\frac{\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})\exp(\eta\sum_{e\in\mathfrak{p}}\mathsf{w}^{t-1}(e))}{\sum_{\mathfrak{p}^{\prime}}\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p}^{\prime})\exp(\eta\sum_{e\in\mathfrak{p}^{\prime}}\mathsf{w}^{t-1}(e))}\,.$
-Now, we show that for any path $\mathfrak{p}$ and $t\in[T]$, $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$. To this end, we first present the following result:
+$\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})=\frac{\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})\exp(\eta\sum_{e\in\mathfrak{p}}\textsf{w}^{t-1}(e))}{\sum_{\mathfrak{p}^{\prime}}\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p}^{\prime})\exp(\eta\sum_{e\in\mathfrak{p}^{\prime}}\textsf{w}^{t-1}(e))}\,.$
+We prove Algorithm 1 is equivalent to the Hedge algorithm by showing that $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p}),\forall\mathfrak{p}\in\mathscr{P},\forall t\in[T]$.
+To this end, we first present the following result:
 Claim B.1.
 For any node $u$ in the graph, let $\mathscr{P}(u)$ be the set of paths starting at $u$ and terminating at $d$. Then,
-$\Gamma^{t-1}(u)=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))$
+$\Gamma^{t-1}(u)=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))$
 Proof.
-We prove the result by induction in a bottom-to-top order.
+We prove the result by strong induction in a bottom-to-top order.
 For the base case, $u=d$, $\Gamma^{t-1}(d)=1$.
-Suppose the result is true for all the nodes in layer $\ell+1$ for some $0\leq\ell\leq m$.
+Suppose the result is true for all the nodes in layer $\ell^{\prime}$ for some $\ell<\ell^{\prime}$.
 By the recursion in Eq.
-12, for any node $u$ in layer $\ell$,
-$\Gamma^{t-1}(u)$ $=\sum_{v:u\to v\in E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(u\to v)\cdot\exp(\eta\mathsf{w}^{t-1}(u\to v))$
-$=\sum_{v:u\to v\in E}\Big{(}\sum_{\mathfrak{p}\in\mathscr{P}(v)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))\Big{)}\varphi^{t-1}(u\to v)\cdot\exp(\eta\mathsf{w}^{t-1}(u\to v))$
-$=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e)),$
+31, for any node $u$ in layer $\ell$,
+$\Gamma^{t-1}(u)$ $=\sum_{v:u\to v\in E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(u\to v)\cdot\exp(\eta\textsf{w}^{t-1}(u\to v))$
+$=\sum_{v:u\to v\in E}\Big(\sum_{\mathfrak{p}\in\mathscr{P}(v)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))\Big)\varphi^{t-1}(u\to v)\cdot\exp(\eta\textsf{w}^{t-1}(u\to v))$
+$=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e)),$
 where the first equality follows from induction hypothesis. ∎
-We will show by induction on $t\in[T]$ that for any path $\mathfrak{p}$, $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$.
-For $t=1$ and any path $\mathfrak{p}$, $\mathbb{P}_{\textsc{hedge}}^{1}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{1}(e)=\mathbb{P}^{1}(\mathfrak{p})$.
-Here, the second equality holds by Eq.
-10.
-Suppose the result holds for round $t-1$, i.e., $\mathbb{P}^{t-1}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)$.
+Claim B.2.
+$\mathbb{P}_{\textsc{hedge}}^{1}(\mathfrak{p})=\frac{1}{|\mathscr{P}|}=\prod_{e\in\mathfrak{p}}\varphi^{1}(e)=\mathbb{P}^{1}(\mathfrak{p})$.
+Proof.
+The first equality holds from the definition of the naïve Hedge algorithm above and the last equality follows from Eq.
+29.
+To complete the proof, we need to show that $\prod_{e\in\mathfrak{p}}\varphi^{1}(e)=\frac{1}{|\mathscr{P}|}$.
+As $\varphi^{0}(e)=1$ and $\textsf{w}^{0}(e)=0$ for all $e\in E$ by initialization in Algorithm 1, for any edge $e=u\to v$, we get that
+[…]
+for any $s$-$d$ path $\mathfrak{p}\in\mathscr{P}$.
+For any node $u\in V$, by B.1, $\Gamma^{0}(u)=\sum_{\mathfrak{p}\in\mathscr{P}(u)}\prod_{e\in\mathfrak{p}}1=|\mathscr{P}(u)|$, the number of paths that start at $u$ and terminate at $d$.
+By initialization, $\Gamma^{0}(d)=1$ and $\Gamma^{0}(s)=|\mathscr{P}(s)|=|\mathscr{P}|$, which proves the result.
+∎
+We will show by induction on $t\in[T]$ that for any path $\mathfrak{p}\in\mathscr{P}$, $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$.
+For $t=1$ and any path $\mathfrak{p}$, the result is true due to B.2.
+Suppose the result holds for round t1t-1, i.e., $\mathbb{P}^{t-1}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)$.
 For round $t$,
+$\mathbb{P}^{t}(\mathfrak{p})$ $=\prod_{e\in\mathfrak{p}}\varphi^{t}(e)$
 […]
+$=\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p})\cdot\exp\Big(\eta\sum_{e\in\mathfrak{p}}\textsf{w}^{t-1}(e)\Big)\cdot\frac{\Gamma^{t-1}(d)}{\Gamma^{t-1}(s)},$
 where the last equality follows by telescoping product and invoking the induction hypothesis. By definition, $\Gamma^{t-1}(d)=1$. By B.1,
-$\Gamma^{t-1}(s)=\sum_{\mathfrak{p}^{\prime}}\prod_{e\in\mathfrak{p}^{\prime}}\varphi^{t-1}(e)\cdot\exp(\eta\mathsf{w}^{t-1}(e))=\sum_{\mathfrak{p}^{\prime}}\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p}^{\prime})\cdot\exp\Big{(}\eta\sum_{e\in\mathfrak{p}^{\prime}}\mathsf{w}^{t-1}(e)\Big{)}$
-Substituting the values gives $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$ implying Algorithm 1 is a correct implementation of the Hedge algorithm.
-B.3.1.  Bandit Feedback Model
-In the bandit setting, the bidder learns about their own allocation only for which $\mathsf{w}^{t}(e)$ can not be exactly computed for all the edges.
-Hence, we use an unbiased estimator, $\widehat{\mathsf{w}}^{t}(e)$, of $\mathsf{w}^{t}(e)$ as follows:
-$\widehat{\mathsf{w}}^{t}(e)$ $=\overline{\mathsf{w}}(e)-\frac{\overline{\mathsf{w}}(e)-\mathsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}\right]}\quad\text{where}\quad p^{t}(e)=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\,.$
-Here, $p^{t}(e)$ is the probability of selecting edge $e$ in round $t$, $\mathbb{P}^{t}$ is the distribution over all $s$-$d$ paths in $\mathcal{G}^{t}(V,E)$ and $\overline{\mathsf{w}}(e)$ is an edge-dependent upper bound of the edge weight $\mathsf{w}^{t}(e)$.
-Formally,
-$\overline{\mathsf{w}}(e)=\begin{cases}j^{\prime}-j,&\text{if }e=(\ell,j)\to(\ell+1,j^{\prime})\\ 0,&\text{if }e=(\ell,j)\to d\,.\end{cases}$
-The algorithm followed by the bidder in the bandit setting is the same as Algorithm 1, except in Line 9, the bidder sets the edge weights as $\widehat{\mathsf{w}}^{t}(e)$ instead of $\mathsf{w}^{t}(e)$.
-Clearly, $\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)]=\mathsf{w}^{t}(e)$.
-The unbiased estimator $\widehat{\mathsf{w}}^{t}(e)$ is different from the natural importance-weighted estimator $\frac{\mathsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}\right]}$ for technical reasons (similar to Lattimore and Szepesvári (2020, Equation 11.6)) which we will clarify shortly.
-B.3.2.  Regret Upper Bound
-Full Information Setting.
-Having shown that Algorithm 1 emulates the Hedge algorithm, we now recall the its regret bound from Brânzei et al.
-(2023, Corollary 1) which states that for learning rate $\eta$:
-$\mathsf{REG}=-\frac{1}{\eta}\max_{\mathfrak{p}}\log(\mathbb{P}^{1}(\mathfrak{p}))+\frac{\eta TM^{2}}{8}\,.$
-Observe that $\mathbb{P}^{1}(\mathfrak{p})\geq M^{-m}$, which implies,
-$\mathsf{REG}\leq\frac{m\log M}{\eta}+\frac{\eta TM^{2}}{8}\,.$
-Setting $\eta=\frac{1}{M}\sqrt{\frac{8m\log M}{T}}$, we get $\mathsf{REG}\leq O(M\sqrt{mT\log M})$.
-Bandit Setting. Using the standard analysis for EXP3 algorithm (see Lattimore and Szepesvári (2020, Chapter 11)), we get
-$\mathsf{REG}=-\frac{1}{\eta}\max_{\mathfrak{p}}\log(\mathbb{P}^{1}(\mathfrak{p}))+\sum_{t=1}^{T}\mathbb{E}\left[\frac{1}{\eta}\log\left(\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})e^{\eta\widehat{\mathsf{w}}^{t}(\mathfrak{p})}\right)-\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\widehat{\mathsf{w}}^{t}(\mathfrak{p})\right]$
-where $\widehat{\mathsf{w}}^{t}(\mathfrak{p})=\sum_{e\in\mathfrak{p}}\widehat{\mathsf{w}}^{t}(e)$.
-Note that $\widehat{\mathsf{w}}^{t}(\mathfrak{p})=\sum_{e\in\mathfrak{p}}\widehat{\mathsf{w}}^{t}(e)\leq\sum_{e\in\mathfrak{p}}\overline{\mathsf{w}}(e)\leq M$ and $\mathbb{P}^{1}(\mathfrak{p})\geq M^{-m}$.
-For $\eta\leq\frac{1}{M}$,
-$\mathsf{REG}$ $\leq\frac{m\log M}{\eta}+\eta\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\mathbb{E}[\widehat{\mathsf{w}}^{t}(\mathfrak{p})^{2}]$
-$\leq\frac{m\log M}{\eta}+\eta m\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]$
-where in the first inequality, we used $e^{x}\leq 1+x+x^{2}$ for $x\leq 1$ and $\log(1+x)\leq x$ for $x\geq 0$ and the second inequality follows from Cauchy-Schwarz. Observe that
-$\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]=\sum_{e}\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})=\sum_{e}p^{t}(e)\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}],$
+$\Gamma^{t-1}(s)=\sum_{\mathfrak{p}^{\prime}}\prod_{e\in\mathfrak{p}^{\prime}}\varphi^{t-1}(e)\cdot\exp(\eta\textsf{w}^{t-1}(e))=\sum_{\mathfrak{p}^{\prime}}\mathbb{P}_{\textsc{hedge}}^{t-1}(\mathfrak{p}^{\prime})\cdot\exp\Big(\eta\sum_{e\in\mathfrak{p}^{\prime}}\textsf{w}^{t-1}(e)\Big)$
+Substituting the values gives $\mathbb{P}^{t}(\mathfrak{p})=\mathbb{P}_{\textsc{hedge}}^{t}(\mathfrak{p})$. So, Algorithm 1 is a correct implementation of the Hedge algorithm.
+Regret Upper Bound and Time Complexity.
+Having shown that Algorithm 1 is equivalent to the Hedge algorithm, we now recall its regret bound from Cesa-Bianchi and Lugosi (2006, Theorem 2.2) which states that for learning rate $\eta$:
+$\textsf{REG}\leq\frac{\log|\mathscr{P}|}{\eta}+\frac{\eta TM^{2}}{8}\,.$
+As $|\mathscr{P}|\leq M^{m}$,
+$\textsf{REG}\leq\frac{m\log M}{\eta}+\frac{\eta TM^{2}}{8}\,.$
+Setting $\eta=\frac{1}{M}\sqrt{\frac{8m\log M}{T}}$, we get $\textsf{REG}\leq O(M\sqrt{mT\log M})$.
+In the full information setting, the running time bottleneck is computing Eq.
+31.
+Thus, Algorithm 1 runs in $|E|=O(mM^{2})$ time per round.
+B.3. Proof of Theorem 4.4
+Regret Analysis.
+In the bandit setting, the bidder follows Algorithm 1 by substituting $\textsf{w}^{t}(e)$ with its unbiased estimator $\widehat{\textsf{w}}^{t}(e)$.
+Thus, in this setting, the updates is similar to that of the EXP3 algorithm (Auer et al., 2002).
+Using the standard analysis for EXP3 algorithm (see Lattimore and Szepesvári (2020, Chapter 11)), we obtain
+$\textsf{REG}=\frac{\log|\mathscr{P}|}{\eta}+\sum_{t=1}^{T}\mathbb{E}\left[\frac{1}{\eta}\log\left(\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})e^{\eta\widehat{\textsf{w}}^{t}(\mathfrak{p})}\right)-\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\widehat{\textsf{w}}^{t}(\mathfrak{p})\right],$
+where $\widehat{\textsf{w}}^{t}(\mathfrak{p})=\sum_{e\in\mathfrak{p}}\widehat{\textsf{w}}^{t}(e)$.
+Note that $\widehat{\textsf{w}}^{t}(\mathfrak{p})=\sum_{e\in\mathfrak{p}}\widehat{\textsf{w}}^{t}(e)\leq\sum_{e\in\mathfrak{p}}\overline{\textsf{w}}(e)\leq M$ and $|\mathscr{P}|\leq M^{m}$.
+Note that if the natural importance-weighted estimator had been used instead, we would not be able to guarantee that $\widehat{\textsf{w}}^{t}(\mathfrak{p})\leq M$.
+For $\eta\leq\frac{1}{M}$, $\eta\widehat{\textsf{w}}^{t}(\mathfrak{p})\leq 1$.
+Hence,
+REG $\leq\frac{m\log M}{\eta}+\eta\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\mathbb{E}[\widehat{\textsf{w}}^{t}(\mathfrak{p})^{2}]$
+$\leq\frac{m\log M}{\eta}+\eta m\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\textsf{w}}^{t}(e)^{2}],$
+where in the first inequality, we used $e^{x}\leq 1+x+x^{2}$ for $x\leq 1$ and $\log(1+x)\leq x$ for $x\geq 0$. The second inequality follows from Cauchy-Schwarz. Observe that
+$\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\textsf{w}}^{t}(e)^{2}]=\sum_{e\in E}\mathbb{E}[\widehat{\textsf{w}}^{t}(e)^{2}]\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})=\sum_{e\in E}p^{t}(e)\mathbb{E}[\widehat{\textsf{w}}^{t}(e)^{2}],$
 where the last equality holds as $p^{t}(e)=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})$.
-$\sum_{e}p^{t}(e)\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]$ $=\sum_{e}p^{t}(e)\Big{[}\overline{\mathsf{w}}(e)^{2}(1-p^{t}(e))+\Big{(}\overline{\mathsf{w}}(e)-\frac{\overline{\mathsf{w}}(e)-\mathsf{w}^{t}(e)}{p^{t}(e)}\Big{)}^{2}p^{t}(e)\Big{]}$
-$=\sum_{e}\mathsf{w}^{t}(e)^{2}p^{t}(e)+(\overline{\mathsf{w}}(e)-\mathsf{w}^{t}(e))^{2}(1-p^{t}(e))\leq\sum_{e}\overline{\mathsf{w}}(e)^{2}\,.$
-We now aim to upper bound $\sum_{e}\overline{\mathsf{w}}(e)^{2}$.
-$\sum_{e}\overline{\mathsf{w}}(e)^{2}=\sum_{\ell=0}^{m-1}\sum_{j<j^{\prime}\leq M}\overline{\mathsf{w}}((\ell,j)\to(\ell+1,j^{\prime}))^{2}=m\sum_{j<j^{\prime}\leq M}(j-j^{\prime})^{2}\leq 4m\sum_{j=1}^{M}j^{2}\leq 4mM^{3}\,.$
-Hence,
+$\sum_{e\in E}p^{t}(e)\mathbb{E}[\widehat{\textsf{w}}^{t}(e)^{2}]$ $=\sum_{e\in E}p^{t}(e)\Big[\overline{\textsf{w}}(e)^{2}(1-p^{t}(e))+\Big(\overline{\textsf{w}}(e)-\frac{\overline{\textsf{w}}(e)-\textsf{w}^{t}(e)}{p^{t}(e)}\Big)^{2}p^{t}(e)\Big]$
+$=\sum_{e\in E}\textsf{w}^{t}(e)^{2}p^{t}(e)+(\overline{\textsf{w}}(e)-\textsf{w}^{t}(e))^{2}(1-p^{t}(e))\leq\sum_{e\in E}\overline{\textsf{w}}(e)^{2}\,.$
+Observe that $\sum_{e}\overline{\mathsf{w}}(e)^{2}\leq|E|M^{2}\lesssim mM^{4}$. Hence,
 $\mathsf{REG}$ $\leq\frac{m\log M}{\eta}+\eta m\sum_{t=1}^{T}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\mathbb{E}[\widehat{\mathsf{w}}^{t}(e)^{2}]$
-$\lesssim\frac{m\log M}{\eta}+\eta Tm^{2}M^{3}$
-Setting $\eta=\frac{1}{M}\sqrt{\frac{\log M}{mMT}}$, we get $\mathsf{REG}\leq O(M\sqrt{m^{3}MT\log M})\,.$  [^10] Recall that the analysis assumes that $\eta\leq\nicefrac{{1}}{{M}}$.
-As $m\geq 1,T\geq 1$ and $\log M\leq M$, the chosen $\eta=\frac{1}{M}\sqrt{\frac{\log M}{mMT}}\leq\frac{1}{M}.$, [^11] The upper bound on $\mathsf{REG}$ has a linear dependence on $M$ which is unavoidable as the reward in each round is in $[0,M]$.
-If normalized regret (Zhang et al., 2024) is considered instead, i.e., the rewards are scaled such that in each round it lies in $[0,1]$, the dependence on $M$ is logarithmic (or sublinear in the bandit setting).
-B.3.3.  Time Complexity
-In the full information setting, the running time bottleneck is Eq.
-12.
-Thus, Algorithm 1 runs in $|E|=O(mM^{2})$ time per round.
+$\lesssim\frac{m\log M}{\eta}+\eta Tm^{2}M^{4}\,.$
+Setting $\eta=\frac{1}{M^{2}}\sqrt{\frac{\log M}{mT}}$, we get $\mathsf{REG}\leq O(M^{2}\sqrt{m^{3}T\log M})$.
+Time Complexity.
 The bottleneck of running Algorithm 1 under the bandit feedback is the efficient computation of the marginal distribution, $p^{t}(e)$.
-Note that $p^{t}(e)$ has a sum-product structure as follows:
-$p^{t}(e)=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\prod_{e^{\prime}\in\mathfrak{p}}\varphi^{t}(e^{\prime})\,.$
-Hence, by the message passing algorithm from Wainwright and Jordan (2008, Section 2.5.1), we can compute the marginal probabilities (and hence run Algorithm 1) in $O(M|E|)=O(mM^{3})$ time per round.
+We claim that this can be done in $O(|E|)=O(mM^{2})$ time per-round.
+We begin by defining a quantity analogous to $\Gamma^{t-1}(\cdot)$.
+Define $\widetilde{\Gamma}^{t-1}(s)=1$ and $\widetilde{\Gamma}^{t-1}(\cdot)$ is computed recursively for every node $u\in V$ as follows:
+$\widetilde{\Gamma}^{t-1}(u)=\sum_{v:v\to u\in E}\widetilde{\Gamma}^{t-1}(v)\cdot\varphi^{t-1}(v\to u)\cdot\exp(\eta\widehat{\textsf{w}}^{t-1}(v\to u))\,.$ (32)
+Claim B.3.
+For any node $u$ in the graph, let $\widetilde{\mathscr{P}}(u)$ be the set of paths starting at $s$ and terminating at $u$. Then,
+$\widetilde{\Gamma}^{t-1}(u)=\sum_{\mathfrak{p}\in\widetilde{\mathscr{P}}(u)}\prod_{e\in\mathfrak{p}}\varphi^{t-1}(e)\cdot\exp(\eta\widehat{\textsf{w}}^{t-1}(e))$
+Proof.
+The proof follows the same structure as B.1, but instead of proceeding bottom-up through the nodes, we use top-down induction. ∎
+Let $e=u\to v$. Recall that
+[…]
+$=\varphi^{t}(e)\left(\sum_{\mathfrak{p}\in\widetilde{\mathscr{P}}(u)}\prod_{e^{\prime}\in\mathfrak{p}}\varphi^{t}(e^{\prime})\right)\left(\sum_{\mathfrak{p}\in\mathscr{P}(v)}\prod_{e^{\prime\prime}\in\mathfrak{p}}\varphi^{t}(e^{\prime\prime})\right)$
+[…]
+$\quad\times\left(\sum_{\mathfrak{p}\in\mathscr{P}(v)}\prod_{e^{\prime\prime}\in\mathfrak{p}}\varphi^{t-1}(e^{\prime\prime})\exp(\eta\widehat{\textsf{w}}^{t-1}(e^{\prime\prime}))\right)\,.$
+By B.1, B.3, Eq. 30 and the fact that $\Gamma^{t-1}(d)=1$, we get
+$p^{t}(e)$ $=\varphi^{t-1}(e)\cdot\exp(\eta\widehat{\textsf{w}}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)}\cdot\frac{\Gamma^{t-1}(u)}{\Gamma^{t-1}(s)}\cdot\frac{1}{\Gamma^{t-1}(v)}\cdot\widetilde{\Gamma}^{t-1}(u)\cdot\Gamma^{t-1}(v)$
+$=\varphi^{t-1}(e)\cdot\exp(\eta\widehat{\textsf{w}}^{t-1}(e))\cdot\frac{\widetilde{\Gamma}^{t-1}(u)\Gamma^{t-1}(v)}{\Gamma^{t-1}(s)}$
+Observe that $\Gamma^{t-1}(\cdot)$ (resp.
+$\widetilde{\Gamma}^{t-1}(\cdot)$) can be computed for all nodes recursively using Eq.
+31 (resp.
+Eq.
+32) in $O(|E|)$ time.
+Hence, $p^{t}(e)$ can be computed for all edges $e\in E$ in $O(|E|)=O(mM^{2})$ time in any round.
 B.4. Proof of Theorem 4.5
-For $m=1$, let $K=M$ be an even integer.
-Let $\mathbf{v}=[1,\dots,1,v,\dots,v]$.
-Here, the first $\frac{M}{2}$ entries are 1 and the remaining $\frac{M}{2}$ entries are $v$.
-Let $v=1-\delta$.
-Define $\delta^{\prime}=\frac{\delta}{2M}$.
+For $m=1$, let $K=M$ be an even integer. Let
+$\mathbf{v}=[1,\dots,1,v,\dots,v]\,.$ (33)
+Here, the first $\frac{M}{2}$ entries are 1 and the remaining $\frac{M}{2}$ entries are $v$. Let $v=1-\delta$. Define $\delta^{\prime}=\frac{\delta}{2M}$,
+$\bm{\beta}_{-}^{\clubsuit}=[1-\delta^{\prime},\dots,1-\delta^{\prime}]\quad\text{and}\quad\bm{\beta}_{-}^{\diamondsuit}=\Big[\frac{1+v}{2}-\delta^{\prime},\dots,\frac{1+v}{2}-\delta^{\prime}\Big]$ (34)
 Consider two scenarios:
 Scenario 1. In this scenario, for every $t\in[T]$, the competing bids $\bm{\beta}_{-}^{t}$ are:
-$\bm{\beta}_{-}^{t}=\begin{cases}[1-\delta^{\prime},\dots,1-\delta^{\prime}],&\text{ w.p.
-}\frac{1}{2}+\delta,\\ [\frac{1+v}{2}-\delta^{\prime},\dots,\frac{1+v}{2}-\delta^{\prime}],&\text{ w.p.
-}\frac{1}{2}-\delta,\\ \end{cases}$
+[…]
 Scenario 2. In this scenario, for every $t\in[T]$, the competing bids $\bm{\beta}_{-}^{t}$ are:
-$\bm{\beta}_{-}^{t}=\begin{cases}[1-\delta^{\prime},\dots,1-\delta^{\prime}],&\text{ w.p.
-}\frac{1}{2}-\delta,\\ [\frac{1+v}{2}-\delta^{\prime},\dots,\frac{1+v}{2}-\delta^{\prime}],&\text{ w.p.
-}\frac{1}{2}+\delta,\\ \end{cases}$
-for some $\delta\in(0,1/4)$ that would be determined shortly. Assume the randomness used in different rounds are independent.
-Let $P$ and $Q$ be the distribution of $[\bm{\beta}_{-}^{t}]_{t\in[T]}$ for scenario 1 and 2 respectively. Then, for $\delta\in(0,\frac{1}{4})$,
+[…]
+for some $\delta\in(0,1/4)$ to be determined shortly. Assume the randomness used in different rounds are independent.
+Let $P$ and $Q$ be the distribution of $\bm{\beta}_{-}^{t}]_{t\in[T]}$ for scenario 1 and 2 respectively. Then, for $\delta\in(0,\frac{1}{4})$,
 $\textsc{KL}(P||Q)=T\cdot\textsc{KL}(\textsc{Bern}(0.5+\delta)||\textsc{Bern}(0.5-\delta))=2T\delta\log\Big(\frac{1+2\delta}{1-2\delta}\Big)\leq\frac{8T\delta^{2}}{1-2\delta}\leq 16T\delta^{2}$
@@
 $w_{\frac{M}{2}+1}=\frac{\frac{M}{2}+1-\delta}{\frac{M}{2}+1}=1-\frac{\delta}{\frac{M}{2}+1}\leq 1-\frac{\delta}{M}<1-\delta^{\prime}\,.$
-Hence, no strategy of form $(w_{q},q)$ for $q\geq\frac{M}{2}+1$ is allocated any unit in when the competing bids are $[1-\delta^{\prime},\dots,1-\delta^{\prime}]$.
+Hence, no strategy of form $(w_{q},q)$ for $q\geq\frac{M}{2}+1$ is allocated any unit in when the competing bid profile is $\bm{\beta}_{-}^{\clubsuit}$.
 The smallest bid value for $\frac{M}{2}+1\leq q\leq M$ is for $q=M$.
 Note that, $w_{M}=\frac{1+v}{2}>\frac{1+v}{2}-\delta^{\prime}$.
-Hence, $(w_{M},M)=(\frac{1+v}{2},M)=(1-\frac{\delta}{2},M)$ is allocated all the units when the competing bids are $[\frac{1+v}{2}-\delta^{\prime},\dots,\frac{1+v}{2}-\delta^{\prime}]$ and, by definition, weakly dominates all the strategies of the form $(w_{q},q)$ for $q\geq\frac{M}{2}+1$.
+Hence, $(w_{M},M)=(\frac{1+v}{2},M)=(1-\frac{\delta}{2},M)$ is allocated all the units when the competing bid profile is $\bm{\beta}_{-}^{\diamondsuit}$ and, by definition, weakly dominates all the strategies of the form $(w_{q},q)$ for $q\geq\frac{M}{2}+1$.
 Hence, we have two undominated (for the constructed competing bids) strategies: $\mathbf{b}_{1}=(1,\frac{M}{2})$ and $\mathbf{b}_{2}=(1-\frac{\delta}{2},M)$.
-Now, we compute the expected value obtained by $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ for scenarios $P$ and $Q$.
+Now, we compute the expected value obtained by $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ for scenarios 1 and 2.
 $\mathbb{E}_{P}[V(\mathbf{b}_{1};\bm{\beta}_{-}^{t})]$ $=\mathbb{E}_{Q}[V(\mathbf{b}_{1};\bm{\beta}_{-}^{t})]=\frac{M}{2}$
@@
 $\mathbb{E}_{Q}[V(\mathbf{b}_{2};\bm{\beta}_{-}^{t})]$ $=M\left(\frac{1}{2}+\delta\right)\left(\frac{1+v}{2}\right)=\frac{M}{2}\left(\frac{1}{2}+\delta\right)(2-\delta)>\frac{M}{2}$
-So, $\mathbf{b}_{1}$ (resp. $\mathbf{b}_{2}$) is optimal for scenario $P$ (resp. scenario $Q$). Now, consider the distribution $\frac{P+Q}{2}$, i.e., $\textsc{Bern}(0.5)$.
-$\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}_{1};\bm{\beta}_{-}^{t})]=\frac{M}{2}~{}~{}\text{ and }~{}~{}$ $\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}_{2};\bm{\beta}_{-}^{t})]=\frac{M}{2}\left(\frac{1+v}{2}\right)\leq\frac{M}{2}$
+So, $\mathbf{b}_{1}$ (resp. $\mathbf{b}_{2}$) is optimal for scenario 1 (resp. scenario 2). Now, consider the distribution $\frac{P+Q}{2}$, i.e., $\textsc{Bern}(0.5)$.
+[…]
 $\implies\max_{\mathbf{b}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b};\bm{\beta}_{-}^{t})]$ $\leq\frac{M}{2}$ (35)
@@
 $\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{P}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})-V(\mathbf{b};\bm{\beta}_{-}^{t})]+\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{Q}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})-V(\mathbf{b};\bm{\beta}_{-}^{t})]$
-$\geq\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{P}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]+\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{Q}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]-2\max_{b^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]$
+$\geq\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{P}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]+\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{Q}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]-2\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]$
 $=\mathbb{E}_{P}[V(\mathbf{b}_{1};\bm{\beta}_{-}^{t})]+\mathbb{E}_{Q}[V(\mathbf{b}_{2};\bm{\beta}_{-}^{t})]-2\max_{\mathbf{b}^{*}\in\mathscr{U}_{1}^{\star}}\mathbb{E}_{\frac{P+Q}{2}}[V(\mathbf{b}^{*};\bm{\beta}_{-}^{t})]$
-(13)${{\geq}}\frac{M}{2}+\frac{M}{2}\left(\frac{1}{2}+\delta\right)(2-\delta)-2\cdot\frac{M}{2}=\frac{M}{2}\left(\frac{3\delta}{2}-\delta^{2}\right)\geq\frac{5M\delta}{8},$
+[…]
 where the last inequality follows as $\delta\in(0,\frac{1}{4})$.
@@
 By two-point method from Tsybakov (2009, Theorem 2.2),
-$\mathbb{E}_{\frac{P+Q}{2}}[\mathsf{REG}]\geq\frac{5MT\delta}{16}\cdot(1-\text{TV}(P,Q))\geq\frac{5MT\delta}{32}\exp{\left(-16T\delta^{2}\right)}$
-Setting $\delta=\frac{1}{4\sqrt{2T}}$, we get $\mathbb{E}_{\frac{P+Q}{2}}[\mathsf{REG}]\geq\frac{M\sqrt{T}}{36\sqrt{e}}$. Hence, $\mathbb{E}_{\frac{P+Q}{2}}[\mathsf{REG}]=\Omega(M\sqrt{T})$.
+$\mathbb{E}_{\frac{P+Q}{2}}[\textsf{REG}]\geq\frac{5MT\delta}{16}\cdot(1-\text{TV}(P,Q))\geq\frac{5MT\delta}{32}\exp{\left(-16T\delta^{2}\right)}$
+Setting $\delta=\frac{1}{4\sqrt{2T}}$, we get $\mathbb{E}_{\frac{P+Q}{2}}[\textsf{REG}]\geq\frac{M\sqrt{T}}{36\sqrt{e}}$. Hence, $\mathbb{E}_{\frac{P+Q}{2}}[\textsf{REG}]=\Omega(M\sqrt{T})$.
 Appendix C Omitted Proofs from Section 5
-C.1.  Proof of Theorem 5.3
-To prove that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}\leq 2$ for any $m\in\mathbb{N}$, recall from the proof sketch that we first defined the following metric for any bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$,
-$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})=\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$ (14)
-Let $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{m}^{*},q_{m}^{*})\rangle$ where $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right):=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ [^12] We assumed $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{m}^{*},q_{m}^{*})\rangle$ without loss of generality.
-The proof also follows if we considered $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b_{1}^{*},q_{1}^{*}),\dots,(b_{k}^{*},q_{k}^{*})\rangle$ for some $k\in[m]$ instead.Define $Q_{\ell}^{*}=\sum_{j\leq\ell}q_{j}^{*}$.
-Suppose $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated $r_{t}^{*}$ units in round $t$.
-For any $j\in[m]$, let $T_{j}$ be the set of rounds in which the least winning bid is $b_{j}^{*}$, i.e.,
-$T_{j}=\big{\{}t\in[T]:Q_{j-1}^{*}<r_{t}^{*}\leq Q_{j}^{*}\big{\}}\,.$ (15)
-For any $j\in[m]$, partition $T_{j}$ into $T_{j,0}$ and $T_{j,1}$ such that $T_{j,0}$ is the set of rounds where the bidder gets strictly less than $Q_{j}^{*}$ units and $T_{j,1}$ is the set of rounds when they get exactly $Q_{j}^{*}$ units:
-$T_{j,0}=\Big{\{}t\in[T_{j}]:r_{t}^{*}<Q_{j}^{*}\Big{\}},\qquad T_{j,1}=\Big{\{}t\in[T_{j}]:r_{t}^{*}=Q_{j}^{*}\Big{\}}\,.$ (16)
-For any $j\in[m]$, let
-$\widehat{Q}_{j}=\begin{cases}\max\{r_{t}^{*}:t\in T_{j,0}\},~{}&\text{ if }T_{j,0}\neq\emptyset\\ Q_{j}^{*}~{}&\text{ if }T_{j,0}=\emptyset\,.\end{cases}$
-In other words, $\widehat{Q}_{j}$ is the $2^{nd}$ highest number of units allocated to $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ over the rounds in $T_{j}$ (or the highest if $T_{j,0}=\emptyset$ or $T_{j,1}=\emptyset$).
-Step 1.
-Constructing a Restricted Class of Safe Bidding Strategies.
-Here, as a crucial part of the proof, we construct a restricted class of safe bidding strategies, denoted by $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, where $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})\subset\mathscr{U}_{m}^{\star}$.
-This construction serves two purposes.
-First, it reduces the search space for the optimal safe bidding strategy.
-Second, and more importantly, it enables us to establish a connection between the optimal safe bidding strategy and $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$.
-In defining the restricted class, we use the quantities, $\{\widehat{Q}_{j},Q_{j}^{*}\}_{j\in[m]}$ as follows:
-$\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})=\Big{\{}\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}},~{}~{}Q_{\ell}\in\{\widehat{Q}_{\ell},Q_{\ell}^{*}\},\quad\forall\ell\in[m]\Big{\}}\,.$ (17)
-Recall that any strategy in $\mathscr{U}_{m}^{\star}$ with $m$ bid-quantity pairs takes the form of $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}},\forall\ell\in[m]$.
-The strategies in $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$ also have the same structure but as an important difference, for any $\ell\in[m]$, we enforce $Q_{\ell}\in\{\widehat{Q}_{\ell},Q_{\ell}^{*}\}$.
-Observe that for any $\ell\in[m-1]$,
-$Q_{\ell}\leq Q_{\ell}^{*}<\widehat{Q}_{\ell+1}\leq Q_{\ell+1},$
-where the first and third inequalities follow directly from the definition of $Q_{\ell}$ and the second one follows from the definition of $\widehat{Q}_{\ell+1}$.
-So, $Q_{\ell}$’s are distinct and ordered.
-Further observe that the number of bidding strategies in $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$ is $O(2^{m})$; significantly smaller than the number of strategies in $\mathscr{U}_{m}^{\star}$, which is $O(M^{m})$.
-With the definition of the restricted class of safe bidding strategies, we have
-$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})=\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\leq\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$ (18)
-Step 2.
-Value Decomposition for $m$-uniform Strategies.
-Let $V_{j,k}$ be the time-average value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in the set of rounds in $T_{j,k}$ (as defined in Eq.
-16).
-Formally, $\forall j\in[m],k\in\{0,1\}$,
-$V_{j,k}=\dfrac{1}{|T_{j,k}|}\sum_{t\in T_{j,k}}V(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right),\bm{\beta}_{-}^{t})\,.$
-Define ${W}_{k}=\sum_{j=1}^{k}v_{j}$.
-Note that $V_{j,1}={W}_{Q_{j}^{*}}$ because for any $t\in T_{j,1}$, we have $r_{t}^{*}=Q_{j}^{*}$.
-For any $\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, where $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{\ell}=w_{Q_{\ell}}$, and $Q_{\ell}\in\{Q_{\ell}^{*},\widehat{Q}_{\ell}\}$ for any $\ell\in[m]$, define $N^{*},\widehat{N}\subseteq[m]$ as follows:
-$N^{*}=\{j:Q_{j}=Q_{j}^{*},j\in[m]\}\quad\text{ and }\quad\widehat{N}=\{j:Q_{j}=\widehat{Q}_{j},j\in[m]\}\,.$
-We now present a crucial result that expresses the value obtained by a $m$-uniform bidding strategy as a function of the value obtained by $m$ 1-uniform strategies.
-Lemma C.1.
-Let $\mathbf{b}=\langle(w_{Q_{1}},q_{1}),\dots,(w_{Q_{m}},q_{m})\rangle$ be a $m$-uniform safe bidding strategy for some $m\in\mathbb{N}$. Then, for any competing bids $\bm{\beta}_{-}$,
-$V(\mathbf{b};\bm{\beta}_{-})=\max_{\ell\in[m]}V((w_{Q_{\ell}},Q_{\ell});\bm{\beta}_{-}),$
-where we recall that $Q_{\ell}=\sum_{j=1}^{\ell}q_{j},\forall\ell\in[m]$.
-A key consequence of Lemma C.1 is that knowing the value obtained by the strategies in $\mathscr{U}_{1}^{\star}$ is sufficient to compute the value obtained any $m$-uniform safe bidding strategy.
-By Lemma C.1, for any round $t\in T_{j}$, and $\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, we get
-$V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\max_{\ell\in[m]}V((w_{Q_{\ell}},Q_{\ell});\bm{\beta}_{-}^{t})\geq V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\,,$ (19)
-$\implies\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $\geq\sum_{j\in N^{*}}\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})+\sum_{j\in\widehat{N}}\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\,.$ (20)
-Step 3. Allocation Lower Bounds for 1-uniform Strategies. For any $j\in[m]$, we now invoke the following lemma to establish lower bound on $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})$
-Lemma C.2.
-Let $\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$, where $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle:b_{j}=w_{Q_{j}}$, and $Q_{j}\in\{Q_{j}^{*},\widehat{Q}_{j}\},\forall j\in[m]$.
-Then, for any $j\in[m]$,
-  • (a)
-    if $Q_{j}=\widehat{Q}_{j}$ (i.e., $j\in\widehat{N}$), we have
-    $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\geq V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\,.$ (21)
-  • (b)
-    If $Q_{j}=Q_{j}^{*}$ (i.e., $j\in N^{*}$), we have
-    $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\geq{W}_{Q_{j}^{*}}|T_{j,1}|\,.$ (22)
-Note that the right hand side of Eq. 20 depends only on the choice of the partitions $N^{*}$ and $\widehat{N}$. Substituting the lower bound from Lemma C.2 in Eq. 20, we establish that,
-$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ (18)${{\leq}}\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}$
-$\leq\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\max_{(N^{*},\widehat{N})}\Big{\{}\sum_{j\in N^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}\Big{\}}}\,.$ (23)
-Let $(N_{0}^{*},\widehat{N}_{0})=\mathop{\mathrm{argmax}}_{(N^{*},\widehat{N})}\Big{\{}\sum_{j\in N^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}\Big{\}}$.
-Then,
-$\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{t=1}^{T}\sum_{\ell=0}^{1}V_{j,\ell}|T_{j,\ell}|$
-$=\sum_{j\in N_{0}^{*}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}\,.$ (24)
-Consider a partition of $[m]$ that differs from the maximizing partition $(N_{0}^{*},\widehat{N}_{0})$ by exactly one element, i.e., for any $a\in\widehat{N}_{0}$, consider the following partition: $(N_{0}^{*}\cup\{a\},\widehat{N}_{0}\setminus\{a\})$.
-By definition,
-$\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$ $\geq\sum_{j\in N_{0}^{*}\cup\{a\}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}\setminus\{a\}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$
-$\implies V_{a,0}|T_{a,0}|+{W}_{\widehat{Q}_{a}}|T_{a,1}|$ $\geq{W}_{Q_{a}^{*}}|T_{a,1}|\,.$ (25)
-Now, for any $b\in N^{*}_{0}$, consider the following partition: $(N_{0}^{*}\setminus\{b\},\widehat{N}_{0}\cup\{b\})$. By definition,
-$\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$ $\geq\sum_{j\in N_{0}^{*}\setminus\{b\}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}\cup\{b\}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}$
-$\implies{W}_{Q_{b}^{*}}|T_{b,1}|-{W}_{\widehat{Q}_{b}}|T_{b,1}|$ $\geq V_{b,0}|T_{b,0}|\,.$ (26)
-Plugging in the values,
-$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ (C.1)${{\leq}}\frac{\sum_{j\in N_{0}^{*}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
-(25)${{\leq}}\frac{\sum_{j\in N_{0}^{*}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}2V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
-(26)${{\leq}}\frac{\sum_{j\in N_{0}^{*}}\Big{(}2{W}_{Q_{j}^{*}}|T_{j,1}|-{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}2V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
-$=\frac{2\Big{\{}\sum_{j\in N_{0}^{*}}{W}_{Q_{j}^{*}}|T_{j,1}|+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}\Big{\}}-\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
-$=2-\theta_{\mathcal{H}_{-}}\,,$
-where
-$0<\theta_{\mathcal{H}_{-}}$ $=\frac{\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big{)}}$
-(25)${{\leq}}\frac{\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in N_{0}^{*}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}+\sum_{j\in\widehat{N}_{0}}\Big{(}{W}_{Q_{j}^{*}}|T_{j,1}|\Big{)}}$
-$=\frac{\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j=1}^{m}{W}_{Q_{j}^{*}}|T_{j,1}|}\stackrel{{\scriptstyle(a)}}{{\leq}}\max_{j\in[m]}\frac{{W}_{\widehat{Q}_{j}}}{{W}_{Q_{j}^{*}}}\leq 1\,,$
-and $(a)$ follows from 1. Finally,
-$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}=\max_{\mathcal{H}_{-}}\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})\leq 2-\min_{\mathcal{H}_{-}}\theta_{\mathcal{H}_{-}}=:2-\theta,\quad\text{where}\quad\theta=\min_{\mathcal{H}_{-}}\theta_{\mathcal{H}_{-}}\in(0,1]\,.$
-Fact 1 (Williamson and Shmoys (2011, pp. 25)).
-For positive numbers $a_{1},\dots,a_{k}$ and $b_{1},\dots,b_{k}$,
-$\frac{\sum_{j=1}^{k}a_{j}}{\sum_{j=1}^{k}b_{j}}\leq\max_{j\in[m]}\frac{a_{j}}{b_{j}}\,.$
-C.2.  Proof of Lemma C.1
-We state and prove a stronger version of the result in Lemma C.1. Formally,
+C.1. Proof of Lemma 5.6
+We state and prove a stronger version of the result in Lemma 5.6. Formally,
 Lemma C.1.
@@
 where we recall that $Q_{\ell}=\sum_{j=1}^{\ell}q_{j},\forall\ell\in[m]$.
-Lemma C.3 states a similar result as Lemma C.1, except a key difference that the bidding strategies are not necessarily safe.
-C.2.1.  Proof of Lemma C.3
+Lemma C.1 states a similar result as Lemma 5.6, except a key difference that the bidding strategies are not necessarily safe.
+C.1.1. Proof of Lemma C.1
 We prove the lemma via induction on $m$.
@@
 $V(\mathbf{b};\bm{\beta}_{-})=V(\mathbf{b}[1:m];\bm{\beta}_{-})\,.$ (36)
-Hence, by C.4 and Eq. 27,
+Hence, by C.2 and Eq. 36,
 $V(\mathbf{b};\bm{\beta}_{-})$ $=\max\Big\{V(\mathbf{b}[1:m];\bm{\beta}_{-}),V((b_{m+1},Q_{m+1});\bm{\beta}_{-})\Big\}\,.$ (37)
@@
 So, $(b_{m+1},Q_{m+1})$ is allocated at least $r$ units which implies $V((b_{m+1},Q_{m+1}),\bm{\beta}_{-})\geq V(\mathbf{b};\bm{\beta}_{-})$.
-By C.4, we also have $V(\mathbf{b};\bm{\beta}_{-})\geq V((b_{m+1},Q_{m+1}),\bm{\beta}_{-})$.
+By C.2, we also have $V(\mathbf{b};\bm{\beta}_{-})\geq V((b_{m+1},Q_{m+1}),\bm{\beta}_{-})$.
 Hence,
@@
 $V(\mathbf{b};\bm{\beta}_{-})$ $=\max\Big\{V(\mathbf{b}[1:m];\bm{\beta}_{-}),V((b_{m+1},Q_{m+1});\bm{\beta}_{-})\Big\}$
-=$\stackrel{{\scriptstyle(a)}}{{=}}\max\Big{\{}\max_{\ell\in[m]}V((b_{\ell},Q_{\ell});\bm{\beta}_{-}),V((b_{m+1},Q_{m+1});\bm{\beta}_{-})\Big{\}}$
+$\stackrel{{\scriptstyle(a)}}{{=}}\max\Big\{\max_{\ell\in[m]}V((b_{\ell},Q_{\ell});\bm{\beta}_{-}),V((b_{m+1},Q_{m+1});\bm{\beta}_{-})\Big\}$
 $=\max_{\ell\in[m+1]}V((b_{\ell},Q_{\ell});\bm{\beta}_{-})\,.$
 Here, $(a)$ holds as $\mathbf{b}[1:m]$ is feasible for $\bm{\beta}_{-}$ allowing us to apply the induction hypothesis for $m$.
-C.3.  Proof of Lemma C.2
+C.2. Proof of Lemma 5.7
 To prove this result, we use the following key lemma that measures the value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in terms of 1-uniform safe bidding strategies.
@@
 Then in any round $t^{\prime}\in T_{j}$ in which $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated at most $r_{t}^{*}$ units, i.e., $r_{t^{\prime}}^{*}\leq r_{t}^{*}$, $(w_{r_{t}^{*}},r_{t}^{*})$ is allocated at least $r_{t^{\prime}}^{*}$ units.
-With this lemma, we are ready to prove Lemma C.2.
-Case I: $Q_{j}=\widehat{Q}_{j}$.
-By assumption, $Q_{j}=\widehat{Q}_{j}\leq Q_{j}^{*}$.
-So, for any $t\in T_{j,1}$, invoking Lemma C.5 (1) with $q=\widehat{Q}_{j}$, we conclude that $(w_{\widehat{Q}_{j}},\widehat{Q}_{j})$ is allocated exactly $\widehat{Q}_{j}$ units.
+With this lemma, we are ready to prove Lemma 5.7.
+Case I: $Q_{j}=\widehat{Q}_{j}<Q_{j}^{*}$.
+For any $t\in T_{j,1}$, invoking Lemma C.3 (1) with $q=\widehat{Q}_{j}$, we conclude that $(w_{\widehat{Q}_{j}},\widehat{Q}_{j})$ is allocated exactly $\widehat{Q}_{j}$ units.
 Summing over all rounds in $T_{j,1}$,
 $\sum_{t\in T_{j,1}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})={W}_{\widehat{Q}_{j}}|T_{j,1}|\,.$ (40)
-Now, suppose $T_{j,0}\neq\emptyset$ and $r_{t}^{*}=\widehat{Q}_{j}$ for some $t\in T_{j,0}$.
+As $\widehat{Q}_{j}<Q_{j}^{*}$, $T_{j,0}\neq\emptyset$.
+Let $r_{t}^{*}=\widehat{Q}_{j}$ for some $t\in T_{j,0}$.
 By definition, $\widehat{Q}_{j}<Q_{j}^{*}$ and $\widehat{Q}_{j}\geq r_{s}^{*}$ for any $s\in T_{j,0}$.
@@
 $\sum_{t\in T_{j,0}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})$ $\geq\sum_{t\in T_{j,0}}V\Big(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right),\bm{\beta}_{-}^{t}\Big)=V_{j,0}|T_{j,0}|\,.$ (41)
-If $T_{j,0}=\emptyset$, the lower bound is trivially true. Combining Eq. 31 and (32), for $Q_{j}=\widehat{Q}_{j}$,
+Combining Eq. 40 and (41), for $Q_{j}=\widehat{Q}_{j}$,
 $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\geq V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\,.$ (42)
@@
 $\sum_{t\in T_{j}}V((w_{Q_{j}},Q_{j});\bm{\beta}_{-}^{t})\geq{W}_{Q_{j}^{*}}|T_{j,1}|\,.$ (45)
-C.3.1.  Proof of Lemma C.5
+C.2.1. Proof of Lemma C.3
 (1) First we show that for any $t\in[T]$ and $q\leq r_{t}^{*}$, the 1-uniform bid $(w_{q},q)$ is allocated exactly $q$ units.
@@
 (2) For any $j\in[m]$, let $T_{j}\subseteq[T]$, defined in Eq.
-15, be the rounds in which $b_{j}^{*}$ is the least winning bid.
+11, be the rounds in which $b_{j}^{*}$ is the least winning bid.
 Suppose that $\exists t\in T_{j}$ such that $r_{t}^{*}<Q_{j}^{*}$.
@@
 Hence, $(w_{r_{t}^{*}},r_{t}^{*})$ is allocated at least $r_{t^{\prime}}^{*}$ units in round $t^{\prime}$.
+C.3. Proof of Lemma 5.8
+Let $(\mathcal{S}_{0}^{*},\widehat{\mathcal{S}}_{0})=\mathop{\mathrm{argmax}}_{(\mathcal{S}^{*},\widehat{\mathcal{S}})}\Big\{\sum_{j\in\mathcal{S}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)\Big\}$.
+Then,
+$\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{t=1}^{T}\sum_{\ell=0}^{1}V_{j,\ell}|T_{j,\ell}|$
+$=\sum_{j\in\mathcal{S}_{0}^{*}}\Big(V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big)\,.$ (46)
+Consider a partition of $[m]$ that differs from the maximizing partition $(\mathcal{S}_{0}^{*},\widehat{\mathcal{S}}_{0})$ by exactly one element, i.e., for any $a\in\widehat{\mathcal{S}}_{0}$, consider the following partition: $(\mathcal{S}_{0}^{*}\cup\{a\},\widehat{\mathcal{S}}_{0}\setminus\{a\})$.
+By definition,
+$\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)$ $\geq\sum_{j\in\mathcal{S}_{0}^{*}\cup\{a\}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}\setminus\{a\}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)$
+$\implies V_{a,0}|T_{a,0}|+{W}_{\widehat{Q}_{a}}|T_{a,1}|$ $\geq{W}_{Q_{a}^{*}}|T_{a,1}|\,.$ (47)
+Now, for any $b\in\mathcal{S}^{*}_{0}$, consider the following partition: $(\mathcal{S}_{0}^{*}\setminus\{b\},\widehat{\mathcal{S}}_{0}\cup\{b\})$. By definition,
+$\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)$ $\geq\sum_{j\in\mathcal{S}_{0}^{*}\setminus\{b\}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}\cup\{b\}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)$
+$\implies{W}_{Q_{b}^{*}}|T_{b,1}|-{W}_{\widehat{Q}_{b}}|T_{b,1}|$ $\geq V_{b,0}|T_{b,0}|\,.$ (48)
+Plugging in the values,
+$\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)}$
+$=\frac{\sum_{j\in\mathcal{S}_{0}^{*}}\Big(V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big)}{\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)}$
+[…]
+[…]
+$=\frac{2\Big\{\sum_{j\in\mathcal{S}_{0}^{*}}{W}_{Q_{j}^{*}}|T_{j,1}|+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)\Big\}-\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)}$
+$=2-\theta_{\mathcal{H}_{-}}\,$,
+where
+$0<\theta_{\mathcal{H}_{-}}$ $=\frac{\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)}$
+[…]
+and $(a)$ follows because for positive $a_{1},\dots,a_{m}$ and $b_{1},\dots,b_{m}$, $\frac{\sum_{j=1}^{m}a_{j}}{\sum_{j=1}^{m}b_{j}}\leq\max_{j\in[m]}\frac{a_{j}}{b_{j}}$.
 C.4. Proof of Theorem 5.4
@@
 For $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, we define
-$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-})=\frac{\max_{\mathbf{b}^{\prime}\in\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$
-Then, maximizing over all bid histories, we get the desired result.
+$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-},\mathbf{v})=\frac{\max_{\mathbf{b}^{\prime}\in\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$
+Then, maximizing over all bid histories and valuation vectors, we get the desired result.
 Let $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$.
-If $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is a $k$-uniform bidding strategy where $k\leq m\leq m^{\prime}$, by definition, $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)\in\mathscr{F}^{\mathcal{H}_{-}}_{m}$ implying $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-})=1\leq\frac{m^{\prime}}{m}$.
-Without loss of generality, let the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$ be $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b^{*}_{1},q^{*}_{1}),\dots,(b^{*}_{m^{\prime}},q^{*}_{m^{\prime}})\rangle$ and $Q_{\ell}^{*}=\sum_{j\leq\ell}q_{j}^{*}$ for all $\ell\in[m^{\prime}]$ [^13] A similar analysis also follows if $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is a $k$-uniform strategy where $m\leq k\leq m^{\prime}$.
+If $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is a $k$-uniform bidding strategy where $k\leq m\leq m^{\prime}$, by definition, $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)\in\mathscr{F}^{\mathcal{H}_{-}}_{m}$ implying $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-},\mathbf{v})=1\leq\frac{m^{\prime}}{m}$.
+Without loss of generality, let the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$ be
+$\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b^{*}_{1},q^{*}_{1}),\dots,(b^{*}_{m^{\prime}},q^{*}_{m^{\prime}})\rangle$
+and $Q_{\ell}^{*}=\sum_{j\leq\ell}q_{j}^{*}$ for all $\ell\in[m^{\prime}]$.
+[^23] A similar analysis also follows if $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is a $k$-uniform strategy where $m\leq k\leq m^{\prime}$.
 Let $r_{t}^{*}$ be the number of units allocated to $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in round $t$.
 Recall that $T_{j}$ is the set of rounds when $b_{j}^{*}$ is the least winning bid when $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is submitted (see Eq.
-15).
+11).
 So,
@@
 Algorithm 2 Constructing $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$
-0:  $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b^{*}_{1},q^{*}_{1}),\dots,(b^{*}_{m^{\prime}},q^{*}_{m^{\prime}})\rangle$.
-The collection, $\mathscr{C}$, of all subsets $\mathcal{S}\subseteq[m^{\prime}]$ such that $|S|=m$.
-Initialize $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}=\emptyset$.
-1:  for $\mathcal{S}\in\mathscr{C}$ do
-2:     Suppose $\mathcal{S}=\{k_{1},\dots,k_{m}\}$ such that ${k_{1}}<{k_{2}}<\dots<{k_{m}}$.
-3:     Construct a $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ where
-$b_{j}=b_{k_{j}}^{*}\quad\text{and}\quad q_{j}=Q_{k_{j}}^{*}-Q_{k_{j-1}}^{*},~{}$ $\forall j\in[m]\,.$ (39)
-4:     $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\leftarrow\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\cup\{\mathbf{b}\}$
-5:  end forReturn: The class of bidding strategies, $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$.
+[…]
+2:Initialize $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\leftarrow\emptyset$.
+3:for $\mathcal{S}\in\mathscr{C}$ do
+4:  Suppose $\mathcal{S}=\{k_{1},\dots,k_{m}\}$ such that ${k_{1}}<{k_{2}}<\dots<{k_{m}}$.
+5:  Construct a $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ where
+$b_{j}=b_{k_{j}}^{*}\qquad\text{and}\qquad q_{j}=Q_{k_{j}}^{*}-Q_{k_{j-1}}^{*},$~ $\forall j\in[m]\,.$ (51)
+6:  $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\leftarrow\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\cup\{\mathbf{b}\}$
+7:end for
+8:return The class of bidding strategies, $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$.
 For example: suppose $m^{\prime}=4$ and $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(w_{3},3),(w_{7},4),(w_{8},1),(w_{10},2)\rangle$.
@@
 where the first inequality follows because entries in bid vector are in non increasing order, the first equality follows as $b_{(Q_{r})}^{*}=b_{r}^{*}$ for all $r\in[m^{\prime}]$ and second equality follows from Eq.
-39.
+51.
 As $j$ and $\ell$ were picked arbitrarily, we get that for all $j\in[Q_{k_{m}}]$, $b_{(j)}\leq b_{(j)}^{*}$.
@@
 because for any $\ell\in[m]$, $b_{\ell}=b_{k_{\ell}}^{*}$ and $Q_{\ell}=\sum_{j=1}^{j}q_{j}=\sum_{j=1}^{j}Q_{k_{j}}^{*}-Q_{k_{j-1}}^{*}=Q_{k_{\ell}}^{*}\,.$
-Recall that $T_{j}$s, the set of rounds in which $b_{j}^{*}$ is the least winning bid, form a partition of $[T]$. So,
-$\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{t=1}^{m^{\prime}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})=\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})+\sum_{j\in[m^{\prime}]\setminus\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})\geq\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (42)
+Recall that $T_{j}$s, the set of rounds in which $b_{j}^{*}$ is the least winning bid, partition the set $[T]$. So,
+$\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{t=1}^{m^{\prime}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})$
+$=\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})+\sum_{j\in[m^{\prime}]\setminus\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (54)
+$\geq\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (55)
 For any $j\in\mathcal{S}$, consider any round $t\in T_{j}$:
-$V(\mathbf{b};\bm{\beta}_{-}^{t})$ =(41)${{=}}\max_{\ell\in\mathcal{S}}V((b_{\ell}^{*},Q_{\ell}^{*});\bm{\beta}_{-}^{t})\geq V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})\implies\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})\geq\sum_{t\in T_{j}}V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})$ ) (43)
+[…]
+$\implies\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $\geq\sum_{t\in T_{j}}V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})$ (56)
 which implies,
@@
 $\max_{\mathbf{b}\in\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\max_{\mathcal{S}\subseteq[m^{\prime}]:|\mathcal{S}|=m}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$
-(44)${{\geq}}\max_{\mathcal{S}\subseteq[m^{\prime}]:|\mathcal{S}|=m}\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})$ )
+[…]
 $\geq\frac{m}{m^{\prime}}\sum_{j=1}^{m^{\prime}}\sum_{t\in T_{j}}V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})\,.$ (58)
@@
 […]
-Maximizing over all bid histories, we get that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}\leq\frac{m^{\prime}}{m}$.
+Maximizing over all bid histories and valuation vectors, we get that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}\leq\frac{m^{\prime}}{m}$.
 Now, we prove that the result also holds for safe bidding strategies.
 Formally, we show that for $m,m^{\prime}\in\mathbb{N}$ and $m^{\prime}\geq m$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}\leq\frac{m^{\prime}}{m}$.
-To this end, we consider the quantity $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ analogous to $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-})$, upper bound it and maximize over all bid histories to get the desired result.
+To this end, we consider the quantity $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ analogous to $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-},\mathbf{v})$, upper bound it and maximize over all bid histories and valuation vectors to get the desired result.
 For any bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, let $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right):=\mathop{\mathrm{argmax}}_{\mathbf{b}^{\prime}\in\mathscr{U}_{m^{\prime}}^{\star}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime};\bm{\beta}_{-}^{t})$.
@@
 The proof to bound $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ is similar to that of Theorem C.4 till Eq.
-38, where we have
+50, where we have
 $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})\leq\frac{\sum_{j=1}^{m^{\prime}}\sum_{t\in T_{j}}V((w_{Q_{j}}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{S}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$
@@
 Lemma C.6.
-For any subset $\mathcal{S}\subseteq[m^{\prime}]$ such that $|S|=m$, suppose $\mathbf{b}$ is the bidding strategy obtained from $\mathcal{S}$.
+For any subset $\mathcal{S}\subseteq[m^{\prime}]$ such that $|\mathcal{S}|=m$, suppose $\mathbf{b}$ is the bidding strategy obtained from $\mathcal{S}$ by applying Algorithm 2.
 Then, $\mathbf{b}$ is a safe bidding strategy.
@@
 Appendix D Tight Lower Bounds for Results in Section 5
-D.1.  Tight Lower Bound for Theorem 5.3 (For $m=1$)
-In this section, we construct a bid history, $\mathcal{H}_{-}$ and valuation curve $\mathbf{v}$ (equivalently the $\mathscr{U}_{m}^{\star}$) for which the upper bound on $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}$, presented in Theorem 5.3, is tight.
-Recall that for any choice of $\mathcal{H}_{-}$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})\leq 2-\theta_{\mathcal{H}_{-}}$, where $\theta_{\mathcal{H}_{-}}\leq\max_{j\in[m]}({W}_{\widehat{Q}j}/{W}_{Q_{j}^{*}})$.
-To minimize $\theta_{\mathcal{H}_{-}}$, we choose a valuation curve that is very weakly decreasing.
-We then set the competing bids such that $\widehat{Q}_{j}\ll Q_{j}^{*}$ for all $j\in[m]$.
-Finally, we determine the values of $|T_{j,0}|$ and $|T_{j,1}|$ for which the upper bound is tight.
-See the definition of these quantities in Section C.1, where the first part of the theorem is proven.
-We present the case when $m=1$ below.
-The case for $m\geq 2$ is deferred to Section D.2.
-Although the main idea for both the cases are the same, for $m\geq 2$, the construction is more involved, and hence presented separately.
-Formally, for any $\delta\in(0,1/2]$, we design a bid history, $\mathcal{H}_{-}$, and valuation vector, $\mathbf{v}$ for which $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{1},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq 2-\delta$.
-Let $M=2\left\lceil\frac{1}{\delta}\right\rceil$.
-Suppose $\mathbf{v}=[1,v,\cdots,v]\in\mathbb{R}^{M}$, target RoI $\gamma=0$ and $v=1-4\epsilon$ where $\epsilon=\frac{\delta-1/M}{4(1-1/M)}<\frac{\delta}{4}$.
-Observe that $\epsilon\in(0,1/8)$ as $\delta\leq 1/2$.
-Let $T=M$ and $K=M+1$.
-The bid history is defined as:
-$\bm{\beta}_{-,t}^{-(j)}=\begin{cases}1-\epsilon,~{}&\text{ if $t\leq M-1$ and $j=1$}\\ C,~{}&\text{ if $t\leq M-1$ and $2\leq j\leq K$}\\ \epsilon,~{}&\text{ if $t=M$ and $1\leq j\leq K$}\\ \end{cases}\,,$
-where $C\gg w_{1}$. The bid history is presented in Table 1.
-Table 1: Bid history, $\mathcal{H}_{-}$, that achieves the $2$-approximation asymptotically for $m=1$. Here $C\gg w_{1}$ and $\epsilon>0$ is a small real number.
-$t=1$ $t=2$cdots $t=M-1$ $t=M$
-$C$ $C$cdots $C$ $\epsilon$
-$\vdots$ $\vdots$ $\cdots$ $\vdots$ $\vdots$
-$C$ $C$cdots $C$vdots
-$1-\epsilon$ $1-\epsilon$cdots $1-\epsilon$ $\epsilon$
-Notice that the constructed $\mathcal{H}_{-}$ contains $M-1$ rounds with high competition (where the bidder can acquire at most 1 unit), while there is one round with minimal competition, allowing the bidder to obtain any desired number of units.
-The 1-uniform safe bidding strategies of the form $(w_{q},q)$ do not perform well universally on all rounds because with increase in $q$, despite the increasing demand (i.e., $q$), the bid value $w_{q}$ decreases, thereby reducing the likelihood of acquiring any units.
-Notations.
-For a given bid history $\mathcal{H}_{-}$, recall that for any $m\in\mathbb{N}$, $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$.
-We define $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)=\sum_{t=1}^{T}V(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})$, the value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ over the bid history $\mathcal{H}_{-}$.
-Similarly, $\mathbf{b}_{m}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right):=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ and $V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)=\sum_{t=1}^{T}V(\mathbf{b}_{m}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})$.
-Computing the optimal strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{1}$.
-We observe that no bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{1}$ can obtain more than 1 unit in the first $M-1$ rounds and $M$ units in the final round.
-We claim that $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=(1,M)$.
-Observe that $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated 1 unit in each of the first $M-1$ rounds and $M$ units in the final round.
-Hence, it is allocated the maximum number of units possible.
-To verify that $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is feasible for the bid history, notice that for $\forall t\in[M-1]$, $p(\bm{\beta}^{t})=1\leq w_{1}=1$.
-For round $t=M$, as $K>M$,
-$p(\bm{\beta}^{t})=\epsilon\leq 1-4\epsilon<w_{M}=\mu\cdot 1+(1-\mu)\cdot(1-4\epsilon),\quad\text{where}\quad\mu=\frac{1}{M},$
-implying that RoI constraint is satisfied by $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in round $t=T$. So, $V^{\mathsf{OPT}}_{1}\left(\mathcal{H}_{-}\right)=M+v(M-1)$.
-Computing the optimal strategy in $\mathscr{U}_{1}^{\star}$.
-Here, we argue that the optimal strategy in $\mathscr{U}_{1}^{\star}$ is $(w_{1},1)=(1,1)$.
-Observe that the strategy $(1,1)$ is allocated 1 unit in each round.
-Hence, it obtains a total value of $M$.
-As $\bm{\beta}_{-,t}^{-(1)}=1-\epsilon>1-2\epsilon=w_{2}$, for $t=1,2,\dots,M-1$, bidding $\left(w_{q},q\right)$ for $q\geq 2$ does not get any value in the first $M-1$ rounds.
-Bidding $\left(w_{q},q\right)$ gets exactly $q$ units in round $M$ as $w_{q}>\epsilon$ for any $q\geq 2$.
-So, for $2\leq q\leq M$, the total value obtained by bidding $(w_{q},q)$ is $1+(q-1)v<q$.
-Hence, $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)=M$, which is the value obtained by the safe strategy $(w_{1},1)=(1,1)$.
-So,
-$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{1};\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})=\frac{V^{\mathsf{OPT}}_{1}\left(\mathcal{H}_{-}\right)}{V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}=\frac{M+v(M-1)}{M}=\frac{2M-1-4\epsilon(M-1)}{M}=2-\delta\,.$
-D.2.  Tight Lower Bound for Theorem 5.3 (For $m\geq 2$)
+D.1. Tight Lower Bound for Theorem 5.3 (For $m\geq 2$)
 In this section, we design a bid history $\mathcal{H}_{-}$ and valuation vector, $\mathbf{v}$ such that for any $\delta\in(0,1/2]$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq 2-\delta$.
@@
 where the second inequality follows because $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}\leq m$ (by Theorem 5.4).
-So, instead of computing $V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)$ directly which requires solving a DP, we obtain $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$ and show that the bound is tight.
-D.2.1.  Construction of $\mathcal{H}_{-}$.
+Surprisingly, we show that instead of computing $V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)$ directly, computing the upper bound of $mV^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$ also gives a tight bound.
+D.1.1. Construction of $\mathcal{H}_{-}$.
 We first decide all the parameters.
@@
 We present an example for such a bid history in Table 3.
-Table 2: Example of bid history, $\mathcal{H}_{-}$, achieving 2-approximation asymptotically for $m=2$.
-Each round in the same partition has identical competing bid profile.
-Total number of units in each auction is $K=N^{3}+1$.
+Table 3: Bid history achieving tight lower bound for $m=2$. Each round in the same partition has identical competing bid profile. Total number of units in each auction is $K=N^{3}+1$.
 Partition 1 Partition 2 Partition 3 Partition 4
@@
 $N^{3}+1$ bids are $w_{N^{3}+1}+\epsilon$ $N^{2}$ bids are $w_{N^{2}+1}+\epsilon$ $N+1$ bids are $w_{N+1}+\epsilon$ 1 bid is $w_{2}+\epsilon$
-D.2.2.  Computing $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$.
-We make the following claim about the optimal $m$-uniform bidding strategy for the constructed $\mathcal{H}_{-}$.
+D.1.2. Computing $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$.
 Lemma D.1.
@@
 $=N^{2m-1}+(2m-1)(N^{2m-1}-N^{2m-2})v\,.$
-D.2.3.  Computing $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$
+D.1.3. Computing $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$
 Recall that we compute $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$ and invoke the bounds on $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}$, instead of directly evaluating $V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)$.
@@
 So, if $N^{2m-j}<q\leq N^{2m-j+1}$, $(w_{q},q)$ gets no units in
-$\sum_{\ell=j}^{2m}N^{\ell-1}-N^{\ell-2}=N^{2m-1}-N^{j-2}\text{ auctions.}$.
+$\sum_{\ell=j}^{2m}N^{\ell-1}-N^{\ell-2}=N^{2m-1}-N^{j-2}\text{ auctions.}$
 In the remaining $N^{j-2}$ auctions it can win at most $q$ units. So, the maximum value obtained by $(w_{q},q)$ for any $q\geq 2$ is
@@
 where the last step follows by substituting the value of $\epsilon^{\prime}$ and thus concludes the proof.
-D.3.  Tight Lower Bound for Theorem 5.4
+D.2. Tight Lower Bound for Theorem 5.4
 In this section, we construct a bid history $\mathcal{H}_{-}$ and valuation vector, $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{m^{\prime}}{m}-\delta$, i.e., the upper bound is tight for $m^{\prime}\geq m+1$ (If $m^{\prime}=m$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}=1$ implying the bound is tight).
@@
 We now establish a lower bound on $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$.
-D.3.1.  Construction of $\mathcal{H}_{-}$.
+D.2.1. Construction of $\mathcal{H}_{-}$.
 We first decide all the parameters.
@@
 We present an example for such a bid history in Table 4.
-Table 3: Example of bid history, $\mathcal{H}_{-}$, achieving 4-approximation asymptotically for $m^{\prime}=4$ (and $m=1$).
+Table 4: Bid history attaining tight lower bound for $m^{\prime}=4$ (and $m=1$).
 Each round in the same partition has identical competing bid profile.
@@
 $N^{3}$ bids are $w_{N^{3}+1}+\epsilon$ $N^{2}+1$ bids are $w_{N^{2}+1}+\epsilon$ $N+1$ bids are $w_{N+1}+\epsilon$ 2 bids are $w_{2}+\epsilon$
-D.3.2.  Computing $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
-We make the following claim about the optimal $m^{\prime}$-uniform safe bidding strategy for $\mathcal{H}_{-}$.
+D.2.2. Computing $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
 Lemma D.3.
@@
 Note that the unique bid values (ignoring the quantity for the sake of brevity) in $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$, provided in Eq.
-49, are $\mathbf{b}=\left\{1,w_{N},\dots,w_{N^{m^{\prime}-1}}\right\}$.
+62, are $\mathbf{b}=\left\{1,w_{N},\dots,w_{N^{m^{\prime}-1}}\right\}$.
 We claim that the winning bid values of $\mathbf{b}_{m^{\prime}}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$ in the $j^{th}$ partition are $\widetilde{\mathbf{b}}=\left\{1,w_{N},\dots,w_{N^{m^{\prime}-j}}\right\}$.
@@
 $=N^{m^{\prime}-1}+(m^{\prime}-1)(N^{m^{\prime}-1}-N^{m^{\prime}-2})v\,.$
-D.3.3.  Computing $\mathbf{b}_{1}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
-In this section, we compute the optimal 1-uniform safe bidding strategy for the bid history, $\mathcal{H}_{-}$.
+D.2.3. Computing $\mathbf{b}_{1}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right)$
 Lemma D.4.
@@
 ∎
-Substituting values from Lemma D.3, Lemma D.4 and Eq. 48
-$\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{1}{m}\cdot\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ $=\frac{V^{\mathsf{SAFE}}_{m^{\prime}}\left(\mathcal{H}_{-}\right)}{mV^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}$
+Substituting values from Lemma D.3, Lemma D.4
+[…]
+$=\frac{V^{\mathsf{SAFE}}_{m^{\prime}}\left(\mathcal{H}_{-}\right)}{mV^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}$
 $=\frac{N^{m^{\prime}-1}+(m^{\prime}-1)(N^{m^{\prime}-1}-N^{m^{\prime}-2})v}{mN^{m^{\prime}-1}}$
@@
 As $\mathscr{U}_{m}^{\star}\subseteq\mathscr{F}^{\mathcal{H}_{-}}_{m}$ for all $m\in\mathbb{N}$, the upper bound in Theorem C.4 is also tight, i.e., for any $m^{\prime}\geq m$ and $\delta\in(0,\frac{1}{2}]$, there exists a bid history $\mathcal{H}_{-}$ and valuation curve $\mathbf{v}$ such that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{m^{\prime}}{m}-\delta$.
-D.4.  Tight Lower Bound for Theorem 5.5
+D.3. Tight Lower Bound for Theorem 5.5
 In this section, we present a bid history $\mathcal{H}_{-}$ and valuation vector, $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{2m^{\prime}}{m}-\delta$, i.e., the upper bound is tight for $m^{\prime}\geq m$.
@@
 Eq.
-46).
+59).
 Thus, we consider a valuation curve and a bid history that has a structure identical to the one presented in Section D.1 but with the following modified parameters,
@@
 […]
-$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{1}{m}\cdot\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ $=\frac{N^{2m^{\prime}-1}+(2m^{\prime}-1)(N^{2m^{\prime}-1}-N^{2m^{\prime}-2})v}{mN^{2m^{\prime}-1}}$
+[…]
+$=\frac{N^{2m^{\prime}-1}+(2m^{\prime}-1)(N^{2m^{\prime}-1}-N^{2m^{\prime}-2})v}{mN^{2m^{\prime}-1}}$
 $=\frac{2m^{\prime}N^{2m^{\prime}-1}-(2m^{\prime}-1)N^{2m^{\prime}-2}-2\epsilon^{\prime}(2m^{\prime}-1)(N^{2m^{\prime}-1}-N^{2m^{\prime}-2})}{mN^{2m^{\prime}-1}}$
@@
 where the last step follows by substituting the value of $\epsilon^{\prime}$ and thus concludes the proof.
-Appendix E Simulation Details
-In Section 5.3, we conducted numerical experiments to estimate the richness ratio (or equivalently $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}$) in practice for the classes of bidding strategies discussed earlier.
-We now discuss a few details regarding the same.
-Dataset.
-We estimate the richness ratio for the EU ETS emission permit auction data for 2022 and 2023.
-However, only aggregate statistics of the submitted bids is publicly available for privacy reasons.
-Hence, we synthesize individual level bid data from these available statistics.
-The exact procedure to reconstruct the bid data is presented in Section E.1.
+Appendix E Omitted Details from Section 6
+E.1. Proof of Lemma 6.1
+We introduce a few additional notations to make the proof easier to parse.
+Specifically, we denote random variables with a hat, e.g.
+$\widehat{\mathbf{b}}^{t}$, and their realizations without hats, e.g.
+$\mathbf{b}^{t}$.
+Let $\mathbb{E}_{t}[\cdot]$ denote the expectation taken only with respect to randomness in selecting $\widehat{\mathbf{b}}^{t}$.
+For an adaptive adversary and any fixed $\mathbf{b}\in\mathscr{U}_{m}^{\star}$,
+[…]
+$=\sup_{\bm{\beta}_{-}^{1},\dots,\bm{\beta}_{-}^{T}}\sum_{t=1}^{T}\mathbb{E}_{t}\Big[V(\mathbf{b};\bm{\beta}_{-}^{t})-V(\widehat{\mathbf{b}}^{t};\bm{\beta}_{-}^{t})\Big]\,.$ (64)
+To see why Section E.1 holds, let $T=2$. Then, the right hand side in the first equation becomes
+$\sup_{\bm{\beta}_{-}^{1}}\mathbb{E}_{1}\Big[V(\mathbf{b};\bm{\beta}_{-}^{1})-V(\widehat{\mathbf{b}}^{1};\bm{\beta}_{-}^{1})\Big]+\sup_{\bm{\beta}_{-}^{2}}\mathbb{E}_{2}\Big[V(\mathbf{b};\bm{\beta}_{-}^{2})-V(\widehat{\mathbf{b}}^{2};\bm{\beta}_{-}^{2})\Big]$
+Note that $\mathbb{E}_{1}\Big[V(\mathbf{b};\bm{\beta}_{-}^{1})-V(\widehat{\mathbf{b}}^{1};\bm{\beta}_{-}^{1})\Big]$ does not depend on $\bm{\beta}_{-}^{2}$ and $\bm{\beta}_{-}^{2}$, due to the forgetful property of the learning algorithm, does not depend on the realization of $\widehat{\bm{\beta}_{-}^{1}}$.
+Hence,
+$\sup_{\bm{\beta}_{-}^{1}}\mathbb{E}_{1}\Big[V(\mathbf{b};\bm{\beta}_{-}^{1})-V(\widehat{\mathbf{b}}^{1};\bm{\beta}_{-}^{1})\Big]+\sup_{\bm{\beta}_{-}^{2}}\mathbb{E}_{2}\Big[V(\mathbf{b};\bm{\beta}_{-}^{2})-V(\widehat{\mathbf{b}}^{2};\bm{\beta}_{-}^{2})\Big]$
+$=\sup_{\bm{\beta}_{-}^{1},\bm{\beta}_{-}^{2}}\Big\{\mathbb{E}_{1}\Big[V(\mathbf{b};\bm{\beta}_{-}^{1})-V(\widehat{\mathbf{b}}^{1};\bm{\beta}_{-}^{1})\Big]+\mathbb{E}_{2}\Big[V(\mathbf{b};\bm{\beta}_{-}^{2})-V(\widehat{\mathbf{b}}^{2};\bm{\beta}_{-}^{2})\Big]\Big\}$
+Repeating the same argument for any $T\geq 2$, we get the equality in Section E.1. Hence,
+[…]
+$=\sup_{\bm{\beta}_{-}^{1},\dots,\bm{\beta}_{-}^{T}}\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\Big\{\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}_{t}[V(\widehat{\mathbf{b}}^{t};\bm{\beta}_{-}^{t})]\Big\}\leq B,$
+where the last inequality holds as $\textsf{REG}\leq B$ for any sequence of competing bids $\bm{\beta}_{-}^{1},\dots,\bm{\beta}_{-}^{T}$ by an oblivious adversary.
+Taking expectations on both sides gives $\textsf{REG}_{\text{adap}}\leq B$.
+E.2. Proof of Lemma 6.2
+We have that $\widehat{\textsf{REG}}_{\text{adap}}\leq B(\delta)$ w. p. at least $1-\delta$. Then,
+$\textsf{REG}_{\text{adap}}=\mathbb{E}[\widehat{\textsf{REG}}_{\text{adap}}]$ $=\mathbb{E}\left[\widehat{\textsf{REG}}_{\text{adap}}\cdot\mathbb{I}{\left[\widehat{\textsf{REG}}_{\text{adap}}\geq B(\delta)\right]}\right]+\mathbb{E}\left[\widehat{\textsf{REG}}_{\text{adap}}\cdot\mathbb{I}{\left[\widehat{\textsf{REG}}_{\text{adap}}\leq B(\delta)\right]}\right]$
+$\leq MT\cdot\mathbb{P}[\widehat{\textsf{REG}}_{\text{adap}}\geq B(\delta)]+B(\delta)\leq MT\delta+B(\delta)\,$,
+where the first inequality holds as $\textsf{REG}_{\text{adap}}\leq MT$ w.p. 1.
+E.3. Proof of Theorem 6.3
+The Algorithm.
+Before proving Theorem 6.3, we present Algorithm 3, which is designed for this setting.
+As stated earlier, our approach builds on ideas from György et al.
+(2007).
+We first define a covering path set $\mathcal{C}$ as a collection of $s$-$d$ paths $\mathfrak{p}$ such that every edge $e\in E$ appears in at least one path $\mathfrak{p}\in\mathcal{C}$.
+Such a set can be constructed with size at most $|E|$ by including, for each edge $e$, at least one path that contains $e$.
+Hence, $|\mathcal{C}|\leq|E|$.
+In each round $t$, similar to Algorithm 1, the bidder constructs a DAG $\mathcal{G}^{t}(V,E)$ without edge weights.
+With probability $\lambda>0$, the bidder selects an $s$-$d$ path uniformly at random from the set $\mathcal{C}$.
+Otherwise, they update edge weights and sample a path as a Markov chain, following Algorithm 1.
+Once a path is selected, the bidder maps it to a safe bidding strategy with at most $m$ bid-quantity pairs.
+Since the true weights cannot be computed $\forall e\in E$ under the bandit feedback structure, we therefore use the following biased estimator $\widehat{\textsf{w}}^{t}(e)$ of $\textsf{w}^{t}(e)$:
+$\widehat{\textsf{w}}^{t}(e)$ $=\frac{\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}+\frac{\theta}{p^{t}(e)}\quad\text{where}\quad p^{t}(e)=\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\,.$ (65)
+Here, $\textsf{w}^{t}(e)$ is as per Eq.
+6 and $\theta\in[0,M]$ is a parameter that will be determined shortly.
+As before, $p^{t}(e)$ denotes the probability of selecting edge $e$ in round $t$, and $\mathbb{P}^{t}(\cdot)$ denotes the distribution over all ssdd paths in the graph $\mathcal{G}^{t}(V,E)$.
+Following the description of the algorithm, the distribution $\mathbb{P}^{t}(\cdot)$ can be expressed as a convex combination of two distributions: the uniform distribution over the paths in the covering path set $\mathcal{C}$, given by $\frac{\mathbb{I}{\left[\cdot\in\mathcal{C}\right]}}{|\mathcal{C}|}$, and $\widehat{\mathbb{P}}^{t}(\cdot)$, which is the distribution over ssdd paths in round $t$ obtained using the path selection step from Algorithm 1.
+Formally,
+$\mathbb{P}^{t}(\mathfrak{p})$ $:=(1-\lambda)\cdot\widehat{\mathbb{P}}^{t}(\mathfrak{p})+\lambda\cdot\frac{\mathbb{I}{\left[\mathfrak{p}\in\mathcal{C}\right]}}{|\mathcal{C}|},\quad\text{and}\quad\widehat{\mathbb{P}}^{t}(\mathfrak{p})=\prod_{e\in\mathfrak{p}}\varphi^{t}(e)=\frac{\exp(\eta\sum_{\tau=1}^{t-1}\widehat{\textsf{w}}^{\tau}(\mathfrak{p}))}{\sum_{\mathfrak{p}^{\prime}}\exp(\eta\sum_{\tau=1}^{t-1}\widehat{\textsf{w}}^{\tau}(\mathfrak{p}^{\prime}))}\,.$ (66)
+Here, the last equality holds because the $s$-$d$ path distribution induced by the path selection step Algorithm 1 matches that of the Hedge algorithm over $s$-$d$ paths (see proof of Theorem 4.3).
+The estimator, $\widehat{\textsf{w}}^{t}(e)$, is well defined as
+[…]
+where the second inequality follows as every edge $e$ is contained in at least one path $\mathfrak{p}\in\mathcal{C}$.
+Algorithm 3 Learning Safe Bidding Strategies (Bandit Setting, Adaptive Adversary)
+1:valuation curve $\mathbf{v}$, time horizon $T$, learning rate $\eta>0$, parameters $\lambda$, $\theta$, covering path set $\mathcal{C}$.
+Initialize $\varphi^{0}(e)=1$ and $\widehat{\textsf{w}}^{0}(e)=0,\forall e\in E$.
+2:for $t=1,2,\dots,T$ do
+3:  Construct $\mathcal{G}^{t}(V,E)$ similar to $\mathcal{G}(V,E)$ without weights.
+4:  Sample $Z_{t}\sim\text{Unif}[0,1]$.
+5:  if $Z_{t}\leq\lambda$ then
+6:    Select path $\mathfrak{p}^{t}$ uniformly at random from the set $\mathcal{C}$.
+7:  else
+8:    Set $\Gamma^{t-1}(d)=1$ and recursively compute in bottom-to-top fashion for every node $u$ in $\mathcal{G}^{t}(V,E)$:
+$\Gamma^{t-1}(u)=\sum_{v:u\to v=e\ni E}\Gamma^{t-1}(v)\cdot\varphi^{t-1}(e)\cdot\exp(\eta\widehat{\textsf{w}}^{t-1}(e))\,.$
+9:    For edge $e=u\to v$ in $\mathcal{G}^{t}(V,E)$, update edge probability:
+$\varphi^{t}(e)=\varphi^{t-1}(e)\cdot\exp(\eta\widehat{\textsf{w}}^{t-1}(e))\cdot\frac{\Gamma^{t-1}(v)}{\Gamma^{t-1}(u)}.$
+   Define initial node $u=s$ and path $\mathfrak{p}^{t}=s$.
+10:    while $u\neq d$ do
+11:     Sample $v$ with probability $\varphi^{t}(u\to v)$.
+12:     Append $v$ to the path $\mathfrak{p}^{t}$; set $u\leftarrow v$.
+13:    end while
+14:  end if
+15:  If $\mathfrak{p}^{t}=s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for some $k\in[m]$, submit $\mathbf{b}^{t}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
+$b_{\ell}=w_{z_{\ell}}\qquad\text{and}\qquad q_{\ell}=z_{\ell}-z_{\ell-1},~\forall\ell\in[k]\,.$
+16:  The bidder sets edge weights $\widehat{\textsf{w}}^{t}(e)$ per Eq. 65.
+17:end for
+Regret Analysis. We begin by stating the following result which is a modification of (György et al., 2007, Lemma 2).
+Lemma E.1.
+For any $\delta\in(0,1)$ and $\theta\in(0,M]$ and $e\in E$,
+$\mathbb{P}\left[\sum_{t=1}^{T}\textsf{w}^{t}(e)>\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(e)+\frac{M^{2}}{\theta}\log\frac{|E|}{\delta}\right]\leq\frac{\delta}{|E|}\,.$
+To upper bound the regret of Algorithm 3, define $\Phi_{t}=\sum_{\mathfrak{p}}\exp(\eta\sum_{\tau=1}^{t}\widehat{\textsf{w}}^{\tau}(\mathfrak{p}))$ where $\widehat{\textsf{w}}^{\tau}(\mathfrak{p})=\sum_{e\in\mathfrak{p}}\widehat{\textsf{w}}^{\tau}(e)$ for any path $\mathfrak{p}$ in the DAG.
+Suppose
+$\eta=\frac{\lambda}{2(m+1)M|E|}\quad\text{and}\quad 0<\lambda\leq\frac{1}{2}\,.$ (67)
+Observe that $\widehat{\textsf{w}}^{t}(e)=\frac{\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}+\frac{\theta}{p^{t}(e)}\implies\widehat{\textsf{w}}^{t}(e)\leq\frac{2M|E|}{\lambda}$ as $\textsf{w}^{t}(e)\leq M,\theta\leq M$ and $p^{t}(e)\geq\frac{\lambda}{|\mathcal{C}|}\geq\frac{\lambda}{|E|}$.
+So,
+[…]
+where the inequality holds as any path $\mathfrak{p}$ in the DAG has at most $m+1$ edges. Now,
+$\frac{\Phi_{t}}{\Phi_{t-1}}$ $=\sum_{\mathfrak{p}}\frac{\exp(\eta\sum_{\tau=1}^{t-1}\widehat{\textsf{w}}^{\tau}(\mathfrak{p}))}{\Phi_{t-1}}\cdot\exp(\eta\widehat{\textsf{w}}^{t}(\mathfrak{p}))$
+[…]
+[…]
+$=1+\sum_{\mathfrak{p}}\widehat{\mathbb{P}}^{t}(\mathfrak{p})\cdot(\eta\widehat{\textsf{w}}^{t}(\mathfrak{p})+\eta^{2}\widehat{\textsf{w}}^{t}(\mathfrak{p})^{2})$
+$\leq 1+\frac{\eta}{1-\lambda}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\widehat{\textsf{w}}^{t}(\mathfrak{p})+\frac{\eta^{2}}{1-\lambda}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\widehat{\textsf{w}}^{t}(\mathfrak{p})^{2}$
+$\leq\exp\left(\frac{\eta}{1-\lambda}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\widehat{\textsf{w}}^{t}(\mathfrak{p})+\frac{\eta^{2}}{1-\lambda}\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\widehat{\textsf{w}}^{t}(\mathfrak{p})^{2}\right)$
+where the first inequality uses $e^{x}\leq 1+x+x^{2}$ for $x\leq 1$, second one follows as $\widehat{\mathbb{P}}^{t}(\mathfrak{p})=\frac{\mathbb{P}^{t}(\mathfrak{p})-\frac{\lambda}{|\mathcal{C}|}\cdot\mathbb{I}{\left[\mathfrak{p}\in\mathcal{C}\right]}}{1-\lambda}\leq\frac{\mathbb{P}^{t}(\mathfrak{p})}{1-\lambda}$ and third one is true as $1+x\leq e^{x},\forall x$.
+Now,
+$\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\widehat{\textsf{w}}^{t}(\mathfrak{p})$ $=\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\widehat{\textsf{w}}^{t}(e)$
+$=\sum_{e\in E}\widehat{\textsf{w}}^{t}(e)\sum_{\mathfrak{p}:e\in\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})$
+[…]
+[…]
+$=\textsf{w}^{t}(\mathfrak{p}^{t})+|E|\theta\,.$
+Similarly,
+$\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\widehat{\textsf{w}}^{t}(\mathfrak{p})^{2}$ $\leq(m+1)\sum_{\mathfrak{p}}\mathbb{P}^{t}(\mathfrak{p})\sum_{e\in\mathfrak{p}}\widehat{\textsf{w}}^{t}(e)^{2}$
+$=(m+1)\sum_{e\in E}\widehat{\textsf{w}}^{t}(e)^{2}p^{t}(e)$
+[…]
+$\leq 2(m+1)M\sum_{e\in E}\widehat{\textsf{w}}^{t}(e),$
+where the first inequality holds due to Cauchy-Schwarz inequality. Combining everything,
+$\frac{\Phi_{t}}{\Phi_{t-1}}$ $\leq\exp\left(\frac{\eta}{1-\lambda}(\textsf{w}^{t}(\mathfrak{p}^{t})+|E|\theta)+\frac{2(m+1)M\eta^{2}}{1-\lambda}\sum_{e\in E}\widehat{\textsf{w}}^{t}(e)\right)$
+Taking logarithms on both sides and adding from $t=1$ till $t=T$, we get
+$\log\Phi_{T}-\log\Phi_{0}$ $\leq\frac{\eta}{1-\lambda}\sum_{t=1}^{T}(\textsf{w}^{t}(\mathfrak{p}^{t})+|E|\theta)+\frac{2(m+1)M\eta^{2}}{1-\lambda}\sum_{t=1}^{T}\sum_{e\in E}\widehat{\textsf{w}}^{t}(e)$
+$\leq\frac{\eta}{1-\lambda}\sum_{t=1}^{T}(\textsf{w}^{t}(\mathfrak{p}^{t})+|E|\theta)+\frac{2(m+1)M\eta^{2}|E|}{1-\lambda}\max_{\mathfrak{p}}\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(\mathfrak{p}),$
+where the last inequality follows from the fact that $\sum_{e\in E}\widehat{\textsf{w}}^{t}(e)\leq\sum_{\mathfrak{p}\in\mathcal{C}}\widehat{\textsf{w}}^{t}(\mathfrak{p})$, since each edge $e$ is contained in at least one path $\mathfrak{p}\in\mathcal{C}$.
+Hence,
+$\sum_{t=1}^{T}\sum_{e\in E}\widehat{\textsf{w}}^{t}(e)\leq\sum_{t=1}^{T}\sum_{\mathfrak{p}\in\mathcal{C}}\widehat{\textsf{w}}^{t}(\mathfrak{p})=\sum_{\mathfrak{p}\in\mathcal{C}}\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(\mathfrak{p})\leq|E|\max_{\mathfrak{p}}\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(\mathfrak{p}),$
+where the last inequality holds as $|\mathcal{C}|\leq|E|$. Observe that, $\Phi_{0}\leq M^{m}$ and $\log\Phi_{T}\geq\eta\max_{\mathfrak{p}}\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(\mathfrak{p})$. Thus,
+$\max_{\mathfrak{p}}\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(\mathfrak{p})-\frac{m\log M}{\eta}$ $\leq\frac{1}{1-\lambda}\sum_{t=1}^{T}(\textsf{w}^{t}(\mathfrak{p}^{t})+|E|\theta)+\frac{2(m+1)M\eta|E|}{1-\lambda}\max_{\mathfrak{p}}\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(\mathfrak{p})\,.$
+Rearranging the terms, we get
+$(1-\lambda-2\eta(m+1)M|E|)\max_{\mathfrak{p}}\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(\mathfrak{p})-\sum_{t=1}^{T}\textsf{w}^{t}(\mathfrak{p}^{t})$ $\leq\frac{m\log M}{\eta}+T|E|\theta\,.$
+Note that $1-\lambda-2\eta(m+1)M|E|=1-2\lambda\geq 0$. By Lemma E.1 and union bound over all edges $e\in E$, we get that with probability at least $1-\delta$,
+$\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(e)\geq\sum_{t=1}^{T}\textsf{w}^{t}(e)-\frac{M^{2}}{\theta}\log\frac{|E|}{\delta},\forall e\in E\,.$
+For any $\mathfrak{p}$, summing over the edges, we get that with probability at least $1-\delta$,
+$\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(\mathfrak{p})\geq\sum_{t=1}^{T}\textsf{w}^{t}(\mathfrak{p})-\frac{(m+1)M^{2}}{\theta}\log\frac{|E|}{\delta}\,.$
+Hence,
+$(1-\lambda-2\eta(m+1)M|E|)\left(\max_{\mathfrak{p}}\sum_{t=1}^{T}\textsf{w}^{t}(\mathfrak{p})-\frac{(m+1)M^{2}}{\theta}\log\frac{|E|}{\delta}\right)-\sum_{t=1}^{T}\textsf{w}^{t}(\mathfrak{p}^{t})$ $\leq\frac{m\log M}{\eta}+T|E|\theta\,.$
+Using the fact that $\lambda=2\eta(m+1)M|E|$ and $\max_{\mathfrak{p}}\sum_{t=1}^{T}\textsf{w}^{t}(\mathfrak{p})\leq MT$, we get
+$\max_{\mathfrak{p}}\sum_{t=1}^{T}\textsf{w}^{t}(\mathfrak{p})-\sum_{t=1}^{T}\textsf{w}^{t}(\mathfrak{p}^{t})$ $\leq\frac{m\log M}{\eta}+4\eta(m+1)M^{2}|E|T+T|E|\theta+\frac{(m+1)M^{2}}{\theta}\log\frac{|E|}{\delta}$
+Setting
+$\eta=\frac{1}{2M}\sqrt{\frac{m\log M}{(m+1)|E|T}},~~\theta=M\sqrt{\frac{(m+1)}{T|E|}\log\frac{|E|}{\delta}},\quad\text{and}\quad\lambda=\sqrt{\frac{m(m+1)|E|\log M}{T}}$
+and $T\geq\max\{4m(m+1)|E|\log M,\log(|E|/\delta)\}$ (to ensure $\theta\in(0,M]$ and $\lambda\in(0,\frac{1}{2}]$) we obtain,
+$\max_{\mathfrak{p}}\sum_{t=1}^{T}\textsf{w}^{t}(\mathfrak{p})-\sum_{t=1}^{T}\textsf{w}^{t}(\mathfrak{p}^{t})$ $\leq 4M\sqrt{m(m+1)|E|T\log M}+2M\sqrt{(m+1)|E|T\log(|E|/\delta)}\,.$
+By Theorem 4.2 and Algorithm 3, $\max_{\mathfrak{p}}\sum_{t=1}^{T}\textsf{w}^{t}(\mathfrak{p})=\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ and $\sum_{t=1}^{T}\textsf{w}^{t}(\mathfrak{p}^{t})=\sum_{t=1}^{T}V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})$.
+Furthermore, $|E|=O(mM^{2})$.
+Hence, with probability at least $1-\delta$,
+$\widehat{\textsf{REG}}_{\text{adap}}=\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})$ $\leq O(m^{3/2}M^{2}\sqrt{T\log M}+mM^{2}\sqrt{T\log(M/\delta)})\,.$
+Time Complexity.
+The computational bottleneck is in computing the edge weight estimator in Eq.
+65.
+Specifically, we need to compute the marginal probabilities $p^{t}(e)$ efficiently.
+From the proof of Theorem 4.4, we know that this can be done efficiently in $O(mM^{2})$ time.
+Thus, the per-round running time of Algorithm 3 is $O(mM^{2})$ time per-round.
+E.3.1. Proof of Lemma E.1
+Fix any $e\in E$. By Chernoff bound, for any $z>0$ and $\zeta>0$,
+$\mathbb{P}\left[\sum_{t=1}^{T}\textsf{w}^{t}(e)>\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(e)+z\right]\leq e^{-\zeta z}\mathbb{E}\left[\exp\left(\zeta\left(\sum_{t=1}^{T}\textsf{w}^{t}(e)-\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(e)\right)\right)\right]$
+Setting $z=\frac{M^{2}}{\theta}\log\frac{|E|}{\delta}$ and $\zeta=\frac{\theta}{M^{2}}$, we get
+$\mathbb{P}\left[\sum_{t=1}^{T}\textsf{w}^{t}(e)>\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(e)+\frac{M^{2}}{\theta}\log\frac{|E|}{\delta}\right]\leq\frac{\delta}{|E|}\mathbb{E}\left[\exp\left(\zeta\left(\sum_{t=1}^{T}\textsf{w}^{t}(e)-\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(e)\right)\right)\right]\,.$
+To complete the proof, we need to show that $\mathbb{E}[\exp(\zeta(\sum_{t=1}^{T}\textsf{w}^{t}(e)-\sum_{t=1}^{T}\widehat{\textsf{w}}^{t}(e)))]\leq 1$. Define
+$X_{t}=\exp\left(\zeta\left(\sum_{\tau=1}^{t}\textsf{w}^{\tau}(e)-\sum_{\tau=1}^{t}\widehat{\textsf{w}}^{\tau}(e)\right)\right)\quad\text{and}\quad X_{0}=1\,.$
+We show that for any $t$, $\mathbb{E}[X_{t}|\mathbf{b}_{1},\dots,\mathbf{b}_{t-1}]:=\mathbb{E}_{t}[X_{t}]\leq X_{t-1}$.
+Computing conditional expectations recursively results in $\mathbb{E}[X_{t}]\leq 1,\forall t$ which proves the result.
+To prove $\mathbb{E}_{t}[X_{t}]\leq X_{t-1}$, note that $X_{t}=X_{t-1}\exp(\zeta(\textsf{w}^{t}(e)-\widehat{\textsf{w}}^{t}(e)))$.
+So,
+$\mathbb{E}_{t}[X_{t}]$ $=X_{t-1}\mathbb{E}_{t}\left[\exp(\zeta(\textsf{w}^{t}(e)-\widehat{\textsf{w}}^{t}(e)))\right]$
+[…]
+$=X_{t-1}e^{-\frac{\theta^{2}}{M^{2}p^{t}(e)}}\mathbb{E}_{t}\left[\exp\left(\frac{\theta}{M^{2}}\left(\textsf{w}^{t}(e)-\frac{\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}\right)\right)\right]$
+$\leq X_{t-1}e^{-\frac{\theta^{2}}{M^{2}p^{t}(e)}}\mathbb{E}_{t}\left[1+\frac{\theta}{M^{2}}\left(\textsf{w}^{t}(e)-\frac{\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}\right)+\frac{\theta^{2}}{M^{4}}\left(\textsf{w}^{t}(e)-\frac{\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}\right)^{2}\right]$
+$=X_{t-1}e^{-\frac{\theta^{2}}{M^{2}p^{t}(e)}}\mathbb{E}_{t}\left[1+\frac{\theta^{2}}{M^{4}}\left(\textsf{w}^{t}(e)-\frac{\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}\right)^{2}\right]$
+$\leq X_{t-1}e^{-\frac{\theta^{2}}{M^{2}p^{t}(e)}}\mathbb{E}_{t}\left[1+\frac{\theta^{2}}{M^{4}}\left(\frac{\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}\right)^{2}\right]$
+$\leq X_{t-1}e^{-\frac{\theta^{2}}{M^{2}p^{t}(e)}}\left(1+\frac{\theta^{2}}{M^{2}p^{t}(e)}\right)$
+$\leq X_{t-1},$
+which is the desired result.
+Here, the first inequality follows as $\frac{\theta}{M^{2}}\left(\textsf{w}^{t}(e)-\frac{\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}\right)\leq\theta/M\leq 1$ and $e^{x}\leq 1+x+x^{2},\forall x\leq 1$, the fourth equality is true as $\mathbb{E}_{t}[\frac{\textsf{w}^{t}(e)}{p^{t}(e)}\cdot\mathbb{I}{\left[e\in\mathfrak{p}^{t}\right]}]=\textsf{w}^{t}(e)$, the third inequality follows as $\textsf{w}^{t}(e)\leq M$ and the final inequality holds true as $1+x\leq e^{x},\forall x$.
+E.4. Proof of Theorem 6.5
+Let $m=M=1$ and $K=2$.
+Fix any $c\in(0,1]$ and choose $\epsilon\in(0,\frac{c}{2})$.
+Assume $\epsilon T\in\mathbb{N}$.
+For the first $(1-\epsilon)T$ rounds, the top $K=2$ competing bids are $3T,1+\frac{\epsilon}{2}]$.
+For the remaining $\epsilon T$ rounds, the competing bids are $\epsilon,\epsilon]$.
+Define $\mathbf{v}=1$.
+Here, $\mathbf{b}^{*}=1+\epsilon$ is the fixed hindsight optimal strategy that satisfies the cumulative RoI constraint.
+The total value obtained by $\mathbf{b}^{*}$ is $T$ (1 in each round).
+Similarly, the total payment is $(1+\epsilon)\cdot(1-\epsilon)T+\epsilon\cdot\epsilon T=T$.
+By 6.4, the bidder follows safe strategies for the first $(1-\epsilon)T$ rounds and does not win any units.
+In the remaining $\epsilon T$ rounds, it can win at most $M=1$ unit per round.
+Thus, the maximum possible value obtained by the bidder is $\epsilon T$.
+Hence, $c$-$\textsf{REG}_{\text{cumul}}\geq c\cdot T-\epsilon T\geq\frac{cT}{2}=\Omega(T)$.
+E.5. Proof of Theorem 6.6
+Before proving the result, we formally present the learning algorithm for stochastic contexts in Algorithm 4.
+Note that the learner does not need to know the distribution $F_{\mathbf{v}}$ from which the contexts are sampled.
+The space complexity and per-round time complexity of Algorithm 4 is $O(mM^{2}|\mathcal{V}|)$.
+Algorithm 4 Learning Safe Bidding Strategies (Full Information, Stochastic Contexts)
+1:Set of valuation vectors, $\mathcal{V}$, time horizon $T$, learning rate $0<\eta\leq\frac{1}{M}$.
+Initialize $\varphi^{0}(e;\mathbf{v})=1$ and $\textsf{w}^{0}(e;\mathbf{v})=0,\forall e\in E,\forall\mathbf{v}\in\mathcal{V}$.
+2:for $t=1,2,\dots,T$ do
+3:  Observe valuation curve $\mathbf{v}^{t}\in\mathcal{V}$ which is sampled from $F_{\mathbf{v}}$.
+4:  Construct $\mathcal{G}^{t}(V,E,\mathbf{v})$ similar to $\mathcal{G}(V,E)$ without weights for all $\mathbf{v}\in\mathcal{V}$.
+5:  $\textsf{UPDATE}:$ Obtain edge probabilities $\varphi^{t}(\cdot;\mathbf{v})$ for all $\mathbf{v}\in\mathcal{V}$ as follows:
+6:  for $\mathbf{v}\in\mathcal{V}$ do
+7:    Set $\Gamma^{t-1}(d;\mathbf{v})=1$ and recursively compute in bottom-to-top fashion for every node $u$ in $\mathcal{G}^{t}(V,E,\mathbf{v})$:
+$\Gamma^{t-1}(u;\mathbf{v})=\sum_{v:u\to v=e\ni E}\Gamma^{t-1}(v;\mathbf{v})\cdot\varphi^{t-1}(e;\mathbf{v})\cdot\exp(\eta\textsf{w}^{t-1}(e;\mathbf{v}))$
+8:    For edge $e=u\to v$ in $\mathcal{G}^{t}(V,E,\mathbf{v})$, update: $\varphi^{t}(e;\mathbf{v})=\varphi^{t-1}(e;\mathbf{v})\cdot\exp(\eta\textsf{w}^{t-1}(e;\mathbf{v}))\cdot\frac{\Gamma^{t-1}(v;\mathbf{v})}{\Gamma^{t-1}(u;\mathbf{v})}.$
+9:  end for
+10:  SAMPLE: Define initial node $u=s$ and path $\mathfrak{p}^{t}=s$.
+11:  while $u\neq d$ do
+12:    Sample $v$ with probability $\varphi^{t}(u\to v;\mathbf{v}^{t})$.
+13:    Append $v$ to the path $\mathfrak{p}^{t}$; set $u\leftarrow v$.
+14:  end while
+15:  MAP: If $\mathfrak{p}^{t}=s\to(1,z_{1})\to\dots\to(k,z_{k})\to d$ for some $k\in[m]$, submit $\mathbf{b}^{t}=\langle(b_{1},q_{1}),\dots,(b_{k},q_{k})\rangle$ where
+$b_{\ell}=w_{z_{\ell}}^{t}\qquad\text{and}\qquad q_{\ell}=z_{\ell}-z_{\ell-1},~\forall\ell\in[k]\,.$
+where $w_{z_{\ell}}^{t}$ is the $z_{\ell}^{th}$ entry of the average cumulative valuation vector for all $\ell\in[k]$ corresponding to $\mathbf{v}^{t}$.
+16:  The bidder observes $\bm{\beta}_{-}^{t}$ and sets the edge weights for $e=x\to y$ as follows for each $\mathbf{v}=[v_{1},\dots,v_{M}]\in\mathcal{V}$: (i) If $x=(\ell-1,j)$ and $y=(\ell,j^{\prime})$ with $\ell\in[m]$ and $j<j^{\prime}$,
+$\textsf{w}^{t}(e;\mathbf{v})=\sum_{k=j+1}^{j^{\prime}}v_{k}\cdot\mathbb{I}{\left[w_{j^{\prime}}\geq\bm{\beta}_{-,t}^{-(k)}\right]}\,.$
+(ii) If $x=(\ell,j)$ and $y=d$, then $\textsf{w}^{t}(e,\mathbf{v})=0,\forall\ell\in[m],j\in[M]$.
+17:end for
+Regret Analysis.
+For any fixed $\mathbf{v}\in\mathcal{V}$, define $\Phi_{t}(\mathbf{v})=\sum_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}\exp(\eta\sum_{\tau=1}^{t}V(\mathbf{b};\bm{\beta}_{-}^{\tau}))$.
+Then,
+$\frac{\Phi_{t}(\mathbf{v})}{\Phi_{t-1}(\mathbf{v})}$ $=\sum_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}\frac{\exp(\eta\sum_{\tau=1}^{t-1}V(\mathbf{b};\bm{\beta}_{-}^{\tau}))}{\Phi_{t-1}(\mathbf{v})}\cdot\exp(\eta V(\mathbf{b};\bm{\beta}_{-}^{t}))$
+$=\sum_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}\mathbb{P}[\mathbf{b}^{t}=\mathbf{b}|\mathbf{v}^{t}=\mathbf{v}]\cdot\exp(\eta V(\mathbf{b};\bm{\beta}_{-}^{t})),$
+where the last line follows because for a fixed context, Algorithm 4 is equivalent to Algorithm 1, which is an equivalent implementation of the Hedge algorithm (see proof of Theorem 4.3).
+For $x\leq 1$, $e^{x}\leq 1+x+x^{2}$ and $\log(1+x)\leq x$ for all $x\geq 0$. As $\eta V(\mathbf{b};\bm{\beta}_{-}^{t})\leq\eta M\leq 1$, we have:
+$\frac{\Phi_{t}(\mathbf{v})}{\Phi_{t-1}(\mathbf{v})}$ $\leq\sum_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}\mathbb{P}[\mathbf{b}^{t}=\mathbf{b}|\mathbf{v}^{t}=\mathbf{v}]\cdot(1+\eta V(\mathbf{b};\bm{\beta}_{-}^{t})+\eta^{2}V(\mathbf{b};\bm{\beta}_{-}^{t})^{2})$
+$\leq\exp\Big(\sum_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}\mathbb{P}[\mathbf{b}^{t}=\mathbf{b}|\mathbf{v}^{t}=\mathbf{v}](\eta V(\mathbf{b};\bm{\beta}_{-}^{t})+\eta^{2}V(\mathbf{b};\bm{\beta}_{-}^{t})^{2})\Big)$
+Taking logarithms on both sides and summing over $t=1$ to $t=T$, we get
+$\log\Phi_{T}(\mathbf{v})-\log\Phi_{0}(\mathbf{v})\leq\sum_{t=1}^{T}\sum_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}\mathbb{P}[\mathbf{b}^{t}=\mathbf{b}|\mathbf{v}^{t}=\mathbf{v}](\eta V(\mathbf{b};\bm{\beta}_{-}^{t})+\eta^{2}V(\mathbf{b};\bm{\beta}_{-}^{t})^{2})$
+Define $\Phi_{0}(\mathbf{v})=\sum_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}1\leq M^{m}$. Moreover, for any stationary policy $\pi\in\Pi$, per Eq. 25,
+$\log\Phi_{T}(\mathbf{v})\geq\eta\sum_{t=1}^{T}V(\pi(\mathbf{v});\bm{\beta}_{-}^{t})\,.$
+Hence,
+$\sum_{t=1}^{T}V(\pi(\mathbf{v});\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\sum_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}\mathbb{P}[\mathbf{b}^{t}=\mathbf{b}|\mathbf{v}^{t}=\mathbf{v}]V(\mathbf{b};\bm{\beta}_{-}^{t})$
+$\leq\frac{m\log M}{\eta}+\eta\sum_{t=1}^{T}\sum_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}\mathbb{P}[\mathbf{b}^{t}=\mathbf{b}|\mathbf{v}^{t}=\mathbf{v}]V(\mathbf{b};\bm{\beta}_{-}^{t})^{2}$ (69)
+Taking expectations with respect to the randomness of the context (valuation vectors) and using the fact that the choice of $\pi$ was arbitrary, we get:
+$\textsf{REG}_{sto}$ $=\max_{\pi\in\Pi}\sum_{t=1}^{T}\mathbb{E}_{\mathbf{v}_{t}\sim F_{\mathbf{v}}}[V(\pi(\mathbf{v}^{t});\bm{\beta}_{-}^{t})]-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]$
+[…]
+$\leq\frac{m\log M}{\eta}+\eta M^{2}\sum_{t=1}^{T}\sum_{\mathbf{v}}\mathbb{P}[\mathbf{v}^{t}=\mathbf{v}]\sum_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}\mathbb{P}[\mathbf{b}^{t}=\mathbf{b}|\mathbf{v}^{t}=\mathbf{v}]=\frac{m\log M}{\eta}+\eta M^{2}T,$
+where the second inequality follows as $V(\mathbf{b};\bm{\beta}_{-}^{t})\leq M$. Setting, $\eta=\frac{1}{M}\sqrt{\frac{m\log M}{T}}$, we get
+$\textsf{REG}_{sto}$ $\leq O(M\sqrt{mT\log M})\,.$
+E.6. Adversarial Contexts
+In this section, we consider the case when the contexts (valuation vectors) are generated adversarially in advance (akin to an oblivious adversary). The regret in this setting is defined as
+$\textsf{REG}_{adv}=\max_{\pi^{*}\in\Pi}\sum_{t=1}^{T}V(\pi^{*}({\mathbf{v}^{t}});\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})],$
+where $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}(\mathbf{v}^{t}),\forall t\in[T]$ and $\Pi$ is defined in Eq.
+25.
+Observe that for any $\pi$, $\sum_{t=1}^{T}V(\pi({\mathbf{v}^{t}});\bm{\beta}_{-}^{t})=\sum_{\mathbf{v}\in\mathcal{V}}\sum_{t=1}^{T}V(\pi({\mathbf{v}});\bm{\beta}_{-}^{t})\cdot\mathbb{I}{\left[\mathbf{v}^{t}=\mathbf{v}\right]}$.
+Hence, the optimal stationary policy in this setting is
+$\pi^{*}(\mathbf{v})=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\mathbf{v})}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})\cdot\mathbb{I}{\left[\mathbf{v}^{t}=\mathbf{v}\right]},\quad\forall\mathbf{v}\in\mathcal{V}.$ (70)
+This benchmark is the natural choice for adversarial rewards and contexts (Lattimore and Szepesvári, 2020, Chapter 18.1).
+Note that the optimal policy $\pi^{*}(\cdot)$ differs between the stochastic and adversarial settings.
+In the stochastic case, the benchmark assumes that the clairvoyant knows the competing bids and distribution over the contexts a priori but does not observe the realized contexts in each round in advance.
+Hence, their goal is to maximize the expected total value over the $T$ rounds ex ante.
+In contrast, under adversarial contexts, the bidder is assumed to know both the contexts and competing bids a priori, and thus aims to maximize the total value ex post over the $T$ rounds.
+In this setting, the bidder maintains $|\mathcal{V}|$ copies of Algorithm 1, each corresponding to a context in $\mathcal{V}$.
+In each round $t\in[T]$, after observing $\mathbf{v}^{t}$, the bidder updates the edge probabilities only in the algorithm corresponding to $\mathbf{v}^{t}$ and submits a bidding strategy in $\mathscr{U}_{m}^{\star}(\mathbf{v}^{t})$.
+The space complexity of Algorithm 5 is $O(mM^{2}|\mathcal{V}|)$ while the per-round time complexity is $O(mM^{2})$.
+Algorithm 5 Learning Safe Bidding Strategies (Full Information, Adversarial Contexts)
+1:Initialize $|\mathcal{V}|$ copies of Algorithm 1, each corresponding to a context in $\mathcal{V}$.
+2:for $t=1,2,\dots,T$ do
+3:  Observe valuation curve $\mathbf{v}^{t}\in\mathcal{V}$.
+4:  Obtain bidding strategy and update edge probabilities by running Algorithm 1 corresponding to the context $\mathbf{v}^{t}$.
+5:end for
+Observe that for stochastic as well as adversarial contexts, a separate DAG is maintained for each context.
+However, in any given round $t$, while the edge probabilities in each DAG are updated under stochastic contexts the edge probabilities in only the DAG corresponding to the context $\mathbf{v}^{t}$ are updated for adversarial contexts.
+Theorem E.2.
+When the contexts are adversarial, there exists an algorithm that runs in $\text{poly}(m,M)$ space and per-round time and achieves $\textsf{REG}_{adv}\leq O(M\sqrt{m|\mathcal{V}|T\log M})$ under full information feedback.
+Proof.
+Recall that when the contexts are adversarial, regret is defined as
+$\textsf{REG}_{adv}=\sum_{t=1}^{T}V(\pi^{*}({\mathbf{v}^{t}});\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]$
+where $\pi^{*}(\cdot)$ is defined in Eq. 70. Let $T_{\mathbf{v}}$ be the set of rounds in which the context is $\mathbf{v}\in\mathcal{V}$. Then,
+$\textsf{REG}_{adv}$ $=\sum_{\mathbf{v}\in\mathcal{V}}\left\{\sum_{t\in T_{\mathbf{v}}}V(\pi^{*}({\mathbf{v}});\bm{\beta}_{-}^{t})-\sum_{t\in T_{\mathbf{v}}}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\right\}$
+[…]
+$\leq\sum_{\mathbf{v}\in\mathcal{V}}O(M\sqrt{m|T_{\mathbf{v}}|\log M})\,$,
+where the last inequality follows as a separate instance of Algorithm 1 is utilized for each context and using the regret upper bound of Algorithm 1 in the full information setting.
+[^24] Although $T_{\mathbf{v}}$ is not known in advance (which is required for tuning the learning rate), the bidder can use the standard doubling trick (Cesa-Bianchi and Lugosi, 2006) to implement Algorithm 1 which yields the same regret guarantee up to constant factors.
+Finally, by Jensen’s inequality, we get that $\textsf{REG}_{adv}\leq O(M\sqrt{m|\mathcal{V}|T\log M})$.
+∎
+We complement the regret upper bound for adversarial contexts with the following lower bound:
+Theorem E.3.
+Let $M\geq 2$, $m=1$ and $|\mathcal{V}|<\infty$.
+There exist competing bids, $\bm{\beta}_{-}^{t}]_{t\in[T]}$, such that for adversarial contexts, under any learning algorithm, $\mathbb{E}[\textsf{REG}_{adv}]=\Omega(M\sqrt{|\mathcal{V}|T})$ in the full information setting.
+Proof.
+We begin by presenting a corollary of Theorem 4.5:
+Corollary E.4.
+Fix any $\theta>0$ and assume that $v_{j}\in[0,\theta],\forall j\in[M]$.
+Then, for $M\geq 2$ and $m=1$, there exist competing bids, $\bm{\beta}_{-}^{t}]_{t\in[T]}$, such that, under any learning algorithm, $\mathbb{E}[\textsf{REG}]\geq c\theta M\sqrt{T}$ for some $c>0$, in the full information setting where REG is defined in Eq.
+4.
+Proof of Corollary E.4.
+The result follows directly by considering the scaled valuation vector $\theta\mathbf{v}$ (as defined in Eq.
+33), the scaled competing bid profiles $\theta\bm{\beta}_{-}^{\clubsuit}$ and $\theta\bm{\beta}_{-}^{\diamondsuit}$ (as defined in Eq.
+34), and applying the analysis of Theorem 4.5.
+∎
+Fix any $N\in\mathbb{N}$ and define $\mathcal{V}=\{\alpha_{i}\mathbf{v}:\alpha_{i}=\frac{i}{N},~~\text{for}~~i\in[N]\}$ where $\mathbf{v}$ is defined in Eq.
+33.
+Note that $\mathbf{v}^{\prime}\in[0,1]^{M}$ for all $\mathbf{v}^{\prime}\in\mathcal{V}$.
+Assume that $T$ is a multiple of $N$.
+Partition the $T$ rounds into $N$ epochs with the $i^{th}$ epoch spanning the rounds $t=\frac{(i-1)T}{N}+1,\dots,\frac{iT}{N}$.
+For all the rounds in the $i^{th}$ epoch, set the context as $\alpha_{i}\mathbf{v}$.
+Then, for some $c^{\prime}>0$,
+$\mathbb{E}[\textsf{REG}_{adv}]=\sum_{i=1}^{N}\mathbb{E}\left[\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}(\alpha_{i}\mathbf{v})}\sum_{t=\frac{(i-1)T}{N}+1}^{\frac{iT}{N}}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=\frac{(i-1)T}{N}+1}^{\frac{iT}{N}}V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})\right]\geq c^{\prime}M\sqrt{\frac{T}{N}}\sum_{i=1}^{N}\alpha_{i}\,.$
+The inequality follows from Corollary E.4.
+As $\sum_{i=1}^{N}\alpha_{i}\geq N/2$, $\mathbb{E}[\textsf{REG}_{adv}]\geq\frac{1}{2}c^{\prime}MN\sqrt{\frac{T}{N}}=\Omega(M\sqrt{|\mathcal{V}|T})$ where the last equality holds as $|\mathcal{V}|=N$.
+∎
+Appendix F Simulation Details
 We sample the values from the $\text{Unif}[0,1]$ distribution.
@@
 4.
-As computing $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ (the value obtained by the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m}$) can be non-trivial, we obtain a uniform upper bound for $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ that is independent of $m$ (see details in Section E.2).
-E.1.  Reconstructing Individual Bid Data
+As computing $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ (the value obtained by the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m}$) can be non-trivial, we obtain a uniform upper bound for $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ that is independent of $m$ (see details in Section F.2).
+F.1. Reconstructing Individual Bid Data
 We obtained the publicly available auction data for $T_{\max}=443$ EU ETS emission permit auctions held in 2022 and 2023 (EEX, 2023).
@@
 We normalized the bids to be in $[0,1]$.
-For all rounds $t$, $b_{\text{avg}}^{t}\approx b_{\text{med}}^{t}$ (linear regression yields coefficient 1.01 and intercept $-0.008$).
+For all rounds $t$, $b_{\text{avg}}^{t}\approx b_{\text{med}}^{t}$ (linear regression yields coefficient 1.01 and intercept 0.008-0.008).
 Upon further investigation, we observed that, except a few, a significant number of auctions had either $b_{\min}^{t}\approx b_{\text{avg}}^{t}\ll b_{\max}^{t}$ (Type I) or $b_{\min}^{t}\ll b_{\text{avg}}^{t}\approx b_{\max}^{t}$ (Type II).
-As $b_{\text{avg}}^{t}\approx b_{\text{med}}^{t},\forall t$, we deduce that for Type I, most of the bids are concentrated in the interval $[b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\min}^{t}]$ whereas for Type II, most of the bids are in the interval $[2b_{\text{avg}}^{t}-b_{\max}^{t},b_{\max}^{t}]$.
+As $b_{\text{avg}}^{t}\approx b_{\text{med}}^{t},\forall t$, we deduce that for Type I, most of the bids are concentrated in the interval $b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\min}^{t}]$ whereas for Type II, most of the bids are in the interval $2b_{\text{avg}}^{t}-b_{\max}^{t},b_{\max}^{t}]$.
 We posit that for Type I (resp.
-Type II) auctions, $f\in(0,1)$ fraction of the bids ($n_{\text{sub}}^{t}$) are in $[b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\min}^{t}]$ (resp.
-$[2b_{\text{avg}}^{t}-b_{\max}^{t},b_{\max}^{t}]$) and the $1-f$ fraction of bids are in $[2b_{\text{avg}}^{t}-b_{\min}^{t},b_{\max}^{t}]$ (resp.
-$[b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\max}^{t}]$).
+Type II) auctions, $f\in(0,1)$ fraction of the bids ($n_{\text{sub}}^{t}$) are in $b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\min}^{t}]$ (resp.
+$2b_{\text{avg}}^{t}-b_{\max}^{t},b_{\max}^{t}]$) and the 1f1-f fraction of bids are in $2b_{\text{avg}}^{t}-b_{\min}^{t},b_{\max}^{t}]$ (resp.
+$b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\max}^{t}]$).
 If for Type I (resp.
 Type II) auctions, $2b_{\text{avg}}^{t}>b_{\min}^{t}+b_{\max}^{t}$ (resp.
-$2b_{\text{avg}}^{t}<b_{\min}^{t}+b_{\max}^{t}$), we assume that all the bids are uniformly present in the interval $[b_{\min}^{t},b_{\max}^{t}]$.
+$2b_{\text{avg}}^{t}<b_{\min}^{t}+b_{\max}^{t}$), we assume that all the bids are uniformly present in the interval $b_{\min}^{t},b_{\max}^{t}]$.
 With these assumptions, we generate individual bid data for each auction by sampling uniformly from these intervals.
@@
 The bids are normalized to be in $[0,1]$.
-E.2.  An Uniform Upper Bound for $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right).$
+F.2. An Uniform Upper Bound for $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right).$
 For any bid history, $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, suppose $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated $r_{t}$ units in any round $t$.
@@
 $\implies V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ $=\sum_{t=1}^{T}V(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})\leq\sum_{t=1}^{T}\max_{\mathbf{b}\in\mathscr{U}_{1}^{\star}}V(\mathbf{b};\bm{\beta}_{-}^{t})\,.$
-Appendix F Resolving Ties
+Appendix G Resolving Ties
 We first emphasize that the effect of resolution of ties on the objective function under the value maximization behavioral model is fundamentally different from the quasilinear utility maximization model.
-To illustrate this, consider a single-item second price auction with two bidders each valuing the item equally at $v$.
+To illustrate this, consider a second price auction with a single indivisible item with two bidders each valuing the item equally at $v$.
 Assume that both the bidders bid $v$ and there exists a definitive tie breaking rule (either randomized or deterministic).
@@
 [^25] Although, the total (liquid) welfare is $v$ in both the cases.
-Hence, we need to carefully analyze the key results of our work in the event of ties.
+So, we need to carefully analyze the key results of our work in the event of ties.
 In the context of multi-unit auctions, we can classify ties into two informal types: (a) ‘good ties’—ties occurring at any bid other than the last accepted bid (LAB) and (b) ‘bad ties’, which are the ties occurring at the LAB.
 All the results in this work are unaffected in case only ‘good ties’ occur.
-Thus, we focus on ‘bad ties’ for the rest of this section.
+Thus, we focus on ‘bad ties’ in this paper.
 For this discussion, we consider a public, deterministic tie breaking rule under which ties are always broken in favor of lower indexed bidder.